### v2

# Avoiding Catastrophe in Continuous Spaces by Asking for Help

Abstract
Most reinforcement learning algorithms with formal regret guarantees assume all mistakes are reversible and essentially rely on trying all possible behaviors. This approach leads to poor outcomes when some mistakes are irreparable or even catastrophic. We propose a variant of the contextual bandit problem where the goal is to minimize the chance of catastrophe. Specifically, we assume that the payoff each round represents the chance of avoiding catastrophe that round, and try to maximize the product of payoffs (the overall chance of avoiding catastrophe). We allow a limited number of queries to a mentor and assume a Lipschitz continuous payoff function. We first show that in general, any algorithm either constantly queries the mentor or is nearly guaranteed to cause catastrophe. However, when the mentor policy class has bounded Natarajan dimension and contains at least some “reasonable” policies, we provide an algorithm whose regret and rate of querying the mentor both approach 0 as the time horizon grows. We also present an alternative algorithm which provides the same regret and query guarantees when the mentor’s action changes a constant number of times in a 1D state space, and can handle adversarially chosen states.
1 Introduction
In reinforcement learning (RL), an agent finds itself in an unknown environment and wants to maximize some payoff function, commonly modeled as a Markov Decision Process (MDP). Typically, the only way for the agent to learn about its environment is by trying actions and observing their effects. This presents a challenge when some actions are inherently dangerous: for example, the agent must learn to avoid self-destructive actions without actually trying those actions.
This problem is hopeless without additional assumptions, since any untried action could lead to either paradise or disaster, and the agent has no way to predict which. Most RL theory evades this issue by either assuming ergodicity, meaning that any state is reachable from any other (e.g., [15]), or by resetting the agent at the start of each “episode” (e.g., [1]). In the real world, however, many actions lead to unrecoverable situations: we call these catastrophes. Examples could include ingesting poison or crashing a robotic vehicle. Resetting the agent may be expensive (if it destroys itself), inadequate (if it irreparably harms others), or impossible (if it is outside of human supervision).
Rather than ignore the possibility of catastrophe, we explore whether access to a mentor might enable an agent to navigate risky (i.e., non-ergodic single-episode) environments. We assume the mentor illustrates the optimal action in the current state, and we would like an algorithm which obtains near-optimal payoff with a limited number of mentor queries. “Near-optimal payoff” is often modeled via regret, defined as the gap between the optimal total payoff and the agent’s actual total payoff. For undiscounted MDPs (i.e., each time step has the same weight in the objective), one typically aims for regret sublinear in the time horizon $T$, or equivalently, for the average regret per time step to approach 0 for large $T$. Similarly, a sublinear number of mentor queries implies that the rate of querying the mentor goes to 0 for large $T$—that is, the agent gradually becomes self-sufficient. Formally, we pose the following open problem:
Is there an algorithm for non-ergodic single-episode undiscounted MDPs which ensures that both the regret and the number of mentor queries are sublinear in $T$?
An additional assumption is necessary, as the agent needs some way to make inferences about unseen states in order to decide when to query. Much past work has used Bayesian inference, which suffers tractability issues in complex environments. [^1] For the curious reader, [3] provides a thorough treatment. See also Section 1.3 for further discussion. In this paper, we use Lipschitz continuity instead, which captures the intuition that an action should have similar effects in similar states. Importantly, the Lipschitz assumption is compatible with any state space and action space with a distance metric.
1.1 Our model
In this paper, we target a more modest goal: can the agent simply avoid catastrophe? This is clearly necessary (but not sufficient) for sublinear regret in general MDPs. Furthermore, avoiding catastrophe is the precise ability which standard RL algorithms lack.
Focusing on catastrophe also allows us to streamline the model. If we redefine the payoff at time $t$ to be the probability of avoiding catastrophe on that time step, then the product of payoffs represents the overall chance of avoiding catastrophe. Representing catastrophe in the objective function eliminates the need to model non-ergodicity through the transition function, as MDPs do. Our multiplicative objective captures the same stringency: a single mistake can ruin the entire endeavor. [^2] An additive objective with unbounded payoffs could also capture the core challenge of needing to avoid catastrophic actions without ever trying them. However, we feel that the multiplicative objective is more intuitive due its interpretation as the probability of avoiding catastrophe.
The resulting model is similar to the contextual bandit problem: on each time step, the agent observes a state (we generally assume i.i.d. states), selects an action (or queries the mentor), and obtains some payoff. To our knowledge, multiplicative objective functions have not been studied for contextual bandits. Lastly, unlike the standard contextual bandit model, we assume that the agent does not observe payoffs. This is because the payoff in our model represents the chance of avoiding catastrophe that round, and in the real world, one only observes whether catastrophe occurred, not its probability. [^3] One may be able to detect “close calls” in some cases, but observing the precise probability seems unrealistic.
1.2 Our results
We first show that in general, any algorithm with sublinear queries to the mentor has arbitrarily poor regret in the worst-case (Theorem 3.1). As a corollary, even when it is possible to avoid catastrophe with certainty, any algorithm either asks for help on a constant fraction of time steps or is nearly guaranteed to cause catastrophe. This holds even for randomized algorithms and simple environments: our construction uses just two actions and assumes states are uniformly random across $[0,1]$.
On the other hand, the class of mentor policies in our negative result is quite complex in the sense of having unbounded VC dimension [31]. What if we restrict the mentor policy to be relatively simple?
We present an intuitive algorithm whose regret and rate of querying the mentor both go to 0 as $T\to\infty$ when the mentor policy class has bounded VC dimension (or Natarajan dimension, more generally [^4] Natarajan dimension [24] generalizes VC dimension (which only handles binary prediction) to multi-class prediction. In our case, each class is an action, and a policy is a multi-class predictor from states to actions.) and contains at least some “reasonable” policies (Theorem 4.5). Our algorithm can handle an arbitrary unknown distribution over states, a multi-dimensional unbounded state space, and many actions. It does not need detailed access to the feature embedding (instead relying on two simple operations) and does not need to know the Lipschitz constant. It does need to know the policy class and (an upper bound on) the Natarajan dimension (as is standard in PAC learning). This is our primary positive result.
Finally, we give another algorithm with the same guarantees of sublinear queries and near-zero regret under a different simplicity assumption (Theorem 5.2). Specifically, we assume that the mentor’s action changes a constant number of times in a 1D state space. This algorithm can also handle adversarially chosen states.
We generally assume that the mentor’s policy is optimal, but our positive results can be generalized to a suboptimal mentor under a slightly different Lipschitz assumption (Remark B.1).
1.3 Related work
See [9; 11] for full surveys of safe RL. Here, we discuss only the work most relevant to our own.
Learning with irreversible costs.
Despite the ubiquity of irreversible costs in the real world, there remains a dearth of theoretical work on this topic. This may be due to the fundamental modeling question of how the agent should learn about novel states or actions without actually trying them.
The most common approach is to allow the agent to ask for help. This alone is insufficient, however: the agent must have some way to decide when to ask for help. A popular solution is to perform Bayesian inference on the world model, but this has two tricky requirements: (1) a prior distribution which contains the true world model (or an approximation), and (2) an environment where computing (or approximating) the posterior is tractable. A finite set of possible environments satisfies both conditions, but is unrealistic in many real-world scenarios. In contrast, our algorithm can handle a continuous unbounded state space, which is crucial for many real-world scenarios in which one never sees the exact same state twice.
Bayesian inference combined with asking for help is studied by [5; 6; 17; 21]. We also mention [12; 22; 30], which utilize Bayesian inference in the context of safe RL without asking for help (and without regret bounds).
We are only aware of two papers which theoretically address irreversibility without Bayesian inference: [10] and [20]. The former proposes to sample trajectories and learn reversibility based on temporal consistency between states: intuitively, if $s_{1}$ always precedes $s_{2}$, we can infer that $s_{1}$ is unreachable from $s_{2}$. Although the paper theoretically grounds this intuition, there is no formal regret guarantee. The latter presents an algorithm which asks for help in the form of rollouts from the current state. However, the regret bound and number of rollouts are both linear in the worst case, due to the dependence on the $\gamma^{*}$ parameter which roughly captures how bad an irreversible action can be. In contrast, our algorithms achieve near-zero regret even when actions can be maximally bad.
To our knowledge, we are the first to provide an algorithm which formally guarantees avoidance of catastrophe (with high probability) without Bayesian inference.
Contextual bandits.
Like MDPs, the contextual bandit model involves an agent who selects an action on each time step based on the current state (or “context”) to maximize payoff. Unlike MDPs, the state on each time step is determined randomly or by an adversary, instead of by a transition function. See [29] for an overview.
The contextual bandit literature often uses Lipschitz continuity of payoff functions as a way to generalize between states, which affirms our use of that same assumption [4; 14; 18; 19; 26; 28]. However, to our knowledge, all prior contextual bandit work aims to maximize the sum of payoffs, which implicitly assumes that no mistake is catastrophic: no matter how poorly you perform, you can always make it up given a sufficiently long time horizon.
Multiplicative objectives.
Perhaps the best-studied example of a multiplicative objective is the Nash welfare function (also known as the Nash bargaining solution), defined as the geometric mean of agent utilities [16; 23]. Inspired by Nash welfare, [2] recently studied algorithms for maximizing the geometric mean of payoffs across time steps in the classic (non-contextual) multi-armed bandit problem. However, the geometric mean of payoffs is much more forgiving of errors compared to the product of payoffs. Imagine that the agent obtains a payoff of $1-1/\sqrt{T}$ per time step: then $\lim_{T\to\infty}(1-1/\sqrt{T})^{T}=0$ but $\lim_{T\to\infty}((1-1/\sqrt{T})^{T})^{1/T}=\lim_{T\to\infty}(1-1/\sqrt{T})=1$. In our setting, it is crucial that the cumulative chance of catastrophe (the product of payoffs) is small, not just the average chance of catastrophe per time step (the geometric mean of payoffs).
Multiplicative objectives also arise in the context of compound growth or repeated gambles, where one’s wealth at time $t$ can be modeled as the initial wealth times a product of growth factors (see, e.g., [25]). Our problem could be viewed as a compound growth problem where one’s “wealth” (the total probability of avoiding catastrophe) never increases. However, we are not aware of any mathematical tools related to compound growth that handle the contextual bandit aspects of our problem.
The rest of the paper is structured as follows. Section 2 formally defines our model. Section 3 presents our negative result for general mentor policies. Section 4 presents our positive result for policy classes with bounded Natarajan dimension. Section 5 presents our second positive result under a different simplicity assumption. Section 6 discusses limitations. All proofs are deferred to the appendix.
2 Model
Lipschitz contextual bandits.
Let $\mathbb{N}$ refer to the strictly positive integers and let $T\in\mathbb{N}$ be the time horizon. Let $S\subseteq\mathbb{R}^{n}$ be the state space, and let $\mathbf{s}=(s_{1},s_{2},\dots s_{T})\in S^{T}$ be a sequence of states. (The contextual bandit literature typically uses “context” instead of “state”, but we have chosen to use “state” to align with MDP terminology.) With the exception of Section 5, we assume that each state is sampled i.i.d. from an unknown distribution $\mathcal{D}$ over $S$.
For $k\in\mathbb{N}$, let $[k]=\{1,2,\dots,k\}$. On each time step $t\in[T]$, the agent must select an action $a_{t}\in A\cup\{\hat{a}\}$, which generates payoff $\mu(s_{t},a_{t})\in[0,1]$. We allow the agent’s choice of action to be randomized. The action $\hat{a}$ constitutes querying the mentor, which we will define shortly. The agent never observes payoffs: the only way to learn about $\mu$ is by querying the mentor. We assume that $|A|\leq T$, since the agent can never observe more than $T$ distinct actions.
We assume that $\mu$ is Lipschitz continuous with respect to $S$: there exists $L>0$ such $\forall s,s^{\prime}\in S,a\in A$, $|\mu(s,a)-\mu(s^{\prime},a)|\leq L||s-s^{\prime}||$, where |||||$\cdot$||| | ⋅ | | denotes Euclidean distance.
Asking for help.
The special action $\hat{a}$ constitutes querying a mentor who knows an optimal policy $\pi^{*}:S\to A$. When action $\hat{a}$ is chosen, the mentor informs the agent of the action $\pi^{*}(s_{t})$ and the agent obtains optimal payoff $\mu(s_{t},\pi^{*}(s_{t}))=\mu^{*}(s_{t})=\max_{a\in A}\mu(s_{t},a)$. Let $Q_{t}=\{i\in[t]:a_{i}=\hat{a}\}$ be the set of time steps up to and including $t$ which involve a query to the mentor.
We would like an algorithm which becomes “self-sufficient” over time: the rate of querying the mentor should go to 0 as $T\to\infty$. Intuitively, this is equivalent to the cumulative number of queries being sublinear in $T$: $|Q_{T}|\in o(T)$. However, interpreted literally, $|Q_{T}|\in o(T)$ is not a rigorous definition. This is because $|Q_{T}|$ is not just a function of $T$: it can also depend on $\mathbf{s}$, $\mu$ and $\pi^{*}$.
To formalize this, let $|Q_{T}(\mathbf{s},\mu,\pi^{*})|$ be the number of queries to the mentor up to time $T$, given payoff function $\mu$ and state sequence $\mathbf{s}$. Then we say that $|Q_{T}|$ is sublinear in $T$, denoted $|Q_{T}|\in o(T)$, if there exists $g:\mathbb{N}\to\mathbb{N}$ such that $g(T)\in o(T)$ and $|Q_{T}(\mathbf{s},\mu,\pi^{*})|\leq g(T)$ for all $\mathbf{s},\mu,\pi^{*}$.
Multiplicative objective and regret.
If $\mu(s_{t},a_{t})\in[0,1]$ is the chance of avoiding catastrophe on time step $t$ (conditioned on no prior catastrophe), then $\prod_{t=1}^{T}\mu(s_{t},a_{t})$ represents the agent’s overall chance of avoiding catastrophe. [^5] Conditioning on no prior catastrophe means we do not need to assume that these probabilities are independent (and if catastrophe has already occurred, this time step does not matter). This is due to the chain rule of probability. For a time horizon $T$, payoff function $\mu$, mentor policy $\pi^{*}$, and state sequence $\mathbf{s}$, the agent’s regret is
$R_{T}(\mathbf{s},\mu,\pi^{*})=\prod_{t=1}^{T}\mu^{*}(s_{t})-\prod_{t=1}^{T}\mu(s_{t},a_{t})$
where $\mathbf{a}=\{a_{1},\dots,a_{T}\}$ is the set of actions selected by the algorithm (which depends on $\pi^{*}$). We focus on a worst-case choice of $\mu$ and $\pi^{*}$. [^6] The choice of $\pi^{*}$ can actually matter: suppose $a$ is optimal everywhere but $a^{\prime}$ is optimal only for this state.
VC and Natarajan dimension.
VC dimension [31] and Natarajan dimension [24] are standard measures of learning difficulty. The former applies to binary classes, while the latter handles multi-class prediction (and reduces to the former in the case of two classes). In our setting, a policy can be viewed as a multi-class predictor, where each action is a class. The precise definitions of these dimensions are not relevant to our paper, as we only invoke Natarajan dimension indirectly through the fundamental theorem of multi-class PAC learning (Theorem 4.4). We direct the reader to [27] for a comprehensive overview.
Misc.
We will often deal with the measure of subsets of $S$. When $n=1$, we use $\operatorname*{len}(X)$ to denote the length of interval $X\subset S$; for general $n$, we use $\operatorname*{vol}(X)$ to denote the $n$-dimensional measure of $X\subseteq S$ [^7] Although we could use the same notation for both cases, we find this to be more intuitive. The convex hull of a set $X$ is denoted $\operatorname*{conv}(X)$. All logarithms and exponents are base $e$ unless otherwise noted.
3 Hardness result for complex mentor policies
We begin by showing that in general, any algorithm with sublinear mentor queries has arbitrary poor regret in the worst-case, even when states are uniformly random on $[0,1]$ and there are only two actions. The result also holds even if the algorithm knows $L$ and $\mathbf{s}$ ahead of time.
Let $U(S)$ denote the uniform distribution on $S$ and let $\operatorname*{\mathbb{E}}_{\mathbf{s}\sim U(S^{T}),\mathbf{a}}$ denote the expectation over both $\mathbf{s}$ and any additional randomness in the algorithm.
Theorem 3.1.
The worst-case expected regret of any algorithm with sublinear queries goes to 1 as $T$ goes to infinity. Formally,
$\lim_{T\to\infty}\ \,\sup_{\mu,\pi^{*}}\ \,\operatorname*{\mathbb{E}}_{\mathbf{s}\sim U(S^{T}),\,\mathbf{a}}\ \,[R_{T}(\mathbf{s},\mu,\pi^{*})]=1$
3.1 Intuition
We partition $S$ into equally-sized sections that are “independent” in the sense that querying a state in section $i$ gives you no information about section $j$. The number of sections is determined by a function $f:\mathbb{N}\to\mathbb{N}$ that we will choose. If $|Q_{T}|\in o(f(T))$, most of these sections will never contain a query. When the agent sees a state in a section not containing a query, it essentially has to guess, meaning it will be wrong a constant fraction of the time. Figure 1 fleshes out this idea.
Picking $\boldsymbol{f(T)}$.
A natural idea is to try $f(T)=T$. Unfortunately, this doesn’t quite work: even if the agent chooses wrong on every time step, the minimum payoff is still at least $1-\frac{L}{2T}$, and $\lim_{T\to\infty}\prod_{t=1}^{T}\left(1-\frac{L}{2T}\right)=\lim_{T\to\infty}\left(1-\frac{L}{2T}\right)^{T}=e^{-L/2}$. In order for the regret to approach 1, we need $f(T)$ to be asymptotically between $|Q_{T}|$ and $T$. Lemma A.1 implies the existence of such an $f(T)$ if $|Q_{T}|$ is bounded by some function $g(T)\in o(T)$, so we get a bound of the following form: $\prod_{t=1}^{T}\mu_{f,\mathbf{x}}(s_{t},a_{t})\leq\big{(}1-\frac{L}{\Theta(f(T))}\big{)}^{\Theta(T)}$. [^8] The bound in the actual proof will be in log space, so it will look slightly different, but the idea is the same. When $f(T)\in o(T)$, the right hand side converges to 0, while $\prod_{t=1}^{T}\mu_{f,\mathbf{x}}^{*}(s_{t})=1$. In words, the agent is nearly guaranteed to cause catastrophe, despite the existence of a policy which is guaranteed to avoid catastrophe.
$s$$\mu$  $\mu(s,0)$$\mu(s,1)$$\cfrac{L}{2f(T)}$$\frac{1}{f(T)}$$\frac{2}{f(T)}$$\frac{3}{f(T)}$$\frac{f(T)-1}{f(T)}$$\dots$1001111
Figure 1: An illustration of the construction we use to prove Theorem 3.1 (not to scale). The horizontal axis indicates the state $s\in[0,1]$ and the vertical axis indicates the payoff $\mu(s,a)\in[0,1]$. The solid line represents $\mu(s,0)$ and the dotted line represents $\mu(s,1)$. In each section, one of the actions has the optimal payoff of 1, and the other action has the worst possible payoff allowed by the Lipschitz constraint, reaching a minimum of $1-\frac{L}{2f(T)}$. Crucially, both actions result in a payoff of 1 at the boundaries between sections: this allows us to “reset” for the next section. As a result, we can freely toggle the optimal action for each section independently.
VC/Natarajan dimension.
The class of mentor policies induced by our construction has VC and Natarajan dimensions of $f(T)$. [^9] Recall that VC and Natarajan dimensions coincide for $|A|=2$. This is expected given our positive result for policy classes with bounded Natarajan dimension (Theorem 4.5).
3.2 Formal definition of construction
Let $S=[0,1]$, $A=\{0,1\}$, and $\mathcal{D}=U(S)$. Assume that $L\leq 1$; this will simplify the math and only makes the problem easier for the agent. We define a family of payoff functions parameterized by a function $f:\mathbb{N}\to\mathbb{N}$ and a bit string $\mathbf{x}=(x_{1},x_{2},\dots,x_{f(T)})\in\{0,1\}^{f(T)}$. The bit $x_{j}$ will denote the optimal action in section $j$. Note that $f(T)\geq 1$ and since we defined $\mathbb{N}$ to exclude 0.
[…]
$\mu_{f,\mathbf{x}}(s,a)=\begin{cases}1&\textnormal{ if }a=x_{j(s)}\\ 1-L\left(\cfrac{1}{2f(T)}-|m_{j(s)}-s|\right)&\textnormal{ if }a\neq x_{j(s)}\\ \end{cases}$
For any $\mathbf{x}\in\{0,1\}^{f(T)}$, $\mu_{f,\mathbf{x}}$ is piecewise linear (trivially) and continuous (because both actions have payoff 1 on the boundary between sections). Since the slope of each piece is in $\{-L,0,L\}$, $\mu_{f,\mathbf{x}}$ is Lipschitz continuous with Lipschitz constant $L$.
Since each $s_{t}$ belongs to exactly one $S_{j}$, there is a unique optimal action for each $s_{t}$, so the choice of $\pi^{*}$ does not matter in this scenario.
4 Near-zero regret assuming bounded Natarajan dimension
In this section, we present a simple algorithm with near-zero regret when the policy class $\Pi$ has bounded Natarajan dimension $d$ and which contains at least some “reasonable” policies. The algorithm needs to know $\Pi$ and (an upper bound on) $d$, as is standard in PAC learning (see, e.g., [27]). The algorithm does not need to know $L$, and can handle an unbounded state space (the number of queries simply scales with the volume of the convex hull of the observed states).
We need a few definitions to formally state our algorithm and result.
Definition 4.1.
A policy $\pi$ is consistent with $X\subset S\times A$ if $\pi(s)=a$ for all $(s,a)\in X$.
Definition 4.2.
Given $X\subset S\times A$ and $r\geq 0$, a policy $\pi$ is $(X,r)$-local if (1) $\pi$ is consistent with $X$ and (2) whenever $(s^{\prime},a^{\prime})\in X$ and $||s-s^{\prime}||\leq r$, there exists $(s^{\prime\prime},a^{\prime\prime})\in X$ such that $||s-s^{\prime\prime}||\leq r$ and $\pi(s)=a^{\prime\prime}$.
A policy which knows $X$ but is not $(X,r)$-local says either (1) “even though I know the optimal action for $s$, I’m going to do something else” or (2) “even though I have a data point near $s$, I’m going to base my action on a data point far from $s$”. [^10] Note that the mentor policy cannot itself query the mentor, so we are only dealing with actions in $A$. Given the Lipschitz nature of the problem, this type of policy seems quite unintuitive. However, we allow $\Pi$ to contain non-local policies, as long as it also contains local policies. Note that knowing $X,r$, and $\Pi$ is sufficient to find a $(X,r)$-local policy.
Definition 4.3.
Given $r\geq 0$, $\Pi$ is $r$-local if for every $\tilde{S}\subset S$ and $X=\{(s,\pi^{*}(s)):s\in\tilde{S}\}$, there exists an $(X,r)$-local policy in $\Pi$.
As an example, let $S=\mathbb{R}$ and suppose $\Pi$ is the set of policies where $\pi(s)$ changes at most $K$ times as $s$ goes from $\infty$- ∞ to +$\infty$+ ∞ (in the language of Section 5, $\pi$ has at most $K$ segments). Then $\Pi$ is $r$-local for any $r$, since for any set of queries $X$, the nearest neighbor policy is $(X,r)$-local and has at most $K$ segments (and thus is in $\Pi$). Note that this $\Pi$ also has bounded Natarajan dimension.
4.1 Intuition behind the algorithm
Our algorithm has a simple structure: gather enough data to choose a solid baseline policy, and then mostly follow that policy, but ask for help if you find yourself out-of-distribution (OOD). The algorithm does not need direct access to the feature embedding of states: the only two operations it uses are (1) update baseline policy and (2) find nearest neighbor distance. This approach seems quite natural for real-world applications, where one can imagine updating the baseline policy via gradient descent and treating an OOD detector score as the nearest neighbor distance.
In our case, the initial dataset $X$ is gathered by querying on the first $g(T)$ time steps. We will choose $g(T)$ to be large enough so that – along with the assumption on Natarajan dimension – we can apply the fundamental theorem of multi-class PAC learning (Theorem 4.4). [^11] Theorem 4.4 is called the “realizable” case of PAC learning, in contrast to the “agnostic” case where $\pi^{*}$ may not be in $\Pi$. Some sources (e.g., [27]) state that $\pi$ should minimize empirical loss, rather than being consistent with $X$. These are equivalent in the realizable case since consistent policies have zero empirical loss. This is the only place in the proof where Natarajan dimension is used.
Theorem 4.4 (Fundamental Theorem of Multi-Class PAC Learning [13; 24; 27]).
Assume $\Pi$ has finite Natarajan dimension $d$ and $\pi^{*}\in\Pi$. Then $\exists\,c>0$ such that $\forall\,\varepsilon,\delta\in(0,1)$, if $X\subset S\times A$ consists of at least $\frac{c}{\varepsilon}\left(d\log(\frac{|A|d}{\varepsilon})+\log\frac{1}{\delta}\right)$ i.i.d. samples $\mathcal{D}$ and the corresponding actions from $\pi^{*}$, then with probability at least $1-\delta$, any policy $\pi\in\Pi$ consistent with $X$ satisfies $\Pr_{s\sim\mathcal{D}}[\pi(s)\neq\pi^{*}(s)]\leq\varepsilon$.
Theorem 4.4 will ensure that $\Pr[a_{t}\neq\pi^{*}(s_{t})]$ is small, but even a single incorrect action can spell defeat if the payoff is bad enough. To handle this, the algorithm continues to selectively ask for help, but only when the state is far from any previously queried state (i.e., out-of-distribution). The policy $\pi$ is then updated based on this new information. The threshold $r$ is chosen such that we can cover the entire state $S$ with a sublinear number of queries.
1:function LookBeforeYouLeap($T\in\mathbb{N},\>r\in\mathbb{R}_{\geq 0},\>g:\mathbb{N}\to\mathbb{N}$)
2:     $X\leftarrow\emptyset$
3:     for $t$ from $1$ to $g(T)$ do $\triangleright$ Gather baseline data set
4:         Query mentor and observe $\pi^{*}(s_{t})$
5:         $X\leftarrow X\cup\{(s_{t},\pi^{*}(s_{t}))\}$
6:     $\pi\leftarrow$ any $(X,r)$-local policy in $\Pi$
7:     for $t$ from $g(T)+1$ to $T$ do $\triangleright$ Mostly follow $\pi$, but ask for help if out-of-distribution
8:         if $\min_{(s,a)\in X}||s_{t}-s||\leq r$ then
9:              Take action $\pi(s_{t})$
10:         else
11:              Query mentor and observe $\pi^{*}(s_{t})$
12:              $X\leftarrow X\cup\{(s_{t},\pi^{*}(s_{t}))\}$
13:              $\pi\leftarrow$ any $(X,r)$-local policy in $\Pi$
Algorithm 1 achieves near-zero regret when $\Pi$ has bounded Natarajan dimension and is $r$-local.
Tractability. Our algorithm can only be useful in practice if finding a $(X,r)$-local policy in $\Pi$ is computationally tractable. Such guarantees are impossible without further assumptions on $\Pi$ (this is again standard for PAC learning). However, if $\Pi$ supports nearest-neighbor policies (which satisfy locality for all values for $r$), those policies are both intuitive and can be computed easily.
We are now ready to state the main result of this section. The constant $c$ in Theorem 4.5 is the same as the constant $c$ in Theorem 4.4.
Theorem 4.5.
[…]
$\operatorname*{\mathbb{E}}_{\mathbf{s}\sim\mathcal{D}^{T}}\left[R_{T}(\mathbf{s},\mu,\pi^{*})\right]\leq\frac{6L}{T^{\frac{0.5}{n+0.5}}}+\frac{1}{T}+\exp(-3T^{\frac{0.5}{n+0.5}}/8)$
Corollary 4.6 makes the near-zero regret and sublinear mentor queries more explicit.
Corollary 4.6.
Assume $\pi^{*}\in\Pi$, where $\Pi$ has finite Natarajan dimension $d$ and is $\varepsilon(T)^{1/n}$-local for $\varepsilon(T)=T^{\frac{-n}{n+0.5}}$. Then Algorithm 1 with $g(T)=2cdT^{\frac{n}{n+0.5}}\log(d\,T)$ and $r=\varepsilon(T)^{1/n}$ makes at most $O\left(T^{\frac{n}{n+0.5}}(d\log(d\,T)+\operatorname*{vol}(\operatorname*{conv}(S)))\right)$ queries and satisfies
$\lim_{T\to\infty}\>\operatorname*{\mathbb{E}}_{\mathbf{s}\sim\mathcal{D}^{T}}\left[R_{T}(\mathbf{s},\mu,\pi^{*})\right]=0$
When $d$ is constant, the number of queries in Corollary 4.6 is $o(T)$. (Actually, $d$ can be as large as $T^{\frac{0.5-\delta}{n+0.5}}$ for $\delta\in(0,1/2)$.) Also, since the expected regret goes to zero, the regret must go to zero with high probability (since regret is always nonnegative).
Finally, note that standard PAC learning lower bounds do not directly yield lower bounds for our setting: even $\Pr[a_{t}\neq\pi^{*}(s_{t})]=1$ is insufficient for high regret, since the payoffs could be nearly optimal. Also, PAC learning lower bounds typically state that there exists a distribution for which the algorithm performs poorly, while our lower bound uses a fixed known distribution.
5 Near-zero regret for simple mentor policies in 1D with adversarial states
Our final result is another algorithm with near-zero regret and sublinear queries, but under different assumptions. In this section, we assume a 1D state space and we allow the state sequence to be adversarially chosen. We consider on the following notion of simplicity:
Definition 5.1.
Given a mentor policy $\pi^{*}$, partition the state space $S$ into intervals such that all states within each interval share the same mentor action. Let $\{S_{1},\dots,S_{k}\}$ be a partition that minimizes the number of intervals. We call each $S_{j}$ a segment. Let $f(\pi^{*})$ denote the number of segments in $\pi^{*}$.
This is similar conceptually to Natarajan dimension in that it limits the ability of the policy class to realize arbitrary combinations of labels (i.e., optimal actions) on $\mathbf{s}$. The number of segments also relates to the minimum depth needed by a neural network to learn the policy [7]. Figure 1 shows why the Lipschitz constant does not bound the number of segments ($f(T)$ can be arbtrarily large).
We prove the following result:
Theorem 5.2.
For any $c\in[0,1]$, any $\mathbf{s}\in S^{T}$, and any $\pi^{*}$ with $f(\pi^{*})\leq K$, Algorithm 2 with $g(T)=T^{c}$ makes at most $(4+\operatorname*{len}(\operatorname*{conv}(\mathbf{s})))g(T)$ queries and satisfies $R_{T}(\mathbf{s},\mu,\pi^{*})\leq 6LKT^{1-2c}$.
Corollary 5.3 makes the sublinear queries and near-zero regret of Theorem 5.2 more explicit when the number of segments is a constant.
Corollary 5.3.
For any $c\in(1/2,1)$, Algorithm 2 with $g(T)=T^{c}$ makes $O(T^{c}\operatorname*{len}(\operatorname*{conv}(S)))$ queries and satisfies
$\lim_{T\to\infty}\ \sup_{\mathbf{s}\in S^{T}}\ \sup_{\mu}\ \sup_{\pi^{*}:f(\pi^{*})\leq K}\ R_{T}(\mathbf{s},\mu,\pi^{*})=0$
Our algorithm does not need to know the Lipschitz constant $L$ or the number of segments; it only needs to know the time horizon $T$.
5.1 Intuition behind the algorithm
The algorithm maintains a set of buckets which partition the observed portion of the state space. Each bucket’s length determines the maximum loss in payoff we will allow from that subset of the state space. As long as the bucket contains a query from a prior time step, the Lipschitz constraint allows us to bound $\mu^{*}(s_{t})-\mu(s_{t},a_{t})$ based on the length of the bucket containing $s_{t}$. We always query if the bucket does not contain a prior query
The granularity of the buckets is controlled by a function $g$, with the initial buckets having length $1/g(T)$. Since we can expect one query per bucket, we need $g(T)\in o(T)$ to ensure sublinear queries.
Regardless of the bucket length, the adversary can still place multiple segments in the same bucket $B$. A single query only tells us the optimal action for one of those segments, so we risk a payoff as bad as $\mu^{*}(s_{t})-O(\operatorname*{len}(B))$ whenever we choose not to query. We can endure a limited number of such payoffs, but if we never query again in that bucket, we may suffer $\Theta(T)$ such payoffs. Letting $\mu^{*}(s_{t})=1$ for simplicity, that would lead to $\prod_{t=1}^{T}\mu(s_{t},a_{t})\leq\big{(}1-\frac{1}{O(g(T))}\big{)}^{\Theta(T)}$, which converges to 0 (i.e., guaranteed catastrophe) when $g(T)\in o(T)$.
This failure mode suggests a natural countermeasure: if we start to suffer significant (potential) losses in the same bucket, then we should probably query there again. One way to structure these supplementary queries is by splitting the bucket in half when enough time steps have involved that bucket. It turns out that splitting after $T/g(T)$ time steps is a sweet spot.
This approach is reminiscent of Algorithm 1 in that it collects an initial dataset (one query per initial bucket), and then performs supplemental queries as needed based on the observed states. We do not think this is a coincidence.
1:function AvoidCatastrophe($T\in\mathbb{N},\>g:\mathbb{N}\to\mathbb{N}$)
2:     $S_{Q}\leftarrow\emptyset$triangleright Previously queried states
3:     $\pi\leftarrow\emptyset$triangleright Records $\pi^{*}(s)$ for each $s\in S_{Q}$
4:     $\mathcal{B}\leftarrow\emptyset$triangleright The set of active buckets
5:     for $t$ from $1$ to $T$ do
6:         EvaluateState$(s_{t})$
7:function EvaluateState($s\in S$)
8:     if $s\not\in B$ for all $B\in\mathcal{B}$ then $\triangleright$ No bucket containing $s$: create a new bucket and try again
9:         $B\leftarrow\left[\frac{j-1}{g(T)},\frac{j}{g(T)}\right]$ for $j\in\mathbb{Z}$ such that $s\in B$
10:         $\mathcal{B}\leftarrow\mathcal{B}\cup\{B\}$
11:         $n_{B}\leftarrow 0$triangleright Number of time steps that have used $B$
12:         EvaluateState$(s)$
13:     else
14:         $B\leftarrow$ any bucket containing $s$
15:         if $S_{Q}\cap B=\emptyset$ then $\triangleright$ No queries in this bucket
16:              Query mentor and observe $\pi^{*}(s)$
17:              $\pi(s)\leftarrow\pi^{*}(s)$
18:              $S_{Q}\leftarrow S_{Q}\cup\{s\}$
19:              $n_{B}\leftarrow n_{B}+1$
20:         else if $n_{B}<T/g(T)$ then $\triangleright$ Bucket has a query and isn’t full: take that action
21:              Let $s^{\prime}\in S_{Q}\cap B$
22:              Take action $\pi(s^{\prime})$
23:              $n_{B}\leftarrow n_{B}+1$
24:         $\triangleright$ Bucket is full: split bucket and try again
25:              $B=[x,y]$
26:              $(B_{1},B_{2})\leftarrow\Big{(}\left[x,\frac{x+y}{2}\right],\left[\frac{x+y}{2},y\right]\Big{)}$
27:              $(s_{B_{1}},s_{B_{2}})\leftarrow(0,0)$
28:              $\mathcal{B}\leftarrow\mathcal{B}\cup\{B_{1},B_{2}\}\setminus B$
29:              EvaluateState$(s)$
Algorithm 2 achieves near-zero regret when the mentor’s policy has a bounded number of segments.
6 Limitations
Our paper contains both positive and negative results. As such, our claim in this paper is not “avoiding catastrophe is possible (or impossible) under assumptions X, Y, Z, and we argue that X, Y, Z hold”. Instead, we aim to understand the conditions under which it is possible to avoid catastrophe. Our claim is “avoiding catastrophe is possible under X, Y, Z, but impossible under A, B, C; we might sometimes expect X, Y, Z and sometimes expect A, B, C”. For example, some real-world instances might exhibit bounded Natarajan dimension, while others might not. Our paper elucidates what to expect from either scenario.
The goal, then, is to make the gap between X, Y, Z and A, B, C as small as possible: to isolate the boundary of tractability. Our results suggest that the key factor determining tractability is the complexity of the mentor policy class. Our negative result (Theorem 3.1) shows that a simple environment (bounded 1D state space, known uniform distribution over states, two actions) is not enough for tractability if the mentor policy class is complex. On the other hand, our primary positive result (Theorem 4.5) holds for quite general environments (unbounded $n$-dimensional state space, unknown arbitrary distribution over states, many actions), as long as the mentor policy class is simple.
One remaining gap between those two results is the assumption of $r$-locality in Theorem 4.5. Although we think $r$-locality is a fairly plausible assumption (as argued in Section 4), that assumption would need to be removed in order for Theorem 3.1 and Theorem 4.5 to fully match.
Another question is how far this positive result generalizes. In particular, the assumptions of (1) only caring about catastrophe and (2) i.i.d. states significantly limit the real-world applicability of our algorithm. In future work, we hope to extend this result to general MDPs.
Our secondary positive result (Theorem 5.2) also suggests that mentor policy complexity could be the key factor determining tractability. However, it has the drawback that the number of segments in the mentor policy is a less established measure of complexity than the well-studied notion of Natarajan dimension. In our opinion, the more serious limitation is the restriction to a 1D state space. Although Theorem 5.2 does permit adversarial state sequences, we feel that Theorem 4.5 does a better overall job of isolating the boundary of tractability.
7 Conclusion and future work
In this paper, we proposed a model of avoiding catastrophe in sequential decision-making. We showed that in general, any algorithm either constantly asks for help or is guaranteed to cause catastrophe. However, the problem becomes tractable when the mentor’s policy is relatively simple, for two different definitions of “simple”.
There remain some technical questions within this paper’s model: (1) Can Theorem 4.5 be generalized to adversarial state sequences? (2) Is $r$-locality necessary? (3) Are there are reasonable alternatives to the Lipschitz assumption? (Indeed, the utility of $r$-locality steps from the Lipschitz assumption.)
As mentioned, we are also excited to move beyond the current model to full MDPs: recall the open problem we posed in Section 1. Most applications are better captured by MDPs than by this paper’s assumptions of i.i.d. states and only caring about catastrophe.
Finally, although we intentionally avoided Bayesian inference in this paper, it seems premature to abandon that tool entirely. Are there tractable approximations to Bayesian approaches?
Acknowledgements
This work was supported in part by a gift from Open Philanthropy to the Center for Human-Compatible AI (CHAI) at UC Berkeley. This paper also benefited from discussions with many researchers at CHAI. We would like to particularly thank Michael Cohen, Bhaskar Mishra, Cam Allen, and Li Dayan for helpful feedback.
Appendix A Proof of Theorem 3.1
A.1 Proof roadmap
[…]
Recall that $\operatorname*{\mathbb{E}}_{\mathbf{s}\sim U(S^{T}),\mathbf{a}}$ denotes the expectation over both the random states and any additional randomness in the algorithm.
The proof proceeds via the following steps:
  1. 1.
    Prove that $f(T)=\sqrt{(|Q_{T}|+1)T}$ is asymptotically between $|Q_{T}|$ and $T$ (Lemma A.1).
  2. 2.
    Provide a simple variant of the Chernoff bound which we will apply multiple times (Lemma A.2).
  3. 3.
    Show that with high probability, $\sum_{j\in A}|V_{j}|$ is adequately large (Lemma A.3).
  4. 4.
    The key lemma is Lemma A.4, which shows that a randomly sampled $\mathbf{x}$ produces poor agent performance with high probability. The central idea is that at least $f(T)-|Q_{T}|$ sections are never queried (which is large, by Lemma A.1), so the agent has no way of knowing the optimal action in those sections. As a result, the agent picks the wrong answer at least half the time on average (and at least a quarter of the time with high probability). Lemma A.3 implies that a constant fraction of those time steps will have quite suboptimal payoffs, again with high probability.
  5. 5.
    To complete the proof, we observe that $\sup_{\mu}\ \operatorname*{\mathbb{E}}_{\mathbf{s}\sim U(S^{T}),\,\mathbf{a}}\ R_{T}(\mathbf{s},\mu,\pi^{*})\geq\operatorname*{\mathbb{E}}_{\mathbf{x}\sim U(\{0,1\}^{f(T)})}\ \operatorname*{\mathbb{E}}_{\mathbf{s}\sim U(S^{T}),\,\mathbf{a}}\ R_{T}(\mathbf{s},\mu_{f,\mathbf{x}},\pi^{*})$. This is essentially an application of the probabilistic method: if a randomly chosen $\mu_{f,\mathbf{x}}$ has high expected regret, then the worst case $\mu$ also has high expected regret.
Note that $\mathbf{s},\mathbf{a}$, and $\mathbf{x}$ are random variables, so all variables defined on top of them (such as $V_{j}$) are also random variables. In contrast, the partition $S=\{S_{1},\dots,S_{f(T)}\}$ and properties thereof (like the midpoints $m_{j}$) are not random variables.
A.2 Proof
Lemma A.1.
Let $a,b:\mathbb{N}\to\mathbb{N}$ be functions such that $a(x)\in o(b(x))$. Then $c(x)=\sqrt{a(x)b(x)}$ satisfies $a(x)\in o(c(x))$ and $c(x)\in o(b(x))$.
Proof.
Since $a$ and $b$ are strictly positive (and thus $c$ is as well), we have
$\frac{a(x)}{c(x)}=\frac{a(x)}{\sqrt{a(x)b(x)}}=\sqrt{\frac{a(x)}{b(x)}}=\frac{\sqrt{a(x)b(x)}}{b(x)}=\frac{c(x)}{b(x)}$
Then $a(x)\in o(b(x))$ implies
$\lim_{x\to\infty}\frac{a(x)}{c(x)}=\lim_{x\to\infty}\frac{c(x)}{b(x)}=\lim_{x\to\infty}\sqrt{\frac{a(x)}{b(x)}}=0$
as required. ∎
Lemma A.2.
Let $z_{1},\dots,z_{n}$ be i.i.d. variables in $\{0,1\}$ and let $Z=\sum_{i=1}^{n}z_{i}$. If $\operatorname*{\mathbb{E}}[Z]\geq M$, then $\Pr\big{[}Z\leq M/2\big{]}\leq\exp(-M/8)$.
Proof.
By the Chernoff bound for i.i.d. binary variables, we have $\Pr[Z\leq\operatorname*{\mathbb{E}}[Z]/2]\leq\exp(-\operatorname*{\mathbb{E}}[Z]/8)$. Since $-\operatorname*{\mathbb{E}}[Z]\leq-M$ and $\exp$ is an increasing function, we have $\exp(-\operatorname*{\mathbb{E}}[Z]/8)\leq\exp(-M/8)$. Also, $M/2\leq E[Z]/2$ implies $\Pr[Z\leq M/2]\leq\Pr[Z\leq\operatorname*{\mathbb{E}}[Z]/2]$. Combining these inequalities proves the lemma. ∎
Lemma A.3.
let $A\subseteq[f(T)]$ be any nonempty subset of sections. Then
$\Pr\left[\sum_{j\in A}|V_{j}|\leq\frac{T|A|}{4f(T)}\right]\leq\exp\left(\frac{-T}{16f(T)}\right)$
Proof.
[…]
$\Pr\left[\sum_{j\in A}|V_{j}|\leq\frac{T|A|}{4f(T)}\right]\leq\exp\left(\frac{-T|A|}{16f(T)}\right)\leq\exp\left(\frac{-T}{16f(T)}\right)$
with the last step due to $|A|\geq 1$. ∎
Lemma A.4.
Independently sample $\mathbf{x}\sim U(\{0,1\}^{f(T)})$ and $s\sim U(S^{T})$. [^13] That is, the entire set $\{x_{1},\dots,x_{f(T)},s_{1},\dots,s_{T}\}$ is mutually independent. Then with probability at least $1-\exp\big{(}\frac{-T}{16f(T)}\big{)}-\exp\big{(}-\frac{f(T)-|Q_{T}|}{16}\big{)}$,
$\prod_{t=1}^{T}\mu_{f,\mathbf{x}}(s_{t},a_{t})\leq\exp\left(-\frac{LT(f(T)-|Q_{T}|)}{2^{7}f(T)^{2}}\right)$
Proof.
Part 1: setup. Let $J_{\neg Q}=\{j\in[f(T)]:s_{t}\not\in S_{j}\ \forall t\in Q_{T}\}$ be the set of sections that are never queried. Since each query appears in exactly one section (because each state appears in exactly one section), $|J_{\neg Q}|\geq f(T)-|Q_{T}|$.
For each $j\in J_{\neg Q}$, let $y_{j}$ be the action taken most frequently among time steps in $V_{j}$:
$y_{j}=\operatorname*{arg\,max}_{a\in\{0,1\}}\Big{|}\{t\in V_{j}:a=a_{t}\}\Big{|}$
Let $\bar{J}=\{j\in J_{\neg Q}:x_{j}\neq y_{j}\}$. For each $j\in\bar{J}$, let $V_{j}^{\prime}=\{t\in V_{j}:a_{t}\neq x_{j}\}$ be the set of time steps where the agent chooses the wrong action (assuming payoff function $\mu_{f,\mathbf{x}}$).
[…]
$\Pr\left[|\bar{J}|\leq\frac{f(T)-|Q_{T}|}{4}\right]\leq\exp\left(-\frac{f(T)-|Q_{T}|}{16}\right)$
Part 3: $V_{j}^{\prime}$ is not too small. Since $j\in J_{\neg Q}$, the mentor is not queried on any time step $t\in V_{j}$, so $a_{t}\in\{0,1\}$ for all $t\in V_{j}$. Since the agent chooses one of two actions for each $t\in V_{j}$, the more frequent action must be chosen chosen at least half of the time: $a_{t}=y_{j}$ for at least half of the time steps in $V_{j}$. Since $x_{j}\neq y_{j}$ for $j\in\bar{J}$, we have $a_{t}=y_{j}\neq x_{j}$ for those time steps, so $|V_{j}^{\prime}|\geq|V_{j}|/2$.
Part 4: a bound in terms of $\bar{J}$ and $V_{j}$. Consider any $j\in\bar{J}$ and $t\in V_{j}^{\prime}\subseteq V_{j}$. By definition of $V_{j}$, we have $|m_{j}-s_{t}|\leq\frac{1}{4f(T)}$. Then by definition of $\mu_{f,\mathbf{x}}$,
$\mu_{f,\mathbf{x}}(s_{t},a_{t})=$ $\ 1-L\left(\frac{1}{2f(T)}-|s_{t}-m_{j}|\right)$
$\leq$ $\ 1-L\left(\frac{1}{2f(T)}-\frac{1}{4f(T)}\right)$
$=$ $\ =1-\frac{L}{4f(T)}$
Now aggregating across time steps,
$\prod_{t=1}^{T}\mu_{f,\mathbf{x}}(s_{t},a_{t})\leq$ $\ \prod_{j\in\bar{J}}\prod_{t\in V_{j}^{\prime}}\mu_{f,\mathbf{x}}(s_{t},a_{t})$
$\leq$ $\ \prod_{j\in\bar{J}}\left(1-\frac{L}{4f(T)}\right)^{|V_{j}^{\prime}|}$
$\leq$ $\ \prod_{j\in\bar{J}}\left(1-\frac{L}{4f(T)}\right)^{|V_{j}|/2}$
[…]
$\log\prod_{t=1}^{T}\mu_{f,\mathbf{x}}(s_{t},a_{t})\leq$ $\ \log\prod_{j\in\bar{J}}\left(1-\frac{L}{4f(T)}\right)^{|V_{j}|/2}$
$=$ $\ \sum_{j\in\bar{J}}\frac{|V_{j}|}{2}\log\left(1-\frac{L}{4f(T)}\right)$
$\leq$ $\ -\sum_{j\in\bar{J}}\frac{L|V_{j}|}{8f(T)}$
[…]
$\log\prod_{t=1}^{T}\mu_{f,\mathbf{x}}(s_{t},a_{t})\leq$ $\ -\sum_{j\in\bar{J}}\frac{L|V_{j}|}{8f(T)}$
$\leq$ $\ -\frac{L}{8f(T)}\cdot\frac{T|\bar{J}|}{4f(T)}$
$\leq$ $\ -\frac{L}{8f(T)}\cdot\frac{T}{4f(T)}\cdot\frac{f(T)-|Q_{T}|}{4}$
$=$ $\ -\frac{LT(f(T)-|Q_{T}|)}{2^{7}f(T)^{2}}$
Exponentiating both sides proves the lemma. ∎
Let $\alpha(T)=\exp\big{(}\frac{-T}{16f(T)}\big{)}+\exp\big{(}-\frac{f(T)-|Q_{T}|}{16}\big{)}$ for brevity. See 3.1
Proof.
If the algorithm has sublinear queries, then there exists $g(T)\in o(T)$ such that $|Q_{T}|\leq g(T)$ for all $\mathbf{s},\mu,\pi^{*}$. Let $f(T)=\sqrt{(g(T)+1)T}$. Then by Lemma A.1, $g(T)\in o(f(T))$ and $f(T)\in o(T)$. Combining this with $|Q_{T}|\leq g(T)$, we get $\lim_{T\to\infty}\alpha(T)=0$. Also, since $|Q_{T}|\in o(f(T))$, there exists $T_{0}$ such that $|Q_{T}|\leq f(T)/2$ for all $T\geq T_{0}$. Combining this with Lemma A.4 and noting that $\prod_{t=1}^{T}\mu_{f,\mathbf{x}}(s_{t},a_{t})\leq 1$, we have
$\ \operatorname*{\mathbb{E}}_{\mathbf{x}\sim U(\{0,1\}^{f(T)})}\ \,\operatorname*{\mathbb{E}}_{\mathbf{s}\sim U(S^{T}),\,\mathbf{a}}\ \>\prod_{t=1}^{T}\mu_{f,\mathbf{x}}(s_{t},a_{t})$
$\leq$ $\ \alpha(T)\cdot 1+\big{(}1-\alpha(T)\big{)}\exp\left(-\frac{LT(f(T)-|Q_{T}|)}{2^{7}f(T)^{2}}\right)$
$\leq$ $\ \alpha(T)+\big{(}1-\alpha(T)\big{)}\exp\left(-\frac{LTf(T)/2}{2^{7}f(T)^{2}}\right)$
$=$ $\ \alpha(T)+\big{(}1-\alpha(T)\big{)}\exp\left(-\frac{LT}{2^{8}f(T)}\right)$
whenever $T\geq T_{0}$. Since $\prod_{t=1}^{T}\mu_{f,\mathbf{x}}^{*}(s_{t})=1$ always, we have [^14] Note that Fubini’s theorem means we need not worry about the order of the expectation operators.
$\sup_{\mu}\ \operatorname*{\mathbb{E}}_{\mathbf{s}\sim U(S^{T}),\,\mathbf{a}}\ R_{T}(\mathbf{s},\mu,\pi^{*})\geq$ $\ \operatorname*{\mathbb{E}}_{\mathbf{x}\sim U(\{0,1\}^{f(T)})}\ \,\operatorname*{\mathbb{E}}_{\mathbf{s}\sim U(S^{T}),\,\mathbf{a}}\ R_{T}(\mathbf{s},\mu_{f,\mathbf{x}},\pi^{*})$
$=\operatorname*{\mathbb{E}}_{\mathbf{x}\sim U(\{0,1\}^{f(T)})}\ \,\operatorname*{\mathbb{E}}_{\mathbf{s}\sim U(S^{T}),\,\mathbf{a}}$ $\ \,\left[\prod_{t=1}^{T}\mu_{f,\mathbf{x}}^{*}(s_{t})-\prod_{t=1}^{T}\mu_{f,\mathbf{x}}(s_{t},a_{t})\right]$
$\geq$ $\ 1-\alpha(T)-\big{(}1-\alpha(T)\big{)}\exp\left(-\frac{LT}{2^{8}f(T)}\right)$
Therefore
$\lim_{T\to\infty}\sup_{\mu}\ \operatorname*{\mathbb{E}}_{\mathbf{s}\sim U(S^{T}),\,\mathbf{a}}\ R_{T}(\mathbf{s},\mu,\pi^{*})\geq$ $\ 1-\lim_{T\to\infty}\alpha(T)-\big{(}1-\lim_{T\to\infty}\alpha(T)\big{)}\cdot\exp\left(\lim_{T\to\infty}-\frac{LT}{2^{8}f(T)}\right)$
$=$ $\ 1-0-(1-0)\cdot\exp(-\infty)$
[…]
as required. ∎
Appendix B Proof of Theorem 4.5
B.1 Proof roadmap
We use the following notation throughout the proof:
  1. 1.
    For each $t\in[T]$, let $X_{t},$ refer to the values of $X$ respectively at the start of time step $t$. Define $\pi_{t}$ similarly, except only for $t>g(T)$.
  2. 2.
    Let $V_{T}=\{t\in[T]:\mu(s_{t},a_{t})<\mu^{*}(s_{t})\}$ be the set of time steps with a suboptimal payoff.
  3. 3.
    Unless specified (e.g., $\Pr_{s\sim\mathcal{D}}$), all probabilities are with respect to $\mathbf{s}\sim\mathcal{D}^{T}$.
The proof proceeds in the following steps:
  1. 1.
    Show that $\pi_{t}$ meets the criteria for the fundamental theorem of multi-class PAC learning and thus has small expected error on a single state (Lemma B.1).
  2. 2.
    Use Lemma B.1 and a non-standard concentration inequality (Lemma B.2) to show that $|V_{T}|$ is small with high probability. This involves two steps (Lemmas B.3 and B.4).
  3. 3.
    Having a small number of mistakes is not sufficient, because a single payoff of $\mu(s_{t},a_{t})=0$ can ruin the entire endeavor. Lemmas B.5 and B.6 bound the suboptimality of a single payoff.
  4. 4.
    Lemma B.7 transforms the multiplicative objective into an additive form that is easier to manipulate.
  5. 5.
    Lemma B.8 bounds the total regret of the algorithm.
  6. 6.
    Lemma B.10 bounds the total number of queries based on a packing number lemma (Lemma B.9).
Lemmas B.8 and B.10 together imply Theorem 4.5.
B.2 Proof
Lemma B.1.
With probability at least $1-1/T$, we have $\Pr_{s\sim\mathcal{D}}[\pi_{t}(s)\neq\pi^{*}(s)]\leq\varepsilon(T)$ $\forall t>g(T)$.
Proof.
Let $\delta=1/T$. Then
$|X_{g(T)+1}|=$ $\ g(T)$
$=$ $\ \frac{cd}{\varepsilon(T)}\log(d\,T)$
$=$ $\ 2\frac{cd}{\varepsilon(T)}\log(d^{2}T^{2})$
$=$ $\ \frac{c}{\varepsilon(T)}(d\log(d^{2}T^{2})+d\log(d^{2}T^{2}))$
$\geq$ $\ \frac{c}{\varepsilon(T)}(d\log(d\,T\cdot T^{\frac{n}{n+0.5}})+\log T)$
$\geq$ $\ \frac{c}{\varepsilon(T)}\left(d\log(\frac{d|A|}{\varepsilon(T)})+\log\frac{1}{\delta}\right)$
Now consider any $t>g(T)$. By construction, $\pi_{t}$ is $(X_{t},r)$-local ($\pi$ and $X$ are not updated on every time step, but are always updated at the same time). Therefore $\pi_{t}$ is consistent with $X_{t}$ and in turn is consistent with $X_{g(T)+1}$ since $X_{g(T)+1}\subseteq X_{t}$. Since $X_{g(T)+1}$ is a set of i.i.d. samples of the requisite size, Theorem 4.4 implies that with at probability at least $1-1/T$, any policy $\pi\in\Pi$ consistent with $X_{g(T)+1}$ satisfies $\Pr_{s\sim\mathcal{D}}[\pi(s)\neq\pi^{*}(s)]\leq\varepsilon(T)$. In particular, this holds for $\pi_{t}$ for all $t>g(T)$, completing the proof. ∎
Next, we wish to prove that $|V_{T}|$ is small. The $\{t\in V_{T}\}$ indicator variables are not independent, and standard martingale concentration bounds like the Azuma-Hoeffding inequality are not tight enough, so we use a specialized inequality from [8]. The statement below is slightly simplified from its original source, since we do not need the more general version.
Lemma B.2 (Corollary 2.1 in [8]).
Let $x_{0},\dots,x_{k}$ be a supermartingale and assume there exists $b\in(0,1]$ such that $x_{i}-x_{i-1}\in[-b,1]$ for all $i\in[k]$. Then for any $x>0$,
$\Pr[x_{k}\geq x]\leq\exp\left(-\frac{x^{2}}{2(kb+x/3)}\right)$
Lemma B.3.
Assume $\Pr_{s\sim\mathcal{D}}[\pi_{t}(s)\neq\pi^{*}(s)]\leq\varepsilon(T)$ for all $t>g(T)$. Then
$\Pr\left[|V_{T}|\geq 2T\varepsilon(T)\right]\leq\exp(-3T\varepsilon(T)/8)$
Proof.
Let $k=T-g(T)$ for brevity. For each $t\in\{g(T)+1,\dots,T\}$, define a random variable $y_{t-g(T)}=\mathbf{1}_{\pi_{t}(s_{t})\neq\pi^{*}(s_{t})}-\varepsilon(T)$. Also let $x_{0}=0$, and for $i\in[k]$, let $x_{i}=x_{i-1}+y_{i}$. This defines random variables $y_{1},\dots,y_{k}$ and $x_{0},\dots,x_{k}$. For each $i\in[k]$, let $y_{<i}=\{y_{1},\dots,y_{i-1}\}$ and $x_{<i}=\{x_{0},\dots,x_{i-1}\}$.
To show that $x_{0},\dots,x_{k}$ form a supermartingale, we first claim that $\operatorname*{\mathbb{E}}[y_{i}\mid y_{<i}]\leq 0$ for all $i\in[k]$. Regardless of what occurs on previous time steps, each state is sampled randomly from $\mathcal{D}$. Thus
$\operatorname*{\mathbb{E}}[y\mid y_{<i}]=\Pr_{s\sim\mathcal{D}}[\pi_{i+g(T)}\neq\pi^{*}(s)\mid y_{<i}]-\varepsilon(T)\leq 0$
Since $y_{<i}$ and $x_{<i}$ provide the same information (i.e., either can be constructed from the other), we have $\operatorname*{\mathbb{E}}[y_{i}\mid x_{<i}]\leq 0$. Therefore
$\operatorname*{\mathbb{E}}[x_{i}\mid x_{<i}]=$ $\ \operatorname*{\mathbb{E}}[x_{i-1}+y_{i}-\varepsilon(T)\mid x_{<i}]$
$=$ $\ \operatorname*{\mathbb{E}}[x_{i-1}\mid x_{<i}]+\operatorname*{\mathbb{E}}[y_{i}\mid x_{<i}]-\varepsilon(T)$
$\leq$ $\ x_{i-1}+\varepsilon(T)-\varepsilon(T)$
$=$ $\ x_{i-1}$
Also, since each $x_{i}$ is bounded, we have $\operatorname*{\mathbb{E}}[x_{i}]<\infty$ for all $i\in[k]$. Therefore $x_{0},\dots,x_{k}$ is indeed a supermartingale. Furthermore, $x_{i}-x_{i-1}=y_{i}\in[-\varepsilon(T),1]$ for all $i\in[k]$, so by Lemma B.2,
$\Pr[x_{k}\geq 2T\varepsilon(T)]\leq$ $\ \exp\left(-\frac{T^{2}\varepsilon(T)^{2}}{2(k\varepsilon(T)+T\varepsilon(T)/3)}\right)$
Since $k\leq T$,
$\Pr[x_{k}\geq 2T\varepsilon(T)]\leq$ $\ \exp\left(-\frac{T^{2}\varepsilon(T)^{2}}{2(T\varepsilon(T)+T\varepsilon(T)/3)}\right)$
$=$ $\ \exp\left(-\frac{T^{2}\varepsilon(T)^{2}}{8T\varepsilon(T)/3}\right)$
$=$ $\ \exp(-3T\varepsilon(T)/8)$
as required. ∎
Lemma B.4.
We have $\Pr[|V_{T}|\geq 2T\varepsilon(T)]\leq\exp(-3T^{\frac{0.5}{n+0.5}}/8)+1/T$.
Proof.
Let $\xi$ be the event that $\Pr_{s\sim\mathcal{D}}[\pi_{t}(s)\neq\pi^{*}(s)]\leq\varepsilon(T)$ for all $t>g(T)$. Lemma B.1 states that $\Pr[\neg\xi]\leq 1/T$ and Lemma B.3 states that $\Pr[|V_{T}|\geq 2T\varepsilon(T)\mid\xi]\leq\exp(-3T\varepsilon(T)/8)$. Then
$\Pr[|V_{T}|\geq 2T\varepsilon(T)]=$ $\ \Pr[|V_{T}|\geq 2T\varepsilon(T)\mid\xi]\Pr[\xi]+\Pr[|V_{T}|\geq 2T\varepsilon(T)\mid\neg\xi]\Pr[\neg\xi]$
$\leq$ $\ \exp(-3T\varepsilon(T)/8)\cdot 1+1\cdot 1/T$
$=$ $\ \exp(-3T^{1-\frac{n}{n+0.5}}/8)+1/T$
$=$ $\ \exp(-3T^{\frac{0.5}{n+0.5}}/8)+1/T$
as claimed. ∎
Remark B.1.
The astute reader note that Lemma B.5 is the only place in the paper where we invoke the assumption that the mentor’s policy is optimal. Both of our positive results (Theorems 4.5 and 5.2) can be generalized to a possibly suboptimal mentor policy $\pi^{m}$ if:
  1. 1.
    We modify the Lipschitz assumption to be $|\mu(s,\pi^{m}(s))-\mu(s,\pi^{m}(s^{\prime}))|\leq L||s-s^{\prime}||$. In words, the value of the mentor’s policy shouldn’t change too much between nearby states.
  2. 2.
    We define replace $\mu^{*}(s)$ with $\mu(s,\pi^{m}(s))$ throughout.
Lemma B.5 follows immediately from these two assumptions.
Also note that Lemmas B.5 and B.7 are not specific to this algorithm and will also be used in the proof of Theorem 5.2.
Lemma B.5.
If $||s-s^{\prime}||\leq\delta$ for some $\delta\geq 0$, then $\mu(s,\pi^{*}(s^{\prime}))\geq\mu^{*}(s)-2L\delta$.
Proof.
We have By the Lipschitz continuity of $\mu$, $\mu(s,\pi^{*}(s^{\prime}))\geq\mu(s^{\prime},\pi^{*}(s^{\prime}))-L||s-s^{\prime}||$. Also, $\mu(s^{\prime},\pi^{*}(s^{\prime}))=\max_{a\in A}\mu(s^{\prime},a)\geq\mu(s^{\prime},\pi^{*}(s))$ by assumption. Thus
$\mu(s,\pi^{*}(s^{\prime}))\geq$ $\ \mu(s^{\prime},\pi^{*}(s^{\prime}))-L||s-s^{\prime}||$ (Lipschitz continuity of $\mu$)
$\geq$ $\ \mu(s^{\prime},\pi^{*}(s))-L||s-s^{\prime}||$ (By assumption, $\pi^{*}$ is optimal)
$\geq$ $\ \mu(s,\pi^{*}(s))-2L||s-s^{\prime}||$ (Lipschitz continuity of $\mu$ again)
$=$ $\ \mu^{*}(s)-2L||s-s^{\prime}||$ (Definition of $\mu^{*}(s)$)
Plugging in $||s-s^{\prime}||\leq\delta$ completes the proof. ∎
Lemma B.6.
Under the conditions of Theorem 4.5, $\mu(s_{t},a_{t})\geq\mu^{*}(s_{t})-2Lr$ for all $t\in[T]$.
Proof.
If we query the mentor at time $t$, then $\mu(s_{t},a_{t})=\mu^{*}(s_{t})$, so assume we do not query the mentor at time $t$. Let $(s^{\prime},a^{\prime})=\operatorname*{arg\,min}_{(s,a)\in X_{t}}||s_{t}-s||$: then $||s_{t}-s^{\prime}||\leq r$. Since $\pi_{t}$ is $(X_{t},r)$-local by construction, there exists $(s^{\prime\prime},a^{\prime\prime})\in X_{t}$ such that $||s_{t}-s^{\prime\prime}||\leq r$ and $a_{t}=\pi_{t}(s_{t})=a^{\prime\prime}$. Then by Lemma B.5, $\mu(s_{t},\pi^{*}(s^{\prime\prime}))\geq\mu^{*}(s_{t})-2Lr$. Furthermore, because $(s^{\prime\prime},a^{\prime\prime})\in X$, we have $\pi^{*}(s^{\prime\prime})=a^{\prime\prime}=a_{t}$. Therefore $\mu(s_{t},a_{t})\geq\mu^{*}(s_{t})-2Lr$. ∎
Lemma B.7.
If $\mu^{*}(s_{t})\geq\mu_{0}>0$ and $c_{t}\in\big{[}0,\mu_{0}/2\big{]}$ for all $t\in[T]$, then
$\prod_{t=1}^{T}\left(\mu^{*}(s_{t})-c_{t}\right)\geq\prod_{t=1}^{T}\mu^{*}(s_{t})-1.5\sum_{t=1}^{T}c_{t}$
Proof.
Recall the standard limit $\lim_{\delta\to 0}\delta^{-1}\log(1+\delta)=1$. Thus for all $\varepsilon>0$, there exists $\delta_{0}>0$ such that whenever $|\delta|\leq\delta_{0}$, we have $|\delta^{-1}\log(1+\delta)-1|\leq\varepsilon$. For our purposes, $\varepsilon=0.5$ is sufficient, we only care about $\delta<0$, and we only need the lower bound. One can verify that $\delta_{0}=0.5$ is sufficient. To summarize, we have $1.5\delta\leq\log(1+\delta)$ whenever $|\delta|\leq 0.5$.
Let $\delta_{t}=-c_{t}/\mu^{*}(s_{t})$ for each $t\in[T]$. Then
$|\delta_{t}|=\frac{c_{t}}{\mu^{*}(s_{t})}\leq\frac{\mu_{0}/2}{\mu_{0}}=0.5$
and so
$-\frac{1.5c_{t}}{\mu^{*}(s_{t})}\leq$ $\ \log\left(1-\frac{c_{t}}{\mu^{*}(s_{t})}\right)$
$\log(\mu^{*}(s_{t}))-\frac{1.5c_{t}}{\mu^{*}(s_{t})}\leq$ $\ \log\left(\mu^{*}(s_{t})-c_{t}\right)$
Now summing over $t$,
$\sum_{t=1}^{T}\log(\mu^{*}(s_{t}))-\sum_{t=1}^{T}\frac{1.5c_{t}}{\mu^{*}(s_{t})}\leq$ $\ \sum_{t=1}^{T}\log\left(\mu^{*}(s_{t})-c_{t}\right)$
$\log\prod_{t=1}^{T}\mu^{*}(s_{t})-\sum_{t=1}^{T}\frac{1.5c_{t}}{\mu^{*}(s_{t})}\leq$ $\ \log\prod_{t=1}^{T}\left(\mu^{*}(s_{t})-c_{t}\right)$
Exponentiating both sides yields
$\prod_{t=1}^{T}\left(\mu^{*}(s_{t})-c_{t}\right)\geq\exp\left(-\sum_{t=1}^{T}\frac{1.5c_{t}}{\mu^{*}(s_{t})}\right)\prod_{t=1}^{T}\mu^{*}(s_{t})$
Applying the standard inequality $e^{x}\geq 1+x$ gives us
$\prod_{t=1}^{T}\left(\mu^{*}(s_{t})-c_{t}\right)\geq$ $\ \left(1-\sum_{t=1}^{T}\frac{1.5c_{t}}{\mu^{*}(s_{t})}\right)\prod_{t=1}^{T}\mu^{*}(s_{t})$
$=$ $\ \prod_{t=1}^{T}\mu^{*}(s_{t})-1.5\sum_{t=1}^{T}c_{t}\frac{\prod_{i=1}^{T}\mu^{*}(s_{i})}{\mu^{*}(s_{t})}$
Since $\mu^{*}(s_{t})\in[0,1]$ for all $t\in[T]$, $\prod_{i=1}^{T}\mu^{*}(s_{i})\leq\mu^{*}(s_{t})$. Thus
$\prod_{t=1}^{T}\left(\mu^{*}(s_{t})-c_{t}\right)\geq$ $\ \prod_{t=1}^{T}\mu^{*}(s_{t})-1.5\sum_{t=1}^{T}c_{t}\frac{\mu^{*}(s_{t})}{\mu^{*}(s_{t})}$
$=$ $\ \prod_{t=1}^{T}\mu^{*}(s_{t})-1.5\sum_{t=1}^{T}c_{t}$
as required. ∎
Lemma B.8.
Under the conditions of Theorem 4.5, Algorithm 1 satisfies
$\operatorname*{\mathbb{E}}_{\mathbf{s}\sim\mathcal{D}^{T}}\left[R_{T}(\mathbf{s},\mu,\pi^{*})\right]\leq\frac{6L}{T^{\frac{0.5}{n+0.5}}}+\frac{1}{T}+\exp(-3T^{\frac{0.5}{n+0.5}}/8)$
Proof.
Let $\mu_{0}=4Lr$. If there exists $t\in[T]$ with $\mu^{*}(s_{t})<\mu_{0}$, then
$R_{T}(\mathbf{s},\mu,\pi^{*})=\prod_{t=1}^{T}\mu^{*}(s_{t})-\prod_{t=1}^{T}\mu(s_{t},a_{t})<\mu_{0}-0=4L\varepsilon(T)^{1/n}=4LT^{\frac{-1}{n+0.5}}\leq 6LT^{\frac{-0.5}{n+0.5}}$
so the claim is satisfied. Thus assume $\mu^{*}(s_{t})\geq\mu_{0}$ for all $t\in[T]$. Let $c_{t}=0$ for $t\not\in V_{T}$ and $c_{t}=2Lr$ for $t\in V_{T}$. Then by Lemma B.6,
$\prod_{t=1}^{T}\mu(s_{t},a_{t})=$ $\ \prod_{t\in V_{T}}\mu(s_{t},a_{t})\prod_{t\not\in V_{T}}\mu(s_{t},a_{t})$
$\geq$ $\ \prod_{t\in V_{T}}\left(\mu^{*}(s_{t})-2Lr\right)\prod_{t\not\in V_{T}}\mu^{*}(s_{t})$
$=$ $\ \prod_{t=1}^{T}\left(\mu^{*}(s_{t})-c_{t}\right)$
Since $c_{t}\in[0,\mu_{0}/2]$ for all $t\in[T]$, Lemma B.7 implies that
$\prod_{t=1}^{T}\mu(s_{t},a_{t})\geq$ $\ \prod_{t=1}^{T}\mu^{*}(s_{t})-1.5\sum_{t=1}^{T}c_{t}$
$=$ $\ \prod_{t=1}^{T}\mu^{*}(s_{t})-3L|V_{T}|r$
If $|V_{T}|<2T\varepsilon(T)$, then
$R_{T}(\mathbf{s},\mu,\pi^{*})=$ $\ \prod_{t=1}^{T}\mu^{*}(s_{t})-\prod_{t=1}^{T}\mu(s_{t},a_{t})$
$\leq$ $\ 6LT\varepsilon(T)r$
$=$ $\ 6LT\varepsilon(T)^{(n+1)/n}$
$=$ $\ 6LT^{(1+\frac{-n}{n+0.5}\cdot\frac{n+1}{n})}$
$=$ $\ 6LT^{(\frac{n+0.5}{n+0.5}-\frac{n+1}{n+0.5})}$
$=$ $\ 6LT^{\frac{-0.5}{n+0.5}}$
Then with the help of Lemma B.4 (and noting that $R_{T}(\mathbf{s},\mu,\pi^{*})\leq 1$ always),
$\operatorname*{\mathbb{E}}_{\mathbf{s}\sim\mathcal{D}^{T}}\left[R_{T}(\mathbf{s},\mu,\pi^{*})\right]\leq$ $\ \Pr[|V_{T}|\geq 2T\varepsilon(T)]\cdot 6LT^{\frac{-0.5}{n+0.5}}+\Pr[|V_{T}|\geq 2T\varepsilon(T)]\cdot 1$
$\leq$ $\ 6LT^{\frac{-0.5}{n+0.5}}+\frac{1}{T}+\exp(-3T^{\frac{0.5}{n+0.5}}/8)$
Definition B.2.
Let $(V,||\cdot||)$ be a normed vector space and let $\delta>0$. Then $W\subseteq V$ is a $\delta$-packing of $V$ if for all $x,y\in W$, $||x-y||>\delta$. The $\delta$-packing number of $V$, denoted $\mathcal{M}(V,||\cdot||,\delta)$, is the maximum cardinality of any $\delta$-packing of $V$.
Lemma B.9 (Theorem 14.2 in [32]).
Let $B$ be the unit ball and let |||||$\cdot$||| | ⋅ | | denote Euclidean distance. Assume $V\subset\mathbb{R}^{n}$ is convex, and $\delta B\subset V$ for some $\delta>0$. Then
$\mathcal{M}(V,||\cdot||,\delta)\leq\left(\frac{3}{\delta}\right)^{n}\frac{\operatorname*{vol}(V)}{\operatorname*{vol}(B)}$
Lemma B.10.
[…]
Proof.
Let $\tilde{S}=\{s_{t}:t\in Q_{T}\text{ and }t>g(T)\}$ be the set of states queried after $t=g(T)$. Since we always query on the first $g(T)$ time steps, the total number of queries is $g(T)+|\tilde{S}|$.
We claim that for any $s,s^{\prime}\in\tilde{S}$, we have $||s-s^{\prime}||>r$. Suppose not, and WLOG assume $s$ was queried after $s^{\prime}$, with $s=s_{t}$. Since $s^{\prime}$ was queried earlier, $(s^{\prime},\pi^{*}(s^{\prime}))\in X_{t}$. Therefore
$\min_{(s^{\prime\prime},a)\in X_{t}}||s-s^{\prime\prime}||\leq||s-s^{\prime}||\leq r$
But then we would have taken the action $\pi_{t}(s_{t})$ instead of querying, which is a contradiction. Hence $\tilde{S}$ is a $r$-packing of $\operatorname*{conv}(\mathbf{s})$, and thus is also an $r$-packing of $\operatorname*{conv}(\mathbf{s}\cup rB)$. Since $rB\subset\operatorname*{conv}(\mathbf{s}\cup rB)$, Lemma B.9 implies that
$|\tilde{S}|\leq\mathcal{M}(\operatorname*{conv}(\mathbf{s}\cup rB),||\cdot||,r)\leq\left(\frac{3}{r}\right)^{n}\frac{\operatorname*{vol}(\operatorname*{conv}(\mathbf{s}\cup rB))}{\operatorname*{vol}(B)}=\frac{3^{n}\operatorname*{vol}(\operatorname*{conv}(\mathbf{s}\cup rB))}{\varepsilon(T)\operatorname*{vol}(B)}$
Theorem 4.5 follows from Lemmas B.8 and B.10.
Appendix C Proof of Theorem 5.2
C.1 Notation for the proof
We will use the following notation throughout the proof of Theorem 5.2:
  • Let $V_{T}=\{t\in[T]:\mu(s_{t},a_{t})<\mu^{*}(s)\}$ be the set of time steps with a suboptimal payoff.
  • Let $B_{t}$ be the bucket that is used on time step $t$ (as defined on line 14 of Algorithm 2).
  • Let $d(B)$ be the depth of bucket $B$
    • Buckets created on line 9 are depth 0.
    • We refer to $B_{1},B_{2}$ created on line 26 as the children of the bucket $B$ defined on line 14.
    • If $B^{\prime}$ is the child of $B$, $d(B^{\prime})=d(B)+1$.
[…]
  • Viewing the set of buckets are a binary tree defined by the “child” relation, we use the terms “ancestor” and ”descendant” in accordance with their standard tree definitions.
  • Let $\mathcal{B}_{V}=\{B:\exists t\in V_{T}\text{ s.t. }B_{t}=B\}$ be the set of buckets that ever produced a suboptimal payoff.
  • Let $\mathcal{B}_{V}^{\prime}=\{B\in\mathcal{B}_{V}:\text{no descendant of $B$ is in $\mathcal{B}_{V}$}\}$.
C.2 Proof roadmap
The proof proceeds in the following steps:
  1. 1.
    Bound the total number of buckets and therefore the total number of queries (Lemma C.1).
  2. 2.
    Bound the suboptimality on a single time step based on the bucket length and the Lipschitz constant (Lemma C.2).
  3. 3.
    Bound the sum of bucket lengths on time steps where we make a mistake (Lemma C.4), with Lemma C.3 as an intermediate step. This captures the total amount of suboptimality.
  4. 4.
    The bound from Lemma C.4 is in additive form, whereas the objective is multiplicative. Lemma B.7 factors the objective into one term for the optimal payoffs and one term for the suboptimality in the form of Lemma C.4.
  5. 5.
    Combine Lemmas C.2, C.4 and B.7 to bound the total regret (Lemma C.5).
  6. 6.
    Prove Theorem 5.2 using Lemma C.1, Lemma C.5, and some light arithmetic.
C.3 Proof
Lemma C.1.
Algorithm 2 performs at most $(4+\operatorname*{len}(\operatorname*{conv}(\mathbf{s})))g(T)$ queries.
Proof.
Algorithm 2 performs at most one query per bucket, so the total number of queries is bounded by the total number of buckets. There are two ways to create a bucket: from scratch (line 9), or by splitting an existing bucket (line 26).
Since depth 0 buckets overlap only at their boundaries, and each depth 0 bucket has length $1/g(T)$, at most $g(T)\max_{t,t^{\prime}\in[T]}|s_{t}-s_{t^{\prime}}|=g(T)\operatorname*{len}(\operatorname*{conv}(\mathbf{s}))$ depth 0 buckets are subsets of the interval $[\min_{t\in[T]}s_{t},\max_{t\in[T]}s_{t}]$. At most two depth 0 buckets are not subsets of that interval (one at each end), so the total number of depth 0 buckets is at most $g(T)\operatorname*{len}(\operatorname*{conv}(\mathbf{s}))+2$.
We split a bucket $B$ when $n_{B}$ reaches $T/g(T)$, which creates two new buckets. Since each time step increments $n_{B}$ for a single bucket $B$, and there are a total of $T$ time steps, the total number of buckets created via splitting is at most
$\frac{2T}{T/g(T)}=2g(T)$
Therefore the total number of buckets ever in existence is $(2+\operatorname*{len}(\operatorname*{conv}(\mathbf{s})))g(T)+2\leq(4+\operatorname*{len}(\operatorname*{conv}(\mathbf{s})))g(T)$, so Algorithm 2 performs at most $(4+\operatorname*{len}(\operatorname*{conv}(\mathbf{s})))g(T)$ queries. ∎
Lemma C.2.
For each $t\in[T]$, $\mu(s_{t},a_{t})\geq\mu^{*}(s_{t})-2L\operatorname*{len}(B_{t})$.
Proof.
If we query the mentor at time $t$, $\mu(s_{t},a_{t})=\mu^{*}(s_{t})$. Thus assume we do not query the mentor at time $t$: then there exists $s^{\prime}\in B_{t}$ (as defined on line 21 of Algorithm 2) such that $a_{t}=\pi(s^{\prime})=\pi^{*}(s^{\prime})$. Since $s_{t}$ and $s^{\prime}$ are both in $B_{t}$, $|s_{t}-s^{\prime}|\leq\operatorname*{len}(B_{t})$. Applying Lemma B.5 completes the proof. ∎
Lemma C.3.
If $\pi^{*}$ has at most $K$ segments, $|\mathcal{B}_{V}^{\prime}|\leq K$.
Proof.
If two buckets $B$ and $B^{\prime}$ overlap by more than at a single point, one must be the descendant of the other. Thus by definition of $\mathcal{B}_{V}^{\prime}$, any two buckets in $\mathcal{B}_{V}^{\prime}$ overlap by at most a single point.
Now consider any $B\in\mathcal{B}_{V}^{\prime}$. By definition, there exists $t\in V_{T}$ such that $s_{t}\in B$. Then there exists $s^{\prime}\in B$ (as defined in Algorithm 2) such that $a_{t}=\pi(s^{\prime})=\pi^{*}(s^{\prime})$. Since $t\in V_{T}$, we have $\pi^{*}(s_{t})\neq a_{t}=\pi^{*}(s^{\prime})$. Thus $s_{t}$ and $s^{\prime}$ are in different segments; let $S_{B}$ be the segment containing $\min(s_{t},s^{\prime})$. Then $\min(s_{t},s^{\prime})\leq\max(S_{B})<\max(s_{t},s^{\prime})\leq\max(B)$.
Consider any $B,B^{\prime}\in\mathcal{B}_{V}^{\prime}$ with $B\neq B^{\prime}$ and WLOG assume $\max(B)<\max(B^{\prime})$. Then $\max(S_{B})<\max(B)\leq\max(S_{B}^{\prime})$, so $S_{B}\neq S_{B^{\prime}}$. Thus the number of segments in $\pi^{*}$ is at least $\mathcal{B}_{V}^{\prime}$. Since the number of segments in $\pi^{*}$ is at most $K$, we must have $|\mathcal{B}_{V}^{\prime}|\leq K$. ∎
Lemma C.4.
[…]
Proof.
For every $t\in V_{T}$, we have $B_{t}=B$ for some $B\in\mathcal{B}_{V}$, so
$\sum_{t\in V_{T}}\operatorname*{len}(B_{t})=\sum_{B\in\mathcal{B}_{V}}\ \sum_{t\in V_{T}:B=B_{t}}\operatorname*{len}(B_{t})$
Next, observe that every $B\in\mathcal{B}_{V}\setminus\mathcal{B}_{V}^{\prime}$ must have a descendent in $\mathcal{B}_{V}^{\prime}$: otherwise we would have $B\in\mathcal{B}_{V}^{\prime}$. Let $\mathcal{A}(B)$ denote the set of ancestors of $B$, plus $B$ itself. Then we can write
$\sum_{t\in V_{T}}$ $\ \operatorname*{len}(B_{t})\leq\sum_{B^{\prime}\in\mathcal{B}_{V}^{\prime}}\ \sum_{B\in\mathcal{A}(B^{\prime})}\ \sum_{t\in V_{T}:B=B_{t}}\operatorname*{len}(B_{t})$
$=$ $\ \sum_{B^{\prime}\in\mathcal{B}_{V}^{\prime}}\ \sum_{B\in\mathcal{A}(B^{\prime})}\ |\{t\in V_{T}:B=B_{t}\}|\cdot\operatorname*{len}(B_{t})$
[…]
$\sum_{B\in\mathcal{A}(B^{\prime})}$ |$\ \frac{|\{t\in V_{T}:B=B_{t}\}|}{g(T)2^{d(B)}}\leq\frac{T}{g(T)^{2}}\sum_{B\in\mathcal{A}(B^{\prime})}\ \frac{1}{2^{d(B)}}$
$=$ $\ \frac{T}{g(T)^{2}}\sum_{d=0}^{d(B^{\prime})}\frac{1}{2^{d}}\leq\frac{T}{g(T)^{2}}\sum_{d=0}^{\infty}\frac{1}{2^{d}}=\frac{2T}{g(T)^{2}}$
Then by Lemma C.3,
$\sum_{t\in V_{T}}\operatorname*{len}(B_{t})\leq\sum_{B^{\prime}\in\mathcal{B}_{V}^{\prime}}\frac{2T}{g(T)^{2}}=\frac{2T|\mathcal{B}_{V}^{\prime}|}{g(T)^{2}}\leq\frac{2KT}{g(T)^{2}}$
as claimed. ∎
Lemma C.5.
For any $\mathbf{s}\in S^{T}$, any Lipschitz $\mu$, any $\pi^{*}$ with at most $K$ segments, and any $g:\mathbb{N}\to\mathbb{N}$, Algorithm 2 satisfies
$R_{T}(\mathbf{s},\mu,\pi^{*})\leq\frac{6L}{g(T)}\max\left(1,\frac{TK}{g(T)}\right)$
Proof.
Let $\mu_{0}=4L/g(T)$. Suppose there exists $t\in[T]$ such that $\mu^{*}(s_{t})<\mu_{0}$. Then
$R_{T}(\mathbf{s},\mu,\pi^{*})=\prod_{t=1}^{T}\mu^{*}(s_{t})-\prod_{t=1}^{T}\mu(s_{t},a_{t})\leq\mu_{0}-0\leq\frac{6L}{g(T)}$
so the claim is satisfied. Thus assume $\mu^{*}(s_{t})\geq\mu_{0}$ for all $t\in[T]$. Let $c_{t}=0$ for $t\not\in V_{T}$ and $c_{t}=2L\operatorname*{len}(B_{t})$ for $t\in V_{T}$. Then by Lemma C.2,
$\prod_{t=1}^{T}\mu(s_{t},a_{t})=$ $\ \prod_{t\in V_{T}}\mu(s_{t},a_{t})\prod_{t\not\in V_{T}}\mu(s_{t},a_{t})$
$\geq$ $\ \prod_{t\in V_{T}}\left(\mu^{*}(s_{t})-2L\operatorname*{len}(B_{t})\right)\prod_{t\not\in V_{T}}\mu^{*}(s_{t})$
$=$ $\ \prod_{t=1}^{T}\left(\mu^{*}(s_{t})-c_{t}\right)$
Note that $\operatorname*{len}(B_{t})\leq 1/g(T)$ for all $t\in[T]$, so $c_{t}\in[0,\mu_{0}/2]$. Thus by Lemma B.7,
$\prod_{t=1}^{T}\mu(s_{t},a_{t})\geq$ $\ \prod_{t=1}^{T}\mu(s_{t},a_{t})-1.5\sum_{t=1}^{T}c_{t}$
$=$ $\ \prod_{t=1}^{T}\mu(s_{t},a_{t})-3L\sum_{t\in V_{T}}\operatorname*{len}(B_{t})$
Then by Lemma C.4,
$\prod_{t=1}^{T}\mu(s_{t},a_{t})\geq$ $\ \prod_{t=1}^{T}\mu^{*}(s_{t})-\frac{6LKT}{g(T)^{2}}$
and so
$R_{T}(\mathbf{s},\mu,\pi^{*})=\prod_{t=1}^{T}\mu^{*}(s_{t})-\prod_{t=1}^{T}\mu(s_{t},a_{t})\leq\frac{6LKT}{g(T)^{2}}$
as required. ∎
See 5.2
Proof.
Lemma C.1 implies the query bound. Since $KT\geq T\geq g(T)$, Lemma C.5 implies that for any such $\mathbf{s},\mu$, and $\pi^{*}$,
$R_{T}(\mathbf{s},\mu,\pi^{*})\leq$ $\ \frac{6L}{g(T)}\max\left(1,\frac{TK}{g(T)}\right)=\frac{6L}{g(T)}\cdot\frac{TK}{g(T)}$
$=$ $\ \frac{6LTK}{T^{2c}}=6LKT^{1-2c}$

### v3

@@
-Avoiding Catastrophe in Continuous Spaces by Asking for Help
+Avoiding Catastrophe in Online Learning by Asking for Help
 Abstract
-Most reinforcement learning algorithms with formal regret guarantees assume all mistakes are reversible and essentially rely on trying all possible behaviors.
-This approach leads to poor outcomes when some mistakes are irreparable or even catastrophic.
-We propose a variant of the contextual bandit problem where the goal is to minimize the chance of catastrophe.
-Specifically, we assume that the payoff each round represents the chance of avoiding catastrophe that round, and try to maximize the product of payoffs (the overall chance of avoiding catastrophe).
-We allow a limited number of queries to a mentor and assume a Lipschitz continuous payoff function.
+Most learning algorithms with formal regret guarantees assume that no mistake is irreparable and essentially rely on trying all possible behaviors.
+This approach is problematic when some mistakes are catastrophic, i.e., irreparable.
+We propose an online learning problem where the goal is to minimize the chance of catastrophe.
+Specifically, we assume that the payoff in each round represents the chance of avoiding catastrophe that round and aim to maximize the product of payoffs (the overall chance of avoiding catastrophe) while allowing a limited number of queries to a mentor.
 We first show that in general, any algorithm either constantly queries the mentor or is nearly guaranteed to cause catastrophe.
-However, when the mentor policy class has bounded Natarajan dimension and contains at least some “reasonable” policies, we provide an algorithm whose regret and rate of querying the mentor both approach 0 as the time horizon grows.
-We also present an alternative algorithm which provides the same regret and query guarantees when the mentor’s action changes a constant number of times in a 1D state space, and can handle adversarially chosen states.
+However, in settings where the mentor policy class is learnable in the standard online learning model, we provide an algorithm whose regret and rate of querying the mentor both approach 0 as the time horizon grows.
+Conceptually, if a policy class is learnable in the absence of catastrophic risk, it is learnable in the presence of catastrophic risk if the agent can ask for help.
 1 Introduction
-In reinforcement learning (RL), an agent finds itself in an unknown environment and wants to maximize some payoff function, commonly modeled as a Markov Decision Process (MDP).
-Typically, the only way for the agent to learn about its environment is by trying actions and observing their effects.
-This presents a challenge when some actions are inherently dangerous: for example, the agent must learn to avoid self-destructive actions without actually trying those actions.
-This problem is hopeless without additional assumptions, since any untried action could lead to either paradise or disaster, and the agent has no way to predict which.
-Most RL theory evades this issue by either assuming ergodicity, meaning that any state is reachable from any other (e.g., [15]), or by resetting the agent at the start of each “episode” (e.g., [1]).
-In the real world, however, many actions lead to unrecoverable situations: we call these catastrophes.
-Examples could include ingesting poison or crashing a robotic vehicle.
-Resetting the agent may be expensive (if it destroys itself), inadequate (if it irreparably harms others), or impossible (if it is outside of human supervision).
-Rather than ignore the possibility of catastrophe, we explore whether access to a mentor might enable an agent to navigate risky (i.e., non-ergodic single-episode) environments.
-We assume the mentor illustrates the optimal action in the current state, and we would like an algorithm which obtains near-optimal payoff with a limited number of mentor queries.
-“Near-optimal payoff” is often modeled via regret, defined as the gap between the optimal total payoff and the agent’s actual total payoff.
-For undiscounted MDPs (i.e., each time step has the same weight in the objective), one typically aims for regret sublinear in the time horizon $T$, or equivalently, for the average regret per time step to approach 0 for large $T$.
-Similarly, a sublinear number of mentor queries implies that the rate of querying the mentor goes to 0 for large $T$—that is, the agent gradually becomes self-sufficient.
-Formally, we pose the following open problem:
-Is there an algorithm for non-ergodic single-episode undiscounted MDPs which ensures that both the regret and the number of mentor queries are sublinear in $T$?
-An additional assumption is necessary, as the agent needs some way to make inferences about unseen states in order to decide when to query.
+There has been mounting concern over catastrophic risk from AI, including but not limited to autonomous weapon accidents [1], bioterrorism [30], and cyberattacks on critical infrastructure [14].
+See [10] and [20] for taxonomies of societal-scale AI risks.
+In this paper, we use “catastrophe” to refer to any kind of irreparable harm.
+This definition also covers smaller-scale (yet still unacceptable) incidents such as serious medical errors [11], crashing a robotic vehicle [24], or discriminatory sentencing [39].
+The gravity of these risks contrasts starkly with the dearth of theoretical understanding of how to avoid them.
+Nearly all of learning theory explicitly or implicitly assumes that no single mistake is too costly.
+We focus on online learning, where an agent repeatedly interacts with an unknown environment and uses its observations to gradually improve its performance.
+Most online learning algorithms essentially try all possible behaviors and see what works well.
+We do not want autonomous weapons or surgical robots to try all possible behaviors.
+More precisely, trial-and-error-style algorithms only work when catastrophe is assumed to be impossible.
+This assumption manifests differently in different subtypes of online learning.
+In the standard online learning model, the agent’s actions have no permanent effect on the environment.
+[^1] More precisely, the state can depend on the agent’s previous actions, but the agent’s performance is always evaluated with respect to the optimal policy on the same sequence of states.
+Online reinforcement learning allows the agent’s actions to permanently affect the environment, but typically assumes that either no action has irreversible effects (e.g., [21]) or that the agent is reset at the start of each “episode” (e.g., [2]).
+One could train an agent entirely in a controlled lab setting where the above assumptions do hold, but we argue that sufficiently general agents will inevitably encounter novel scenarios when deployed in the real world.
+Machine learning models often behave unpredictably in unfamiliar environments (see, e.g., [32]), and we do not want AI biologists or robotic vehicles to behave unpredictably.
+The goal of this paper is to understand the conditions under which it is possible to formally guarantee avoidance of catastrophe in online learning.
+Certainly some conditions are necessary, because if the agent can only learn by trying actions directly, the problem is hopeless: any untried action could lead to paradise or disaster and the agent has no way to predict which.
+In the real world, however, one needn’t learn through pure trial-and-error: one can also ask for help.
+1.1 Our model
+We propose an online learning model of avoiding catastrophe with mentor help.
+On each time step, the agent observes a state, selects an action (or queries the mentor), and obtains a payoff.
+[^2] Depending on the field, states are also known as inputs, instances, or contexts, actions are also known as outputs, labels, or arms, and payoffs are also known as rewards or losses.
+Each payoff represents the probability of avoiding catastrophe on that time step, and the agent’s goal is to maximize the product of payoffs, which corresponds to minimizing the total chance of catastrophe.
+We consider the product of payoffs obtained while learning, not the product of payoffs of some final policy: finding a catastrophe-avoiding policy is futile if catastrophe occurs while learning.
+The (possibly suboptimal) mentor has a fixed policy, and when queried, the mentor illustrates their policy’s action in the current state.
+We desire an agent whose regret – defined as the gap between the mentor’s performance and the agent’s performance – approaches zero as the time horizon $T$ grows.
+In other words, with enough time, the agent should avoid catastrophe nearly as well as the mentor.
+We also expect the agent to become self-sufficient over time: formally, the number of queries to the mentor should be sublinear in $T$, or equivalently, the rate of querying the mentor should go to zero.
+1.2 Our assumptions
+The agent needs some way to make inferences about unseen states in order to decide when to ask for help.
 Much past work has used Bayesian inference, which suffers tractability issues in complex environments.
 [^3] For the curious reader, [4] provides a thorough treatment.
-See also Section 1.3 for further discussion.
-In this paper, we use Lipschitz continuity instead, which captures the intuition that an action should have similar effects in similar states.
-Importantly, the Lipschitz assumption is compatible with any state space and action space with a distance metric.
-1.1 Our model
-In this paper, we target a more modest goal: can the agent simply avoid catastrophe?
-This is clearly necessary (but not sufficient) for sublinear regret in general MDPs.
-Furthermore, avoiding catastrophe is the precise ability which standard RL algorithms lack.
-Focusing on catastrophe also allows us to streamline the model.
-If we redefine the payoff at time $t$ to be the probability of avoiding catastrophe on that time step, then the product of payoffs represents the overall chance of avoiding catastrophe.
-Representing catastrophe in the objective function eliminates the need to model non-ergodicity through the transition function, as MDPs do.
-Our multiplicative objective captures the same stringency: a single mistake can ruin the entire endeavor.
-[^2] An additive objective with unbounded payoffs could also capture the core challenge of needing to avoid catastrophic actions without ever trying them.
-However, we feel that the multiplicative objective is more intuitive due its interpretation as the probability of avoiding catastrophe.
-The resulting model is similar to the contextual bandit problem: on each time step, the agent observes a state (we generally assume i.i.d.
-states), selects an action (or queries the mentor), and obtains some payoff.
-To our knowledge, multiplicative objective functions have not been studied for contextual bandits.
-Lastly, unlike the standard contextual bandit model, we assume that the agent does not observe payoffs.
-This is because the payoff in our model represents the chance of avoiding catastrophe that round, and in the real world, one only observes whether catastrophe occurred, not its probability.
+See also Section 2.
+In this paper, we instead assume that the mentor policy satisfies a Lipschitz-flavored condition we call “local generalization”: informally, if the mentor told us that action $a$ was safe in a nearby state, then $a$ is probably (but not certainly) also safe in the current state.
+Local generalization captures the intuition that one can transfer knowledge between similar states but the generalization becomes progressively less reliable as the states become less similar.
+Unlike Bayesian inference, local generalization only requires computing distances and is compatible with any state space which admits a distance metric.
+Unlike the standard online learning model, we assume that the agent does not observe payoffs.
+This is because the payoff in our model represents the chance of avoiding catastrophe on that time step.
+In the real world, one only observes whether catastrophe occurred, not its probability.
 [^4] One may be able to detect “close calls” in some cases, but observing the precise probability seems unrealistic.
+1.3 Standard online learning
+An overview of standard online learning is in order before discussing our results.
+In the standard model, the agent observes a state on each time step and must choose an action.
+An adversary then reveals the correct action, which results in some payoff for the agent.
+The goal is sublinear regret, i.e., the average regret per time step should go to 0 as $T\to\infty$.
+Figure 1 delineates the precise differences between the standard model and our model.
+If the adversary’s choices are unconstrained, the problem is hopeless: if the adversary determines the correct action on each time step randomly and independently, the agent can do no better than random guessing.
+However, sublinear regret becomes possible if (1) the hypothesis class has finite Littlestone dimension [26], or (2) the hypothesis class has finite VC dimension [38] and the input is $\sigma$-smooth [^5] Informally, the adversarial chooses a distribution over states instead of a precise state.
+See Section 3.
+[17].
+The goal of sublinear regret in online learning implicitly assumes catastrophe is impossible: the agent can make arbitrarily many (and arbitrarily costly) mistakes as long as the average regret per time step goes to 0.
+In contrast, we demand subconstant regret: the total probability of catastrophe should go to 0.
+Furthermore, standard online learning allows the agent to observe payoffs on every time step, while our agent only receives feedback on time steps with queries.
+However, access to a mentor (and local generalization) allows our agent to learn without trying actions directly, which is enough to offset all of the above disadvantages.
+Objective Regret goal Feedback Mentor Local gen.
+Standard model Sum of payoffs Sublinear Every time step No No
+Our model Product of payoffs Subconstant Only from queries Yes Yes
+Figure 1: Comparison between the standard online learning model and our model.
+On a technical level, sublinear vs subconstant regret is a more important distinction than the sum vs product of payoffs: for example, taking a logarithm can transform the product into a sum.
+However, the multiplicative objective and the goal of subconstant regret are tightly related, since the former’s interpretation as the chance of catastrophe is what gives rise to the latter: we want the total (not average) chance of catastrophe to go to 0.
 1.4 Our results
-We first show that in general, any algorithm with sublinear queries to the mentor has arbitrarily poor regret in the worst-case (Theorem 3.1).
+Informally, we show that avoiding catastrophe with the help of a mentor and local generalization is no harder than online learning without catastrophic risk.
+More precisely, we first show that in general, any algorithm with sublinear queries to the mentor has arbitrarily poor regret in the worst-case (Theorem 4.1).
 As a corollary, even when it is possible to avoid catastrophe with certainty, any algorithm either asks for help on a constant fraction of time steps or is nearly guaranteed to cause catastrophe.
-This holds even for randomized algorithms and simple environments: our construction uses just two actions and assumes states are uniformly random across $[0,1]$.
-On the other hand, the class of mentor policies in our negative result is quite complex in the sense of having unbounded VC dimension [31].
-What if we restrict the mentor policy to be relatively simple?
-We present an intuitive algorithm whose regret and rate of querying the mentor both go to 0 as $T\to\infty$ when the mentor policy class has bounded VC dimension (or Natarajan dimension, more generally [^4] Natarajan dimension [24] generalizes VC dimension (which only handles binary prediction) to multi-class prediction.
-In our case, each class is an action, and a policy is a multi-class predictor from states to actions.) and contains at least some “reasonable” policies (Theorem 4.5).
-Our algorithm can handle an arbitrary unknown distribution over states, a multi-dimensional unbounded state space, and many actions.
-It does not need detailed access to the feature embedding (instead relying on two simple operations) and does not need to know the Lipschitz constant.
-It does need to know the policy class and (an upper bound on) the Natarajan dimension (as is standard in PAC learning).
-This is our primary positive result.
-Finally, we give another algorithm with the same guarantees of sublinear queries and near-zero regret under a different simplicity assumption (Theorem 5.2).
-Specifically, we assume that the mentor’s action changes a constant number of times in a 1D state space.
-This algorithm can also handle adversarially chosen states.
-We generally assume that the mentor’s policy is optimal, but our positive results can be generalized to a suboptimal mentor under a slightly different Lipschitz assumption (Remark B.1).
-1.3 Related work
-See [9; 11] for full surveys of safe RL. Here, we discuss only the work most relevant to our own.
+Unlike online learning where the general impossibility result is trivial (the agent might as well guess randomly given an unconstrained adversary), local generalization significantly limits the adversary’s power and necessitates a careful construction.
+Next, we present a simple algorithm whose total regret and rate of querying the mentor both go to 0 as $T\to\infty$ when either (1) the mentor policy class has finite Littlestone dimension or (2) the mentor policy class has finite VC dimension and the state sequence is $\sigma$-smooth.
+Our algorithm can handle a multi-dimensional unbounded state space and does not need detailed access to the feature embedding, instead using two simple operations.
+It does need to know the mentor policy class, as is standard in online learning.
+We initially prove the theorem for binary actions (Theorem 5.2) and then reduce learning with (finitely) many actions to the binary action case (Theorem C.1).
+Finally, Appendix D presents another algorithm with the same guarantees of subconstant regret and sublinear queries under different assumptions.
+Most importantly, this result applies to the class of thresholds on $[0,1]$, which is known to be hard for standard online learning (see Section 2).
+This shows that our problem and standard online learning do not have exactly the same difficulty, but rather our problem is no harder than online learning.
+The rest of the paper is structured as follows.
+Section 2 discusses related work.
+Section 3 formally defines our model.
+Section 4 presents our negative result for general mentor policies.
+Section 5 presents our positive result for simple mentor policy classes.
+Proofs are deferred to the appendix.
+2 Related work
 Learning with irreversible costs.
-Despite the ubiquity of irreversible costs in the real world, there remains a dearth of theoretical work on this topic.
+Despite the ubiquity of irreparable/irreversible costs in the real world, theoretical work on this topic remains limited.
 This may be due to the fundamental modeling question of how the agent should learn about novel states or actions without actually trying them.
@@
 A finite set of possible environments satisfies both conditions, but is unrealistic in many real-world scenarios.
-In contrast, our algorithm can handle a continuous unbounded state space, which is crucial for many real-world scenarios in which one never sees the exact same state twice.
-Bayesian inference combined with asking for help is studied by [5; 6; 17; 21].
-We also mention [12; 22; 30], which utilize Bayesian inference in the context of safe RL without asking for help (and without regret bounds).
+In contrast, our algorithm can handle an uncountable policy class and a continuous unbounded state space, which is crucial for many real-world scenarios in which one never sees the exact same state twice.
+Bayesian inference combined with asking for help is studied by [8; 9; 25; 28].
+We also mention [15; 29; 37], which utilize Bayesian inference in the context of safe (online) reinforcement learning without asking for help (and without regret bounds).
 We are only aware of two papers which theoretically address irreversibility without Bayesian inference: [13] and [27].
@@
 However, the regret bound and number of rollouts are both linear in the worst case, due to the dependence on the $\gamma^{*}$ parameter which roughly captures how bad an irreversible action can be.
-In contrast, our algorithms achieve near-zero regret even when actions can be maximally bad.
+In contrast, our algorithm achieves good regret even when actions are maximally bad.
 To our knowledge, we are the first to provide an algorithm which formally guarantees avoidance of catastrophe (with high probability) without Bayesian inference.
-Contextual bandits.
-Like MDPs, the contextual bandit model involves an agent who selects an action on each time step based on the current state (or “context”) to maximize payoff.
-Unlike MDPs, the state on each time step is determined randomly or by an adversary, instead of by a transition function.
-See [29] for an overview.
-The contextual bandit literature often uses Lipschitz continuity of payoff functions as a way to generalize between states, which affirms our use of that same assumption [4; 14; 18; 19; 26; 28].
-However, to our knowledge, all prior contextual bandit work aims to maximize the sum of payoffs, which implicitly assumes that no mistake is catastrophic: no matter how poorly you perform, you can always make it up given a sufficiently long time horizon.
+Online learning.
+See [7] and Chapter 21 of [34] for introductions to online learning.
+A classical result states that sublinear regret is possible iff the hypothesis class has finite Littlestone dimension [26].
+However, even some simple hypothesis classes have infinite Littlestone dimension, such as the class of thresholds on $[0,1]$ (Example 21.4 in [34]).
+Recently, [17] showed that if the adversary only chooses a distribution over states rather than the precise state, only the weaker assumption of finite VC dimension [38] is needed for sublinear regret.
+Specifically, they assume that each state is sampled from a distribution whose concentration is upper bounded by $\frac{1}{\sigma}$ times the uniform distribution.
+This framework is known as smoothed analysis.
+Smoothed analysis.
+Smoothed analysis was originally proposed by [36] to explain the fact that some algorithms almost always perform well in practice, despite the theoretical existence of pathological inputs.
+Informally, algorithms which perform well under smoothed analysis are guaranteed to perform well for nearly all – but not necessarily all – inputs.
+Active learning.
+Our assumption that the agent only receives feedback in response to queries falls under the umbrella of active learning [18].
+This contrasts with passive learning, where the agent receives feedback automatically.
+Although ideas from active learning could be useful in our domain, we are not aware of any results from that literature which account for irreversible costs.
 Multiplicative objectives.
-Perhaps the best-studied example of a multiplicative objective is the Nash welfare function (also known as the Nash bargaining solution), defined as the geometric mean of agent utilities [16; 23].
-Inspired by Nash welfare, [2] recently studied algorithms for maximizing the geometric mean of payoffs across time steps in the classic (non-contextual) multi-armed bandit problem.
+Perhaps the best-studied example of a multiplicative objective is the Nash welfare function (also known as the Nash bargaining solution), defined as the geometric mean of agent utilities [23; 31].
+Inspired by Nash welfare, [3] recently studied algorithms for maximizing the geometric mean of payoffs across time steps in the multi-armed bandit problem.
 However, the geometric mean of payoffs is much more forgiving of errors compared to the product of payoffs.
 Imagine that the agent obtains a payoff of $1-1/\sqrt{T}$ per time step: then $\lim_{T\to\infty}(1-1/\sqrt{T})^{T}=0$ but $\lim_{T\to\infty}((1-1/\sqrt{T})^{T})^{1/T}=\lim_{T\to\infty}(1-1/\sqrt{T})=1$.
-In our setting, it is crucial that the cumulative chance of catastrophe (the product of payoffs) is small, not just the average chance of catastrophe per time step (the geometric mean of payoffs).
-Multiplicative objectives also arise in the context of compound growth or repeated gambles, where one’s wealth at time $t$ can be modeled as the initial wealth times a product of growth factors (see, e.g., [25]).
-Our problem could be viewed as a compound growth problem where one’s “wealth” (the total probability of avoiding catastrophe) never increases.
-However, we are not aware of any mathematical tools related to compound growth that handle the contextual bandit aspects of our problem.
-The rest of the paper is structured as follows.
-Section 2 formally defines our model.
-Section 3 presents our negative result for general mentor policies.
-Section 4 presents our positive result for policy classes with bounded Natarajan dimension.
-Section 5 presents our second positive result under a different simplicity assumption.
-Section 6 discusses limitations.
-All proofs are deferred to the appendix.
+In our setting, it is crucial that the total chance of catastrophe (the product of payoffs) is small, not just the average chance of catastrophe per time step (the geometric mean of payoffs).
+This is closely related to the difference between sublinear and subconstant regret discussed in Section 1.3.
 3 Model
-Lipschitz contextual bandits.
-Let $\mathbb{N}$ refer to the strictly positive integers and let $T\in\mathbb{N}$ be the time horizon.
-Let $S\subseteq\mathbb{R}^{n}$ be the state space, and let $\mathbf{s}=(s_{1},s_{2},\dots s_{T})\in S^{T}$ be a sequence of states.
-(The contextual bandit literature typically uses “context” instead of “state”, but we have chosen to use “state” to align with MDP terminology.) With the exception of Section 5, we assume that each state is sampled i.i.d.
-from an unknown distribution $\mathcal{D}$ over $S$.
+States.
+Let $\mathbb{N}$ refer to the set of strictly positive integers and let $T\in\mathbb{N}$ be the time horizon.
+Let $S\subseteq\mathbb{R}^{n}$ be the state space and let $\mathbf{s}=(s_{1},s_{2},\dots s_{T})\in S^{T}$ be the sequence of states.
+In the adversarial setting, each $s_{t}$ can have arbitrary dependence on the events of prior time steps.
+In the smoothed setting, the adversary only chooses the distribution $s_{t}$ from which $s_{t}$ is sampled.
+Formally, a distribution $\mathcal{D}$ over $S$ is $\sigma$-smooth if for any $X\subseteq S$, $\mathcal{D}(X)\leq\frac{1}{\sigma}U(X)$.
+(In the smoothed setting, we assume that $S$ supports a uniform distribution $U$ [^6] For example, $S$ having finite Lebesgue measure is sufficient.
+Note that this does not imply boundedness.
+Alternatively, $\sigma$-smoothness can be defined with respect to a different distribution, as long as the Radon-Nikodym derivative is uniformly bounded; see Definition 1 of [5].) If each $s_{t}$ is sampled from a $\sigma$-smooth $\mathcal{D}_{t}$, we say that $\mathbf{s}$ is $\sigma$-smooth.
+The sequence $\boldsymbol{\mathcal{D}}=\mathcal{D}_{1},\dots,\mathcal{D}_{T}$ can still be adaptive, i.e., the choice of $\mathcal{D}_{t}$ can depend on the events of prior time steps.
+Actions.
+Let $A$ be a finite set of actions.
+There also exists a special action $\hat{a}$ which corresponds to querying the mentor.
 For $k\in\mathbb{N}$, let $[k]=\{1,2,\dots,k\}$.
 On each time step $t\in[T]$, the agent must select an action $a_{t}\in\ A\cup\{\hat{a}\}$, which generates payoff $\mu(s_{t},a_{t})\in[0,1]$.
-We allow the agent’s choice of action to be randomized.
-The action $\hat{a}$ constitutes querying the mentor, which we will define shortly.
+Asking for help.
+The mentor is endowed with a (not necessarily optimal) policy $\pi^{m}:S\to A$.
+When action $\hat{a}$ is chosen, the mentor informs the agent of the action $\pi^{m}(s_{t})$ and the agent obtains payoff $\mu(s_{t},\pi^{m}(s_{t}))$.
+For brevity, let $\mu^{m}(s)=\mu(s,\pi^{m}(s))$.
 The agent never observes payoffs: the only way to learn about $\mu$ is by querying the mentor.
-We assume that $|A|\leq T$, since the agent can never observe more than $T$ distinct actions.
-We assume that $\mu$ is Lipschitz continuous with respect to $S$: there exists $L>0$ such $\forall s,s^{\prime}\in S,a\in A$, $|\mu(s,a)-\mu(s^{\prime},a)|\leq L||s-s^{\prime}||$, where |||||$\cdot$||| | ⋅ | | denotes Euclidean distance.
-Asking for help.
-The special action $\hat{a}$ constitutes querying a mentor who knows an optimal policy $\pi^{*}:S\to A$.
-When action $\hat{a}$ is chosen, the mentor informs the agent of the action $\pi^{*}(s_{t})$ and the agent obtains optimal payoff $\mu(s_{t},\pi^{*}(s_{t}))=\mu^{*}(s_{t})=\max_{a\in A}\mu(s_{t},a)$.
-Let $Q_{t}=\{i\in[t]:a_{i}=\hat{a}\}$ be the set of time steps up to and including $t$ which involve a query to the mentor.
 We would like an algorithm which becomes “self-sufficient” over time: the rate of querying the mentor should go to 0 as $T\to\infty$.
-Intuitively, this is equivalent to the cumulative number of queries being sublinear in $T$: $|Q_{T}|\in o(T)$.
-However, interpreted literally, $|Q_{T}|\in o(T)$ is not a rigorous definition.
-This is because $|Q_{T}|$ is not just a function of $T$: it can also depend on $\mathbf{s}$, $\mu$ and $\pi^{*}$.
-To formalize this, let $|Q_{T}(\mathbf{s},\mu,\pi^{*})|$ be the number of queries to the mentor up to time $T$, given payoff function $\mu$ and state sequence $\mathbf{s}$.
-Then we say that $|Q_{T}|$ is sublinear in $T$, denoted $|Q_{T}|\in o(T)$, if there exists $g:\mathbb{N}\to\mathbb{N}$ such that $g(T)\in o(T)$ and $|Q_{T}(\mathbf{s},\mu,\pi^{*})|\leq g(T)$ for all $\mathbf{s},\mu,\pi^{*}$.
+Intuitively, this is equivalent to the cumulative number of queries being sublinear in $T$: $|Q_{T}|\in o(T)$, where $Q_{T}=\{t\in[T]:a_{t}=\hat{a}\}$.
+However, interpreted literally, $|Q_{T}|\in o(T)$ is not a rigorous statement.
+This is because $|Q_{T}|$ is not just a function of $T$: it can also depend on $\mathbf{s}$, $\mu$, $\pi^{m}$, and any randomness in the agent’s actions.
+To formalize this, let $Q_{T}(\mathbf{s},\mu,\pi^{m})$ be the set of time steps with queries given $\mathbf{s},\mu$, and $\pi^{m}$.
+Then we say that $|Q_{T}|$ is sublinear in $T$, denoted $|Q_{T}|\in o(T)$, if there exists $g:\mathbb{N}\to\mathbb{N}$ such that $g(T)\in o(T)$ and $|Q_{T}(\mathbf{s},\mu,\pi^{m})|\leq g(T)$ for all $\mathbf{s},\mu,\pi^{m}$ (and for all possible realizations of the agent’s randomization).
+However, we will mostly write $Q_{T}$ and $|Q_{T}|\in o(T)$ for brevity.
+Local generalization.
+We assume that the mentor policy permits “local generalization”.
+Formally, we assume there exists $L>0$ such that for all $s,s^{\prime}\in S$, $|\mu^{m}(s)-\mu(s,\pi^{m}(s^{\prime}))|\leq L||s-s^{\prime}||$, where |||||$\cdot$||| | ⋅ | | denotes Euclidean distance.
+Informally, if the agent is in state $s$, taking the mentor action from a nearby state $s^{\prime}$ is almost as good as taking the mentor action for $s$.
+In other words, it’s okay to use an action you learned from a nearby state.
+This assumption might initially feel less intuitive than a standard Lipschitz continuity assumption of $|\mu(s,a)-\mu(s^{\prime},a)|\leq L||s-s^{\prime}||$.
+However, Lipschitz continuity is strictly stronger than local generalization when the mentor is optimal (Proposition E.1).
+The properties are incomparable in general, but Lipschitz continuity allows odd situations where the mentor provides an action which is only acceptable in that specific state.
+This sets the agent up for failure: even if such actions exist, a thoughtful mentor should avoid those actions when possible.
+See Figure 2 for an illustration.
+[…]
+[…]
+[…]
+Figure 2: The difference between local generalization and Lipschitz continuity.
+The x-axis corresponds to a one-dimensional state space, with $s_{1}$ and $s_{2}$ denoting two nearby states.
+The y-axis is the payoff.
+In all figures, $\pi^{m}(s_{1})=0$ and $\pi^{m}(s_{2})=1$.
+Left: What one might consider the “typical case” with a near-optimal mentor.
+The mentor takes action 0 in $s_{1}$, so $|\mu(s_{2},0)-\mu(s_{2},\pi^{m}(s_{2})|$ is small (local generalization) and $|\mu(s_{2},0)-\mu(s_{1},0)|$ is small (Lipschitz continiuty).
+Middle: Local generalization without Lipschitz continuity.
+Action 0 is much worse in $s_{2}$, but it is still nearly as good as the mentor’s action in $s_{2}$, which is also quite bad.
+We primarily care about the agent’s performance relative to the mentor.
+Right: Lipschitz continuity without local generalization.
+The payoff of each action is similar in nearby states.
+However, despite the mentor recommending action 0 in $s_{1}$, action 0 is much worse than the mentor’s action in $s_{2}$.
+The mentor is setting up the agent for failure by recommending action 0 in $s_{1}$.
 Multiplicative objective and regret.
-If $\mu(s_{t},a_{t})\in[0,1]$ is the chance of avoiding catastrophe on time step $t$ (conditioned on no prior catastrophe), then $\prod_{t=1}^{T}\mu(s_{t},a_{t})$ represents the agent’s overall chance of avoiding catastrophe.
+If $\mu(s_{t},a_{t})\in[0,1]$ is the chance of avoiding catastrophe on time step $t$ (conditioned on no prior catastrophe), then $\prod_{t=1}^{T}\mu(s_{t},a_{t})$ is the agent’s overall chance of avoiding catastrophe.
 [^7] Conditioning on no prior catastrophe means we do not need to assume that these probabilities are independent (and if catastrophe has already occurred, this time step does not matter).
 This is due to the chain rule of probability.
-For a time horizon $T$, payoff function $\mu$, mentor policy $\pi^{*}$, and state sequence $\mathbf{s}$, the agent’s regret is
-$R_{T}(\mathbf{s},\mu,\pi^{*})=\prod_{t=1}^{T}\mu^{*}(s_{t})-\prod_{t=1}^{T}\mu(s_{t},a_{t})$
-where $\mathbf{a}=\{a_{1},\dots,a_{T}\}$ is the set of actions selected by the algorithm (which depends on $\pi^{*}$).
-We focus on a worst-case choice of $\mu$ and $\pi^{*}$.
-[^6] The choice of $\pi^{*}$ can actually matter: suppose $a$ is optimal everywhere but $a^{\prime}$ is optimal only for this state.
-VC and Natarajan dimension.
-VC dimension [31] and Natarajan dimension [24] are standard measures of learning difficulty.
-The former applies to binary classes, while the latter handles multi-class prediction (and reduces to the former in the case of two classes).
-In our setting, a policy can be viewed as a multi-class predictor, where each action is a class.
-The precise definitions of these dimensions are not relevant to our paper, as we only invoke Natarajan dimension indirectly through the fundamental theorem of multi-class PAC learning (Theorem 4.4).
-We direct the reader to [27] for a comprehensive overview.
+For a fixed state sequence $\mathbf{s}$ and agent actions $\mathbf{a}=(a_{1},\dots,a_{T})$, the agent’s regret is
+$R_{T}(\mathbf{s},\mathbf{a},\mu,\pi^{m})=\prod_{t=1}^{T}\mu^{m}(s_{t})-\prod_{t=1}^{T}\mu(s_{t},a_{t})$
+We will usually write $R_{T}=R_{T}(\mathbf{s},\mathbf{a},\mu,\pi^{m})$ for brevity.
+We will study the expected regret over any randomness in $\mathbf{s}$ and/or $\mathbf{a}$.
+We desire subconstant worst-case regret: the total (not average) expected regret should go to 0 for any $\mu$ and $\pi^{m}$.
+Formally, we want $\lim_{T\to\infty}\sup_{\mu,\pi^{m}}\operatorname*{\mathbb{E}}[R_{T}]=0$.
+VC and Littlestone dimensions.
+VC dimension [38] and Littlestone dimension [26] are standard measures of learning difficulty which capture the ability of a hypothesis class (in our case, a policy class) to realize arbitrary combinations of labels (in our case, actions).
+We omit the precise dimensions since we only utilize these concept via existing results.
+See [34] for a comprehensive overview.
 Misc.
-We will often deal with the measure of subsets of $S$.
-When $n=1$, we use $\operatorname*{len}(X)$ to denote the length of interval $X\subset S$; for general $n$, we use $\operatorname*{vol}(X)$ to denote the $n$-dimensional measure of $X\subseteq S$ [^7] Although we could use the same notation for both cases, we find this to be more intuitive.
-The convex hull of a set $X$ is denoted $\operatorname*{conv}(X)$.
-All logarithms and exponents are base $e$ unless otherwise noted.
-3 Hardness result for complex mentor policies
-We begin by showing that in general, any algorithm with sublinear mentor queries has arbitrary poor regret in the worst-case, even when states are uniformly random on $[0,1]$ and there are only two actions.
+The diameter of $X\subseteq S$ is defined by $\operatorname*{diam}(X)=\max_{x,y\in X}||x-y||$. All logarithms and exponents are base $e$ unless otherwise noted.
+4 Avoiding catastrophe is impossible in general
+We begin by showing that in general, any algorithm with sublinear mentor queries has arbitrary poor regret in the worst-case, even when states are i.i.d.
+on $[0,1]$.
 The result also holds even if the algorithm knows $L$ and $\mathbf{s}$ ahead of time.
-Let $U(S)$ denote the uniform distribution on $S$ and let $\operatorname*{\mathbb{E}}_{\mathbf{s}\sim U(S^{T}),\mathbf{a}}$ denote the expectation over both $\mathbf{s}$ and any additional randomness in the algorithm.
+We use $\operatorname*{\mathbb{E}}_{\mathbf{s}\sim U^{T},\mathbf{a}}$ denote the expectation over both $\mathbf{s}$ and any additional randomness in the algorithm.
 Theorem 4.1.
-The worst-case expected regret of any algorithm with sublinear queries goes to 1 as $T$ goes to infinity. Formally,
-$\lim_{T\to\infty}\ \,\sup_{\mu,\pi^{*}}\ \,\operatorname*{\mathbb{E}}_{\mathbf{s}\sim U(S^{T}),\,\mathbf{a}}\ \,[R_{T}(\mathbf{s},\mu,\pi^{*})]=1$
+The worst-case expected regret of any algorithm with sublinear queries goes to 1 as $T$ goes to infinity.
+Formally, $\lim_{T\to\infty}\ \,\sup_{\mu,\pi^{m}}\ \,\operatorname*{\mathbb{E}}_{\mathbf{s}\sim U^{T},\mathbf{a}}\ \,[R_{T}]=1$.
 4.1 Intuition
@@
 Figure 3 fleshes out this idea.
-Picking $\boldsymbol{f(T)}$.
-A natural idea is to try $f(T)=T$.
-Unfortunately, this doesn’t quite work: even if the agent chooses wrong on every time step, the minimum payoff is still at least $1-\frac{L}{2T}$, and $\lim_{T\to\infty}\prod_{t=1}^{T}\left(1-\frac{L}{2T}\right)=\lim_{T\to\infty}\left(1-\frac{L}{2T}\right)^{T}=e^{-L/2}$.
-In order for the regret to approach 1, we need $f(T)$ to be asymptotically between $|Q_{T}|$ and $T$.
-Lemma A.1 implies the existence of such an $f(T)$ if $|Q_{T}|$ is bounded by some function $g(T)\in o(T)$, so we get a bound of the following form: $\prod_{t=1}^{T}\mu_{f,\mathbf{x}}(s_{t},a_{t})\leq\big{(}1-\frac{L}{\Theta(f(T))}\big{)}^{\Theta(T)}$.
-[^8] The bound in the actual proof will be in log space, so it will look slightly different, but the idea is the same.
-When $f(T)\in o(T)$, the right hand side converges to 0, while $\prod_{t=1}^{T}\mu_{f,\mathbf{x}}^{*}(s_{t})=1$.
-In words, the agent is nearly guaranteed to cause catastrophe, despite the existence of a policy which is guaranteed to avoid catastrophe.
 $s$$\mu$  $\mu(s,0)$$\mu(s,1)$$\cfrac{L}{2f(T)}$$\frac{1}{f(T)}$$\frac{2}{f(T)}$$\frac{3}{f(T)}$$\frac{f(T)-1}{f(T)}$$\dots$1001111
@@
 The solid line represents $\mu(s,0)$ and the dotted line represents $\mu(s,1)$.
-In each section, one of the actions has the optimal payoff of 1, and the other action has the worst possible payoff allowed by the Lipschitz constraint, reaching a minimum of $1-\frac{L}{2f(T)}$.
+In each section, one of the actions has the optimal payoff of 1, and the other action has the worst possible payoff allowed by $L$, reaching a minimum of $1-\frac{L}{2f(T)}$.
 Crucially, both actions result in a payoff of 1 at the boundaries between sections: this allows us to “reset” for the next section.
 As a result, we can freely toggle the optimal action for each section independently.
-VC/Natarajan dimension.
-The class of mentor policies induced by our construction has VC and Natarajan dimensions of $f(T)$.
-[^9] Recall that VC and Natarajan dimensions coincide for $|A|=2$.
-This is expected given our positive result for policy classes with bounded Natarajan dimension (Theorem 4.5).
+Picking $\boldsymbol{f(T)}$.
+A natural idea is to try $f(T)=T$, but this doesn’t quite work: even if the agent chooses wrong on every time step, the minimum payoff is still at least $1-\frac{L}{2T}$, and $\lim_{T\to\infty}\prod_{t=1}^{T}\left(1-\frac{L}{2T}\right)=\lim_{T\to\infty}\left(1-\frac{L}{2T}\right)^{T}=e^{-L/2}$.
+In order for the regret to approach 1, we need $f(T)$ to be asymptotically between $|Q_{T}|$ and $T$ (such $f$ must exist since $|Q_{T}|\leq g(T)\in o(T)$).
+This leads to the following bound: $\prod_{t=1}^{T}\mu(s_{t},a_{t})\leq\big{(}1-\frac{L}{\Theta(f(T))}\big{)}^{\Theta(T)}$.
+When $f(T)\in o(T)$, the right hand side converges to 0, while $\prod_{t=1}^{T}\mu^{m}(s_{t})=1$.
+In words, the agent is nearly guaranteed to cause catastrophe, despite the existence of a policy which is guaranteed to avoid catastrophe.
+VC dimension.
+The class of mentor policies induced by our construction has VC dimension $f(T)$; considered over all possible values of $T$, this implies infinite VC (and Littlestone) dimension.
+This is necessary given our positive results in Section 5.
 4.2 Formal definition of construction
-Let $S=[0,1]$, $A=\{0,1\}$, and $\mathcal{D}=U(S)$.
+Let $S=[0,1]$ and $\mathcal{D}_{t}=U$ for each $t\in[T]$.
 Assume that $L\leq 1$; this will simplify the math and only makes the problem easier for the agent.
@@
 $\mu_{f,\mathbf{x}}(s,a)=\begin{cases}1&\textnormal{ if }a=x_{j(s)}\\ 1-L\left(\cfrac{1}{2f(T)}-|m_{j(s)}-s|\right)&\textnormal{ if }a\neq x_{j(s)}\\ \end{cases}$
+Let $\pi^{m}$ be any policy which is optimal for $\mu_{f,\mathbf{x}}$.
+Note that there is a unique optimal action for each $s_{t}$, since each $s_{t}$ belongs to exactly one $S_{j}$; formally, $\pi^{m}(s_{t})=x_{j(s_{t})}$.
 For any $\mathbf{x}\in\{0,1\}^{f(T)}$, $\mu_{f,\mathbf{x}}$ is piecewise linear (trivially) and continuous (because both actions have payoff 1 on the boundary between sections).
-Since the slope of each piece is in $\{-L,0,L\}$, $\mu_{f,\mathbf{x}}$ is Lipschitz continuous with Lipschitz constant $L$.
-Since each $s_{t}$ belongs to exactly one $S_{j}$, there is a unique optimal action for each $s_{t}$, so the choice of $\pi^{*}$ does not matter in this scenario.
-4 Near-zero regret assuming bounded Natarajan dimension
-In this section, we present a simple algorithm with near-zero regret when the policy class $\Pi$ has bounded Natarajan dimension $d$ and which contains at least some “reasonable” policies.
-The algorithm needs to know $\Pi$ and (an upper bound on) $d$, as is standard in PAC learning (see, e.g., [27]).
-The algorithm does not need to know $L$, and can handle an unbounded state space (the number of queries simply scales with the volume of the convex hull of the observed states).
-We need a few definitions to formally state our algorithm and result.
+Since the slope of each piece is in $\{-L,0,L\}$, $\mu_{f,\mathbf{x}}$ is Lipschitz continuous.
+Thus by Proposition E.1, $\pi^{m}$ satisfies local generalization.
+5 Avoiding catastrophe given finite VC or Littlestone dimension
+Theorem 4.1 shows that avoiding catastrophe is impossible in general, which is also true in online learning.
+What if we restrict ourselves to settings where standard online learning is possible?
+Specifically, we assume that $\pi^{m}$ belongs to a policy class $\Pi$ where either (1) $\Pi$ has finite VC dimension $d$ and $\mathbf{s}$ is $\sigma$-smooth or (2) $\Pi$ has finite Littlestone dimension $d$ [^8] Recall from Section 1.3 that standard online learning becomes tractable under either of these assumptions.
+This section presents a simple algorithm which guarantees subconstant regret and sublinear queries under either of those assumptions.
+Our algorithm needs to know $\Pi$, as is standard in online learning.
+The algorithm does not need to know $\sigma$ (in the smooth case) or $L$, and can handle an unbounded state space (the number of queries simply scales with the maximum distance between observed states).
+For simplicity, we initially prove our result for $A=\{0,1\}$.
+Appendix C extends our result to many actions using the“one versus rest” reduction.
+[^9] For each action $a$, we learn a binary classifier which predicts whether $\pi^{m}(s)=a$.
+If every binary classifier is correct, we can correctly determine $\pi^{m}(s)$.
+See, e.g., Chapter 29 of [34].
+5.1 Intuition behind the algorithm
+Algorithm 1 has two simple components: (1) run a modified version of the Hedge algorithm for online learning, but (2) ask for help in unfamiliar states (specifically, when the current state is far from any queried state with the same action under the proposed policy).
+Hedge ensures that the number of time steps where the agent’s action doesn’t match the mentor’s is small, and asking for help in unfamiliar states ensures that when we do make a mistake, the cost isn’t too high.
+This algorithmic structure seems quite natural: mostly follow a baseline strategy, but ask for help when out-of-distribution.
+Simple operations.
+The algorithm does not require detailed access to the state embedding, instead relying on two simple operations: evaluating a policy on a particular state, and computing a nearest neighbor distance.
+The former seems necessary for any algorithm.
+The latter could be modeled as an out-of-distribution detector score, for which many methods are available (see e.g., [41]).
+Hedge.
+Hedge [12] is a standard online learning algorithm which ensures sublinear regret when the number of hypotheses (in our case, the number of policies in $\Pi$) is finite.
+[^10] See Chapter 5 of [35] and Chapter 21 of [34] for modern introductions to Hedge.
+We would prefer not to assume that $\Pi$ is finite.
+Luckily, any policy $\Pi$ can be approximated within $\varepsilon$ when either (1) $\Pi$ has finite VC dimension and $\mathbf{s}$ $\sigma$-smooth or (2) $\Pi$ has finite Littlestone dimension.
+Thus we can run Hedge on this approximative policy class instead.
+One other modification is necessary.
+In standard online learning, losses are observed on every time step, but our agent only receives feedback in response to queries.
+To handle this, we modify Hedge to only perform updates on time steps with queries and to issue a query with probability $p$ on each time step.
+Continuing our lucky streak, [33] analyzes exactly this modification of Hedge.
+1:function AvoidCatastrophe($T\in\mathbb{N},\>\varepsilon\in\mathbb{R}_{>0},\>d\in\mathbb{N},\,$ policy class $\Pi$)
+2:     if $\Pi$ has VC dimension $d$ then
+3:         $\tilde{\Pi}\leftarrow$ any smooth $\varepsilon$-cover of $\Pi$ of size at most $(41/\varepsilon)^{d}$triangleright See Definition 5.3
+4:     else
+5:         $\tilde{\Pi}\leftarrow$ any adversarial cover of size at most $(eT/d)^{d}$triangleright See Definition 5.4
+6:     $X\leftarrow\emptyset$
+7:     $w(\pi)\leftarrow 1$ for all $\pi\in\tilde{\Pi}$
+8:     $p\leftarrow 1/\sqrt{\varepsilon T}$
+9:     $\eta\leftarrow\max\big{(}\sqrt{\frac{p\log|\tilde{\Pi}|}{2T}},\>\frac{p^{2}}{\sqrt{2}}\big{)}$
+10:     for $t$ from $1$ to $T$ do $\triangleright$ Run one step of Hedge, which selects policy $\pi_{t}$
+[…]
+12:         if hedgeQuery then
+13:              Query mentor and observe $\pi^{m}(s_{t})$
+14:              $\ell(t,\pi)\leftarrow\mathbf{1}(\pi(s_{t})\neq\pi^{m}(s_{t}))$ for all $\pi\in\tilde{\Pi}$
+15:              $\ell^{*}\leftarrow\min_{\pi\in\tilde{\Pi}}\ell(t,\pi)$
+16:              $w(\pi)\leftarrow w(\pi)\cdot\exp(-\eta(\ell(t,\pi)-\ell^{*}))$ for all $\pi\in\tilde{\Pi}$
+17:              $\pi_{t}\leftarrow\operatorname*{arg\,min}_{\pi\in\tilde{\Pi}}\ell(t,\pi)$
+18:         else
+19:              $P(\pi)\leftarrow w(\pi)/\sum_{\pi^{\prime}\in\tilde{\Pi}}w(\pi^{\prime})$ for all $\pi\in\tilde{\Pi}$
+20:              Sample $\pi_{t}\sim P$
+21:         if $\min_{(s,a)\in X:a=\pi_{t}(s_{t})}||s_{t}-s||>\varepsilon^{1/n}$ then $\triangleright$ Ask for help if out-of-distribution
+22:              Query mentor and observe $\pi^{m}(s_{t})$
+23:              $X\leftarrow X\cup\{(s_{t},\pi^{m}(s_{t}))\}$
+24:         $\triangleright$ Otherwise, follow Hedge’s chosen policy
+25:              Take action $\pi_{t}(s_{t})$
+Algorithm 1 successfully avoids catastrophe assuming finite VC or Littlestone dimension.
+We prove the following theorem parametrized by $\varepsilon$:
+Theorem 5.1.
+Let $A=\{0,1\}$.
+Assume $\pi^{m}\in\Pi$ where either (1) $\Pi$ has finite VC dimension $d$, $\mathbf{s}$ is $\sigma$-smooth, and $\varepsilon T\log T>12\sigma d\log(4e^{2}/\varepsilon)$ or (2) $\Pi$ has finite Littlestone dimension $d$.
+Then for any $T\in\mathbb{N}$ and $\varepsilon>0$, Algorithm 1 satisfies
+$\operatorname*{\mathbb{E}}\left[R_{T}\right]\in$ $\ O\left(\frac{dL}{\sigma}T\varepsilon^{1+1/n}\log(1/\varepsilon)\log T\right)$
+$\operatorname*{\mathbb{E}}[|Q_{T}|]\in$ $\ O\left(\sqrt{\frac{T}{\varepsilon}}+\frac{d}{\sigma}T\varepsilon\log(1/\varepsilon)\log T+\frac{\operatorname*{diam}(\mathbf{s})^{n}}{\varepsilon}\right)$
+In Case 1, the expectation is over the randomness of both $\mathbf{s}$ and the algorithm, while in Case 2, the expectation is over only the randomness of the algorithm.
+Also, $R_{T}$ and $Q_{T}$ clearly have no dependence on $\sigma$ in Case 2, but we include $\sigma$ anyway to avoid writing two separate bounds.
+To obtain subconstant regret and sublinear queries, we can choose $\varepsilon=T^{\frac{-2n}{2n+1}}$.
+This also satisfies the requirement of $\varepsilon T\log T>12\sigma d\log(4e^{2}/\varepsilon)$ for large enough $T$.
+Theorem 5.2.
+Let $A=\{0,1\}$.
+Assume $\pi^{m}\in\Pi$ where either (1) $\Pi$ has finite VC dimension $d$ and $\mathbf{s}$ is $\sigma$-smooth or (2) $\Pi$ has finite Littlestone dimension $d$.
+Then for any $T\in\mathbb{N}$, Algorithm 1 with $\varepsilon=T^{\frac{-2n}{2n+1}}$ satisfies
+$\operatorname*{\mathbb{E}}\left[R_{T}\right]\in$ $\ O\left(\frac{dL}{\sigma}T^{\frac{-1}{2n+1}}\log T\right)$
+$\operatorname*{\mathbb{E}}[|Q_{T}|]\in$ $\ O\left(T^{\frac{4n+1}{4n+2}}\left(\frac{d}{\sigma}\log T+\operatorname*{diam}(\mathbf{s})^{n}\right)\right)$
+We can also make the query bound unconditional in exchange for slightly worse bounds.
+The probability that $|Q_{T}|$ asymptotically exceeds the bound from Theorem 5.2 goes to 0, so the query threshold in Corollary 5.2.1 is almost never reached.
+Thus we still obtain subconstant regret.
+Corollary 5.2.1.
+If Algorithm 1 is modified to stop querying after $T^{\frac{4n+1.5}{4n+2}}(\frac{d}{\sigma}\log T+\operatorname*{diam}(\mathbf{s})^{n})$ queries, then under the conditions of Theorem 5.2, we have $\operatorname*{\mathbb{E}}[R_{T}]\in O\big{(}\frac{dL}{\sigma}T^{\frac{-1}{8n+4}}\log T\big{)}$.
+5.2 Proof sketch
+The formal proof of Theorem 5.1 can be found in Appendix B, but we outline the key elements here.
+The regret analysis consists of two ingredients: analyzing the Hedge component, and analyzing the “ask for help when out-of-distrubtion” component.
+The former will bound the number of mistakes made by the algorithm (i.e., the number of time steps where the agent’s action doesn’t match the mentor’s), and the latter will bound the cost of any single mistake.
+We must also carefully show that the latter does not result in excessively many queries, which we do via a novel packing argument.
+We begin by formalizing two notion of approximating a policy class:
 Definition 5.3.
-A policy $\pi$ is consistent with $X\subset S\times A$ if $\pi(s)=a$ for all $(s,a)\in X$.
+Let $U$ be the uniform distribution over $S$.
+For $\varepsilon>0$, a policy class $\tilde{\Pi}$ is a smooth $\varepsilon$-cover of a policy class $\Pi$ is for every $\pi\in\Pi$, there exists $\tilde{\pi}\in\tilde{\Pi}$ such that $\Pr_{s\sim U}[\pi(s)\neq\tilde{\pi}(s)]\leq\varepsilon$.
 Definition 5.4.
-Given $X\subset S\times A$ and $r\geq 0$, a policy $\pi$ is $(X,r)$-local if (1) $\pi$ is consistent with $X$ and (2) whenever $(s^{\prime},a^{\prime})\in X$ and $||s-s^{\prime}||\leq r$, there exists $(s^{\prime\prime},a^{\prime\prime})\in X$ such that $||s-s^{\prime\prime}||\leq r$ and $\pi(s)=a^{\prime\prime}$.
-A policy which knows $X$ but is not $(X,r)$-local says either (1) “even though I know the optimal action for $s$, I’m going to do something else” or (2) “even though I have a data point near $s$, I’m going to base my action on a data point far from $s$”.
-[^10] Note that the mentor policy cannot itself query the mentor, so we are only dealing with actions in $A$.
-Given the Lipschitz nature of the problem, this type of policy seems quite unintuitive.
-However, we allow $\Pi$ to contain non-local policies, as long as it also contains local policies.
-Note that knowing $X,r$, and $\Pi$ is sufficient to find a $(X,r)$-local policy.
-Definition 4.3.
-Given $r\geq 0$, $\Pi$ is $r$-local if for every $\tilde{S}\subset S$ and $X=\{(s,\pi^{*}(s)):s\in\tilde{S}\}$, there exists an $(X,r)$-local policy in $\Pi$.
-As an example, let $S=\mathbb{R}$ and suppose $\Pi$ is the set of policies where $\pi(s)$ changes at most $K$ times as $s$ goes from $\infty$- ∞ to +$\infty$+ ∞ (in the language of Section 5, $\pi$ has at most $K$ segments).
-Then $\Pi$ is $r$-local for any $r$, since for any set of queries $X$, the nearest neighbor policy is $(X,r)$-local and has at most $K$ segments (and thus is in $\Pi$).
-Note that this $\Pi$ also has bounded Natarajan dimension.
-4.1 Intuition behind the algorithm
-Our algorithm has a simple structure: gather enough data to choose a solid baseline policy, and then mostly follow that policy, but ask for help if you find yourself out-of-distribution (OOD).
-The algorithm does not need direct access to the feature embedding of states: the only two operations it uses are (1) update baseline policy and (2) find nearest neighbor distance.
-This approach seems quite natural for real-world applications, where one can imagine updating the baseline policy via gradient descent and treating an OOD detector score as the nearest neighbor distance.
-In our case, the initial dataset $X$ is gathered by querying on the first $g(T)$ time steps.
-We will choose $g(T)$ to be large enough so that – along with the assumption on Natarajan dimension – we can apply the fundamental theorem of multi-class PAC learning (Theorem 4.4).
-[^11] Theorem 4.4 is called the “realizable” case of PAC learning, in contrast to the “agnostic” case where $\pi^{*}$ may not be in $\Pi$.
-Some sources (e.g., [27]) state that $\pi$ should minimize empirical loss, rather than being consistent with $X$.
-These are equivalent in the realizable case since consistent policies have zero empirical loss.
-This is the only place in the proof where Natarajan dimension is used.
-Theorem 4.4 (Fundamental Theorem of Multi-Class PAC Learning [13; 24; 27]).
-Assume $\Pi$ has finite Natarajan dimension $d$ and $\pi^{*}\in\Pi$.
-Then $\exists\,c>0$ such that $\forall\,\varepsilon,\delta\in(0,1)$, if $X\subset S\times A$ consists of at least $\frac{c}{\varepsilon}\left(d\log(\frac{|A|d}{\varepsilon})+\log\frac{1}{\delta}\right)$ i.i.d.
-samples $\mathcal{D}$ and the corresponding actions from $\pi^{*}$, then with probability at least $1-\delta$, any policy $\pi\in\Pi$ consistent with $X$ satisfies $\Pr_{s\sim\mathcal{D}}[\pi(s)\neq\pi^{*}(s)]\leq\varepsilon$.
-Theorem 4.4 will ensure that $\Pr[a_{t}\neq\pi^{*}(s_{t})]$ is small, but even a single incorrect action can spell defeat if the payoff is bad enough.
-To handle this, the algorithm continues to selectively ask for help, but only when the state is far from any previously queried state (i.e., out-of-distribution).
-The policy $\pi$ is then updated based on this new information.
-The threshold $r$ is chosen such that we can cover the entire state $S$ with a sublinear number of queries.
-1:function LookBeforeYouLeap($T\in\mathbb{N},\>r\in\mathbb{R}_{\geq 0},\>g:\mathbb{N}\to\mathbb{N}$)
-2:     $X\leftarrow\emptyset$
-3:     for $t$ from $1$ to $g(T)$ do $\triangleright$ Gather baseline data set
-4:         Query mentor and observe $\pi^{*}(s_{t})$
-5:         $X\leftarrow X\cup\{(s_{t},\pi^{*}(s_{t}))\}$
-6:     $\pi\leftarrow$ any $(X,r)$-local policy in $\Pi$
-7:     for $t$ from $g(T)+1$ to $T$ do $\triangleright$ Mostly follow $\pi$, but ask for help if out-of-distribution
-8:         if $\min_{(s,a)\in X}||s_{t}-s||\leq r$ then
-9:              Take action $\pi(s_{t})$
-10:         else
-11:              Query mentor and observe $\pi^{*}(s_{t})$
-12:              $X\leftarrow X\cup\{(s_{t},\pi^{*}(s_{t}))\}$
-13:              $\pi\leftarrow$ any $(X,r)$-local policy in $\Pi$
-Algorithm 1 achieves near-zero regret when $\Pi$ has bounded Natarajan dimension and is $r$-local.
-Tractability.
-Our algorithm can only be useful in practice if finding a $(X,r)$-local policy in $\Pi$ is computationally tractable.
-Such guarantees are impossible without further assumptions on $\Pi$ (this is again standard for PAC learning).
-However, if $\Pi$ supports nearest-neighbor policies (which satisfy locality for all values for $r$), those policies are both intuitive and can be computed easily.
-We are now ready to state the main result of this section. The constant $c$ in Theorem 4.5 is the same as the constant $c$ in Theorem 4.4.
-Theorem 4.5.
-[…]
-$\operatorname*{\mathbb{E}}_{\mathbf{s}\sim\mathcal{D}^{T}}\left[R_{T}(\mathbf{s},\mu,\pi^{*})\right]\leq\frac{6L}{T^{\frac{0.5}{n+0.5}}}+\frac{1}{T}+\exp(-3T^{\frac{0.5}{n+0.5}}/8)$
-Corollary 4.6 makes the near-zero regret and sublinear mentor queries more explicit.
-Corollary 4.6.
-Assume $\pi^{*}\in\Pi$, where $\Pi$ has finite Natarajan dimension $d$ and is $\varepsilon(T)^{1/n}$-local for $\varepsilon(T)=T^{\frac{-n}{n+0.5}}$.
-Then Algorithm 1 with $g(T)=2cdT^{\frac{n}{n+0.5}}\log(d\,T)$ and $r=\varepsilon(T)^{1/n}$ makes at most $O\left(T^{\frac{n}{n+0.5}}(d\log(d\,T)+\operatorname*{vol}(\operatorname*{conv}(S)))\right)$ queries and satisfies
-$\lim_{T\to\infty}\>\operatorname*{\mathbb{E}}_{\mathbf{s}\sim\mathcal{D}^{T}}\left[R_{T}(\mathbf{s},\mu,\pi^{*})\right]=0$
-When $d$ is constant, the number of queries in Corollary 4.6 is $o(T)$.
-(Actually, $d$ can be as large as $T^{\frac{0.5-\delta}{n+0.5}}$ for $\delta\in(0,1/2)$.) Also, since the expected regret goes to zero, the regret must go to zero with high probability (since regret is always nonnegative).
-Finally, note that standard PAC learning lower bounds do not directly yield lower bounds for our setting: even $\Pr[a_{t}\neq\pi^{*}(s_{t})]=1$ is insufficient for high regret, since the payoffs could be nearly optimal.
-Also, PAC learning lower bounds typically state that there exists a distribution for which the algorithm performs poorly, while our lower bound uses a fixed known distribution.
-5 Near-zero regret for simple mentor policies in 1D with adversarial states
-Our final result is another algorithm with near-zero regret and sublinear queries, but under different assumptions.
-In this section, we assume a 1D state space and we allow the state sequence to be adversarially chosen.
-We consider on the following notion of simplicity:
-Definition 5.1.
-Given a mentor policy $\pi^{*}$, partition the state space $S$ into intervals such that all states within each interval share the same mentor action.
-Let $\{S_{1},\dots,S_{k}\}$ be a partition that minimizes the number of intervals.
-We call each $S_{j}$ a segment.
-Let $f(\pi^{*})$ denote the number of segments in $\pi^{*}$.
-This is similar conceptually to Natarajan dimension in that it limits the ability of the policy class to realize arbitrary combinations of labels (i.e., optimal actions) on $\mathbf{s}$.
-The number of segments also relates to the minimum depth needed by a neural network to learn the policy [7].
-Figure 1 shows why the Lipschitz constant does not bound the number of segments ($f(T)$ can be arbtrarily large).
-We prove the following result:
-Theorem 5.2.
-For any $c\in[0,1]$, any $\mathbf{s}\in S^{T}$, and any $\pi^{*}$ with $f(\pi^{*})\leq K$, Algorithm 2 with $g(T)=T^{c}$ makes at most $(4+\operatorname*{len}(\operatorname*{conv}(\mathbf{s})))g(T)$ queries and satisfies $R_{T}(\mathbf{s},\mu,\pi^{*})\leq 6LKT^{1-2c}$.
-Corollary 5.3 makes the sublinear queries and near-zero regret of Theorem 5.2 more explicit when the number of segments is a constant.
-Corollary 5.3.
-For any $c\in(1/2,1)$, Algorithm 2 with $g(T)=T^{c}$ makes $O(T^{c}\operatorname*{len}(\operatorname*{conv}(S)))$ queries and satisfies
-$\lim_{T\to\infty}\ \sup_{\mathbf{s}\in S^{T}}\ \sup_{\mu}\ \sup_{\pi^{*}:f(\pi^{*})\leq K}\ R_{T}(\mathbf{s},\mu,\pi^{*})=0$
-Our algorithm does not need to know the Lipschitz constant $L$ or the number of segments; it only needs to know the time horizon $T$.
-5.1 Intuition behind the algorithm
-The algorithm maintains a set of buckets which partition the observed portion of the state space.
-Each bucket’s length determines the maximum loss in payoff we will allow from that subset of the state space.
-As long as the bucket contains a query from a prior time step, the Lipschitz constraint allows us to bound $\mu^{*}(s_{t})-\mu(s_{t},a_{t})$ based on the length of the bucket containing $s_{t}$.
-We always query if the bucket does not contain a prior query
-The granularity of the buckets is controlled by a function $g$, with the initial buckets having length $1/g(T)$.
-Since we can expect one query per bucket, we need $g(T)\in o(T)$ to ensure sublinear queries.
-Regardless of the bucket length, the adversary can still place multiple segments in the same bucket $B$.
-A single query only tells us the optimal action for one of those segments, so we risk a payoff as bad as $\mu^{*}(s_{t})-O(\operatorname*{len}(B))$ whenever we choose not to query.
-We can endure a limited number of such payoffs, but if we never query again in that bucket, we may suffer $\Theta(T)$ such payoffs.
-Letting $\mu^{*}(s_{t})=1$ for simplicity, that would lead to $\prod_{t=1}^{T}\mu(s_{t},a_{t})\leq\big{(}1-\frac{1}{O(g(T))}\big{)}^{\Theta(T)}$, which converges to 0 (i.e., guaranteed catastrophe) when $g(T)\in o(T)$.
-This failure mode suggests a natural countermeasure: if we start to suffer significant (potential) losses in the same bucket, then we should probably query there again.
-One way to structure these supplementary queries is by splitting the bucket in half when enough time steps have involved that bucket.
-It turns out that splitting after $T/g(T)$ time steps is a sweet spot.
-This approach is reminiscent of Algorithm 1 in that it collects an initial dataset (one query per initial bucket), and then performs supplemental queries as needed based on the observed states.
-We do not think this is a coincidence.
-1:function AvoidCatastrophe($T\in\mathbb{N},\>g:\mathbb{N}\to\mathbb{N}$)
-2:     $S_{Q}\leftarrow\emptyset$triangleright Previously queried states
-3:     $\pi\leftarrow\emptyset$triangleright Records $\pi^{*}(s)$ for each $s\in S_{Q}$
-4:     $\mathcal{B}\leftarrow\emptyset$triangleright The set of active buckets
-5:     for $t$ from $1$ to $T$ do
-6:         EvaluateState$(s_{t})$
-7:function EvaluateState($s\in S$)
-8:     if $s\not\in B$ for all $B\in\mathcal{B}$ then $\triangleright$ No bucket containing $s$: create a new bucket and try again
-9:         $B\leftarrow\left[\frac{j-1}{g(T)},\frac{j}{g(T)}\right]$ for $j\in\mathbb{Z}$ such that $s\in B$
-10:         $\mathcal{B}\leftarrow\mathcal{B}\cup\{B\}$
-11:         $n_{B}\leftarrow 0$triangleright Number of time steps that have used $B$
-12:         EvaluateState$(s)$
-13:     else
-14:         $B\leftarrow$ any bucket containing $s$
-15:         if $S_{Q}\cap B=\emptyset$ then $\triangleright$ No queries in this bucket
-16:              Query mentor and observe $\pi^{*}(s)$
-17:              $\pi(s)\leftarrow\pi^{*}(s)$
-18:              $S_{Q}\leftarrow S_{Q}\cup\{s\}$
-19:              $n_{B}\leftarrow n_{B}+1$
-20:         else if $n_{B}<T/g(T)$ then $\triangleright$ Bucket has a query and isn’t full: take that action
-21:              Let $s^{\prime}\in S_{Q}\cap B$
-22:              Take action $\pi(s^{\prime})$
-23:              $n_{B}\leftarrow n_{B}+1$
-24:         $\triangleright$ Bucket is full: split bucket and try again
-25:              $B=[x,y]$
-26:              $(B_{1},B_{2})\leftarrow\Big{(}\left[x,\frac{x+y}{2}\right],\left[\frac{x+y}{2},y\right]\Big{)}$
-27:              $(s_{B_{1}},s_{B_{2}})\leftarrow(0,0)$
-28:              $\mathcal{B}\leftarrow\mathcal{B}\cup\{B_{1},B_{2}\}\setminus B$
-29:              EvaluateState$(s)$
-Algorithm 2 achieves near-zero regret when the mentor’s policy has a bounded number of segments.
-6 Limitations
-Our paper contains both positive and negative results.
-As such, our claim in this paper is not “avoiding catastrophe is possible (or impossible) under assumptions X, Y, Z, and we argue that X, Y, Z hold”.
-Instead, we aim to understand the conditions under which it is possible to avoid catastrophe.
-Our claim is “avoiding catastrophe is possible under X, Y, Z, but impossible under A, B, C; we might sometimes expect X, Y, Z and sometimes expect A, B, C”.
-For example, some real-world instances might exhibit bounded Natarajan dimension, while others might not.
-Our paper elucidates what to expect from either scenario.
-The goal, then, is to make the gap between X, Y, Z and A, B, C as small as possible: to isolate the boundary of tractability.
-Our results suggest that the key factor determining tractability is the complexity of the mentor policy class.
-Our negative result (Theorem 3.1) shows that a simple environment (bounded 1D state space, known uniform distribution over states, two actions) is not enough for tractability if the mentor policy class is complex.
-On the other hand, our primary positive result (Theorem 4.5) holds for quite general environments (unbounded $n$-dimensional state space, unknown arbitrary distribution over states, many actions), as long as the mentor policy class is simple.
-One remaining gap between those two results is the assumption of $r$-locality in Theorem 4.5.
-Although we think $r$-locality is a fairly plausible assumption (as argued in Section 4), that assumption would need to be removed in order for Theorem 3.1 and Theorem 4.5 to fully match.
-Another question is how far this positive result generalizes.
-In particular, the assumptions of (1) only caring about catastrophe and (2) i.i.d.
-states significantly limit the real-world applicability of our algorithm.
-In future work, we hope to extend this result to general MDPs.
-Our secondary positive result (Theorem 5.2) also suggests that mentor policy complexity could be the key factor determining tractability.
-However, it has the drawback that the number of segments in the mentor policy is a less established measure of complexity than the well-studied notion of Natarajan dimension.
-In our opinion, the more serious limitation is the restriction to a 1D state space.
-Although Theorem 5.2 does permit adversarial state sequences, we feel that Theorem 4.5 does a better overall job of isolating the boundary of tractability.
+A policy class $\tilde{\Pi}$ is an adversarial cover of a policy class $\Pi$ is for every $\mathbf{s}\in S^{T}$ and $\pi\in\Pi$, there exists $\tilde{\pi}\in\tilde{\Pi}$ such that $\pi(s_{t})=\tilde{\pi}(s_{t})$ for all $t\in[T]$.
+The existence of small covers is crucial:
+Lemma 5.1 (Lemma 7.3.2 in [16] [^11] See also [19] or Lemma 13.6 in [6] for variants which are less convenient for our purposes.).
+For all $\varepsilon>0$, any policy class of VC dimension $d$ admits a smooth $\varepsilon$-cover of size at most $(41/\varepsilon)^{d}$.
+Lemma 5.2 (Lemmas 21.13 and A.5 in [34]).
+Any policy class of Littlestone dimension $d$ admits an adversarial cover of size at most $(eT/d)^{d}$.
+An adversarial cover is a perfect cover by definition. The following lemma establishes that a smooth $\varepsilon$-cover is a good approximation for any sequence of $\sigma$-smooth distributions.
+Lemma 5.3 (Equation 2 and Lemma 3.3 in [17]).
+Let $\tilde{\Pi}$ be a finite smooth $\varepsilon$-cover of $\Pi$ and let $\boldsymbol{\mathcal{D}}=\mathcal{D}_{1},\dots,\mathcal{D}_{T}$ be a sequence of $\sigma$-smooth distributions.
+If $\varepsilon T\log T>12\sigma d\log(4e^{2}/\varepsilon)$, then
+$\operatorname*{\mathbb{E}}_{\mathbf{s}\sim\boldsymbol{\mathcal{D}}}\left[\sup\limits_{\pi\in\Pi}\min\limits_{\tilde{\pi}\in\tilde{\Pi}}\sum_{t=1}^{T}\mathbf{1}(\pi(s_{t})\neq\tilde{\pi}(s_{t}))\right]\in O\left(\sqrt{\frac{d\varepsilon T}{\sigma}\log T\log(1/\varepsilon)}+\frac{\varepsilon T}{\sigma}\log T\right)$
+We will run a variant of Hedge on $\tilde{\Pi}$.
+The vanilla Hedge algorithm operates in the standard online learning model where on each time step, the agent selects a policy (or more generally, a hypothesis), and observes the loss of every policy.
+In general the loss function can depend arbitrarily on the time step, the policy, and prior events, but we will only use the indicator loss function $\ell(t,\pi)=\mathbf{1}(\pi(s_{t})\neq\pi^{m}(s_{t}))$.
+Crucially, whenever we query and learn $\pi^{m}(s_{t})$, we can compute $\ell(t,\pi)$ for every $\pi\in\tilde{\Pi}$.
+We cannot afford to query on every time step, however.
+Recently, [33] analyzed a variant of Hedge where losses are observed only in response to queries, which they call “label-efficient feedback”.
+They proved a regret bound when a query is issued on each time step with fixed probability $p$.
+Lemma 5.4 restates their result in a form that is more convenient for us (see Appendix B for details).
+Although their result is stated for non-adaptive adversaries, we explain in Section B.3 how their argument easily generalizes to adaptive adversaries.
+Full pseudocode for HedgeWithQueries can also be found in the appendix (Algorithm 2).
+Lemma 5.4 (Lemma 3.5 in [33]).
+Assume $\tilde{\Pi}$ is finite. Then for any loss function $\ell:[T]\times\tilde{\Pi}\to[0,1]$ and query probability $p$, HedgeWithQueries enjoys the regret bound
+$\sum_{t=1}^{T}\operatorname*{\mathbb{E}}[\ell(t,\pi_{t})]-\min_{\tilde{\pi}\in\tilde{\Pi}}\sum_{t=1}^{T}\ell(t,\pi)\leq\frac{2\log|\tilde{\Pi}|}{p^{2}}$
+where $\pi_{t}$ is the policy chosen at time $t$ and the expectation is over the randomness of the algorithm.
+We apply Lemma 5.4 with $\ell(t,\pi)=\mathbf{1}(\pi(s_{t})\neq\pi^{m}(s_{t}))$ and combine this with Lemmas 5.1 and 5.3 (in the $\sigma$-smooth case) and with Lemma 5.2 (in the adversarial case).
+This yields a $O\left(\frac{d}{\sigma}T\varepsilon\log(1/\varepsilon)\log T\right)$ bound on the number of mistakes made by Algorithm 1 (Lemma B.1).
+The other key ingredient of the proof is analyzing the “ask for help when out-of-distribution” component.
+Combined with the local generalization assumption, this allows us to fairly easily bound the cost of a single mistake (Lemma B.2).
+The trickier part is bounding the number of resulting queries.
+It is tempting to claim that the states queried in the out-of-distribution case must all be separated by at least $\varepsilon^{1/n}$ and thus form an $\varepsilon^{1/n}$-packing, but this is actually not true.
+Instead, the packing analysis must be performed on a per-action basis, and we reuse Lemma B.1 to crucially show that there isn’t too much “spillover” between actions.
+The resulting overall query bound is given by Lemma B.7.
 6 Conclusion and future work
-In this paper, we proposed a model of avoiding catastrophe in sequential decision-making.
-We showed that in general, any algorithm either constantly asks for help or is guaranteed to cause catastrophe.
-However, the problem becomes tractable when the mentor’s policy is relatively simple, for two different definitions of “simple”.
-There remain some technical questions within this paper’s model: (1) Can Theorem 4.5 be generalized to adversarial state sequences?
-(2) Is $r$-locality necessary?
-(3) Are there are reasonable alternatives to the Lipschitz assumption?
-(Indeed, the utility of $r$-locality steps from the Lipschitz assumption.)
-As mentioned, we are also excited to move beyond the current model to full MDPs: recall the open problem we posed in Section 1.
-Most applications are better captured by MDPs than by this paper’s assumptions of i.i.d.
-states and only caring about catastrophe.
-Finally, although we intentionally avoided Bayesian inference in this paper, it seems premature to abandon that tool entirely. Are there tractable approximations to Bayesian approaches?
+In this paper, we proposed a model of avoiding catastrophe in online learning.
+We showed that achieving subconstant regret in our problem (with the help of a mentor and local generalization) is no harder than achieving sublinear regret in standard online learning.
+There remain some technical questions within this paper’s model.
+One question is whether the time complexity of Algorithm 1 be improved, which currently stands at $\Omega(|\tilde{\Pi}|\cdot T)$ plus the time to compute the $\varepsilon$-cover.
+Also, we have not resolved whether our problem is tractable for finite VC dimension and fully adversarial inputs (although Appendix D shows that the problem is tractable for at least some classes with finite VC but infinite Littlestone dimension).
+We are also interested in alternatives to the local generalization assumption, since an action which is safe in one state may not always be safe in a nearby state (depending on the definition of “nearby”).
+Some assumption is necessary, or there is no way to avoid costly mistakes without querying on every time step.
+One possibility is Bayesian inference.
+We intentionally avoided Bayesian approaches in this paper due to tractability concerns, but it seems premature to abandon those ideas entirely.
+Finally, we are excited to apply the ideas in this paper to Markov Decision Processes (MDPs): specifically, MDPs where some actions are irreversible (“non-communicating”) and the agent only gets one attempt (“single-episode”).
+In such MDPs, the agent must not only avoid catastrophe but also obtain high reward.
+As discussed in Section 2, very little theory exists for RL in non-communicating single-episode MDPs.
+Can an agent learn near-optimal behavior in high-stakes environments while becoming self-sufficient over time?
+Formally, we pose the following open problem:
+Is there an algorithm for non-communicating single-episode undiscounted MDPs which ensures that both the regret and the number of mentor queries are sublinear in $T$?
 Acknowledgements
 This work was supported in part by a gift from Open Philanthropy to the Center for Human-Compatible AI (CHAI) at UC Berkeley.
-This paper also benefited from discussions with many researchers at CHAI.
-We would like to particularly thank Michael Cohen, Bhaskar Mishra, Cam Allen, and Li Dayan for helpful feedback.
+This paper also benefited from discussions with many other researchers.
+We would like to especially thank (in alphabetical order) Cameron Allen, Michael Cohen, Scott Emmons, Nika Haghtalab, Aly Lidayan, Juan Liévano-Karim, Bhaskar Mishra, and Matteo Russo.
 Appendix A Proof of Theorem 4.1
@@
 […]
-Recall that $\operatorname*{\mathbb{E}}_{\mathbf{s}\sim U(S^{T}),\mathbf{a}}$ denotes the expectation over both the random states and any additional randomness in the algorithm.
+Recall that $\operatorname*{\mathbb{E}}_{\mathbf{s}\sim U^{T},\mathbf{a}}$ denotes the expectation over both the random states and any additional randomness in the algorithm.
 The proof proceeds via the following steps:
@@
   5. 5.
-    To complete the proof, we observe that $\sup_{\mu}\ \operatorname*{\mathbb{E}}_{\mathbf{s}\sim U(S^{T}),\,\mathbf{a}}\ R_{T}(\mathbf{s},\mu,\pi^{*})\geq\operatorname*{\mathbb{E}}_{\mathbf{x}\sim U(\{0,1\}^{f(T)})}\ \operatorname*{\mathbb{E}}_{\mathbf{s}\sim U(S^{T}),\,\mathbf{a}}\ R_{T}(\mathbf{s},\mu_{f,\mathbf{x}},\pi^{*})$.
+    To complete the proof, we observe that $\sup_{\mu}\ \operatorname*{\mathbb{E}}_{\mathbf{s}\sim U^{T},\,\mathbf{a}}\ R_{T}(\mathbf{s},\mathbf{a},\mu,\pi^{m})\geq\operatorname*{\mathbb{E}}_{\mathbf{x}\sim U(\{0,1\}^{f(T)})}\ \operatorname*{\mathbb{E}}_{\mathbf{s}\sim U^{T},\,\mathbf{a}}\ R_{T}(\mathbf{s},\mathbf{a},\mu_{f,\mathbf{x}},\pi^{m})$.
 This is essentially an application of the probabilistic method: if a randomly chosen $\mu_{f,\mathbf{x}}$ has high expected regret, then the worst case $\mu$ also has high expected regret.
@@
 In contrast, the partition $S=\{S_{1},\dots,S_{f(T)}\}$ and properties thereof (like the midpoints $m_{j}$) are not random variables.
+Lastly, while the intuition provided in Section 4.1 is accurate, the analysis will mostly occur in log space, so the bounds will look different.
+However, bounds of the form discussed in Section 4.1 can still be found as an intermediate step in Part 4 of the proof of Lemma A.4.
 A.2 Proof
@@
 Lemma A.4.
-Independently sample $\mathbf{x}\sim U(\{0,1\}^{f(T)})$ and $s\sim U(S^{T})$.
+Independently sample $\mathbf{x}\sim U(\{0,1\}^{f(T)})$ and $\mathbf{s}\sim U^{T}$.
 [^12] That is, the entire set $\{x_{1},\dots,x_{f(T)},s_{1},\dots,s_{T}\}$ is mutually independent.
@@
 $\Pr\left[|\bar{J}|\leq\frac{f(T)-|Q_{T}|}{4}\right]\leq\exp\left(-\frac{f(T)-|Q_{T}|}{16}\right)$
-Part 3: $V_{j}^{\prime}$ is not too small.
+Part 3: $|V_{j}^{\prime}|\geq|V_{j}|/2$.
 Since $j\in J_{\neg Q}$, the mentor is not queried on any time step $t\in V_{j}$, so $a_{t}\in\{0,1\}$ for all $t\in V_{j}$.
@@
 Now aggregating across time steps,
-$\prod_{t=1}^{T}\mu_{f,\mathbf{x}}(s_{t},a_{t})\leq$ $\ \prod_{j\in\bar{J}}\prod_{t\in V_{j}^{\prime}}\mu_{f,\mathbf{x}}(s_{t},a_{t})$
-$\leq$ $\ \prod_{j\in\bar{J}}\left(1-\frac{L}{4f(T)}\right)^{|V_{j}^{\prime}|}$
-$\leq$ $\ \prod_{j\in\bar{J}}\left(1-\frac{L}{4f(T)}\right)^{|V_{j}|/2}$
+[…]
+[…]
+[…]
 […]
@@
 Exponentiating both sides proves the lemma. ∎
-Let $\alpha(T)=\exp\big{(}\frac{-T}{16f(T)}\big{)}+\exp\big{(}-\frac{f(T)-|Q_{T}|}{16}\big{)}$ for brevity. See 3.1
+Let $\alpha(T)=\exp\big{(}\frac{-T}{16f(T)}\big{)}+\exp\big{(}-\frac{f(T)-|Q_{T}|}{16}\big{)}$ for brevity. See 4.1
 Proof.
-If the algorithm has sublinear queries, then there exists $g(T)\in o(T)$ such that $|Q_{T}|\leq g(T)$ for all $\mathbf{s},\mu,\pi^{*}$.
+If the algorithm has sublinear queries, then there exists $g(T)\in o(T)$ such that $|Q_{T}|\leq g(T)$ for all $\mathbf{s},\mu,\pi^{m}$.
 Let $f(T)=\sqrt{(g(T)+1)T}$.
@@
 Combining this with Lemma A.4 and noting that $\prod_{t=1}^{T}\mu_{f,\mathbf{x}}(s_{t},a_{t})\leq 1$, we have
-$\ \operatorname*{\mathbb{E}}_{\mathbf{x}\sim U(\{0,1\}^{f(T)})}\ \,\operatorname*{\mathbb{E}}_{\mathbf{s}\sim U(S^{T}),\,\mathbf{a}}\ \>\prod_{t=1}^{T}\mu_{f,\mathbf{x}}(s_{t},a_{t})$
+$\ \operatorname*{\mathbb{E}}_{\mathbf{x}\sim U(\{0,1\}^{f(T)})}\ \,\operatorname*{\mathbb{E}}_{\mathbf{s}\sim U^{T},\,\mathbf{a}}\ \>\prod_{t=1}^{T}\mu_{f,\mathbf{x}}(s_{t},a_{t})$
 $\leq$ $\ \alpha(T)\cdot 1+\big{(}1-\alpha(T)\big{)}\exp\left(-\frac{LT(f(T)-|Q_{T}|)}{2^{7}f(T)^{2}}\right)$
@@
 $=$ $\ \alpha(T)+\big{(}1-\alpha(T)\big{)}\exp\left(-\frac{LT}{2^{8}f(T)}\right)$
-whenever $T\geq T_{0}$. Since $\prod_{t=1}^{T}\mu_{f,\mathbf{x}}^{*}(s_{t})=1$ always, we have [^14] Note that Fubini’s theorem means we need not worry about the order of the expectation operators.
-$\sup_{\mu}\ \operatorname*{\mathbb{E}}_{\mathbf{s}\sim U(S^{T}),\,\mathbf{a}}\ R_{T}(\mathbf{s},\mu,\pi^{*})\geq$ $\ \operatorname*{\mathbb{E}}_{\mathbf{x}\sim U(\{0,1\}^{f(T)})}\ \,\operatorname*{\mathbb{E}}_{\mathbf{s}\sim U(S^{T}),\,\mathbf{a}}\ R_{T}(\mathbf{s},\mu_{f,\mathbf{x}},\pi^{*})$
-$=\operatorname*{\mathbb{E}}_{\mathbf{x}\sim U(\{0,1\}^{f(T)})}\ \,\operatorname*{\mathbb{E}}_{\mathbf{s}\sim U(S^{T}),\,\mathbf{a}}$ $\ \,\left[\prod_{t=1}^{T}\mu_{f,\mathbf{x}}^{*}(s_{t})-\prod_{t=1}^{T}\mu_{f,\mathbf{x}}(s_{t},a_{t})\right]$
+whenever $T\geq T_{0}$. Since $\prod_{t=1}^{T}\mu_{f,\mathbf{x}}^{m}(s_{t})=1$ always, we have [^13] Fubini’s theorem means we need not worry about the order of the expectation operators.
+$\sup_{\mu}\ \operatorname*{\mathbb{E}}_{\mathbf{s}\sim U^{T},\,\mathbf{a}}\ R_{T}(\mathbf{s},\mathbf{a},\mu,\pi^{m})\geq$ $\ \operatorname*{\mathbb{E}}_{\mathbf{x}\sim U(\{0,1\}^{f(T)})}\ \,\operatorname*{\mathbb{E}}_{\mathbf{s}\sim U^{T},\,\mathbf{a}}\ R_{T}(\mathbf{s},\mathbf{a},\mu_{f,\mathbf{x}},\pi^{m})$
+$=\operatorname*{\mathbb{E}}_{\mathbf{x}\sim U(\{0,1\}^{f(T)})}\ \,\operatorname*{\mathbb{E}}_{\mathbf{s}\sim U^{T},\,\mathbf{a}}$ $\ \,\left[\prod_{t=1}^{T}\mu_{f,\mathbf{x}}^{m}(s_{t})-\prod_{t=1}^{T}\mu_{f,\mathbf{x}}(s_{t},a_{t})\right]$
 $\geq$ $\ 1-\alpha(T)-\big{(}1-\alpha(T)\big{)}\exp\left(-\frac{LT}{2^{8}f(T)}\right)$
 Therefore
-$\lim_{T\to\infty}\sup_{\mu}\ \operatorname*{\mathbb{E}}_{\mathbf{s}\sim U(S^{T}),\,\mathbf{a}}\ R_{T}(\mathbf{s},\mu,\pi^{*})\geq$ $\ 1-\lim_{T\to\infty}\alpha(T)-\big{(}1-\lim_{T\to\infty}\alpha(T)\big{)}\cdot\exp\left(\lim_{T\to\infty}-\frac{LT}{2^{8}f(T)}\right)$
+$\lim_{T\to\infty}\sup_{\mu}\ \operatorname*{\mathbb{E}}_{\mathbf{s}\sim U^{T},\,\mathbf{a}}\ R_{T}(\mathbf{s},\mathbf{a},\mu,\pi^{m})\geq$ $\ 1-\lim_{T\to\infty}\alpha(T)-\big{(}1-\lim_{T\to\infty}\alpha(T)\big{)}\cdot\exp\left(\lim_{T\to\infty}-\frac{LT}{2^{8}f(T)}\right)$
 $=$ $\ 1-0-(1-0)\cdot\exp(-\infty)$
@@
 Appendix B Proof of Theorem 5.2
-B.1 Proof roadmap
+B.1 Context on Lemma 5.4
+Before diving into the main proof, we provide some context on Lemma 5.4 from Section 5:
+See 5.4
+Lemma 5.4 is a restatement and simplification of Lemma 3.5 in [33].
+First, [33] parametrize their algorithm by the expected number of queries $\hat{k}$ instead of the query probability $p=\hat{k}/T$.
+Second, [33] include a second parameter $k$, which is the eventual target number of queries for their unconditional query bound.
+In our case, an expected query bound is sufficient, so we simply set $k=\hat{k}$.
+Third, [33] provide a second bound which is tighter for small $k$; that bound is less useful for us so we omit it.
+Fourth, their number of actions $n$ is equal to $|\tilde{\Pi}|$ in our setting.
+(Their actions correspond to policies in $\tilde{\Pi}$, not our actions in $A$.) Since [33] set $\eta=\max\Big{(}\frac{1}{T}\sqrt{\frac{\hat{k}\log n}{2}},\frac{k\hat{k}}{\sqrt{2}T^{2}}\Big{)}$, we end up with $\eta=\max\Big{(}\sqrt{\frac{p\log|\tilde{\Pi}|}{2T}},\>\frac{p^{2}}{\sqrt{2}}\Big{)}$.
+Algorithm 2 provides precise pseudocode for the HedgeWithQueries algorithm to which Lemma 5.4 refers.
+1:function HedgeWithQueries($p\in(0,1],\>\text{finite policy class }\tilde{\Pi},\>\text{unknown }\ell:[T]\times\tilde{\Pi}\to[0,1]$)
+2:     $w(\pi)\leftarrow 1$ for all $\pi\in\tilde{\Pi}$
+3:     $\eta\leftarrow\max\big{(}\sqrt{\frac{p\log|\tilde{\Pi}|}{2T}},\>\frac{p^{2}}{\sqrt{2}}\big{)}$
+4:     for $t$ from $1$ to $T$ do
+[…]
+6:         if hedgeQuery then
+7:              Query and observe $\ell(t,\pi)$ for all $\pi\in\tilde{\Pi}$
+8:              $\ell^{*}\leftarrow\min_{\pi\in\tilde{\Pi}}\ell(t,\pi)$
+9:              $w(\pi)\leftarrow w(\pi)\cdot\exp(-\eta(\ell(t,\pi)-\ell^{*}))$ for all $\pi\in\tilde{\Pi}$
+10:              Select policy $\arg\min_{\pi\in\tilde{\Pi}}\ell(t,\pi)$
+11:         else
+12:              $P(\pi)\leftarrow w(\pi)/\sum_{\pi^{\prime}\in\tilde{\Pi}}w(\pi^{\prime})$ for all $\pi\in\tilde{\Pi}$
+13:              Sample $\pi_{t}\sim P$
+14:              Select policy $\pi_{t}$
+Algorithm 2 A variant of the Hedge algorithm which only observes losses in response to queries.
+B.2 Main proof
 We use the following notation throughout the proof:
   1. 1.
-    For each $t\in[T]$, let $X_{t},$ refer to the values of $X$ respectively at the start of time step $t$. Define $\pi_{t}$ similarly, except only for $t>g(T)$.
+    For each $t\in[T]$, let $X_{t}$ refer to the value of $X$ at the start of time step $t$.
   2. 2.
-    Let $V_{T}=\{t\in[T]:\mu(s_{t},a_{t})<\mu^{*}(s_{t})\}$ be the set of time steps with a suboptimal payoff.
+    Let $V_{T}=\{t\in[T]:\pi_{t}(s_{t})\neq\pi^{m}(s_{t})\}$ be the set of time steps where Hedge’s proposed action doesn’t match the mentor’s.
+Note that $|V_{T}|$ upper bounds the number of mistakes the algorithm makes (the number of mistakes could be smaller, since the algorithm sometimes queries instead of taking action $\pi_{t}(s_{t})$).
   3. 3.
-    Unless specified (e.g., $\Pr_{s\sim\mathcal{D}}$), all probabilities are with respect to $\mathbf{s}\sim\mathcal{D}^{T}$.
-The proof proceeds in the following steps:
-  1. 1.
-    Show that $\pi_{t}$ meets the criteria for the fundamental theorem of multi-class PAC learning and thus has small expected error on a single state (Lemma B.1).
-  2. 2.
-    Use Lemma B.1 and a non-standard concentration inequality (Lemma B.2) to show that $|V_{T}|$ is small with high probability. This involves two steps (Lemmas B.3 and B.4).
-  3. 3.
-    Having a small number of mistakes is not sufficient, because a single payoff of $\mu(s_{t},a_{t})=0$ can ruin the entire endeavor. Lemmas B.5 and B.6 bound the suboptimality of a single payoff.
+    For $K\subseteq S$, let $\operatorname*{vol}(K)$ denote the $n$-dimensional Lebesgue measure of $K$.
   4. 4.
-    Lemma B.7 transforms the multiplicative objective into an additive form that is easier to manipulate.
+    With slight abuse of notation, we will use inequalities of the form $f(T)\leq g(T)+O(h(T))$ to mean that there exists a constant $C$ such that $f(T)\leq g(T)+Ch(T)$.
   5. 5.
-    Lemma B.8 bounds the total regret of the algorithm.
-  6. 6.
-    Lemma B.10 bounds the total number of queries based on a packing number lemma (Lemma B.9).
-Lemmas B.8 and B.10 together imply Theorem 4.5.
-B.2 Proof
+    We will use “Case 1” to refer to finite VC dimension and $\sigma$-smooth $\mathbf{s}$ and “Case 2” to refer to finite Littlestone dimension.
+In Case 1, expectations are over the randomness of both $\mathbf{s}$ and the algorithm, while in Case 2, expectations are over just the randomness of the algorithm.
+When we need to distinction, we use $\operatorname*{\mathbb{E}}_{\mathbf{a}}$ to denote the expectation over randomness of the algorithm and $\operatorname*{\mathbb{E}}_{\mathbf{s}\sim\boldsymbol{\mathcal{D}}}$ to denote the expectation over $\mathbf{s}$.
 Lemma B.1.
-With probability at least $1-1/T$, we have $\Pr_{s\sim\mathcal{D}}[\pi_{t}(s)\neq\pi^{*}(s)]\leq\varepsilon(T)$ $\forall t>g(T)$.
+Under the conditions of Theorem 5.1, Algorithm 1 satisfies
+$\operatorname*{\mathbb{E}}[|V_{T}|]\in O\left(\frac{d}{\sigma}T\varepsilon\log(1/\varepsilon)\log T\right)$
 Proof.
-Let $\delta=1/T$. Then
-$|X_{g(T)+1}|=$ $\ g(T)$
-$=$ $\ \frac{cd}{\varepsilon(T)}\log(d\,T)$
-$=$ $\ 2\frac{cd}{\varepsilon(T)}\log(d^{2}T^{2})$
-$=$ $\ \frac{c}{\varepsilon(T)}(d\log(d^{2}T^{2})+d\log(d^{2}T^{2}))$
-$\geq$ $\ \frac{c}{\varepsilon(T)}(d\log(d\,T\cdot T^{\frac{n}{n+0.5}})+\log T)$
-$\geq$ $\ \frac{c}{\varepsilon(T)}\left(d\log(\frac{d|A|}{\varepsilon(T)})+\log\frac{1}{\delta}\right)$
-Now consider any $t>g(T)$.
-By construction, $\pi_{t}$ is $(X_{t},r)$-local ($\pi$ and $X$ are not updated on every time step, but are always updated at the same time).
-Therefore $\pi_{t}$ is consistent with $X_{t}$ and in turn is consistent with $X_{g(T)+1}$ since $X_{g(T)+1}\subseteq X_{t}$.
-Since $X_{g(T)+1}$ is a set of i.i.d.
-samples of the requisite size, Theorem 4.4 implies that with at probability at least $1-1/T$, any policy $\pi\in\Pi$ consistent with $X_{g(T)+1}$ satisfies $\Pr_{s\sim\mathcal{D}}[\pi(s)\neq\pi^{*}(s)]\leq\varepsilon(T)$.
-In particular, this holds for $\pi_{t}$ for all $t>g(T)$, completing the proof.
-∎
-Next, we wish to prove that $|V_{T}|$ is small.
-The $\{t\in V_{T}\}$ indicator variables are not independent, and standard martingale concentration bounds like the Azuma-Hoeffding inequality are not tight enough, so we use a specialized inequality from [8].
-The statement below is slightly simplified from its original source, since we do not need the more general version.
-Lemma B.2 (Corollary 2.1 in [8]).
-Let $x_{0},\dots,x_{k}$ be a supermartingale and assume there exists $b\in(0,1]$ such that $x_{i}-x_{i-1}\in[-b,1]$ for all $i\in[k]$. Then for any $x>0$,
-$\Pr[x_{k}\geq x]\leq\exp\left(-\frac{x^{2}}{2(kb+x/3)}\right)$
-Lemma B.3.
-Assume $\Pr_{s\sim\mathcal{D}}[\pi_{t}(s)\neq\pi^{*}(s)]\leq\varepsilon(T)$ for all $t>g(T)$. Then
-$\Pr\left[|V_{T}|\geq 2T\varepsilon(T)\right]\leq\exp(-3T\varepsilon(T)/8)$
-Proof.
-Let $k=T-g(T)$ for brevity.
-For each $t\in\{g(T)+1,\dots,T\}$, define a random variable $y_{t-g(T)}=\mathbf{1}_{\pi_{t}(s_{t})\neq\pi^{*}(s_{t})}-\varepsilon(T)$.
-Also let $x_{0}=0$, and for $i\in[k]$, let $x_{i}=x_{i-1}+y_{i}$.
-This defines random variables $y_{1},\dots,y_{k}$ and $x_{0},\dots,x_{k}$.
-For each $i\in[k]$, let $y_{<i}=\{y_{1},\dots,y_{i-1}\}$ and $x_{<i}=\{x_{0},\dots,x_{i-1}\}$.
-To show that $x_{0},\dots,x_{k}$ form a supermartingale, we first claim that $\operatorname*{\mathbb{E}}[y_{i}\mid y_{<i}]\leq 0$ for all $i\in[k]$.
-Regardless of what occurs on previous time steps, each state is sampled randomly from $\mathcal{D}$.
-Thus
-$\operatorname*{\mathbb{E}}[y\mid y_{<i}]=\Pr_{s\sim\mathcal{D}}[\pi_{i+g(T)}\neq\pi^{*}(s)\mid y_{<i}]-\varepsilon(T)\leq 0$
-Since $y_{<i}$ and $x_{<i}$ provide the same information (i.e., either can be constructed from the other), we have $\operatorname*{\mathbb{E}}[y_{i}\mid x_{<i}]\leq 0$. Therefore
-$\operatorname*{\mathbb{E}}[x_{i}\mid x_{<i}]=$ $\ \operatorname*{\mathbb{E}}[x_{i-1}+y_{i}-\varepsilon(T)\mid x_{<i}]$
-$=$ $\ \operatorname*{\mathbb{E}}[x_{i-1}\mid x_{<i}]+\operatorname*{\mathbb{E}}[y_{i}\mid x_{<i}]-\varepsilon(T)$
-$\leq$ $\ x_{i-1}+\varepsilon(T)-\varepsilon(T)$
-$=$ $\ x_{i-1}$
-Also, since each $x_{i}$ is bounded, we have $\operatorname*{\mathbb{E}}[x_{i}]<\infty$ for all $i\in[k]$.
-Therefore $x_{0},\dots,x_{k}$ is indeed a supermartingale.
-Furthermore, $x_{i}-x_{i-1}=y_{i}\in[-\varepsilon(T),1]$ for all $i\in[k]$, so by Lemma B.2,
-$\Pr[x_{k}\geq 2T\varepsilon(T)]\leq$ $\ \exp\left(-\frac{T^{2}\varepsilon(T)^{2}}{2(k\varepsilon(T)+T\varepsilon(T)/3)}\right)$
-Since $k\leq T$,
-$\Pr[x_{k}\geq 2T\varepsilon(T)]\leq$ $\ \exp\left(-\frac{T^{2}\varepsilon(T)^{2}}{2(T\varepsilon(T)+T\varepsilon(T)/3)}\right)$
-$=$ $\ \exp\left(-\frac{T^{2}\varepsilon(T)^{2}}{8T\varepsilon(T)/3}\right)$
-$=$ $\ \exp(-3T\varepsilon(T)/8)$
+Define $\ell:[T]\times\tilde{\Pi}\to[0,1]$ by $\ell(t,\pi)=\mathbf{1}(\pi(s_{t})\neq\pi^{m}(s_{t}))$, and let $w^{h}$ and $\pi_{t}^{h}$ denote the values of $w$ and $\pi_{t}$ respectively in HedgeWithQueries, while $w$ and $\pi_{t}$ refer to the variables in Algorithm 1.
+Then $w$ and $w^{h}$ evolve in the exact same way, so the distributions of $\pi_{t}$ and $\pi_{t}^{h}$ coincide.
+Thus by Lemma 5.4,
+$\operatorname*{\mathbb{E}}_{\mathbf{a}}\left[\sum_{t=1}^{T}\ell(t,\pi_{t})\right]-\min_{\tilde{\pi}\in\tilde{\Pi}}\sum_{t=1}^{T}\ell(t,\tilde{\pi})\leq\frac{2\log|\tilde{\Pi}|}{p^{2}}=2T\varepsilon\log|\tilde{\Pi}|$
+Since Lemma 5.4 holds for any loss function, the bound above holds for any $\mathbf{s}\in S^{T}$, so the bound also holds in expectation over $\mathbf{s}\sim\boldsymbol{\mathcal{D}}$ (which is needed for Case 1).
+Next, observe that $|V_{T}|=\sum_{t=1}^{T}\mathbf{1}(\pi_{t}(s_{t})\neq\pi^{m}(s_{t}))=\sum_{t=1}^{T}\ell(t,\pi_{t})$, so
+$\operatorname*{\mathbb{E}}_{\mathbf{a}}[|V_{T}|]\leq 2T\varepsilon\log|\tilde{\Pi}|+\min_{\tilde{\pi}\in\tilde{\Pi}}\sum_{t=1}^{T}\mathbf{1}(\tilde{\pi}(s_{t})\neq\pi^{m}(s_{t}))$
+Case 1: Since $\tilde{\Pi}$ is a smooth $\varepsilon$-cover of $\Pi$, we have
+$\operatorname*{\mathbb{E}}_{\mathbf{s}\sim\boldsymbol{\mathcal{D}}}\left[\min_{\tilde{\pi}\in\tilde{\Pi}}\sum_{t=1}^{T}\mathbf{1}(\tilde{\pi}(s_{t})\neq\pi^{m}(s_{t}))\right]\leq$ $\ \operatorname*{\mathbb{E}}_{\mathbf{s}\sim\boldsymbol{\mathcal{D}}}\left[\sup_{\pi\in\Pi}\min_{\tilde{\pi}\in\tilde{\Pi}}\sum_{t=1}^{T}\mathbf{1}(\tilde{\pi}(s_{t})\neq\pi(s_{t}))\right]$
+$\in$ $\ O\left(\frac{1}{\sigma}T\varepsilon\log T\sqrt{d\log(1/\varepsilon)}\right)$
+with the first step due to $\pi^{m}\in\Pi$ and the second step due to Lemma 5.3.
+The last component we need is that $|\tilde{\Pi}|\leq(41/\varepsilon)^{d}$ by construction (and such a $\tilde{\Pi}$ is guaranteed to exist by Lemma 5.1).
+Combining the above inequalities and taking the expectation over $\mathbf{s}\sim\boldsymbol{\mathcal{D}}$ (in addition to the randomness of the algorithm), we get
+$\operatorname*{\mathbb{E}}_{\mathbf{s}\sim\boldsymbol{\mathcal{D}},\,\mathbf{a}}\,[|V_{T}|]\leq$ $\ 2T\varepsilon\log|\tilde{\Pi}|+\operatorname*{\mathbb{E}}_{\mathbf{s}\sim\boldsymbol{\mathcal{D}}}\left[\min_{\tilde{\pi}\in\tilde{\Pi}}\sum_{t=1}^{T}\mathbf{1}(\tilde{\pi}(s_{t})\neq\pi^{m}(s_{t}))\right]$
+$\leq$ $\ 2dT\varepsilon\log(41/\varepsilon)+O\left(\frac{1}{\sigma}T\varepsilon\log T\sqrt{d\log(1/\varepsilon)}\right)$
+$\in$ $\ O\left(\frac{d}{\sigma}T\varepsilon\log(1/\varepsilon)\log T\right)$
+Case 2: Since $\tilde{\Pi}$ is an adversarial cover of $\Pi$ and $\pi^{m}\in\Pi$, there exists $\tilde{\pi}\in\tilde{\Pi}$ such that $\sum_{t=1}^{T}\mathbf{1}(\tilde{\pi}(s_{t})\neq\pi^{m}(s_{t}))=0$.
+Since $|\tilde{\Pi}|\leq(eT/d)^{d}$ (with such a $\tilde{\Pi}$ guaranteed to exist by Lemma 5.2),
+$\operatorname*{\mathbb{E}}_{\mathbf{a}}[|V_{T}|]\leq$ $\ 2T\varepsilon\log|\tilde{\Pi}|+\min_{\tilde{\pi}\in\tilde{\Pi}}\sum_{t=1}^{T}\mathbf{1}(\tilde{\pi}(s_{t})\neq\pi^{m}(s_{t}))$
+$\leq$ $\ 2T\varepsilon d\ln(eT/d)$
+$\in$ $\ O\left(\frac{d}{\sigma}T\varepsilon\log(1/\varepsilon)\log T\right)$
 as required. ∎
 Lemma B.2.
-We have $\Pr[|V_{T}|\geq 2T\varepsilon(T)]\leq\exp(-3T^{\frac{0.5}{n+0.5}}/8)+1/T$.
+For all $t\in[T]$, $\mu(s_{t},a_{t})\geq\mu^{m}(s_{t})-L\varepsilon^{1/n}$.
 Proof.
-Let $\xi$ be the event that $\Pr_{s\sim\mathcal{D}}[\pi_{t}(s)\neq\pi^{*}(s)]\leq\varepsilon(T)$ for all $t>g(T)$.
-Lemma B.1 states that $\Pr[\neg\xi]\leq 1/T$ and Lemma B.3 states that $\Pr[|V_{T}|\geq 2T\varepsilon(T)\mid\xi]\leq\exp(-3T\varepsilon(T)/8)$.
-Then
-$\Pr[|V_{T}|\geq 2T\varepsilon(T)]=$ $\ \Pr[|V_{T}|\geq 2T\varepsilon(T)\mid\xi]\Pr[\xi]+\Pr[|V_{T}|\geq 2T\varepsilon(T)\mid\neg\xi]\Pr[\neg\xi]$
-$\leq$ $\ \exp(-3T\varepsilon(T)/8)\cdot 1+1\cdot 1/T$
-$=$ $\ \exp(-3T^{1-\frac{n}{n+0.5}}/8)+1/T$
-$=$ $\ \exp(-3T^{\frac{0.5}{n+0.5}}/8)+1/T$
-as claimed. ∎
-Remark B.1.
-The astute reader note that Lemma B.5 is the only place in the paper where we invoke the assumption that the mentor’s policy is optimal.
-Both of our positive results (Theorems 4.5 and 5.2) can be generalized to a possibly suboptimal mentor policy $\pi^{m}$ if:
-  1. 1.
-    We modify the Lipschitz assumption to be $|\mu(s,\pi^{m}(s))-\mu(s,\pi^{m}(s^{\prime}))|\leq L||s-s^{\prime}||$.
-In words, the value of the mentor’s policy shouldn’t change too much between nearby states.
-  2. 2.
-    We define replace $\mu^{*}(s)$ with $\mu(s,\pi^{m}(s))$ throughout.
-Lemma B.5 follows immediately from these two assumptions.
-Also note that Lemmas B.5 and B.7 are not specific to this algorithm and will also be used in the proof of Theorem 5.2.
-Lemma B.5.
-If $||s-s^{\prime}||\leq\delta$ for some $\delta\geq 0$, then $\mu(s,\pi^{*}(s^{\prime}))\geq\mu^{*}(s)-2L\delta$.
-Proof.
-We have By the Lipschitz continuity of $\mu$, $\mu(s,\pi^{*}(s^{\prime}))\geq\mu(s^{\prime},\pi^{*}(s^{\prime}))-L||s-s^{\prime}||$.
-Also, $\mu(s^{\prime},\pi^{*}(s^{\prime}))=\max_{a\in A}\mu(s^{\prime},a)\geq\mu(s^{\prime},\pi^{*}(s))$ by assumption.
-Thus
-$\mu(s,\pi^{*}(s^{\prime}))\geq$ $\ \mu(s^{\prime},\pi^{*}(s^{\prime}))-L||s-s^{\prime}||$ (Lipschitz continuity of $\mu$)
-$\geq$ $\ \mu(s^{\prime},\pi^{*}(s))-L||s-s^{\prime}||$ (By assumption, $\pi^{*}$ is optimal)
-$\geq$ $\ \mu(s,\pi^{*}(s))-2L||s-s^{\prime}||$ (Lipschitz continuity of $\mu$ again)
-$=$ $\ \mu^{*}(s)-2L||s-s^{\prime}||$ (Definition of $\mu^{*}(s)$)
-Plugging in $||s-s^{\prime}||\leq\delta$ completes the proof. ∎
-Lemma B.6.
-Under the conditions of Theorem 4.5, $\mu(s_{t},a_{t})\geq\mu^{*}(s_{t})-2Lr$ for all $t\in[T]$.
-Proof.
-If we query the mentor at time $t$, then $\mu(s_{t},a_{t})=\mu^{*}(s_{t})$, so assume we do not query the mentor at time $t$.
-Let $(s^{\prime},a^{\prime})=\operatorname*{arg\,min}_{(s,a)\in X_{t}}||s_{t}-s||$: then $||s_{t}-s^{\prime}||\leq r$.
-Since $\pi_{t}$ is $(X_{t},r)$-local by construction, there exists $(s^{\prime\prime},a^{\prime\prime})\in X_{t}$ such that $||s_{t}-s^{\prime\prime}||\leq r$ and $a_{t}=\pi_{t}(s_{t})=a^{\prime\prime}$.
-Then by Lemma B.5, $\mu(s_{t},\pi^{*}(s^{\prime\prime}))\geq\mu^{*}(s_{t})-2Lr$.
-Furthermore, because $(s^{\prime\prime},a^{\prime\prime})\in X$, we have $\pi^{*}(s^{\prime\prime})=a^{\prime\prime}=a_{t}$.
-Therefore $\mu(s_{t},a_{t})\geq\mu^{*}(s_{t})-2Lr$.
-∎
-Lemma B.7.
-If $\mu^{*}(s_{t})\geq\mu_{0}>0$ and $c_{t}\in\big{[}0,\mu_{0}/2\big{]}$ for all $t\in[T]$, then
-$\prod_{t=1}^{T}\left(\mu^{*}(s_{t})-c_{t}\right)\geq\prod_{t=1}^{T}\mu^{*}(s_{t})-1.5\sum_{t=1}^{T}c_{t}$
-Proof.
-Recall the standard limit $\lim_{\delta\to 0}\delta^{-1}\log(1+\delta)=1$.
-Thus for all $\varepsilon>0$, there exists $\delta_{0}>0$ such that whenever $|\delta|\leq\delta_{0}$, we have $|\delta^{-1}\log(1+\delta)-1|\leq\varepsilon$.
-For our purposes, $\varepsilon=0.5$ is sufficient, we only care about $\delta<0$, and we only need the lower bound.
-One can verify that $\delta_{0}=0.5$ is sufficient.
-To summarize, we have $1.5\delta\leq\log(1+\delta)$ whenever $|\delta|\leq 0.5$.
-Let $\delta_{t}=-c_{t}/\mu^{*}(s_{t})$ for each $t\in[T]$. Then
-$|\delta_{t}|=\frac{c_{t}}{\mu^{*}(s_{t})}\leq\frac{\mu_{0}/2}{\mu_{0}}=0.5$
-and so
-$-\frac{1.5c_{t}}{\mu^{*}(s_{t})}\leq$ $\ \log\left(1-\frac{c_{t}}{\mu^{*}(s_{t})}\right)$
-$\log(\mu^{*}(s_{t}))-\frac{1.5c_{t}}{\mu^{*}(s_{t})}\leq$ $\ \log\left(\mu^{*}(s_{t})-c_{t}\right)$
-Now summing over $t$,
-$\sum_{t=1}^{T}\log(\mu^{*}(s_{t}))-\sum_{t=1}^{T}\frac{1.5c_{t}}{\mu^{*}(s_{t})}\leq$ $\ \sum_{t=1}^{T}\log\left(\mu^{*}(s_{t})-c_{t}\right)$
-$\log\prod_{t=1}^{T}\mu^{*}(s_{t})-\sum_{t=1}^{T}\frac{1.5c_{t}}{\mu^{*}(s_{t})}\leq$ $\ \log\prod_{t=1}^{T}\left(\mu^{*}(s_{t})-c_{t}\right)$
-Exponentiating both sides yields
-$\prod_{t=1}^{T}\left(\mu^{*}(s_{t})-c_{t}\right)\geq\exp\left(-\sum_{t=1}^{T}\frac{1.5c_{t}}{\mu^{*}(s_{t})}\right)\prod_{t=1}^{T}\mu^{*}(s_{t})$
-Applying the standard inequality $e^{x}\geq 1+x$ gives us
-$\prod_{t=1}^{T}\left(\mu^{*}(s_{t})-c_{t}\right)\geq$ $\ \left(1-\sum_{t=1}^{T}\frac{1.5c_{t}}{\mu^{*}(s_{t})}\right)\prod_{t=1}^{T}\mu^{*}(s_{t})$
-$=$ $\ \prod_{t=1}^{T}\mu^{*}(s_{t})-1.5\sum_{t=1}^{T}c_{t}\frac{\prod_{i=1}^{T}\mu^{*}(s_{i})}{\mu^{*}(s_{t})}$
-Since $\mu^{*}(s_{t})\in[0,1]$ for all $t\in[T]$, $\prod_{i=1}^{T}\mu^{*}(s_{i})\leq\mu^{*}(s_{t})$. Thus
-$\prod_{t=1}^{T}\left(\mu^{*}(s_{t})-c_{t}\right)\geq$ $\ \prod_{t=1}^{T}\mu^{*}(s_{t})-1.5\sum_{t=1}^{T}c_{t}\frac{\mu^{*}(s_{t})}{\mu^{*}(s_{t})}$
-$=$ $\ \prod_{t=1}^{T}\mu^{*}(s_{t})-1.5\sum_{t=1}^{T}c_{t}$
+Fix any $t\in[T]$.
+If $t\in Q_{T}$, then $\mu(s_{t},a_{t})=\mu^{m}(s_{t})$, so assume $t\not\in Q_{T}$.
+Let $(s^{\prime},a^{\prime})=\operatorname*{arg\,min}_{(s,a)\in X_{t}:\pi_{t}(s_{t})=a}||s_{t}-s||$.
+Since $t\not\in Q_{T}$, we must have $||s_{t}-s^{\prime}||\leq\varepsilon^{1/n}$.
+We have $a^{\prime}=\pi^{m}(s^{\prime})$ by construction of $X_{t}$ and $\pi_{t}(s_{t})=a^{\prime}$ by construction of $a^{\prime}$. Combining these with the local generalization assumption, we get
+$\mu(s_{t},a_{t})=\mu(s_{t},\pi_{t}(s_{t}))=\mu(s_{t},\pi^{m}(s^{\prime}))\geq\mu^{m}(s_{t})-L||s_{t}-s^{\prime}||\geq\mu^{m}(s_{t})-L\varepsilon^{1/n}$
 as required. ∎
 Lemma B.3.
+Assume $x_{1},\dots,x_{T},y_{1},\dots,y_{T}\in[0,1]$ and $x_{t}\geq y_{t}$ for all $t\in[T]$. Then
+$\prod_{t=1}^{T}x_{t}-\prod_{t=1}^{T}y_{t}\leq\sum_{t=1}^{T}x_{t}-\sum_{t=1}^{T}y_{t}$
+Proof.
+We proceed by induction on $T$.
+The claim is trivially satisfied for $T=1$, so suppose $T>1$ and assume that $\prod_{t=1}^{T-1}x_{t}-\prod_{t=1}^{T-1}y_{t}\leq\sum_{t=1}^{T-1}x_{t}-\sum_{t=1}^{T-1}y_{t}$.
+Then
+$\sum_{t=1}^{T}x_{t}-\sum_{t=1}^{T}y_{t}-\prod_{t=1}^{T}x_{t}+\prod_{t=1}^{T}y_{t}=$ $\ x_{T}\sum_{t=1}^{T-1}x_{t}-y_{T}\sum_{t=1}^{T-1}y_{t}-x_{T}\prod_{t=1}^{T-1}x_{t}+y_{T}\prod_{t=1}^{T-1}y_{t}$
+$=$ $\ x_{T}\left(\sum_{t=1}^{T-1}x_{t}-\prod_{t=1}^{T-1}x_{t}\right)-y_{T}\left(\sum_{t=1}^{T-1}y_{t}-\prod_{t=1}^{T-1}y_{t}\right)$
+Since $T>1$ and $x_{t}\in[0,1]$ for all $t\in[T]$, we have $\sum_{t=1}^{T-1}x_{t}\geq x_{1}\geq\sum_{t=1}^{T-1}x_{t}$.
+Thus $\sum_{t=1}^{T-1}x_{t}-\prod_{t=1}^{T-1}x_{t}\geq 0$.
+Combining this with $x_{T}\geq y_{T}$, we get
+$\sum_{t=1}^{T}x_{t}-\sum_{t=1}^{T}y_{t}-\prod_{t=1}^{T}x_{t}+\prod_{t=1}^{T}y_{t}=$ $\ x_{T}\left(\sum_{t=1}^{T-1}x_{t}-\prod_{t=1}^{T-1}x_{t}\right)-y_{T}\left(\sum_{t=1}^{T-1}y_{t}-\prod_{t=1}^{T-1}y_{t}\right)$
+$\geq$ $\ y_{T}\left(\sum_{t=1}^{T-1}x_{t}-\prod_{t=1}^{T-1}x_{t}\right)-y_{T}\left(\sum_{t=1}^{T-1}y_{t}-\prod_{t=1}^{T-1}y_{t}\right)$
+$=$ $\ y_{T}\left(\sum_{t=1}^{T-1}x_{t}-\prod_{t=1}^{T-1}x_{t}-\sum_{t=1}^{T-1}y_{t}+\prod_{t=1}^{T-1}y_{t}\right)$
+[…]
+The last step is due to $y_{T}\geq 0$ and our assumption of $\prod_{t=1}^{T-1}x_{t}-\prod_{t=1}^{T-1}y_{t}\leq\sum_{t=1}^{T-1}x_{t}-\sum_{t=1}^{T-1}y_{t}$. ∎
+Lemma B.4.
 Under the conditions of Theorem 5.1, Algorithm 1 satisfies
-$\operatorname*{\mathbb{E}}_{\mathbf{s}\sim\mathcal{D}^{T}}\left[R_{T}(\mathbf{s},\mu,\pi^{*})\right]\leq\frac{6L}{T^{\frac{0.5}{n+0.5}}}+\frac{1}{T}+\exp(-3T^{\frac{0.5}{n+0.5}}/8)$
+$\operatorname*{\mathbb{E}}\left[R_{T}\right]\in O\left(\frac{dL}{\sigma}T\varepsilon^{1+1/n}\log(1/\varepsilon)\log T\right)$
 Proof.
-Let $\mu_{0}=4Lr$. If there exists $t\in[T]$ with $\mu^{*}(s_{t})<\mu_{0}$, then
-$R_{T}(\mathbf{s},\mu,\pi^{*})=\prod_{t=1}^{T}\mu^{*}(s_{t})-\prod_{t=1}^{T}\mu(s_{t},a_{t})<\mu_{0}-0=4L\varepsilon(T)^{1/n}=4LT^{\frac{-1}{n+0.5}}\leq 6LT^{\frac{-0.5}{n+0.5}}$
-so the claim is satisfied. Thus assume $\mu^{*}(s_{t})\geq\mu_{0}$ for all $t\in[T]$. Let $c_{t}=0$ for $t\not\in V_{T}$ and $c_{t}=2Lr$ for $t\in V_{T}$. Then by Lemma B.6,
-$\prod_{t=1}^{T}\mu(s_{t},a_{t})=$ $\ \prod_{t\in V_{T}}\mu(s_{t},a_{t})\prod_{t\not\in V_{T}}\mu(s_{t},a_{t})$
-$\geq$ $\ \prod_{t\in V_{T}}\left(\mu^{*}(s_{t})-2Lr\right)\prod_{t\not\in V_{T}}\mu^{*}(s_{t})$
-$=$ $\ \prod_{t=1}^{T}\left(\mu^{*}(s_{t})-c_{t}\right)$
-Since $c_{t}\in[0,\mu_{0}/2]$ for all $t\in[T]$, Lemma B.7 implies that
-$\prod_{t=1}^{T}\mu(s_{t},a_{t})\geq$ $\ \prod_{t=1}^{T}\mu^{*}(s_{t})-1.5\sum_{t=1}^{T}c_{t}$
-$=$ $\ \prod_{t=1}^{T}\mu^{*}(s_{t})-3L|V_{T}|r$
-If $|V_{T}|<2T\varepsilon(T)$, then
-$R_{T}(\mathbf{s},\mu,\pi^{*})=$ $\ \prod_{t=1}^{T}\mu^{*}(s_{t})-\prod_{t=1}^{T}\mu(s_{t},a_{t})$
-$\leq$ $\ 6LT\varepsilon(T)r$
-$=$ $\ 6LT\varepsilon(T)^{(n+1)/n}$
-$=$ $\ 6LT^{(1+\frac{-n}{n+0.5}\cdot\frac{n+1}{n})}$
-$=$ $\ 6LT^{(\frac{n+0.5}{n+0.5}-\frac{n+1}{n+0.5})}$
-$=$ $\ 6LT^{\frac{-0.5}{n+0.5}}$
-Then with the help of Lemma B.4 (and noting that $R_{T}(\mathbf{s},\mu,\pi^{*})\leq 1$ always),
-$\operatorname*{\mathbb{E}}_{\mathbf{s}\sim\mathcal{D}^{T}}\left[R_{T}(\mathbf{s},\mu,\pi^{*})\right]\leq$ $\ \Pr[|V_{T}|\geq 2T\varepsilon(T)]\cdot 6LT^{\frac{-0.5}{n+0.5}}+\Pr[|V_{T}|\geq 2T\varepsilon(T)]\cdot 1$
-$\leq$ $\ 6LT^{\frac{-0.5}{n+0.5}}+\frac{1}{T}+\exp(-3T^{\frac{0.5}{n+0.5}}/8)$
+We first claim that $a_{t}=\pi^{m}(s_{t})$ for all $t\not\in V_{T}$.
+If $t\in Q_{T}$, the claim is immediate; if not, we have $a_{t}=\pi_{t}(s_{t})$, and $\pi_{t}(s_{t})=\pi^{m}(s_{t})$ due to $t\not\in V_{T}$.
+Thus $\min(\mu^{m}(s_{t}),\mu(s_{t},a_{t}))=\mu^{m}(s_{t})$ for $t\not\in V_{T}$.
+We next claim that $\mu^{m}(s_{t})-\min(\mu^{m}(s_{t}),\mu(s_{t},a_{t}))\leq L\varepsilon^{1/n}$ for all $t\in[T]$.
+If $\mu(s_{t},a_{t})\leq\mu^{m}(s_{t})$, this follows from Lemma B.2.
+If $\mu(s_{t},a_{t})>\mu^{m}(s_{t})$, then $\mu^{m}(s_{t})-\min(\mu^{m}(s_{t}),\mu(s_{t},a_{t}))=0\leq L\varepsilon^{1/n}$.
+Therefore
+$\sum_{t=1}^{T}\Big{(}\mu^{m}(s_{t})-\min(\mu^{m}(s_{t}),\mu(s_{t},a_{t}))\Big{)}\leq$ $\ \sum_{t\in V_{T}}\Big{(}\mu^{m}(s_{t})-\min(\mu^{m}(s_{t}),\mu(s_{t},a_{t}))\Big{)}$
+$\leq$ $\ \sum_{t\in V_{T}}L\varepsilon^{1/n}$
+$=$ $\ |V_{T}|L\varepsilon^{1/n}$
+Now let $x_{t}=\mu^{m}(s_{t})$ and $y_{t}=\min(\mu^{m}(s_{t}),\mu(s_{t},a_{t}))$ for all $t\in[T]$. Then by Lemma B.3,
+$\prod_{t=1}^{T}\mu^{m}(s_{t})-\prod_{t=1}^{T}\min(\mu^{m}(s_{t}),\mu(s_{t},a_{t}))\leq$ $\ \sum_{t=1}^{T}\Big{(}\mu^{m}(s_{t})-\min(\mu^{m}(s_{t}),\mu(s_{t},a_{t}))\Big{)}$
+Since $\mu(s_{t},a_{t})\geq\min(\mu^{m}(s_{t}),\mu(s_{t},a_{t}))$ for all $t\in[T]$, we have
+$R_{T}=$ $\ \prod_{t=1}^{T}\mu^{m}(s_{t})-\prod_{t=1}^{T}\mu(s_{t},a_{t})$
+$\leq$ $\ \sum_{t=1}^{T}\Big{(}\mu^{m}(s_{t})-\min(\mu^{m}(s_{t}),\mu(s_{t},a_{t}))\Big{)}$
+$\leq$ $\ |V_{T}|L\varepsilon^{1/n}$
+Taking the expectation and applying Lemma B.1 completes the proof. ∎
 Definition B.1.
-Let $(V,||\cdot||)$ be a normed vector space and let $\delta>0$.
-Then $W\subseteq V$ is a $\delta$-packing of $V$ if for all $x,y\in W$, $||x-y||>\delta$.
-The $\delta$-packing number of $V$, denoted $\mathcal{M}(V,||\cdot||,\delta)$, is the maximum cardinality of any $\delta$-packing of $V$.
+Let $(K,||\cdot||)$ be a normed vector space and let $\delta>0$.
+Then $W\subseteq K$ is a $\delta$-packing of $K$ if for all $x,y\in W$, $||x-y||>\delta$.
+The $\delta$-packing number of $K$, denoted $\mathcal{M}(K,||\cdot||,\delta)$, is the maximum cardinality of any $\delta$-packing of $K$.
+In this paper, we only consider the Euclidean distance norm, so we just write $M(K,||\cdot||,\delta)=M(K,\delta)$.
 Lemma B.5 (Theorem 14.2 in [40]).
-Let $B$ be the unit ball and let |||||$\cdot$||| | ⋅ | | denote Euclidean distance. Assume $V\subset\mathbb{R}^{n}$ is convex, and $\delta B\subset V$ for some $\delta>0$. Then
-$\mathcal{M}(V,||\cdot||,\delta)\leq\left(\frac{3}{\delta}\right)^{n}\frac{\operatorname*{vol}(V)}{\operatorname*{vol}(B)}$
+If $K\subset\mathbb{R}^{n}$ is convex, bounded, and contains a ball with radius $\delta>0$, then
+$\mathcal{M}(K,\delta)\leq\frac{3^{n}\operatorname*{vol}(K)}{\delta^{n}\operatorname*{vol}(B)}$
+where $B$ is a unit ball.
+Lemma B.6 (Jung’s Theorem [22]).
+[…]
 Lemma B.7.
+Under the conditions of Theorem 5.1, Algorithm 1 satisfies
+$\operatorname*{\mathbb{E}}[|Q_{T}|]\in O\left(\sqrt{\frac{T}{\varepsilon}}+\frac{d}{\sigma}T\varepsilon\log(1/\varepsilon)\log T+\frac{\operatorname*{diam}(\mathbf{s})^{n}}{\varepsilon}\right)$
+Proof.
 […]
+Next, fix an $a\in A$ and let $S_{a}=\{s\in\mathbf{s}:\pi^{m}(s)=a\}$ be the set of observed states which share a mentor action.
+We claim that $\hat{S}_{2}\cap S_{a}$ is a packing of $S_{a}$.
+Suppose instead that there exists $s,s^{\prime}\in\hat{S}_{2}\cap S_{a}$, with $||s-s^{\prime}||\leq\varepsilon^{1/n}$.
+WLOG assume $s$ was queried after $s^{\prime}$ and let $t$ be the time step on which $s$ was queried.
+Then $(s^{\prime},\pi^{m}(s^{\prime}))\in X_{t}$.
+Also, $s,s^{\prime}\in\hat{S}_{2}\cap S_{a}$ implies that and $\pi_{t}(s_{t})=\pi^{m}(s_{t})=a=\pi^{m}(s^{\prime})$.
+Therefore
+$\min_{(s^{\prime\prime},a^{\prime\prime})\in X_{t}:a^{\prime\prime}=\pi_{t}(s_{t})}||s_{t}-s^{\prime\prime}||\leq||s_{t}-s^{\prime}||\leq\varepsilon^{1/n}$
+which contradicts $s_{t}\in\hat{S}$. Thus $\hat{S}_{2}\cap S_{a}$ is a $\varepsilon^{1/n}$-packing of $S_{a}$.
+By Lemma B.6, there exists a ball $B_{1}$ of diameter $\operatorname*{diam}(\mathbf{s})\sqrt{\frac{n}{2(n+1)}}$ which contains $\mathbf{s}$.
+Let $R=\operatorname*{diam}(\mathbf{s})\sqrt{\frac{n}{8(n+1)}}$ denote the radius of $B_{1}$.
+Let $B_{2}$ be the ball with the same center as $B_{1}$ but with radius $\max(R,\varepsilon^{1/n})$.
+Since $S_{a}\subset\mathbf{s}\subset B_{1}\subset B_{2}$, $\hat{S}_{2}\cap S_{a}$ is also a $\varepsilon^{1/n}$-packing of $B_{2}$.
+Also, $B_{2}$ must contain a ball of radius $\varepsilon^{1/n}$, so Lemma B.5 implies that
+$|\hat{S}_{2}\cap S_{a}|\leq$ $\ \mathcal{M}(B_{2},\varepsilon^{1/n})$
+$\leq$ $\ \frac{3^{n}\operatorname*{vol}(B_{2})}{\varepsilon\operatorname*{vol}(B)}$
+$=$ $\ \big{(}\max(R,\varepsilon^{1/n})\big{)}^{n}\,\frac{3^{n}\operatorname*{vol}(B)}{\varepsilon\operatorname*{vol}(B)}$
+$=$ $\ \max\left(\operatorname*{diam}(\mathbf{s})^{n}\left(\frac{n}{8(n+1)}\right)^{n/2},\>\varepsilon\right)\frac{3^{n}}{\varepsilon}$
+$\leq$ $\ O\left(\frac{\operatorname*{diam}(\mathbf{s})^{n}}{\varepsilon}+1\right)$
+(The $+1$ is necessary for now since $\operatorname*{diam}(\mathbf{s})$ could theoretically be zero.) Therefore
+$\operatorname*{\mathbb{E}}[|Q_{T}|]=$ $\ \sqrt{\frac{T}{\varepsilon}}+\operatorname*{\mathbb{E}}[|\hat{S}|]$
+$=$ $\ \sqrt{\frac{T}{\varepsilon}}+\operatorname*{\mathbb{E}}[|\hat{S}_{1}|]+\operatorname*{\mathbb{E}}\left[\sum_{a\in A}|\hat{S}_{2}\cap S_{a}|\right]$
+$\leq$ $\ \sqrt{\frac{T}{\varepsilon}}+O\left(\frac{d}{\sigma}T\varepsilon\log(1/\varepsilon)\log T\right)+\sum_{a\in A}O\left(\frac{\operatorname*{diam}(\mathbf{s})^{n}}{\varepsilon}+1\right)$
+$\leq$ $\ \sqrt{\frac{T}{\varepsilon}}+O\left(\frac{d}{\sigma}T\varepsilon\log(1/\varepsilon)\log T\right)+|A|\cdot O\left(\frac{\operatorname*{diam}(\mathbf{s})^{n}}{\varepsilon}+1\right)$
+$\leq$ $\ O\left(\sqrt{\frac{T}{\varepsilon}}+\frac{d}{\sigma}T\varepsilon\log(1/\varepsilon)\log T+\frac{\operatorname*{diam}(\mathbf{s})^{n}}{\varepsilon}\right)$
+as required. ∎
+Theorem 5.1 follows from Lemmas B.4 and B.7:
+See 5.1
+We then perform some arithmetic to get Theorem 5.2:
+See 5.2
 Proof.
-Let $\tilde{S}=\{s_{t}:t\in Q_{T}\text{ and }t>g(T)\}$ be the set of states queried after $t=g(T)$.
-Since we always query on the first $g(T)$ time steps, the total number of queries is $g(T)+|\tilde{S}|$.
-We claim that for any $s,s^{\prime}\in\tilde{S}$, we have $||s-s^{\prime}||>r$.
-Suppose not, and WLOG assume $s$ was queried after $s^{\prime}$, with $s=s_{t}$.
-Since $s^{\prime}$ was queried earlier, $(s^{\prime},\pi^{*}(s^{\prime}))\in X_{t}$.
-Therefore
-$\min_{(s^{\prime\prime},a)\in X_{t}}||s-s^{\prime\prime}||\leq||s-s^{\prime}||\leq r$
-But then we would have taken the action $\pi_{t}(s_{t})$ instead of querying, which is a contradiction.
-Hence $\tilde{S}$ is a $r$-packing of $\operatorname*{conv}(\mathbf{s})$, and thus is also an $r$-packing of $\operatorname*{conv}(\mathbf{s}\cup rB)$.
-Since $rB\subset\operatorname*{conv}(\mathbf{s}\cup rB)$, Lemma B.9 implies that
-$|\tilde{S}|\leq\mathcal{M}(\operatorname*{conv}(\mathbf{s}\cup rB),||\cdot||,r)\leq\left(\frac{3}{r}\right)^{n}\frac{\operatorname*{vol}(\operatorname*{conv}(\mathbf{s}\cup rB))}{\operatorname*{vol}(B)}=\frac{3^{n}\operatorname*{vol}(\operatorname*{conv}(\mathbf{s}\cup rB))}{\varepsilon(T)\operatorname*{vol}(B)}$
-Theorem 4.5 follows from Lemmas B.8 and B.10.
-Appendix C Proof of Theorem 5.2
-C.1 Notation for the proof
-We will use the following notation throughout the proof of Theorem 5.2:
-  • Let $V_{T}=\{t\in[T]:\mu(s_{t},a_{t})<\mu^{*}(s)\}$ be the set of time steps with a suboptimal payoff.
+We have
+$\operatorname*{\mathbb{E}}\left[R_{T}\right]\in$ $\ O\left(\frac{dL}{\sigma}T^{1-\frac{2n}{2n+1}-\frac{2}{2n+1}}\log(1/\varepsilon)\log T\right)$
+$=$ $\ O\left(\frac{dL}{\sigma}T^{\frac{-1}{2n+1}}\log T\right)$
+and
+$\operatorname*{\mathbb{E}}[|Q_{T}|]\in$ $\ O\left(\sqrt{T^{1+\frac{2n}{2n+1}}}+\frac{d}{\sigma}T^{1-\frac{-2n}{2n+1}}\log(T^{\frac{2n}{2n+1}})\log T+T^{\frac{2n}{2n+1}}\operatorname*{diam}(\mathbf{s})^{n}\right)$
+$=$ $\ O\left(T^{\frac{2n+0.5}{2n+1}}+\frac{d}{\sigma}T^{\frac{1}{2n+1}}\log T+T^{\frac{2n}{2n+1}}\operatorname*{diam}(\mathbf{s})^{n}\right)$
+$\leq$ $\ O\left(T^{\frac{4n+1}{4n+2}}\Big{(}\frac{d}{\sigma}\log T+\operatorname*{diam}(\mathbf{s})^{n}\Big{)}\right)$
+See 5.2.1
+Proof.
+Let $\xi$ be the event that Algorithm 1 exceeds $T^{\frac{4n+1.5}{4n+2}}(\frac{d}{\sigma}\log T+\operatorname*{diam}(\mathbf{s})^{n})$ queries. Noting that $|Q_{T}|\leq T$ always, we have
+$\operatorname*{\mathbb{E}}\big{[}|Q_{T}|\big{]}=$ $\ \Pr[\xi]\,\operatorname*{\mathbb{E}}\big{[}|Q_{T}|\mid\xi\big{]}+\Pr[\neg\xi]\,\operatorname*{\mathbb{E}}\big{[}|Q_{T}|\mid\neg\xi\big{]}$
+$\geq$ $\ \Pr[\xi]\,T^{\frac{4n+1.5}{4n+2}}\left(\frac{d}{\sigma}\log T+\operatorname*{diam}(\mathbf{s})^{n}\right)$
+Therefore by Theorem 5.2,
+$\Pr[\xi]\leq\frac{O\left(T^{\frac{4n+1}{4n+2}}\left(\frac{d}{\sigma}\log T+\operatorname*{diam}(\mathbf{s})^{n}\right)\right)}{T^{\frac{4n+1.5}{4n+2}}\left(\frac{d}{\sigma}\log T+\operatorname*{diam}(\mathbf{s})^{n}\right)}\leq O\left(T^{\frac{-0.5}{4n+2}}\right)$
+If $\xi$ does not occur, then the regret bound from Theorem 5.2 holds. If not, we still have $R_{T}\leq 1$ unconditionally. Thus the regret of the modified version of Algorithm 1 satisfies
+$\operatorname*{\mathbb{E}}[R_{T}]\leq$ $\ \Pr[\xi]\,\operatorname*{\mathbb{E}}[R_{T}\mid\xi]+\Pr[\neg\xi]\,\operatorname*{\mathbb{E}}[R_{T}\mid\neg\xi]$
+$\leq$ $\ O\left(T^{\frac{-0.5}{4n+2}}\right)\cdot 1+1\cdot O\left(\frac{dL}{\sigma}T^{\frac{-1}{2n+1}}\log T\right)$
+$\leq$ $\ O\left(\frac{dL}{\sigma}T^{\frac{-1}{8n+4}}\log T\right)$
+B.3 Adaptive adversaries
+If $s_{t}$ is allowed to depend on the events of prior time steps, we say that the adversary is adaptive.
+In contrast, a non-adaptive or “oblivious” adversary must choose the entire input upfront.
+This distinction is not relevant for deterministic algorithms, since an adversary knows exactly how the algorithm will behave for any input.
+In other words, the adversary gains no new information during the execution of the algorithm.
+For randomized algorithms, an adaptive adversary can base the choice of $s_{t}$ on the results of randomization on previous time steps (but not on the current time step), while an oblivious adversary cannot.
+In the standard online learning model, Hedge guarantees sublinear regret against both oblivious and adaptive adversaries (Chapter 5 of [35] or Chapter 21 of [34]).
+However, [33] state their result only for oblivious adversaries.
+In order for our overall proof of Theorem 5.1 to hold for adaptive adversaries, Lemma 5.4 (Lemma 3.5 in [33]) must also hold for adaptive adversaries.
+In this section, we argue why the proof of Lemma 5.4 (Lemma 3.5 in their paper) goes through for adaptive adversaries as well.
+For this rest of Section B.3, lemma numbers refer to the numbering in [33].
+The importance of independent queries.
+Recall from Section B.1 that [33] allow two separate parameters $k$ and $\hat{k}$, which we unify for simplicity.
+Recall also that Lemma 3.5 refers to the variant of Hedge which queries with probability $p=\hat{k}/T=k/T$ independently on each time step (Algorithm 2.
+More precisely, on each time step $t$, the algorithm samples a Bernoulli random variable $X_{t}\sim\text{Ber}(p)$ and queries if $X_{t}=1$.
+The key idea is that $X_{t}$ is independent of events on previous time steps.
+Thus even conditioning on the history up to time $t$, for any for any random variable $Y_{t}$ we can write
+$\operatorname*{\mathbb{E}}[Y_{t}]=(1-p)\operatorname*{\mathbb{E}}[Y_{t}\mid X_{t}=0]+p\operatorname*{\mathbb{E}}[Y_{t}\mid X_{t}=1]$
+This insight immediately extends Observation 3.3 to adaptive adversaries (with the minor modification that queries are now issued independently with probability $p$ on each time step instead of issuing $k$ uniformly distributed queries).
+Specifically, using the notation from [33] where $i_{t}$ is the action chosen at time $t$, $i_{t}^{0}$ is the action chosen at time $t$ if a query is not issued, and $i_{t}^{*}$ is the optimal action at time $t$, we have
+$\operatorname*{\mathbb{E}}[\ell_{t}(i_{t})]=(1-p)\operatorname*{\mathbb{E}}[\ell_{t}(i_{t}^{0})]+p\operatorname*{\mathbb{E}}[\ell_{t}(i_{t}^{*})]=\left(1-\frac{k}{T}\right)\operatorname*{\mathbb{E}}[\ell_{t}(i_{t}^{0})]+\frac{k}{T}\operatorname*{\mathbb{E}}[\ell_{t}(i_{t}^{*})]$
+The same logic applies to other statements like $\operatorname*{\mathbb{E}}[\hat{\ell}_{t}(i)\mid X_{\leq t-1},I_{\leq t-1}]=\ell_{t}(i)-\ell_{t}(i_{t}^{*})$ and immediately extends those statements to adaptive adversaries as well.
+Applying Observation 3.3.
+The other tricky part of the proof is applying Observation 3.3 using a new loss function $\hat{\ell}$ defined by ^$\hat{\ell}_{t}=\frac{T}{\hat{k}}(\ell_{t}(i)-\ell_{t}(i_{t}^{*}))\mathbf{1}(X_{t}=1)$.
+To do so, we must argue that standard Hedge run on $\hat{\ell}$ is the “counterpart without queries” of HedgeWithQueries.
+Specifically, both algorithms must have the same weight vectors on every time step, and the only difference should be that HedgeWithQueries takes the optimal action on each time step independently with probability $p$ (and otherwise behaves the same as standard Hedge).
+On time steps with $X_{t}=0$, standard Hedge observes ^$\hat{\ell}_{t}(i)=0$ for all actions $i$ and thus makes no updates, and HedgeWithQueries makes no updates by definition.
+On time steps with $X_{t}=1$, both algorithms perform the typical updates $w_{t+1}(i)=w_{t}(i)\cdot\exp(-\eta(\hat{\ell}_{t}(i)-\hat{\ell}_{t}(i_{t}^{*})))$.
+Thus the weight vectors are the same for both algorithms on every time step.
+Furthermore, HedgeWithQueries takes the optimal action at time $t$ iff $X_{t}=1$, which occurs independently with probability $p$ on each time step.
+Thus standard Hedge run on $\hat{\ell}$ is the “counterpart without queries” of HedgeWithQueries.
+The rest of the proof.
+The other elements of the proof of Lemma 3.5 are as follows:
+  1. 1.
+    Lemma 3.1, which analyzes the standard version of Hedge (i.e., no queries and losses are observed on every time step).
+  2. 2.
+    Applying Lemma 3.1 to a $\hat{\ell}$.
+  3. 3.
+    Arithmetic and rearranging terms.
+The proof of Lemma 3.1 relies on simple arithmetic properties of the Hedge weights.
+Regardless of the adversary’s behavior, $\hat{\ell}$ is a well-defined loss function, so Lemma 3.1 can be applied.
+Step 3 clearly has no dependence on the type of adversary.
+Thus we conclude that Lemma 3.5 extends to adaptive adversaries.
+Appendix C Generalizing Theorem 5.2 to many actions
+We use the standard “one versus rest” reduction (see, e.g., Chapter 29 of [34]).
+For each action $a$, we will learn a binary classifier which predicts whether action $a$ is the mentor’s action.
+Formally, for each $a\in A$, define the policy class $\Pi_{a}=\{\pi_{a}:\pi\in\Pi\text{ and }\pi_{a}(s)=\mathbf{1}(\pi(s)=a))\ \forall s\in S\}$.
+Informally, for each policy $\pi:S\to A$ in $\Pi$, there exists a policy $\pi_{a}:S\to\{0,1\}$ in $\Pi_{a}$ such that $\pi_{a}(s)=\mathbf{1}(\pi(s)=a)$ for all $s\in S$.
+Algorithm 3 runs one copy of our binary-action algorithm Algorithm 1 for each action $a\in A$.
+At each time step $t$, the copy for action $a$ returns an action $b_{t}^{a}$, with $b_{t}^{a}=1$ indicating a belief that $a=\pi^{m}(s_{t})$ and $b_{t}^{a}=0$ indicating a belief that $a\neq\pi^{m}(s_{t})$.
+(Note that $b_{t}^{a}=\hat{a}$ is also possible, indicating that the mentor was queried.)
+The key idea is that if $b_{t}^{a}$ is correct for each action $a$, there will be exactly one $a$ such that $b_{t}^{a}=1$, and specifically it will be $a=\pi^{m}(s_{t})$.
+Thus we are guaranteed to take the mentor’s action on such time steps.
+The analysis for Theorem 5.2 (specifically, Lemma B.1) bounds the number of time steps when a given copy of Algorithm 1 is incorrect, so by the union bound, the number of time steps where any copy is incorrect is $|A|$ times that bound.
+That in turn bounds the number of time steps where Algorithm 3 takes an action other than the mentor’s.
+Similarly, the number of queries made by Algorithm 3 is at most $|A|$ times the bound from Theorem 5.2.
+The result is the following theorem:
+1:function AvoidCatastropheManyActions($T\in\mathbb{N},\>\varepsilon\in\mathbb{R}_{>0},\>d\in\mathbb{N},\,$ policy class $\Pi$)
+2:     for $a\in A$ do
+3:         if $\Pi$ has VC dimension $d$ then
+4:              $\tilde{\Pi}_{a}\leftarrow$ any smooth $\varepsilon$-cover of $\Pi$ of size at most $(41/\varepsilon)^{d}$
+5:         else if $\Pi$ has Littlestone dimension $d$ then
+6:              $\tilde{\Pi}_{a}\leftarrow$ any adversarial $\varepsilon$-cover of size at most $(eT/d)^{d}$
+7:     for $t$ from $1$ to $T$ do
+8:         for $a\in A$ do
+9:              $b_{t}^{a}\leftarrow$ action from running one step of Algorithm 1 on $\Pi_{a}$ (with the same $T,\varepsilon,d$)
+10:         if $b_{t}^{a}\neq\hat{a}$ $\forall a\in A$ and $\exists a\in A:b_{t}^{a}=1$ then
+11:              Take any action $a$ with $b_{t}^{a}=1$
+12:         else
+13:              Query the mentor
+Algorithm 3 extends Algorithm 1 to many actions.
+Theorem C.1.
+Assume $\pi^{m}\in\Pi$ where either (1) $\Pi_{a}$ has finite VC dimension $d$ and $\mathbf{s}$ is $\sigma$-smooth or (2) $\Pi_{a}$ has finite Littlestone dimension $d$ for all $a\in A$.
+Then for any $T\in\mathbb{N}$, Algorithm 3 with $T$ and $\varepsilon=T^{\frac{-2n}{2n+1}}$ satisfies
+$\operatorname*{\mathbb{E}}\left[R_{T}\right]\in$ $\ O\left(\frac{|A|dL}{\sigma}T^{\frac{-1}{2n+1}}\log T\right)$
+$\operatorname*{\mathbb{E}}[|Q_{T}|]\in$ $\ O\left(|A|T^{\frac{4n+1}{4n+2}}\left(\frac{d}{\sigma}\log T+\operatorname*{diam}(\mathbf{s})^{n}\right)\right)$
+We use the following terminology and notation in the proof of Theorem C.1:
+  1. 1.
+    We refer to the copy of Algorithm 1 running on $\Pi_{a}$ as “copy $a$ of Algorithm 1”.
+  2. 2.
+    Let $\pi_{t}^{a}$ and $X_{t}^{a}$ refer to the values of $\pi_{t}$ and $X_{t}$ for copy $a$ of Algorithm 1.
+  3. 3.
+    Let $\pi^{ma}:S\to\{0,1\}$ be the policy defined by $\pi^{ma}(s)=\mathbf{1}(\pi^{m}(s_{t})=a)$.
+Note that querying the mentor tells the agent $\pi^{m}(s_{t})$, which allows the agent to compute $\pi^{ma}(s_{t})$: this is necessary when Algorithm 1 queries while running on some $\Pi_{a}$.
+  4. 4.
+    Let $V_{T}^{a}=\{t\in[T]:b_{t}^{a}\neq\pi^{ma}(s_{t})\}$ be the set of time steps where $\pi_{t}^{a}$ does not correctly determine whether the mentor would take action $a$ and let $V_{T}=\{t\in[T]:a_{t}\neq\pi^{m}(s_{t})\}$ be the set of time steps where the agent’s action doesn’t match the mentor’s.
+Lemma C.1.
+We have $|V_{T}|\leq\sum_{a\in A}|V_{T}^{a}|$.
+Proof.
+We claim that $V_{T}\subseteq\cup_{a\in A}V_{T}^{a}$.
+Suppose the opposite: then there exists $t\in V_{T}$ such that $b_{t}^{a}=\pi^{ma}(s_{t})$ for all $a\in A$.
+Since $\pi^{m}(s_{t})\in A$, there is exactly one $a\in A$ such that $\mathbf{1}(\pi^{m}(s_{t})=a)=\pi^{ma}(s_{t})=b_{t}^{a}=1$.
+Specifically, this holds for $a=\pi^{m}(s_{t})$.
+But then Algorithm 3 takes action $\min\{a\in A:b_{t}^{a}=1\}=\pi^{m}(s_{t})$, which contradicts $t\in V_{T}$.
+Therefore $V_{T}\subseteq\cup_{a\in A}V_{T}^{a}$, and applying the union bound completes the proof.
+∎
+Lemma C.2.
+For all $t\in[T]$, $\mu^{m}(s_{t})-\mu(s_{t},a_{t})\leq L\varepsilon^{1/n}$.
+Proof.
+The argument is similar to the proof of Lemma B.2.
+If $\mu^{m}(s_{t})\neq\mu(s_{t},a_{t})$, then $a_{t}=a$ for some $a\in A$ where $b_{t}^{a}=1$.
+Therefore copy $a$ of Algorithm 1 did not query at time $t$ and $\pi_{t}^{a}(s_{t})=1$.
+Let $(s^{\prime},a^{\prime})=\operatorname*{arg\,min}_{(s,a)\in X_{t}^{a}:\pi_{t}^{a}(s_{t})=a}||s_{t}-s||$.
+Then $||s_{t}-s^{\prime}||\leq\varepsilon^{1/n}$ and $a^{\prime}=\pi_{t}^{a}(s_{t})=1$.
+By construction of $X_{t}^{a}$, $a^{\prime}=\pi^{ma}(s^{\prime})$ so $\pi^{ma}(s^{\prime})=1$ which implies $\pi^{m}(s^{\prime})=a$. Then by the local generalization assumption,
+$\mu(s_{t},a_{t})=\mu(s_{t},a)=\mu(s_{t},\pi^{m}(s^{\prime}))\geq\mu^{m}(s_{t})-L||s_{t}-s^{\prime}||\geq\mu^{m}(s_{t})-L\varepsilon^{1/n}$
+as required. ∎
+See C.1
+Proof.
+Theorem 5.2 implies that each copy of Algorithm 1 makes $O\big{(}T^{\frac{4n+1}{4n+2}}\big{(}\frac{d}{\sigma}\log T+\operatorname*{diam}(\mathbf{s})^{n}\big{)}\big{)}$ queries in expectation, so by linearity of expectation, the expected number of queries made by Algorithm 3 is $O\big{(}|A|T^{\frac{4n+1}{4n+2}}\big{(}\frac{d}{\sigma}\log T+\operatorname*{diam}(\mathbf{s})^{n}\big{)}\big{)}$.
+Using the same argument as in the proof of Lemma B.7 (with Lemma C.2 replacing Lemma B.2), we get
+$R_{T}=$ $\ \prod_{t=1}^{T}\mu^{m}(s_{t})-\prod_{t=1}^{T}\mu(s_{t},a_{t})$
+$\leq$ $\ \sum_{t=1}^{T}\Big{(}\mu^{m}(s_{t})-\min(\mu^{m}(s_{t}),\mu(s_{t},a_{t}))\Big{)}$
+$\leq$ $\ |V_{T}|L\varepsilon^{1/n}$
+Then by Lemma C.1, $R_{T}\leq L\varepsilon^{1/n}\sum_{a\in A}|V_{T}^{a}|$. Taking the expectation and applying Lemma B.1 to each $V_{T}^{a}$ gives us
+$\operatorname*{\mathbb{E}}[R_{T}]\leq L\varepsilon^{1/n}\sum_{a\in A}O\left(\frac{d}{\sigma}T\varepsilon\log(1/\varepsilon)\log T\right)=O\left(|A|L\varepsilon^{1/n}\frac{d}{\sigma}T\varepsilon\log(1/\varepsilon)\log T\right)$
+as required. ∎
+Appendix D There exist policy classes which are learnable in our setting but not in the standard online model
+This section presents another algorithm with subconstant regret and sublinear queries, but under different assumptions.
+The most important aspect of our algorithm is that it can handle the class of thresholds on $[0,1]$, which is known to have infinite Littlestone dimension and thus be hard in the standard online learning model.
+(Example 21.4 in [34]).
+Specifically, we assume a 1D state space and we allow the state sequence to be fully adversarial chosen. Instead of VC/Littlestone dimension, we consider the following notion of simplicity:
+Definition D.1.
+Given a mentor policy $\pi^{m}$, partition the state space $S$ into intervals such that all states within each interval share the same mentor action.
+Let $\{S_{1},\dots,S_{k}\}$ be a partition that minimizes the number of intervals.
+We call each $S_{j}$ a segment.
+Let $f(\pi^{m})$ denote the number of segments in $\pi^{m}$.
+Bounding the number of segments is similar conceptually to VC dimension in that it limits the ability of the policy class to realize arbitrary combinations of labels (i.e., mentor actions) on $\mathbf{s}$.
+For example, if $\Pi$ is the class of thresholds on $[0,1]$, every $\pi\in\Pi$ has at most two segments, and thus the positive result in this section will apply.
+This demonstrates the existence of policy classes which are learnable in our setting but not learnable in the standard online learning model, meaning that the two settings do not exactly coincide.
+We prove the following result:
+Theorem D.2.
+[…]
+Choosing $g(T)=T^{c}$ for $c\in(1/2,1)$ is sufficient to subconstant regret and sublinear queries:
+Theorem D.3.
+For any $c\in(1/2,1)$, Algorithm 4 with $g(T)=T^{c}$ makes $O(T^{c}(\operatorname*{diam}(\mathbf{s})+1))$ queries and satisfies
+$\lim_{T\to\infty}\ \sup_{\mathbf{s}\in S^{T}}\ \sup_{\mu}\ \sup_{\pi^{m}:f(\pi^{m})\leq K}\ R_{T}=0$
+Our algorithm does not need to know $L$ or the number of segments; it only needs to know $T$.
+D.1 Intuition behind the algorithm
+The algorithm maintains a set of buckets which partition the observed portion of the state space.
+Each bucket’s length determines the maximum loss in payoff we will allow from that subset of the state space.
+As long as the bucket contains a query from a prior time step, local generalization allows us to bound $\mu^{m}(s_{t})-\mu(s_{t},a_{t})$ based on the length of the bucket containing $s_{t}$.
+We always query if the bucket does not contain a prior query
+The granularity of the buckets is controlled by a function $g$, with the initial buckets having length $1/g(T)$.
+Since we can expect one query per bucket, we need $g(T)\in o(T)$ to ensure sublinear queries.
+Regardless of the bucket length, the adversary can still place multiple segments in the same bucket $B$.
+A single query only tells us the optimal action for one of those segments, so we risk a payoff as bad as $\mu^{m}(s_{t})-O(\operatorname*{len}(B))$ whenever we choose not to query.
+We can endure a limited number of such payoffs, but if we never query again in that bucket, we may suffer $\Theta(T)$ such payoffs.
+Letting $\mu^{m}(s_{t})=1$ for simplicity, that would lead to $\prod_{t=1}^{T}\mu(s_{t},a_{t})\leq\big{(}1-\frac{1}{O(g(T))}\big{)}^{\Theta(T)}$, which converges to 0 (i.e., guaranteed catastrophe) when $g(T)\in o(T)$.
+This failure mode suggests a natural countermeasure: if we start to suffer significant (potential) losses in the same bucket, then we should probably query there again.
+One way to structure these supplementary queries is by splitting the bucket in half when enough time steps have involved that bucket.
+It turns out that splitting after $T/g(T)$ time steps is a sweet spot.
+1:function AvoidCatastrophe($T\in\mathbb{N},\>g:\mathbb{N}\to\mathbb{N}$)
+2:     $S_{Q}\leftarrow\emptyset$triangleright Previously queried states
+3:     $\pi\leftarrow\emptyset$triangleright Records $\pi^{m}(s)$ for each $s\in S_{Q}$
+4:     $\mathcal{B}\leftarrow\emptyset$triangleright The set of active buckets
+5:     for $t$ from $1$ to $T$ do
+6:         EvaluateState$(s_{t})$
+7:function EvaluateState($s\in S$)
+8:     if $s\not\in B$ for all $B\in\mathcal{B}$ then $\triangleright$ No bucket containing $s$: create a new bucket and try again
+9:         $B\leftarrow\left[\frac{j-1}{g(T)},\frac{j}{g(T)}\right]$ for $j\in\mathbb{Z}$ such that $s\in B$
+10:         $\mathcal{B}\leftarrow\mathcal{B}\cup\{B\}$
+11:         $n_{B}\leftarrow 0$triangleright Number of time steps that have used $B$
+12:         EvaluateState$(s)$
+13:     else
+14:         $B\leftarrow$ any bucket containing $s$
+15:         if $S_{Q}\cap B=\emptyset$ then $\triangleright$ No queries in this bucket
+16:              Query mentor and observe $\pi^{m}(s)$
+17:              $\pi(s)\leftarrow\pi^{m}(s)$
+18:              $S_{Q}\leftarrow S_{Q}\cup\{s\}$
+19:              $n_{B}\leftarrow n_{B}+1$
+20:         else if $n_{B}<T/g(T)$ then $\triangleright$ Bucket has a query and isn’t full: take that action
+21:              Let $s^{\prime}\in S_{Q}\cap B$
+22:              Take action $\pi(s^{\prime})$
+23:              $n_{B}\leftarrow n_{B}+1$
+24:         $\triangleright$ Bucket is full: split bucket and try again
+25:              $B=[x,y]$
+26:              $(B_{1},B_{2})\leftarrow\Big{(}\left[x,\frac{x+y}{2}\right],\left[\frac{x+y}{2},y\right]\Big{)}$
+27:              $(s_{B_{1}},s_{B_{2}})\leftarrow(0,0)$
+28:              $\mathcal{B}\leftarrow\mathcal{B}\cup\{B_{1},B_{2}\}\setminus B$
+29:              EvaluateState$(s)$
+Algorithm 4 achieves subconstant regret when the mentor’s policy has a bounded number of segments.
+D.2 Notation for the proof
+We will use the following notation throughout the proof of Theorem D.2:
+  • Let $V_{T}=\{t\in[T]:\mu(s_{t},a_{t})<\mu^{m}(s_{t})\}$ be the set of time steps with a suboptimal payoff.
   • Let $B_{t}$ be the bucket that is used on time step $t$ (as defined on line 14 of Algorithm 4).
@@
   • Let $\mathcal{B}_{V}^{\prime}=\{B\in\mathcal{B}_{V}:\text{no descendant of $B$ is in $\mathcal{B}_{V}$}\}$.
-C.2 Proof roadmap
+D.3 Proof roadmap
 The proof proceeds in the following steps:
   1. 1.
-    Bound the total number of buckets and therefore the total number of queries (Lemma C.1).
+    Bound the total number of buckets and therefore the total number of queries (Lemma D.1).
   2. 2.
-    Bound the suboptimality on a single time step based on the bucket length and the Lipschitz constant (Lemma C.2).
+    Bound the suboptimality on a single time step based on the bucket length and $L$ (Lemma D.2).
   3. 3.
-    Bound the sum of bucket lengths on time steps where we make a mistake (Lemma C.4), with Lemma C.3 as an intermediate step. This captures the total amount of suboptimality.
+    Bound the sum of bucket lengths on time steps where we make a mistake (Lemma D.4), with Lemma D.3 as an intermediate step. This captures the total amount of suboptimality.
   4. 4.
-    The bound from Lemma C.4 is in additive form, whereas the objective is multiplicative.
-Lemma B.7 factors the objective into one term for the optimal payoffs and one term for the suboptimality in the form of Lemma C.4.
+    As in the proof of Theorem 5.2, Lemma B.3 transforms the multiplicative objective into an additive form. Lemma D.5 bounds the additive objective using Lemmas D.2 and D.4.
   5. 5.
-    Combine Lemmas C.2, C.4 and B.7 to bound the total regret (Lemma C.5).
+    Combining Lemmas D.5 and B.3 bounds the regret (Lemma D.6).
   6. 6.
-    Prove Theorem 5.2 using Lemma C.1, Lemma C.5, and some light arithmetic.
-C.3 Proof
-Lemma C.1.
-Algorithm 2 performs at most $(4+\operatorname*{len}(\operatorname*{conv}(\mathbf{s})))g(T)$ queries.
+    Theorem D.2 directly follows from Lemmas D.1 and D.6.
+D.4 Proof
+Lemma D.1.
+Algorithm 4 performs at most $(\operatorname*{diam}(\mathbf{s})+4)g(T)$ queries.
 Proof.
@@
 There are two ways to create a bucket: from scratch (line 9), or by splitting an existing bucket (line 26).
-Since depth 0 buckets overlap only at their boundaries, and each depth 0 bucket has length $1/g(T)$, at most $g(T)\max_{t,t^{\prime}\in[T]}|s_{t}-s_{t^{\prime}}|=g(T)\operatorname*{len}(\operatorname*{conv}(\mathbf{s}))$ depth 0 buckets are subsets of the interval $[\min_{t\in[T]}s_{t},\max_{t\in[T]}s_{t}]$.
-At most two depth 0 buckets are not subsets of that interval (one at each end), so the total number of depth 0 buckets is at most $g(T)\operatorname*{len}(\operatorname*{conv}(\mathbf{s}))+2$.
+Since depth 0 buckets overlap only at their boundaries, and each depth 0 bucket has length $1/g(T)$, at most $g(T)\max_{t,t^{\prime}\in[T]}|s_{t}-s_{t^{\prime}}|=g(T)\operatorname*{diam}(\mathbf{s})$ depth 0 buckets are subsets of the interval $[\min_{t\in[T]}s_{t},\max_{t\in[T]}s_{t}]$.
+At most two depth 0 buckets are not subsets of that interval (one at each end), so the total number of depth 0 buckets is at most $g(T)\operatorname*{diam}(\mathbf{s})+2$.
 We split a bucket $B$ when $n_{B}$ reaches $T/g(T)$, which creates two new buckets.
@@
 $\frac{2T}{T/g(T)}=2g(T)$
-Therefore the total number of buckets ever in existence is $(2+\operatorname*{len}(\operatorname*{conv}(\mathbf{s})))g(T)+2\leq(4+\operatorname*{len}(\operatorname*{conv}(\mathbf{s})))g(T)$, so Algorithm 2 performs at most $(4+\operatorname*{len}(\operatorname*{conv}(\mathbf{s})))g(T)$ queries.
+Therefore the total number of buckets ever in existence is $(\operatorname*{diam}(\mathbf{s})+2)g(T)+2\leq(\operatorname*{diam}(\mathbf{s})+4)g(T)$, so Algorithm 4 performs at most $(\operatorname*{diam}(\mathbf{s})+4)g(T)$ queries.
 ∎
-Lemma C.2.
-For each $t\in[T]$, $\mu(s_{t},a_{t})\geq\mu^{*}(s_{t})-2L\operatorname*{len}(B_{t})$.
+Lemma D.2.
+For each $t\in[T]$, $\mu(s_{t},a_{t})\geq\mu^{m}(s_{t})-L\operatorname*{len}(B_{t})$.
 Proof.
-If we query the mentor at time $t$, $\mu(s_{t},a_{t})=\mu^{*}(s_{t})$.
-Thus assume we do not query the mentor at time $t$: then there exists $s^{\prime}\in B_{t}$ (as defined on line 21 of Algorithm 2) such that $a_{t}=\pi(s^{\prime})=\pi^{*}(s^{\prime})$.
+If we query the mentor at time $t$, $\mu(s_{t},a_{t})=\mu^{m}(s_{t})$.
+Thus assume we do not query the mentor at time $t$: then there exists $s^{\prime}\in B_{t}$ (as defined on line 21 of Algorithm 4) such that $a_{t}=\pi(s^{\prime})=\pi^{m}(s^{\prime})$.
 Since $s_{t}$ and $s^{\prime}$ are both in $B_{t}$, $|s_{t}-s^{\prime}|\leq\operatorname*{len}(B_{t})$.
-Applying Lemma B.5 completes the proof.
-∎
-Lemma C.3.
-If $\pi^{*}$ has at most $K$ segments, $|\mathcal{B}_{V}^{\prime}|\leq K$.
+Then by the local generalization assumption,
+$\mu(s_{t},a_{t})=\mu(s_{t},\pi^{m}(s^{\prime}))\geq\mu^{m}(s_{t})-L||s_{t}-s^{\prime}||\geq\mu^{m}(s_{t})-L\operatorname*{len}(B_{t})$
+as required. ∎
+Lemma D.3.
+If $\pi^{m}$ has at most $K$ segments, $|\mathcal{B}_{V}^{\prime}|\leq K$.
 Proof.
@@
 By definition, there exists $t\in V_{T}$ such that $s_{t}\in B$.
-Then there exists $s^{\prime}\in B$ (as defined in Algorithm 2) such that $a_{t}=\pi(s^{\prime})=\pi^{*}(s^{\prime})$.
-Since $t\in V_{T}$, we have $\pi^{*}(s_{t})\neq a_{t}=\pi^{*}(s^{\prime})$.
+Then there exists $s^{\prime}\in B$ (as defined in Algorithm 4) such that $a_{t}=\pi(s^{\prime})=\pi^{m}(s^{\prime})$.
+Since $t\in V_{T}$, we have $\pi^{m}(s_{t})\neq a_{t}=\pi^{m}(s^{\prime})$.
 Thus $s_{t}$ and $s^{\prime}$ are in different segments; let $S_{B}$ be the segment containing $\min(s_{t},s^{\prime})$.
@@
 Then $\max(S_{B})<\max(B)\leq\max(S_{B^{\prime}})$, so $S_{B}\neq S_{B^{\prime}}$.
-Thus the number of segments in $\pi^{*}$ is at least $\mathcal{B}_{V}^{\prime}$.
-Since the number of segments in $\pi^{*}$ is at most $K$, we must have $|\mathcal{B}_{V}^{\prime}|\leq K$.
+Thus the number of segments in $\pi^{m}$ is at least $\mathcal{B}_{V}^{\prime}$.
+Since the number of segments in $\pi^{m}$ is at most $K$, we must have $|\mathcal{B}_{V}^{\prime}|\leq K$.
 ∎
-Lemma C.4.
+Lemma D.4.
 […]
@@
 $=$ $\ \frac{T}{g(T)^{2}}\sum_{d=0}^{d(B^{\prime})}\frac{1}{2^{d}}\leq\frac{T}{g(T)^{2}}\sum_{d=0}^{\infty}\frac{1}{2^{d}}=\frac{2T}{g(T)^{2}}$
-Then by Lemma C.3,
+Then by Lemma D.3,
 $\sum_{t\in V_{T}}\operatorname*{len}(B_{t})\leq\sum_{B^{\prime}\in\mathcal{B}_{V}^{\prime}}\frac{2T}{g(T)^{2}}=\frac{2T|\mathcal{B}_{V}^{\prime}|}{g(T)^{2}}\leq\frac{2KT}{g(T)^{2}}$
 as claimed. ∎
-Lemma C.5.
-For any $\mathbf{s}\in S^{T}$, any Lipschitz $\mu$, any $\pi^{*}$ with at most $K$ segments, and any $g:\mathbb{N}\to\mathbb{N}$, Algorithm 2 satisfies
-$R_{T}(\mathbf{s},\mu,\pi^{*})\leq\frac{6L}{g(T)}\max\left(1,\frac{TK}{g(T)}\right)$
+Lemma D.5.
+Under the conditions of Theorem D.2, Algorithm 4 satisfies
+$\sum_{t=1}^{T}\Big{(}\mu^{m}(s_{t})-\min(\mu^{m}(s_{t}),\mu(s_{t},a_{t}))\Big{)}\leq\frac{2LKT}{g(T)^{2}}$
 Proof.
-Let $\mu_{0}=4L/g(T)$. Suppose there exists $t\in[T]$ such that $\mu^{*}(s_{t})<\mu_{0}$. Then
-$R_{T}(\mathbf{s},\mu,\pi^{*})=\prod_{t=1}^{T}\mu^{*}(s_{t})-\prod_{t=1}^{T}\mu(s_{t},a_{t})\leq\mu_{0}-0\leq\frac{6L}{g(T)}$
-so the claim is satisfied. Thus assume $\mu^{*}(s_{t})\geq\mu_{0}$ for all $t\in[T]$. Let $c_{t}=0$ for $t\not\in V_{T}$ and $c_{t}=2L\operatorname*{len}(B_{t})$ for $t\in V_{T}$. Then by Lemma C.2,
-$\prod_{t=1}^{T}\mu(s_{t},a_{t})=$ $\ \prod_{t\in V_{T}}\mu(s_{t},a_{t})\prod_{t\not\in V_{T}}\mu(s_{t},a_{t})$
-$\geq$ $\ \prod_{t\in V_{T}}\left(\mu^{*}(s_{t})-2L\operatorname*{len}(B_{t})\right)\prod_{t\not\in V_{T}}\mu^{*}(s_{t})$
-$=$ $\ \prod_{t=1}^{T}\left(\mu^{*}(s_{t})-c_{t}\right)$
-Note that $\operatorname*{len}(B_{t})\leq 1/g(T)$ for all $t\in[T]$, so $c_{t}\in[0,\mu_{0}/2]$. Thus by Lemma B.7,
-$\prod_{t=1}^{T}\mu(s_{t},a_{t})\geq$ $\ \prod_{t=1}^{T}\mu(s_{t},a_{t})-1.5\sum_{t=1}^{T}c_{t}$
-$=$ $\ \prod_{t=1}^{T}\mu(s_{t},a_{t})-3L\sum_{t\in V_{T}}\operatorname*{len}(B_{t})$
-Then by Lemma C.4,
-$\prod_{t=1}^{T}\mu(s_{t},a_{t})\geq$ $\ \prod_{t=1}^{T}\mu^{*}(s_{t})-\frac{6LKT}{g(T)^{2}}$
-and so
-$R_{T}(\mathbf{s},\mu,\pi^{*})=\prod_{t=1}^{T}\mu^{*}(s_{t})-\prod_{t=1}^{T}\mu(s_{t},a_{t})\leq\frac{6LKT}{g(T)^{2}}$
+For $t\not\in V_{T}$ we have $\min(\mu^{m}(s_{t}),\mu(s_{t},a_{t}))=\mu^{m}(s_{t})$ by definition, and Lemma D.2 implies that $\min(\mu^{m}(s_{t}),\mu(s_{t},a_{t}))\geq L\operatorname*{len}(B_{t})$ for all $t\in[T]$.
+Thus
+$\sum_{t=1}^{T}\Big{(}\mu^{m}(s_{t})-\min(\mu^{m}(s_{t}),\mu(s_{t},a_{t}))\Big{)}\leq\sum_{t\in V_{T}}\Big{(}\mu^{m}(s_{t})-\min(\mu^{m}(s_{t}),\mu(s_{t},a_{t}))\Big{)}\leq L\sum_{t\in V_{T}}\operatorname*{len}(B_{t})$
+Then by Lemma D.4,
+$\sum_{t=1}^{T}\Big{(}\mu^{m}(s_{t})-\min(\mu^{m}(s_{t}),\mu(s_{t},a_{t}))\Big{)}\leq\frac{2LKT}{g(T)^{2}}$
 as required. ∎
-See 5.2
+Lemma D.6.
+Under the conditions of Theorem D.2, Algorithm 4 satisfies $R_{T}\leq\frac{2LKT}{g(T)^{2}}$.
 Proof.
-Lemma C.1 implies the query bound. Since $KT\geq T\geq g(T)$, Lemma C.5 implies that for any such $\mathbf{s},\mu$, and $\pi^{*}$,
-$R_{T}(\mathbf{s},\mu,\pi^{*})\leq$ $\ \frac{6L}{g(T)}\max\left(1,\frac{TK}{g(T)}\right)=\frac{6L}{g(T)}\cdot\frac{TK}{g(T)}$
-$=$ $\ \frac{6LTK}{T^{2c}}=6LKT^{1-2c}$
+Let $x_{t}=\mu^{m}(s_{t})$ and $y_{t}=\min(\mu^{m}(s_{t}),\mu(s_{t},a_{t}))$ for all $t\in[T]$. Then by Lemma B.3,
+$\prod_{t=1}^{T}\mu^{m}(s_{t})-\prod_{t=1}^{T}\min(\mu^{m}(s_{t}),\mu(s_{t},a_{t}))\leq$ $\ \sum_{t=1}^{T}\Big{(}\mu^{m}(s_{t})-\min(\mu^{m}(s_{t}),\mu(s_{t},a_{t}))\Big{)}$
+Since $\mu(s_{t},a_{t})\geq\min(\mu^{m}(s_{t}),\mu(s_{t},a_{t}))$ for all $t\in[T]$, we have
+$R_{T}=\prod_{t=1}^{T}\mu^{m}(s_{t})-\prod_{t=1}^{T}\mu(s_{t},a_{t})\leq\sum_{t=1}^{T}\Big{(}\mu^{m}(s_{t})-\min(\mu^{m}(s_{t}),\mu(s_{t},a_{t}))\Big{)}$
+Applying Lemma D.5 completes the proof. ∎
+Theorem D.2 follows from Lemma D.1 and Lemma D.6.
+Appendix E Other proofs
+Proposition E.1.
+Suppose $\mu$ is $L$-Lipschitz continuous and $\pi^{m}$ is optimal, i.e., $\mu(s,\pi^{m}(s))=\max_{a\in A}\mu(s,a)$ for all $s\in S$. Then $\pi^{m}$ satisfies local generalization with constant $2L$.
+Proof.
+For any $s,s^{\prime}\in S$, we have
+$\mu(s,\pi^{m}(s^{\prime}))\geq$ $\ \mu(s^{\prime},\pi^{m}(s^{\prime}))-L||s-s^{\prime}||$ (Lipschitz continuity of $\mu$)
+$\geq$ $\ \mu(s^{\prime},\pi^{m}(s))-L||s-s^{\prime}||$ ($\pi^{m}$ is optimal for $s^{\prime}$)
+$\geq$ $\ \mu(s,\pi^{m}(s))-2L||s-s^{\prime}||$ (Lipschitz continuity of $\mu$ again)
+$=$ $\ \mu^{m}(s)-2L||s-s^{\prime}||$ (Definition of $\mu^{m}(s)$)
+Since $\pi^{m}$ is optimal for $s$, we have
+$\mu^{m}(s)+2L||s-s^{\prime}||\geq\mu^{m}(s)\geq\mu(s,\pi^{m}(s^{\prime}))$
+Thus $-2L||s-s^{\prime}||\leq\mu(s,\pi^{m}(s^{\prime}))-\mu^{m}(s)\leq 2L||s-s^{\prime}||$.
+This is equivalent to $|\mu(s,\pi^{m}(s^{\prime}))-\mu^{m}(s)|\leq 2L||s-s^{\prime}||$, completing the proof.

### v4

@@
 Abstract
-Most learning algorithms with formal regret guarantees assume that no mistake is irreparable and essentially rely on trying all possible behaviors.
+Most learning algorithms with formal regret guarantees assume that all mistakes are recoverable and essentially rely on trying all possible behaviors.
 This approach is problematic when some mistakes are catastrophic, i.e., irreparable.
 We propose an online learning problem where the goal is to minimize the chance of catastrophe.
-Specifically, we assume that the payoff in each round represents the chance of avoiding catastrophe that round and aim to maximize the product of payoffs (the overall chance of avoiding catastrophe) while allowing a limited number of queries to a mentor.
+Specifically, we assume that the payoff in each round represents the chance of avoiding catastrophe that round and try to maximize the product of payoffs (the overall chance of avoiding catastrophe) while allowing a limited number of queries to a mentor.
 We first show that in general, any algorithm either constantly queries the mentor or is nearly guaranteed to cause catastrophe.
-However, in settings where the mentor policy class is learnable in the standard online learning model, we provide an algorithm whose regret and rate of querying the mentor both approach 0 as the time horizon grows.
+However, in settings where the mentor policy class is learnable in the standard online model, we provide an algorithm whose regret and rate of querying the mentor both approach 0 as the time horizon grows.
 Conceptually, if a policy class is learnable in the absence of catastrophic risk, it is learnable in the presence of catastrophic risk if the agent can ask for help.
 1 Introduction
-There has been mounting concern over catastrophic risk from AI, including but not limited to autonomous weapon accidents [1], bioterrorism [30], and cyberattacks on critical infrastructure [14].
-See [10] and [20] for taxonomies of societal-scale AI risks.
+There has been mounting concern over catastrophic risk from AI, including but not limited to autonomous weapon accidents (Abaimov & Martellini, 2020), bioterrorism (Mouton et al., 2024), and cyberattacks on critical infrastructure (Guembe et al., 2022).
+See Critch & Russell (2023) and Hendrycks et al.
+(2023) for taxonomies of societal-scale AI risks.
 In this paper, we use “catastrophe” to refer to any kind of irreparable harm.
-This definition also covers smaller-scale (yet still unacceptable) incidents such as serious medical errors [11], crashing a robotic vehicle [24], or discriminatory sentencing [39].
+This definition also covers smaller-scale (yet still unacceptable) incidents such as serious medical errors (Rajpurkar et al., 2022), crashing a robotic vehicle (Kohli & Chadha, 2020), and discriminatory sentencing (Villasenor & Foggo, 2020).
 The gravity of these risks contrasts starkly with the dearth of theoretical understanding of how to avoid them.
@@
 In the standard online learning model, the agent’s actions have no permanent effect on the environment.
-[^1] More precisely, the state can depend on the agent’s previous actions, but the agent’s performance is always evaluated with respect to the optimal policy on the same sequence of states.
-Online reinforcement learning allows the agent’s actions to permanently affect the environment, but typically assumes that either no action has irreversible effects (e.g., [21]) or that the agent is reset at the start of each “episode” (e.g., [2]).
+[^1] More precisely, the input can depend on the agent’s previous actions, but the agent’s performance is always evaluated relative to the optimal policy’s performance on the same sequence of inputs.
+Online reinforcement learning allows the agent’s actions to permanently affect the environment, but typically assumes that either no action has irreversible effects (e.g., Jaksch et al.
+(2010)) or that the agent is reset at the start of each “episode” (e.g., Azar et al.
+(2017)).
 One could train an agent entirely in a controlled lab setting where the above assumptions do hold, but we argue that sufficiently general agents will inevitably encounter novel scenarios when deployed in the real world.
-Machine learning models often behave unpredictably in unfamiliar environments (see, e.g., [32]), and we do not want AI biologists or robotic vehicles to behave unpredictably.
+Machine learning models often behave unpredictably in unfamiliar environments (see, e.g., Quionero-Candela et al.
+(2009)), and we do not want AI biologists or robotic vehicles to behave unpredictably.
 The goal of this paper is to understand the conditions under which it is possible to formally guarantee avoidance of catastrophe in online learning.
@@
 In the real world, however, one needn’t learn through pure trial-and-error: one can also ask for help.
+We think it is critical for high-stakes AI applications to employ a designated supervisor who can be asked for help.
+Examples include a human doctor supervising AI doctors, a robotic vehicle with a human driver who can take over in emergencies, autonomous weapons with a human operator, and many more.
+We hope that our work constitutes a small step in the direction of practical safety guarantees for such applications.
 1.1 Our model
 We propose an online learning model of avoiding catastrophe with mentor help.
-On each time step, the agent observes a state, selects an action (or queries the mentor), and obtains a payoff.
-[^2] Depending on the field, states are also known as inputs, instances, or contexts, actions are also known as outputs, labels, or arms, and payoffs are also known as rewards or losses.
-Each payoff represents the probability of avoiding catastrophe on that time step, and the agent’s goal is to maximize the product of payoffs, which corresponds to minimizing the total chance of catastrophe.
-We consider the product of payoffs obtained while learning, not the product of payoffs of some final policy: finding a catastrophe-avoiding policy is futile if catastrophe occurs while learning.
-The (possibly suboptimal) mentor has a fixed policy, and when queried, the mentor illustrates their policy’s action in the current state.
+On each time step, the agent observes an input, selects an action (or queries the mentor), and obtains a payoff.
+Each payoff represents the probability of avoiding catastrophe on that time step (conditioned on no prior catastrophe).
+The agent’s goal is to maximize the product of payoffs, which is equal to the overall probability of avoiding catastrophe.
+[^2] Conditioning on no prior catastrophe means we do not need to assume that these probabilities are independent (and if catastrophe has already occurred, this time step does not matter).
+This is due to the chain rule of probability.
+As is standard in online learning, we consider the product of payoffs obtained while learning, not the product of payoffs of some final policy.
+The (possibly suboptimal) mentor has a fixed policy, and when queried, the mentor illustrates their policy’s action for the current input.
 We desire an agent whose regret – defined as the gap between the mentor’s performance and the agent’s performance – approaches zero as the time horizon $T$ grows.
@@
 1.2 Our assumptions
-The agent needs some way to make inferences about unseen states in order to decide when to ask for help.
+The agent needs some way to make inferences about unqueried inputs in order to decide when to ask for help.
 Much past work has used Bayesian inference, which suffers tractability issues in complex environments.
-[^3] For the curious reader, [4] provides a thorough treatment.
+[^3] For the curious reader, Betancourt (2018) provides a thorough treatment.
 See also Section 2.
-In this paper, we instead assume that the mentor policy satisfies a Lipschitz-flavored condition we call “local generalization”: informally, if the mentor told us that action $a$ was safe in a nearby state, then $a$ is probably (but not certainly) also safe in the current state.
-Local generalization captures the intuition that one can transfer knowledge between similar states but the generalization becomes progressively less reliable as the states become less similar.
-Unlike Bayesian inference, local generalization only requires computing distances and is compatible with any state space which admits a distance metric.
+We instead assume that the mentor policy satisfies what we call local generalization: informally, if the mentor told us that an action was safe for a similar input, then that action is probably also safe for the current input (see Section 3 for a formal definition and further discussion).
+This captures the intuition that one can transfer knowledge between similar situations.
+Unlike Bayesian inference, local generalization only requires computing distances and is compatible with any input space which admits a distance metric.
 Unlike the standard online learning model, we assume that the agent does not observe payoffs.
@@
 [^4] One may be able to detect “close calls” in some cases, but observing the precise probability seems unrealistic.
+One consequence of this is that the payoff function does not affect the behavior of the algorithm, so our positive results hold simultaneously for all payoff functions which satisfy local generalization.
 1.3 Standard online learning
 An overview of standard online learning is in order before discussing our results.
-In the standard model, the agent observes a state on each time step and must choose an action.
-An adversary then reveals the correct action, which results in some payoff for the agent.
-The goal is sublinear regret, i.e., the average regret per time step should go to 0 as $T\to\infty$.
+In the standard model, the agent observes an input on each time step and must choose an action.
+An adversary then reveals the correct action, which results in some payoff to the agent.
+The goal is sublinear regret with respect to the sum of payoffs, or equivalently, the average regret per time step should go to 0 as $T\to\infty$.
 Figure 1 delineates the precise differences between the standard model and our model.
 If the adversary’s choices are unconstrained, the problem is hopeless: if the adversary determines the correct action on each time step randomly and independently, the agent can do no better than random guessing.
-However, sublinear regret becomes possible if (1) the hypothesis class has finite Littlestone dimension [26], or (2) the hypothesis class has finite VC dimension [38] and the input is $\sigma$-smooth [^5] Informally, the adversarial chooses a distribution over states instead of a precise state.
+However, sublinear regret becomes possible if (1) the hypothesis class has finite Littlestone dimension (Littlestone, 1988), or (2) the hypothesis class has finite VC dimension (Vapnik & Chervonenkis, 1971) and the input is $\sigma$-smooth [^5] Informally, the adversary chooses a distribution over inputs instead of a precise input.
 See Section 3.
-[17].
+(Haghtalab et al., 2024).
 The goal of sublinear regret in online learning implicitly assumes catastrophe is impossible: the agent can make arbitrarily many (and arbitrarily costly) mistakes as long as the average regret per time step goes to 0.
@@
 However, access to a mentor (and local generalization) allows our agent to learn without trying actions directly, which is enough to offset all of the above disadvantages.
-Objective Regret goal Feedback Mentor Local gen.
+Objective Regret goal Feedback Mentor Local generalization
 Standard model Sum of payoffs Sublinear Every time step No No
@@
 Figure 1: Comparison between the standard online learning model and our model.
-On a technical level, sublinear vs subconstant regret is a more important distinction than the sum vs product of payoffs: for example, taking a logarithm can transform the product into a sum.
-However, the multiplicative objective and the goal of subconstant regret are tightly related, since the former’s interpretation as the chance of catastrophe is what gives rise to the latter: we want the total (not average) chance of catastrophe to go to 0.
 1.4 Our results
-Informally, we show that avoiding catastrophe with the help of a mentor and local generalization is no harder than online learning without catastrophic risk.
+At a high level, we show that avoiding catastrophe with the help of a mentor and local generalization is no harder than online learning without catastrophic risk.
 More precisely, we first show that in general, any algorithm with sublinear queries to the mentor has arbitrarily poor regret in the worst-case (Theorem 4.1).
-As a corollary, even when it is possible to avoid catastrophe with certainty, any algorithm either asks for help on a constant fraction of time steps or is nearly guaranteed to cause catastrophe.
-Unlike online learning where the general impossibility result is trivial (the agent might as well guess randomly given an unconstrained adversary), local generalization significantly limits the adversary’s power and necessitates a careful construction.
-Next, we present a simple algorithm whose total regret and rate of querying the mentor both go to 0 as $T\to\infty$ when either (1) the mentor policy class has finite Littlestone dimension or (2) the mentor policy class has finite VC dimension and the state sequence is $\sigma$-smooth.
-Our algorithm can handle a multi-dimensional unbounded state space and does not need detailed access to the feature embedding, instead using two simple operations.
+This means that even when the mentor can avoid catastrophe with certainty, any algorithm either needs excessive supervision or is nearly guaranteed to cause catastrophe.
+Unlike online learning where the general impossibility result is trivial (the agent might as well guess randomly given an unconstrained adversary), local generalization significantly limits the adversary’s power and necessitates a careful analysis.
+Next, we present a simple algorithm whose total regret and rate of querying the mentor both go to 0 as $T\to\infty$ when either (1) the mentor policy class has finite Littlestone dimension or (2) the mentor policy class has finite VC dimension and the input sequence is $\sigma$-smooth.
+Our algorithm can handle a unbounded input space and does not need detailed access to the feature embedding, instead relying on two simple operations.
 It does need to know the mentor policy class, as is standard in online learning.
-We initially prove the theorem for binary actions (Theorem 5.2) and then reduce learning with (finitely) many actions to the binary action case (Theorem C.1).
-Finally, Appendix D presents another algorithm with the same guarantees of subconstant regret and sublinear queries under different assumptions.
-Most importantly, this result applies to the class of thresholds on $[0,1]$, which is known to be hard for standard online learning (see Section 2).
-This shows that our problem and standard online learning do not have exactly the same difficulty, but rather our problem is no harder than online learning.
-The rest of the paper is structured as follows.
-Section 2 discusses related work.
-Section 3 formally defines our model.
-Section 4 presents our negative result for general mentor policies.
-Section 5 presents our positive result for simple mentor policy classes.
-Proofs are deferred to the appendix.
+We initially prove the theorem for binary actions (Theorem 5.2) and then reduce learning with many actions to the binary action case (Theorem C.1).
+Along the way, we prove that the same subconstant bound holds for standard additive regret (Theorem 5.3).
+Essentially, our techniques are equally effective for maximizing the sum of payoffs and the product of payoffs.
+We emphasize the multiplicative objective due to our motivation of avoiding catastrophe, but our subconstant additive regret bound may also be of interest.
+In summary, the combination of a mentor and local generalization allows us to reduce the regret by an entire factor of $T$, resulting in subconstant regret (multiplicative or additive) instead of the typical sublinear regret.
 2 Related work
@@
 Despite the ubiquity of irreparable/irreversible costs in the real world, theoretical work on this topic remains limited.
-This may be due to the fundamental modeling question of how the agent should learn about novel states or actions without actually trying them.
+This may be due to the fundamental modeling question of how the agent should learn about novel inputs or actions without actually trying them.
 The most common approach is to allow the agent to ask for help.
@@
 A popular solution is to perform Bayesian inference on the world model, but this has two tricky requirements: (1) a prior distribution which contains the true world model (or an approximation), and (2) an environment where computing (or approximating) the posterior is tractable.
-A finite set of possible environments satisfies both conditions, but is unrealistic in many real-world scenarios.
-In contrast, our algorithm can handle an uncountable policy class and a continuous unbounded state space, which is crucial for many real-world scenarios in which one never sees the exact same state twice.
-Bayesian inference combined with asking for help is studied by [8; 9; 25; 28].
-We also mention [15; 29; 37], which utilize Bayesian inference in the context of safe (online) reinforcement learning without asking for help (and without regret bounds).
-We are only aware of two papers which theoretically address irreversibility without Bayesian inference: [13] and [27].
+A finite set of possible environments satisfies both conditions but is unrealistic in many real-world scenarios.
+In contrast, our algorithm can handle an uncountable policy class and a continuous unbounded input space, which is crucial for many real-world scenarios in which one never sees the exact same input twice.
+Bayesian inference combined with asking for help is studied by Cohen et al.
+(2021); Cohen & Hutter (2020); Kosoy (2019); Mindermann et al.
+(2018).
+It is also worth mentioning Hadfield-Menell et al.
+(2017); Moldovan & Abbeel (2012); Turchetta et al.
+(2016), who utilize Bayesian inference in the context of safe (online) reinforcement learning without asking for help (and without regret bounds).
+We are only aware of two papers which theoretically address irreversibility without Bayesian inference: Grinsztajn et al.
+(2021) and Maillard et al.
+(2019).
 The former proposes to sample trajectories and learn reversibility based on temporal consistency between states: intuitively, if $s_{1}$ always precedes $s_{2}$, we can infer that $s_{1}$ is unreachable from $s_{2}$.
@@
 To our knowledge, we are the first to provide an algorithm which formally guarantees avoidance of catastrophe (with high probability) without Bayesian inference.
+We are also not aware of prior results comparable to our negative result, including in the Bayesian regime.
+Constrained Markov Decision Processes (CMDPs).
+CMDPs (Altman, 2021) require the agent to maximize reward while also satisfying safety constraints.
+See Gu et al.
+(2023); Zhao et al.
+(2023); Wachi et al.
+(2024) for surveys.
+The two most relevant papers are Liu et al.
+(2021) and Stradi et al.
+(2024), both of which provide algorithms guaranteed to satisfy initially unknown safety constraints.
+Since neither paper allows external help, they require strong assumptions to make the problem tractable: the aforementioned results assume that the agent (1) knows a strictly safe policy upfront (i.e., a policy which satisfies the safety constraints with slack), (2) is periodically reset, and (3) observes the safety costs.
+In contrast, our agent has no prior knowledge, is never reset, and never observes payoffs.
 Online learning.
-See [7] and Chapter 21 of [34] for introductions to online learning.
-A classical result states that sublinear regret is possible iff the hypothesis class has finite Littlestone dimension [26].
-However, even some simple hypothesis classes have infinite Littlestone dimension, such as the class of thresholds on $[0,1]$ (Example 21.4 in [34]).
-Recently, [17] showed that if the adversary only chooses a distribution over states rather than the precise state, only the weaker assumption of finite VC dimension [38] is needed for sublinear regret.
-Specifically, they assume that each state is sampled from a distribution whose concentration is upper bounded by $\frac{1}{\sigma}$ times the uniform distribution.
-This framework is known as smoothed analysis.
-Smoothed analysis.
-Smoothed analysis was originally proposed by [36] to explain the fact that some algorithms almost always perform well in practice, despite the theoretical existence of pathological inputs.
-Informally, algorithms which perform well under smoothed analysis are guaranteed to perform well for nearly all – but not necessarily all – inputs.
-Active learning.
-Our assumption that the agent only receives feedback in response to queries falls under the umbrella of active learning [18].
+See Cesa-Bianchi & Lugosi (2006) and Chapter 21 of Shalev-Shwartz & Ben-David (2014) for introductions to online learning.
+A classical result states that sublinear regret is possible iff the hypothesis class has finite Littlestone dimension (Littlestone, 1988).
+However, even some simple hypothesis classes have infinite Littlestone dimension, such as the class of thresholds on $[0,1]$ (Example 21.4 in Shalev-Shwartz & Ben-David (2014)).
+Recently, Haghtalab et al.
+(2024) showed that if the adversary only chooses a distribution over inputs rather than the precise input, only finite VC dimension (Vapnik & Chervonenkis, 1971) is needed for sublinear regret.
+Specifically, they assume that each input is sampled from a distribution whose concentration is upper bounded by $\frac{1}{\sigma}$ times the uniform distribution.
+This framework is known as smoothed analysis, originally due to Spielman & Teng (2004).
+Multiplicative objectives.
+Although online learning traditionally studies the sum of payoffs, there is some work on maximizing the product of payoffs, or equivalently the sum of logarithms (Chapter 9 of Cesa-Bianchi et al.
+(2017)).
+However, these regret bounds are still sublinear in $T$, in comparison to our subconstant regret bounds.
+Also, that work still assumes that payoffs are observed on every time step, while our agent only receives feedback in response to queries (Figure 1).
+Barman et al.
+(2023) recently provided regret bounds for a multiplicative objective in a multi-armed bandit problem, but their objective is the geometric mean of payoffs instead of the product.
+Interpreted in our context, their regret bounds imply that the average chance of catastrophe goes to zero, while we guarantee that the total chance of catastrophe goes to zero.
+This distinction is closely related to the difference between subconstant and sublinear regret discussed in Section 1.3.
+Active learning and imitation learning.
+Our assumption that the agent only receives feedback in response to queries falls under the umbrella of active learning (Hanneke, 2014).
 This contrasts with passive learning, where the agent receives feedback automatically.
 Although ideas from active learning could be useful in our domain, we are not aware of any results from that literature which account for irreversible costs.
-Multiplicative objectives.
-Perhaps the best-studied example of a multiplicative objective is the Nash welfare function (also known as the Nash bargaining solution), defined as the geometric mean of agent utilities [23; 31].
-Inspired by Nash welfare, [3] recently studied algorithms for maximizing the geometric mean of payoffs across time steps in the multi-armed bandit problem.
-However, the geometric mean of payoffs is much more forgiving of errors compared to the product of payoffs.
-Imagine that the agent obtains a payoff of $1-1/\sqrt{T}$ per time step: then $\lim_{T\to\infty}(1-1/\sqrt{T})^{T}=0$ but $\lim_{T\to\infty}((1-1/\sqrt{T})^{T})^{1/T}=\lim_{T\to\infty}(1-1/\sqrt{T})=1$.
-In our setting, it is crucial that the total chance of catastrophe (the product of payoffs) is small, not just the average chance of catastrophe per time step (the geometric mean of payoffs).
-This is closely related to the difference between sublinear and subconstant regret discussed in Section 1.3.
+The process of the agent learning from a mentor is also reminiscent of imitation learning (Osa et al., 2018), but we are not aware of any relevant technical implications.
 3 Model
-States.
+Inputs.
 Let $\mathbb{N}$ refer to the set of strictly positive integers and let $T\in\mathbb{N}$ be the time horizon.
-Let $S\subseteq\mathbb{R}^{n}$ be the state space and let $\mathbf{s}=(s_{1},s_{2},\dots s_{T})\in S^{T}$ be the sequence of states.
-In the adversarial setting, each $s_{t}$ can have arbitrary dependence on the events of prior time steps.
-In the smoothed setting, the adversary only chooses the distribution $s_{t}$ from which $s_{t}$ is sampled.
-Formally, a distribution $\mathcal{D}$ over $S$ is $\sigma$-smooth if for any $X\subseteq S$, $\mathcal{D}(X)\leq\frac{1}{\sigma}U(X)$.
-(In the smoothed setting, we assume that $S$ supports a uniform distribution $U$ [^6] For example, $S$ having finite Lebesgue measure is sufficient.
+Let $\boldsymbol{x}=(x_{1},x_{2},\dots x_{T})\in\mathcal{X}^{T}$ be the sequence of inputs.
+In the adversarial setting, each $x_{t}$ can have arbitrary dependence on the events of prior time steps.
+In the smoothed setting, the adversary only chooses the distribution $x_{t}$ from which $x_{t}$ is sampled.
+Formally, a distribution $\mathcal{D}$ over $\mathcal{X}$ is $\sigma$-smooth if for any $X\subseteq\mathcal{X}$, $\mathcal{D}(X)\leq\frac{1}{\sigma}U(X)$.
+(In the smoothed setting, we assume that $\mathcal{X}$ supports a uniform distribution $U$ [^6] For example, $\mathcal{X}$ having finite Lebesgue measure is sufficient.
 Note that this does not imply boundedness.
-Alternatively, $\sigma$-smoothness can be defined with respect to a different distribution, as long as the Radon-Nikodym derivative is uniformly bounded; see Definition 1 of [5].) If each $s_{t}$ is sampled from a $\sigma$-smooth $\mathcal{D}_{t}$, we say that $\mathbf{s}$ is $\sigma$-smooth.
+Alternatively, $\sigma$-smoothness can be defined with respect to a different distribution, as long as the Radon-Nikodym derivative is uniformly bounded; see Definition 1 of Block et al.
+(2022).) If each $x_{t}$ is sampled from a $\sigma$-smooth $\mathcal{D}_{t}$, we say that $\boldsymbol{x}$ is $\sigma$-smooth.
 The sequence $\boldsymbol{\mathcal{D}}=\mathcal{D}_{1},\dots,\mathcal{D}_{T}$ can still be adaptive, i.e., the choice of $\mathcal{D}_{t}$ can depend on the events of prior time steps.
-Actions.
-Let $A$ be a finite set of actions.
-There also exists a special action $\hat{a}$ which corresponds to querying the mentor.
+Actions and payoffs.
+Let $\mathcal{Y}$ be a finite set of actions.
+There also exists a special action $\tilde{y}$ which corresponds to querying the mentor.
 For $k\in\mathbb{N}$, let $[k]=\{1,2,\dots,k\}$.
-On each time step $t\in[T]$, the agent must select an action $a_{t}\in\ A\cup\{\hat{a}\}$, which generates payoff $\mu(s_{t},a_{t})\in[0,1]$.
+On each time step $t\in[T]$, the agent must select action $y_{t}\in\ \mathcal{Y}\cup\{\tilde{y}\}$, which generates a payoff.
+Let $\boldsymbol{y}=(y_{1},\dots,y_{T})$.
+We allow the payoff function to vary between time steps: let $\boldsymbol{\mu}=(\mu_{1},\dots,\mu_{T})\in(\mathcal{X}\times\mathcal{Y}\to[0,1])^{T}$ be the sequence of payoff functions.
+[^7] Variation in $\mu_{t}$ is not a core part of our problem, but we allow it because the exact same proof holds.
+Since the mentor policy is fixed, local generalization limits how much $\mu_{t}$ can vary between time steps.
+Then $\mu_{t}(x_{t},y_{t})\in[0,1]$ is the agent’s payoff at time $t$.
+Like $\boldsymbol{\mathcal{D}}$, we allow $\boldsymbol{\mu}$ to be adaptive.
+Unless otherwise noted, all expectations are over the agent’s randomization (if any) and the randomization in $\boldsymbol{x}$ (if any).
 Asking for help.
-The mentor is endowed with a (not necessarily optimal) policy $\pi^{m}:S\to A$.
-When action $\hat{a}$ is chosen, the mentor informs the agent of the action $\pi^{m}(s_{t})$ and the agent obtains payoff $\mu(s_{t},\pi^{m}(s_{t}))$.
-For brevity, let $\mu^{m}(s)=\mu(s,\pi^{m}(s))$.
-The agent never observes payoffs: the only way to learn about $\mu$ is by querying the mentor.
-We would like an algorithm which becomes “self-sufficient” over time: the rate of querying the mentor should go to 0 as $T\to\infty$.
-Intuitively, this is equivalent to the cumulative number of queries being sublinear in $T$: $|Q_{T}|\in o(T)$, where $Q_{T}=\{t\in[T]:a_{t}=\hat{a}\}$.
-However, interpreted literally, $|Q_{T}|\in o(T)$ is not a rigorous statement.
-This is because $|Q_{T}|$ is not just a function of $T$: it can also depend on $\mathbf{s}$, $\mu$, $\pi^{m}$, and any randomness in the agent’s actions.
-To formalize this, let $Q_{T}(\mathbf{s},\mu,\pi^{m})$ be the set of time steps with queries given $\mathbf{s},\mu$, and $\pi^{m}$.
-Then we say that $|Q_{T}|$ is sublinear in $T$, denoted $|Q_{T}|\in o(T)$, if there exists $g:\mathbb{N}\to\mathbb{N}$ such that $g(T)\in o(T)$ and $|Q_{T}(\mathbf{s},\mu,\pi^{m})|\leq g(T)$ for all $\mathbf{s},\mu,\pi^{m}$ (and for all possible realizations of the agent’s randomization).
-However, we will mostly write $Q_{T}$ and $|Q_{T}|\in o(T)$ for brevity.
+The mentor is endowed with a (not necessarily optimal) policy $\pi^{m}:\mathcal{X}\to\mathcal{Y}$.
+When action $\tilde{y}$ is chosen, the mentor informs the agent of the action $\pi^{m}(x_{t})$ and the agent obtains payoff $\mu_{t}(x_{t},\pi^{m}(x_{t}))$.
+For brevity, let $\mu_{t}^{m}(x)=\mu_{t}(x,\pi^{m}(x))$.
+The agent never observes payoffs: the only way to learn about $\boldsymbol{\mu}$ is by querying the mentor.
+We would like an algorithm which becomes “self-sufficient” over time: the rate of querying the mentor should go to 0 as $T\to\infty$, or equivalently, the cumulative number of queries should be sublinear in $T$.
+Formally, let $Q_{T}(\boldsymbol{\mu},\pi^{m})=\{t\in[T]:y_{t}=\tilde{y}\}$ be the random variable denoting the set of time steps with queries.
+Then we say that the (expected) number of queries is sublinear in $T$ if $\sup_{\boldsymbol{\mu},\pi^{m}}\operatorname*{\mathbb{E}}[|Q_{T}(\boldsymbol{\mu},\pi^{m})|]\in o(T)$.
+In other words, there must exist $g:\mathbb{N}\to\mathbb{N}$ such that $g(T)\in o(T)$ and $\sup_{\boldsymbol{\mu},\pi^{m}}\operatorname*{\mathbb{E}}[|Q_{T}(\boldsymbol{\mu},\pi^{m})|]\leq g(T)$.
+[^8] One could instead consider the worst-case number of queries, but this distinction does not affect whether subconstant regret is achievable (Proposition E.2).
+For brevity, we will usually write $Q_{T}=Q_{T}(\boldsymbol{\mu},\pi^{m})$.
 Local generalization.
 We assume that the mentor policy permits “local generalization”.
-Formally, we assume there exists $L>0$ such that for all $s,s^{\prime}\in S$, $|\mu^{m}(s)-\mu(s,\pi^{m}(s^{\prime}))|\leq L||s-s^{\prime}||$, where |||||$\cdot$||| | ⋅ | | denotes Euclidean distance.
-Informally, if the agent is in state $s$, taking the mentor action from a nearby state $s^{\prime}$ is almost as good as taking the mentor action for $s$.
-In other words, it’s okay to use an action you learned from a nearby state.
-This assumption might initially feel less intuitive than a standard Lipschitz continuity assumption of $|\mu(s,a)-\mu(s^{\prime},a)|\leq L||s-s^{\prime}||$.
-However, Lipschitz continuity is strictly stronger than local generalization when the mentor is optimal (Proposition E.1).
-The properties are incomparable in general, but Lipschitz continuity allows odd situations where the mentor provides an action which is only acceptable in that specific state.
-This sets the agent up for failure: even if such actions exist, a thoughtful mentor should avoid those actions when possible.
-See Figure 2 for an illustration.
-[…]
-[…]
-[…]
-Figure 2: The difference between local generalization and Lipschitz continuity.
-The x-axis corresponds to a one-dimensional state space, with $s_{1}$ and $s_{2}$ denoting two nearby states.
-The y-axis is the payoff.
-In all figures, $\pi^{m}(s_{1})=0$ and $\pi^{m}(s_{2})=1$.
-Left: What one might consider the “typical case” with a near-optimal mentor.
-The mentor takes action 0 in $s_{1}$, so $|\mu(s_{2},0)-\mu(s_{2},\pi^{m}(s_{2})|$ is small (local generalization) and $|\mu(s_{2},0)-\mu(s_{1},0)|$ is small (Lipschitz continiuty).
-Middle: Local generalization without Lipschitz continuity.
-Action 0 is much worse in $s_{2}$, but it is still nearly as good as the mentor’s action in $s_{2}$, which is also quite bad.
-We primarily care about the agent’s performance relative to the mentor.
-Right: Lipschitz continuity without local generalization.
-The payoff of each action is similar in nearby states.
-However, despite the mentor recommending action 0 in $s_{1}$, action 0 is much worse than the mentor’s action in $s_{2}$.
-The mentor is setting up the agent for failure by recommending action 0 in $s_{1}$.
+Informally, if the agent is given an input $x$, taking the mentor action for a similar input $x^{\prime}$ is almost as good as taking the mentor action for $x$.
+Formally, we assume $\mathcal{X}\subseteq\mathbb{R}^{n}$ and there exists $L>0$ such that for all $x,x^{\prime}\in\mathcal{X}$ and $t\in[T]$, $|\mu_{t}^{m}(x)-\mu_{t}(x,\pi^{m}(x^{\prime}))|\leq L||x-x^{\prime}||$, where |||||$\cdot$||| | ⋅ | | denotes Euclidean distance.
+[^9] One could also allow a generic metric space; our assumption of $\mathcal{X}\subseteq\mathbb{R}^{n}$ is only for convenience.
+This represents the ability to transfer knowledge between similar inputs:
+$|\underbrace{\mu_{t}(x,\pi^{m}(x))}_{\text{Taking the right action}}-\underbrace{\mu_{t}(x,\pi^{m}(x^{\prime}))}_{\text{Using what you learned in $x^{\prime}$}}|\ \ \ \leq\underbrace{L||x-x^{\prime}||}_{\text{Similarity between $x$ and $x^{\prime}$}}$
+Borrowing knowledge from similar experiences seems fundamental to learning and is well-understood in the psychology literature (Esser et al., 2023) and education literature (Hajian, 2019).
+Crucially, our input space can be any feature embedding of the agent’s situation, not just its physical positioning.
+Our algorithms will not require knowledge of the feature embedding and do not need to know $L$, so it suffices that there exists some feature embedding which satisfies local generalization.
+The agent does not even need to know which embedding it is.
+Finally, local generalization implies the more familiar Lipschitz continuity for an optimal mentor (Proposition E.1).
 Multiplicative objective and regret.
-If $\mu(s_{t},a_{t})\in[0,1]$ is the chance of avoiding catastrophe on time step $t$ (conditioned on no prior catastrophe), then $\prod_{t=1}^{T}\mu(s_{t},a_{t})$ is the agent’s overall chance of avoiding catastrophe.
-[^7] Conditioning on no prior catastrophe means we do not need to assume that these probabilities are independent (and if catastrophe has already occurred, this time step does not matter).
-This is due to the chain rule of probability.
-For a fixed state sequence $\mathbf{s}$ and agent actions $\mathbf{a}=(a_{1},\dots,a_{T})$, the agent’s regret is
-$R_{T}(\mathbf{s},\mathbf{a},\mu,\pi^{m})=\prod_{t=1}^{T}\mu^{m}(s_{t})-\prod_{t=1}^{T}\mu(s_{t},a_{t})$
-We will usually write $R_{T}=R_{T}(\mathbf{s},\mathbf{a},\mu,\pi^{m})$ for brevity.
-We will study the expected regret over any randomness in $\mathbf{s}$ and/or $\mathbf{a}$.
-We desire subconstant worst-case regret: the total (not average) expected regret should go to 0 for any $\mu$ and $\pi^{m}$.
-Formally, we want $\lim_{T\to\infty}\sup_{\mu,\pi^{m}}\operatorname*{\mathbb{E}}[R_{T}]=0$.
+If $\mu_{t}(x_{t},y_{t})\in[0,1]$ is the chance of avoiding catastrophe on time step $t$ (conditioned on no prior catastrophe), then $\prod_{t=1}^{T}\mu_{t}(x_{t},y_{t})$ is the agent’s overall chance of avoiding catastrophe by the chain rule of probability.
+For given $\boldsymbol{x},\boldsymbol{y},\boldsymbol{\mu},\pi^{m}$, the agent’s regret is
+$R_{T}(\boldsymbol{x},\boldsymbol{y},\boldsymbol{\mu},\pi^{m})=\prod_{t=1}^{T}\mu_{t}^{m}(x_{t})-\prod_{t=1}^{T}\mu_{t}(x_{t},y_{t})$
+We write $R_{T}=R_{T}(\boldsymbol{x},\boldsymbol{y},\boldsymbol{\mu},\pi^{m})$ for brevity.
+We desire subconstant worst-case regret: the total (not average) expected regret should go to 0 for any $\boldsymbol{\mu}$ and $\pi^{m}$; formally, $\lim_{T\to\infty}\sup_{\boldsymbol{\mu},\pi^{m}}\operatorname*{\mathbb{E}}[R_{T}]=0$.
+The interpretation of a bound on $\operatorname*{\mathbb{E}}[R_{T}]$ depends on the quality of the mentor.
+For example, subconstant regret becomes trivial if $\lim_{T\to\infty}\operatorname*{\mathbb{E}}\big{[}\prod_{t=1}^{T}\mu_{t}^{m}(x_{t})\big{]}=0$.
+However, we think that high-stakes AI applications should ensure the presence of a mentor who is almost always safe, i.e., $\operatorname*{\mathbb{E}}\big{[}\prod_{t=1}^{T}\mu_{t}^{m}(x_{t})\big{]}\approx 1$.
+If no such mentor exists for some application, perhaps it is better to avoid the application altogether.
+Also, our regret bounds include rates of convergence, so even if the mentor policy will eventually cause catastrophe, we can still bound how quickly the agent becomes unsafe.
 VC and Littlestone dimensions.
-VC dimension [38] and Littlestone dimension [26] are standard measures of learning difficulty which capture the ability of a hypothesis class (in our case, a policy class) to realize arbitrary combinations of labels (in our case, actions).
-We omit the precise dimensions since we only utilize these concept via existing results.
-See [34] for a comprehensive overview.
+VC dimension (Vapnik & Chervonenkis, 1971) and Littlestone dimension (Littlestone, 1988) are standard measures of learning difficulty which capture the ability of a hypothesis class (in our case, a policy class) to realize arbitrary combinations of labels (in our case, actions).
+We omit the precise dimensions since we only utilize these concepts via existing results.
+See Shalev-Shwartz & Ben-David (2014) for a comprehensive overview.
 Misc.
-The diameter of $X\subseteq S$ is defined by $\operatorname*{diam}(X)=\max_{x,y\in X}||x-y||$. All logarithms and exponents are base $e$ unless otherwise noted.
+The diameter of a set $X\subseteq\mathcal{X}$ is defined by $\operatorname*{diam}(X)=\max_{x,x^{\prime}\in X}||x-x^{\prime}||$. All logarithms and exponents are base $e$ unless otherwise noted.
 4 Avoiding catastrophe is impossible in general
-We begin by showing that in general, any algorithm with sublinear mentor queries has arbitrary poor regret in the worst-case, even when states are i.i.d.
+We begin by showing that in general, any algorithm with sublinear mentor queries has arbitrary poor regret in the worst-case, even when inputs are i.i.d.
 on $[0,1]$.
-The result also holds even if the algorithm knows $L$ and $\mathbf{s}$ ahead of time.
-We use $\operatorname*{\mathbb{E}}_{\mathbf{s}\sim U^{T},\mathbf{a}}$ denote the expectation over both $\mathbf{s}$ and any additional randomness in the algorithm.
+The result also holds even if the algorithm knows $L$ and $\boldsymbol{x}$ ahead of time.
+The proof can be found in Appendix A.
 Theorem 4.1.
 The worst-case expected regret of any algorithm with sublinear queries goes to 1 as $T$ goes to infinity.
-Formally, $\lim_{T\to\infty}\ \,\sup_{\mu,\pi^{m}}\ \,\operatorname*{\mathbb{E}}_{\mathbf{s}\sim U^{T},\mathbf{a}}\ \,[R_{T}]=1$.
+Formally, $\lim_{T\to\infty}\ \,\sup_{\boldsymbol{\mu},\pi^{m}}\,\operatorname*{\mathbb{E}}[R_{T}]=1$.
 4.1 Intuition
-We partition $S$ into equally-sized sections that are “independent” in the sense that querying a state in section $i$ gives you no information about section $j$.
+We partition $\mathcal{X}$ into equally-sized sections that are “independent” in the sense that querying an input in section $i$ gives you no information about section $j$.
 The number of sections is determined by a function $f:\mathbb{N}\to\mathbb{N}$ that we will choose.
 If $|Q_{T}|\in o(f(T))$, most of these sections will never contain a query.
-When the agent sees a state in a section not containing a query, it essentially has to guess, meaning it will be wrong a constant fraction of the time.
+When the agent sees an input in a section not containing a query, it essentially has to guess, meaning it will be wrong a constant fraction of the time.
+We then choose a payoff function (which is the same for all time steps) which makes the wrong guesses as costly as possible, subject to the local generalization constraint.
 Figure 2 fleshes out this idea.
-$s$$\mu$  $\mu(s,0)$$\mu(s,1)$$\cfrac{L}{2f(T)}$$\frac{1}{f(T)}$$\frac{2}{f(T)}$$\frac{3}{f(T)}$$\frac{f(T)-1}{f(T)}$$\dots$1001111
+Picking $\boldsymbol{f(T)}$.
+A natural idea is to try $f(T)=T$, but this doesn’t quite work: even if the agent chooses wrong on every time step, the minimum payoff is still at least $1-\frac{L}{2T}$, and $\lim_{T\to\infty}\prod_{t=1}^{T}\left(1-\frac{L}{2T}\right)=\lim_{T\to\infty}\left(1-\frac{L}{2T}\right)^{T}=e^{-L/2}$.
+In order for the regret to approach 1, we need $f(T)$ to be asymptotically between $|Q_{T}|$ and $T$ (such $f$ must exist since $|Q_{T}|\leq g(T)\in o(T)$).
+This leads to the following bound: $\prod_{t=1}^{T}\mu(x_{t},y_{t})\leq\big{(}1-\frac{L}{\Theta(f(T))}\big{)}^{\Theta(T)}$.
+When $f(T)\in o(T)$, the right hand side converges to 0, while $\prod_{t=1}^{T}\mu^{m}(x_{t})=1$.
+In words, the agent is nearly guaranteed to cause catastrophe, despite the existence of a policy which is guaranteed to avoid catastrophe.
+$x$$\mu$  $\mu(x,0)$$\mu(x,1)$$\cfrac{L}{2f(T)}$$\frac{1}{f(T)}$$\frac{2}{f(T)}$$\frac{3}{f(T)}$$\frac{f(T)-1}{f(T)}$$\dots$1001111
 Figure 2: An illustration of the construction we use to prove Theorem 4.1 (not to scale).
-The horizontal axis indicates the state $s\in[0,1]$ and the vertical axis indicates the payoff $\mu(s,a)\in[0,1]$.
-The solid line represents $\mu(s,0)$ and the dotted line represents $\mu(s,1)$.
+The horizontal axis indicates the input $x\in[0,1]$ and the vertical axis indicates the payoff $\mu(x,y)\in[0,1]$.
+The solid line represents $\mu(x,0)$ and the dotted line represents $\mu(x,1)$.
 In each section, one of the actions has the optimal payoff of 1, and the other action has the worst possible payoff allowed by $L$, reaching a minimum of $1-\frac{L}{2f(T)}$.
@@
 As a result, we can freely toggle the optimal action for each section independently.
-Picking $\boldsymbol{f(T)}$.
-A natural idea is to try $f(T)=T$, but this doesn’t quite work: even if the agent chooses wrong on every time step, the minimum payoff is still at least $1-\frac{L}{2T}$, and $\lim_{T\to\infty}\prod_{t=1}^{T}\left(1-\frac{L}{2T}\right)=\lim_{T\to\infty}\left(1-\frac{L}{2T}\right)^{T}=e^{-L/2}$.
-In order for the regret to approach 1, we need $f(T)$ to be asymptotically between $|Q_{T}|$ and $T$ (such $f$ must exist since $|Q_{T}|\leq g(T)\in o(T)$).
-This leads to the following bound: $\prod_{t=1}^{T}\mu(s_{t},a_{t})\leq\big{(}1-\frac{L}{\Theta(f(T))}\big{)}^{\Theta(T)}$.
-When $f(T)\in o(T)$, the right hand side converges to 0, while $\prod_{t=1}^{T}\mu^{m}(s_{t})=1$.
-In words, the agent is nearly guaranteed to cause catastrophe, despite the existence of a policy which is guaranteed to avoid catastrophe.
 VC dimension.
@@
 4.2 Formal definition of construction
-Let $S=[0,1]$ and $\mathcal{D}_{t}=U$ for each $t\in[T]$.
+Let $\mathcal{X}=[0,1]$ and $\mathcal{D}_{t}=U$ for each $t\in[T]$.
 Assume that $L\leq 1$; this will simplify the math and only makes the problem easier for the agent.
-We define a family of payoff functions parameterized by a function $f:\mathbb{N}\to\mathbb{N}$ and a bit string $\mathbf{x}=(x_{1},x_{2},\dots,x_{f(T)})\in\{0,1\}^{f(T)}$.
-The bit $x_{j}$ will denote the optimal action in section $j$.
+We define a family of payoff functions parameterized by a function $f:\mathbb{N}\to\mathbb{N}$ and a bit string $\boldsymbol{a}=(a_{1},a_{2},\dots,a_{f(T)})\in\{0,1\}^{f(T)}$.
+The bit $a_{j}$ will denote the optimal action in section $j$.
 Note that $f(T)\geq 1$ and since we defined $\mathbb{N}$ to exclude 0.
 […]
-$\mu_{f,\mathbf{x}}(s,a)=\begin{cases}1&\textnormal{ if }a=x_{j(s)}\\ 1-L\left(\cfrac{1}{2f(T)}-|m_{j(s)}-s|\right)&\textnormal{ if }a\neq x_{j(s)}\\ \end{cases}$
-Let $\pi^{m}$ be any policy which is optimal for $\mu_{f,\mathbf{x}}$.
-Note that there is a unique optimal action for each $s_{t}$, since each $s_{t}$ belongs to exactly one $S_{j}$; formally, $\pi^{m}(s_{t})=x_{j(s_{t})}$.
-For any $\mathbf{x}\in\{0,1\}^{f(T)}$, $\mu_{f,\mathbf{x}}$ is piecewise linear (trivially) and continuous (because both actions have payoff 1 on the boundary between sections).
-Since the slope of each piece is in $\{-L,0,L\}$, $\mu_{f,\mathbf{x}}$ is Lipschitz continuous.
+$\mu_{f,\boldsymbol{a}}(x,y)=\begin{cases}1&\textnormal{ if }y=a_{j(x)}\\ 1-L\left(\cfrac{1}{2f(T)}-|m_{j(x)}-x|\right)&\textnormal{ if }y\neq a_{j(x)}\\ \end{cases}$
+We will use this payoff function for all time steps: $\mu_{t}=\mu_{f,\boldsymbol{a}}$ for all $t\in[T]$.
+Let $\pi^{m}$ be any policy which is optimal for $\mu_{f,\boldsymbol{a}}$.
+Note that there is a unique optimal action for each $x_{t}$, since each $x_{t}$ belongs to exactly one $X_{j}$; formally, $\pi^{m}(x_{t})=a_{j(x_{t})}$.
+For any $\boldsymbol{a}\in\{0,1\}^{f(T)}$, $\mu_{f,\boldsymbol{a}}$ is piecewise linear (trivially) and continuous (because both actions have payoff 1 on the boundary between sections).
+Since the slope of each piece is in $\{-L,0,L\}$, $\mu_{f,\boldsymbol{a}}$ is Lipschitz continuous.
 Thus by Proposition E.1, $\pi^{m}$ satisfies local generalization.
@@
 What if we restrict ourselves to settings where standard online learning is possible?
-Specifically, we assume that $\pi^{m}$ belongs to a policy class $\Pi$ where either (1) $\Pi$ has finite VC dimension $d$ and $\mathbf{s}$ is $\sigma$-smooth or (2) $\Pi$ has finite Littlestone dimension $d$ [^8] Recall from Section 1.3 that standard online learning becomes tractable under either of these assumptions.
+Specifically, we assume that $\pi^{m}$ belongs to a policy class $\Pi$ where either (1) $\Pi$ has finite VC dimension $d$ and $\boldsymbol{x}$ is $\sigma$-smooth or (2) $\Pi$ has finite Littlestone dimension $d$ [^10] Recall from Section 1.3 that standard online learning becomes tractable under either of these assumptions.
 This section presents a simple algorithm which guarantees subconstant regret and sublinear queries under either of those assumptions.
 Our algorithm needs to know $\Pi$, as is standard in online learning.
-The algorithm does not need to know $\sigma$ (in the smooth case) or $L$, and can handle an unbounded state space (the number of queries simply scales with the maximum distance between observed states).
-For simplicity, we initially prove our result for $A=\{0,1\}$.
-Appendix C extends our result to many actions using the“one versus rest” reduction.
-[^9] For each action $a$, we learn a binary classifier which predicts whether $\pi^{m}(s)=a$.
-If every binary classifier is correct, we can correctly determine $\pi^{m}(s)$.
-See, e.g., Chapter 29 of [34].
+The algorithm does not need to know $\sigma$ (in the smooth case) or $L$, and can handle an unbounded input space (the number of queries simply scales with the maximum distance between observed inputs).
+For simplicity, we initially prove our result for $\mathcal{Y}=\{0,1\}$ (the proof can be found in Appendix B).
+Appendix C extends our result to many actions using the standard “one versus rest” reduction.
+[^11] For each action $y$, we learn a binary classifier which predicts whether $\pi^{m}(x)=y$.
+If every binary classifier is correct, we can correctly determine $\pi^{m}(x)$.
+See, e.g., Chapter 29 of Shalev-Shwartz & Ben-David (2014).
 5.1 Intuition behind the algorithm
-Algorithm 1 has two simple components: (1) run a modified version of the Hedge algorithm for online learning, but (2) ask for help in unfamiliar states (specifically, when the current state is far from any queried state with the same action under the proposed policy).
-Hedge ensures that the number of time steps where the agent’s action doesn’t match the mentor’s is small, and asking for help in unfamiliar states ensures that when we do make a mistake, the cost isn’t too high.
+Algorithm 1 has two simple components: (1) run a modified version of the Hedge algorithm for online learning, but (2) ask for help for unfamiliar inputs (specifically, when the input is very different from any queried input with the same action under the proposed policy).
+Hedge ensures that the number of time steps where the agent’s action doesn’t match the mentor’s is small, and asking for help for unfamiliar inputs ensures that when we do make a mistake, the cost isn’t too high.
 This algorithmic structure seems quite natural: mostly follow a baseline strategy, but ask for help when out-of-distribution.
-Simple operations.
-The algorithm does not require detailed access to the state embedding, instead relying on two simple operations: evaluating a policy on a particular state, and computing a nearest neighbor distance.
-The former seems necessary for any algorithm.
-The latter could be modeled as an out-of-distribution detector score, for which many methods are available (see e.g., [41]).
+Agnostic to feature embedding.
+The algorithm does not need to know how inputs are represented in $\mathbb{R}^{n}$, since the only nontrivial operation it performs on the input space is computing a nearest neighbor distance (line 21 of Algorithm 1).
+This could be modeled as an out-of-distribution detector score, for which many methods are available (see e.g., Yang et al.
+(2024)).
+This agnosticism reduces the burden on the local generalization assumption: it suffices that there exists some feature embedding satisfying this assumption.
 Hedge.
-Hedge [12] is a standard online learning algorithm which ensures sublinear regret when the number of hypotheses (in our case, the number of policies in $\Pi$) is finite.
-[^10] See Chapter 5 of [35] and Chapter 21 of [34] for modern introductions to Hedge.
+Hedge (Freund & Schapire, 1997) is a standard online learning algorithm which ensures sublinear regret when the number of hypotheses (in our case, the number of policies in $\Pi$) is finite.
+[^12] See Chapter 5 of Slivkins et al.
+(2019) and Chapter 21 of Shalev-Shwartz & Ben-David (2014) for modern introductions to Hedge.
 We would prefer not to assume that $\Pi$ is finite.
-Luckily, any policy $\Pi$ can be approximated within $\varepsilon$ when either (1) $\Pi$ has finite VC dimension and $\mathbf{s}$ $\sigma$-smooth or (2) $\Pi$ has finite Littlestone dimension.
+Luckily, any policy $\Pi$ can be approximated within $\varepsilon$ when either (1) $\Pi$ has finite VC dimension and $\boldsymbol{x}$ is $\sigma$-smooth or (2) $\Pi$ has finite Littlestone dimension.
 Thus we can run Hedge on this approximative policy class instead.
@@
 To handle this, we modify Hedge to only perform updates on time steps with queries and to issue a query with probability $p$ on each time step.
-Continuing our lucky streak, [33] analyzes exactly this modification of Hedge.
+Continuing our lucky streak, Russo et al.
+(2024) analyze exactly this modification of Hedge.
 1:function AvoidCatastrophe($T\in\mathbb{N},\>\varepsilon\in\mathbb{R}_{>0},\>d\in\mathbb{N},\,$ policy class $\Pi$)
@@
 9:     $\eta\leftarrow\max\big{(}\sqrt{\frac{p\log|\tilde{\Pi}|}{2T}},\>\frac{p^{2}}{\sqrt{2}}\big{)}$
-10:     for $t$ from $1$ to $T$ do $\triangleright$ Run one step of Hedge, which selects policy $\pi_{t}$
+10:     for $t$
+from
+$1$
+to
+$T$ do $\triangleright$ Run one step of Hedge, which selects policy $\pi_{t}$
 […]
 12:         if hedgeQuery then
-13:              Query mentor and observe $\pi^{m}(s_{t})$
-14:              $\ell(t,\pi)\leftarrow\mathbf{1}(\pi(s_{t})\neq\pi^{m}(s_{t}))$ for all $\pi\in\tilde{\Pi}$
+13:              Query mentor and observe $\pi^{m}(x_{t})$
+14:              $\ell(t,\pi)\leftarrow\mathbf{1}(\pi(x_{t})\neq\pi^{m}(x_{t}))$ for all $\pi\in\tilde{\Pi}$
 15:              $\ell^{*}\leftarrow\min_{\pi\in\tilde{\Pi}}\ell(t,\pi)$
@@
 20:              Sample $\pi_{t}\sim P$
-21:         if $\min_{(s,a)\in X:a=\pi_{t}(s_{t})}||s_{t}-s||>\varepsilon^{1/n}$ then $\triangleright$ Ask for help if out-of-distribution
-22:              Query mentor and observe $\pi^{m}(s_{t})$
-23:              $X\leftarrow X\cup\{(s_{t},\pi^{m}(s_{t}))\}$
+21:         if $\min_{(x,y)\in X:y=\pi_{t}(x_{t})}||x_{t}-x||>\varepsilon^{1/n}$ then $\triangleright$ Ask for help if out-of-distribution
+22:              Query mentor and observe $\pi^{m}(x_{t})$
+23:              $X\leftarrow X\cup\{(x_{t},\pi^{m}(x_{t}))\}$
 24:         $\triangleright$ Otherwise, follow Hedge’s chosen policy
-25:              Take action $\pi_{t}(s_{t})$
+25:              Take action $\pi_{t}(x_{t})$
 Algorithm 1 successfully avoids catastrophe assuming finite VC or Littlestone dimension.
@@
 Theorem 5.1.
-Let $A=\{0,1\}$.
-Assume $\pi^{m}\in\Pi$ where either (1) $\Pi$ has finite VC dimension $d$, $\mathbf{s}$ is $\sigma$-smooth, and $\varepsilon T\log T>12\sigma d\log(4e^{2}/\varepsilon)$ or (2) $\Pi$ has finite Littlestone dimension $d$.
+Let $\mathcal{Y}=\{0,1\}$.
+Assume $\pi^{m}\in\Pi$ where either (1) $\Pi$ has finite VC dimension $d$, $\boldsymbol{x}$ is $\sigma$-smooth, and $\varepsilon T\log T>12\sigma d\log(4e^{2}/\varepsilon)$ or (2) $\Pi$ has finite Littlestone dimension $d$.
 Then for any $T\in\mathbb{N}$ and $\varepsilon>0$, Algorithm 1 satisfies
 $\operatorname*{\mathbb{E}}\left[R_{T}\right]\in$ $\ O\left(\frac{dL}{\sigma}T\varepsilon^{1+1/n}\log(1/\varepsilon)\log T\right)$
-$\operatorname*{\mathbb{E}}[|Q_{T}|]\in$ $\ O\left(\sqrt{\frac{T}{\varepsilon}}+\frac{d}{\sigma}T\varepsilon\log(1/\varepsilon)\log T+\frac{\operatorname*{diam}(\mathbf{s})^{n}}{\varepsilon}\right)$
-In Case 1, the expectation is over the randomness of both $\mathbf{s}$ and the algorithm, while in Case 2, the expectation is over only the randomness of the algorithm.
+$\operatorname*{\mathbb{E}}[|Q_{T}|]\in$ $\ O\left(\sqrt{\frac{T}{\varepsilon}}+\frac{d}{\sigma}T\varepsilon\log(1/\varepsilon)\log T+\frac{\operatorname*{diam}(\boldsymbol{x})^{n}}{\varepsilon}\right)$
+In Case 1, the expectation is over the randomness of both $\boldsymbol{x}$ and the algorithm, while in Case 2, the expectation is over only the randomness of the algorithm.
 Also, $R_{T}$ and $Q_{T}$ clearly have no dependence on $\sigma$ in Case 2, but we include $\sigma$ anyway to avoid writing two separate bounds.
@@
 Theorem 5.2.
-Let $A=\{0,1\}$.
-Assume $\pi^{m}\in\Pi$ where either (1) $\Pi$ has finite VC dimension $d$ and $\mathbf{s}$ is $\sigma$-smooth or (2) $\Pi$ has finite Littlestone dimension $d$.
+Let $\mathcal{Y}=\{0,1\}$.
+Assume $\pi^{m}\in\Pi$ where either (1) $\Pi$ has finite VC dimension $d$ and $\boldsymbol{x}$ is $\sigma$-smooth or (2) $\Pi$ has finite Littlestone dimension $d$.
 Then for any $T\in\mathbb{N}$, Algorithm 1 with $\varepsilon=T^{\frac{-2n}{2n+1}}$ satisfies
 $\operatorname*{\mathbb{E}}\left[R_{T}\right]\in$ $\ O\left(\frac{dL}{\sigma}T^{\frac{-1}{2n+1}}\log T\right)$
-$\operatorname*{\mathbb{E}}[|Q_{T}|]\in$ $\ O\left(T^{\frac{4n+1}{4n+2}}\left(\frac{d}{\sigma}\log T+\operatorname*{diam}(\mathbf{s})^{n}\right)\right)$
-We can also make the query bound unconditional in exchange for slightly worse bounds.
-The probability that $|Q_{T}|$ asymptotically exceeds the bound from Theorem 5.2 goes to 0, so the query threshold in Corollary 5.2.1 is almost never reached.
-Thus we still obtain subconstant regret.
-Corollary 5.2.1.
-If Algorithm 1 is modified to stop querying after $T^{\frac{4n+1.5}{4n+2}}(\frac{d}{\sigma}\log T+\operatorname*{diam}(\mathbf{s})^{n})$ queries, then under the conditions of Theorem 5.2, we have $\operatorname*{\mathbb{E}}[R_{T}]\in O\big{(}\frac{dL}{\sigma}T^{\frac{-1}{8n+4}}\log T\big{)}$.
+$\operatorname*{\mathbb{E}}[|Q_{T}|]\in$ $\ O\left(T^{\frac{4n+1}{4n+2}}\left(\frac{d}{\sigma}\log T+\operatorname*{diam}(\boldsymbol{x})^{n}\right)\right)$
+Although our focus is the product of payoffs, Algorithm 1 also guarantees subconstant additive regret:
+Theorem 5.3.
+Let $\mathcal{Y}=\{0,1\}$.
+Assume $\pi^{m}\in\Pi$ where either (1) $\Pi$ has finite VC dimension $d$ and $\boldsymbol{x}$ is $\sigma$-smooth or (2) $\Pi$ has finite Littlestone dimension $d$.
+Then for any $T\in\mathbb{N}$, Algorithm 1 with $\varepsilon=T^{\frac{-2n}{2n+1}}$ satisfies
+$\operatorname*{\mathbb{E}}\left[\sum_{t=1}^{T}\mu_{t}^{m}(x_{t})-\sum_{t=1}^{T}\mu_{t}(x_{t},y_{t})\right]\in O\left(\frac{dL}{\sigma}T^{\frac{-1}{2n+1}}\log T\right)$
 5.2 Proof sketch
@@
 Definition 5.4.
-Let $U$ be the uniform distribution over $S$.
-For $\varepsilon>0$, a policy class $\tilde{\Pi}$ is a smooth $\varepsilon$-cover of a policy class $\Pi$ is for every $\pi\in\Pi$, there exists $\tilde{\pi}\in\tilde{\Pi}$ such that $\Pr_{s\sim U}[\pi(s)\neq\tilde{\pi}(s)]\leq\varepsilon$.
+Let $U$ be the uniform distribution over $\mathcal{X}$.
+For $\varepsilon>0$, a policy class $\tilde{\Pi}$ is a smooth $\varepsilon$-cover of a policy class $\Pi$ is for every $\pi\in\Pi$, there exists $\tilde{\pi}\in\tilde{\Pi}$ such that $\Pr_{x\sim U}[\pi(x)\neq\tilde{\pi}(x)]\leq\varepsilon$.
 Definition 5.5.
-A policy class $\tilde{\Pi}$ is an adversarial cover of a policy class $\Pi$ is for every $\mathbf{s}\in S^{T}$ and $\pi\in\Pi$, there exists $\tilde{\pi}\in\tilde{\Pi}$ such that $\pi(s_{t})=\tilde{\pi}(s_{t})$ for all $t\in[T]$.
+A policy class $\tilde{\Pi}$ is an adversarial cover of a policy class $\Pi$ is for every $\boldsymbol{x}\in\mathcal{X}^{T}$ and $\pi\in\Pi$, there exists $\tilde{\pi}\in\tilde{\Pi}$ such that $\pi(x_{t})=\tilde{\pi}(x_{t})$ for all $t\in[T]$.
 The existence of small covers is crucial:
-Lemma 5.1 (Lemma 7.3.2 in [16] [^11] See also [19] or Lemma 13.6 in [6] for variants which are less convenient for our purposes.).
+Lemma 5.1 (Lemma 7.3.2 in Haghtalab (2018) [^13] See also Haussler & Long (1995) or Lemma 13.6 in Boucheron et al. (2013) for variants which are less convenient for our purposes.).
 For all $\varepsilon>0$, any policy class of VC dimension $d$ admits a smooth $\varepsilon$-cover of size at most $(41/\varepsilon)^{d}$.
-Lemma 5.2 (Lemmas 21.13 and A.5 in [34]).
+Lemma 5.2 (Lemmas 21.13 and A.5 in Shalev-Shwartz & Ben-David (2014)).
 Any policy class of Littlestone dimension $d$ admits an adversarial cover of size at most $(eT/d)^{d}$.
 An adversarial cover is a perfect cover by definition. The following lemma establishes that a smooth $\varepsilon$-cover is a good approximation for any sequence of $\sigma$-smooth distributions.
-Lemma 5.3 (Equation 2 and Lemma 3.3 in [17]).
+Lemma 5.3 (Equation 2 and Lemma 3.3 in Haghtalab et al. (2024)).
 Let $\tilde{\Pi}$ be a finite smooth $\varepsilon$-cover of $\Pi$ and let $\boldsymbol{\mathcal{D}}=\mathcal{D}_{1},\dots,\mathcal{D}_{T}$ be a sequence of $\sigma$-smooth distributions.
-If $\varepsilon T\log T>12\sigma d\log(4e^{2}/\varepsilon)$, then
-$\operatorname*{\mathbb{E}}_{\mathbf{s}\sim\boldsymbol{\mathcal{D}}}\left[\sup\limits_{\pi\in\Pi}\min\limits_{\tilde{\pi}\in\tilde{\Pi}}\sum_{t=1}^{T}\mathbf{1}(\pi(s_{t})\neq\tilde{\pi}(s_{t}))\right]\in O\left(\sqrt{\frac{d\varepsilon T}{\sigma}\log T\log(1/\varepsilon)}+\frac{\varepsilon T}{\sigma}\log T\right)$
+If $\varepsilon T\log T>12\sigma d\log(4e^{2}/\varepsilon)$, then $\operatorname*{\mathbb{E}}\limits_{\boldsymbol{x}\sim\boldsymbol{\mathcal{D}}}\left[\sup\limits_{\pi\in\Pi}\min\limits_{\tilde{\pi}\in\tilde{\Pi}}\sum_{t=1}^{T}\mathbf{1}(\pi(x_{t})\neq\tilde{\pi}(x_{t}))\right]\in O\left(\frac{1}{\sigma}T\varepsilon\log T\sqrt{d\log(1/\varepsilon)}\right)$.
 We will run a variant of Hedge on $\tilde{\Pi}$.
 The vanilla Hedge algorithm operates in the standard online learning model where on each time step, the agent selects a policy (or more generally, a hypothesis), and observes the loss of every policy.
-In general the loss function can depend arbitrarily on the time step, the policy, and prior events, but we will only use the indicator loss function $\ell(t,\pi)=\mathbf{1}(\pi(s_{t})\neq\pi^{m}(s_{t}))$.
-Crucially, whenever we query and learn $\pi^{m}(s_{t})$, we can compute $\ell(t,\pi)$ for every $\pi\in\tilde{\Pi}$.
+In general the loss function can depend arbitrarily on the time step, the policy, and prior events, but we will only use the indicator loss function $\ell(t,\pi)=\mathbf{1}(\pi(x_{t})\neq\pi^{m}(x_{t}))$.
+Crucially, whenever we query and learn $\pi^{m}(x_{t})$, we can compute $\ell(t,\pi)$ for every $\pi\in\tilde{\Pi}$.
 We cannot afford to query on every time step, however.
-Recently, [33] analyzed a variant of Hedge where losses are observed only in response to queries, which they call “label-efficient feedback”.
+Recently, Russo et al.
+(2024) analyzed a variant of Hedge where losses are observed only in response to queries, which they call “label-efficient feedback”.
 They proved a regret bound when a query is issued on each time step with fixed probability $p$.
@@
 Full pseudocode for HedgeWithQueries can also be found in the appendix (Algorithm 2).
-Lemma 5.4 (Lemma 3.5 in [33]).
+Lemma 5.4 (Lemma 3.5 in Russo et al. (2024)).
 Assume $\tilde{\Pi}$ is finite. Then for any loss function $\ell:[T]\times\tilde{\Pi}\to[0,1]$ and query probability $p$, HedgeWithQueries enjoys the regret bound
@@
 where $\pi_{t}$ is the policy chosen at time $t$ and the expectation is over the randomness of the algorithm.
-We apply Lemma 5.4 with $\ell(t,\pi)=\mathbf{1}(\pi(s_{t})\neq\pi^{m}(s_{t}))$ and combine this with Lemmas 5.1 and 5.3 (in the $\sigma$-smooth case) and with Lemma 5.2 (in the adversarial case).
+We apply Lemma 5.4 with $\ell(t,\pi)=\mathbf{1}(\pi(x_{t})\neq\pi^{m}(x_{t}))$ and combine this with Lemmas 5.1 and 5.3 (in the $\sigma$-smooth case) and with Lemma 5.2 (in the adversarial case).
 This yields a $O\left(\frac{d}{\sigma}T\varepsilon\log(1/\varepsilon)\log T\right)$ bound on the number of mistakes made by Algorithm 1 (Lemma B.1).
@@
 The trickier part is bounding the number of resulting queries.
-It is tempting to claim that the states queried in the out-of-distribution case must all be separated by at least $\varepsilon^{1/n}$ and thus form an $\varepsilon^{1/n}$-packing, but this is actually not true.
-Instead, the packing analysis must be performed on a per-action basis, and we reuse Lemma B.1 to crucially show that there isn’t too much “spillover” between actions.
-The resulting overall query bound is given by Lemma B.7.
+It is tempting to claim that the inputs queried in the out-of-distribution case must all be separated by at least $\varepsilon^{1/n}$ and thus form an $\varepsilon^{1/n}$-packing, but this is actually not true.
+Instead, we provide a novel method for bounding the number of data points (i.e., queries) needed to cover a set with respect to the realized actions of the algorithm (Lemma B.7).
+This is in contrast to vanilla packing arguments which consider all data points in aggregate.
+Our method may be useful in other contexts where a more refined packing argument is needed.
 6 Conclusion and future work
@@
 Also, we have not resolved whether our problem is tractable for finite VC dimension and fully adversarial inputs (although Appendix D shows that the problem is tractable for at least some classes with finite VC but infinite Littlestone dimension).
-We are also interested in alternatives to the local generalization assumption, since an action which is safe in one state may not always be safe in a nearby state (depending on the definition of “nearby”).
-Some assumption is necessary, or there is no way to avoid costly mistakes without querying on every time step.
-One possibility is Bayesian inference.
+We are also interested in alternatives to the local generalization assumption.
+We should expect some assumption to be necessary: if not, the payoff function $\mu(x,y)=\mathbf{1}(\pi^{m}(x)=y)$ means that the agent must make zero mistakes (with high probability), which turns out to be impossible even for $\sigma$-smooth $\boldsymbol{x}$ and finite VC dimension (Theorem E.3).
+One possible alternative is Bayesian inference.
 We intentionally avoided Bayesian approaches in this paper due to tractability concerns, but it seems premature to abandon those ideas entirely.
@@
 Is there an algorithm for non-communicating single-episode undiscounted MDPs which ensures that both the regret and the number of mentor queries are sublinear in $T$?
-Acknowledgements
-This work was supported in part by a gift from Open Philanthropy to the Center for Human-Compatible AI (CHAI) at UC Berkeley.
-This paper also benefited from discussions with many other researchers.
-We would like to especially thank (in alphabetical order) Cameron Allen, Michael Cohen, Scott Emmons, Nika Haghtalab, Aly Lidayan, Juan Liévano-Karim, Bhaskar Mishra, and Matteo Russo.
 Appendix A Proof of Theorem 4.1
@@
 […]
-Recall that $\operatorname*{\mathbb{E}}_{\mathbf{s}\sim U^{T},\mathbf{a}}$ denotes the expectation over both the random states and any additional randomness in the algorithm.
-The proof proceeds via the following steps:
+The proof considers an arbitrary algorithm with sublinear queries, and proceeds via the following steps:
   1. 1.
@@
   2. 2.
-    Provide a simple variant of the Chernoff bound which we will apply multiple times (Lemma A.2).
+    Prove a simple variant of the Chernoff bound which we will apply multiple times (Lemma A.2).
   3. 3.
-    Show that with high probability, $\sum_{j\in A}|V_{j}|$ is adequately large (Lemma A.3).
+    Show that with high probability, $\sum_{j\in A}|M_{j}|$ is adequately large (Lemma A.3).
   4. 4.
-    The key lemma is Lemma A.4, which shows that a randomly sampled $\mathbf{x}$ produces poor agent performance with high probability.
+    The key lemma is Lemma A.4, which shows that a randomly sampled $\boldsymbol{a}$ produces poor agent performance with high probability.
 The central idea is that at least $f(T)-|Q_{T}|$ sections are never queried (which is large, by Lemma A.1), so the agent has no way of knowing the optimal action in those sections.
@@
   5. 5.
-    To complete the proof, we observe that $\sup_{\mu}\ \operatorname*{\mathbb{E}}_{\mathbf{s}\sim U^{T},\,\mathbf{a}}\ R_{T}(\mathbf{s},\mathbf{a},\mu,\pi^{m})\geq\operatorname*{\mathbb{E}}_{\mathbf{x}\sim U(\{0,1\}^{f(T)})}\ \operatorname*{\mathbb{E}}_{\mathbf{s}\sim U^{T},\,\mathbf{a}}\ R_{T}(\mathbf{s},\mathbf{a},\mu_{f,\mathbf{x}},\pi^{m})$.
-This is essentially an application of the probabilistic method: if a randomly chosen $\mu_{f,\mathbf{x}}$ has high expected regret, then the worst case $\mu$ also has high expected regret.
-Note that $\mathbf{s},\mathbf{a}$, and $\mathbf{x}$ are random variables, so all variables defined on top of them (such as $V_{j}$) are also random variables.
-In contrast, the partition $S=\{S_{1},\dots,S_{f(T)}\}$ and properties thereof (like the midpoints $m_{j}$) are not random variables.
-Lastly, while the intuition provided in Section 4.1 is accurate, the analysis will mostly occur in log space, so the bounds will look different.
+    Finally, $\sup\limits_{\boldsymbol{\mu}}\ \operatorname*{\mathbb{E}}\limits_{\boldsymbol{x}\sim U^{T},\,\boldsymbol{y}}\ R_{T}(\boldsymbol{x},\boldsymbol{y},\boldsymbol{\mu},\pi^{m})\geq\operatorname*{\mathbb{E}}\limits_{\boldsymbol{a}\sim U(\{0,1\}^{f(T)})}\ \operatorname*{\mathbb{E}}\limits_{\boldsymbol{x}\sim U^{T},\,\boldsymbol{y}}\ R_{T}(\boldsymbol{x},\boldsymbol{y},\mu_{f,\boldsymbol{a}},\pi^{m})$, where $U(\{0,1\}^{f(T)})$ is the uniform distribution over bit strings of length $f(T)$.
+This is essentially an application of the probabilistic method: if a randomly chosen $\mu_{f,\boldsymbol{a}}$ has high expected regret, then the worst case $\boldsymbol{\mu}$ also has high expected regret.
+While reading the proof, one should keep in mind the following:
+  1. 1.
+    Since $\boldsymbol{x},\boldsymbol{y}$, and $\boldsymbol{a}$ are random variables, all variables defined on top of them (such as $M_{j}$) are also random variables.
+In contrast, the partition $\mathcal{X}=\{X_{1},\dots,X_{f(T)}\}$ and properties thereof (like the midpoints $m_{j}$) are not random variables.
+  2. 2.
+    We will assume that the worst-case (not expected) number of queries is sublinear in $T$. Proposition E.2 shows that relaxing this to expected sublinear queries would not affect our conclusion.
+  3. 3.
+    While the intuition provided in Section 4.1 is accurate, the analysis will mostly occur in log space, so the bounds will look different.
 However, bounds of the form discussed in Section 4.1 can still be found as an intermediate step in Part 4 of the proof of Lemma A.4.
@@
 Lemma A.2.
-Let $z_{1},\dots,z_{n}$ be i.i.d. variables in $\{0,1\}$ and let $Z=\sum_{i=1}^{n}z_{i}$. If $\operatorname*{\mathbb{E}}[Z]\geq M$, then $\Pr\big{[}Z\leq M/2\big{]}\leq\exp(-M/8)$.
+Let $z_{1},\dots,z_{n}$ be i.i.d. variables in $\{0,1\}$ and let $Z=\sum_{i=1}^{n}z_{i}$. If $\operatorname*{\mathbb{E}}[Z]\geq W$, then $\Pr\big{[}Z\leq W/2\big{]}\leq\exp(-W/8)$.
 Proof.
@@
 binary variables, we have $\Pr[Z\leq\operatorname*{\mathbb{E}}[Z]/2]\leq\exp(-\operatorname*{\mathbb{E}}[Z]/8)$.
-Since $-\operatorname*{\mathbb{E}}[Z]\leq-M$ and $\exp$ is an increasing function, we have $\exp(-\operatorname*{\mathbb{E}}[Z]/8)\leq\exp(-M/8)$.
-Also, $M/2\leq E[Z]/2$ implies $\Pr[Z\leq M/2]\leq\Pr[Z\leq\operatorname*{\mathbb{E}}[Z]/2]$.
+Since $-\operatorname*{\mathbb{E}}[Z]\leq-W$ and $\exp$ is an increasing function, we have $\exp(-\operatorname*{\mathbb{E}}[Z]/8)\leq\exp(-W/8)$.
+Also, $W/2\leq E[Z]/2$ implies $\Pr[Z\leq W/2]\leq\Pr[Z\leq\operatorname*{\mathbb{E}}[Z]/2]$.
 Combining these inequalities proves the lemma.
@@
 let $A\subseteq[f(T)]$ be any nonempty subset of sections. Then
-$\Pr\left[\sum_{j\in A}|V_{j}|\leq\frac{T|A|}{4f(T)}\right]\leq\exp\left(\frac{-T}{16f(T)}\right)$
+$\Pr\left[\sum_{j\in A}|M_{j}|\leq\frac{T|A|}{4f(T)}\right]\leq\exp\left(\frac{-T}{16f(T)}\right)$
 Proof.
 […]
-$\Pr\left[\sum_{j\in A}|V_{j}|\leq\frac{T|A|}{4f(T)}\right]\leq\exp\left(\frac{-T|A|}{16f(T)}\right)\leq\exp\left(\frac{-T}{16f(T)}\right)$
+$\Pr\left[\sum_{j\in A}|M_{j}|\leq\frac{T|A|}{4f(T)}\right]\leq\exp\left(\frac{-T|A|}{16f(T)}\right)\leq\exp\left(\frac{-T}{16f(T)}\right)$
 with the last step due to $|A|\geq 1$. ∎
 Lemma A.4.
-Independently sample $\mathbf{x}\sim U(\{0,1\}^{f(T)})$ and $\mathbf{s}\sim U^{T}$.
-[^12] That is, the entire set $\{x_{1},\dots,x_{f(T)},s_{1},\dots,s_{T}\}$ is mutually independent.
-Then with probability at least $1-\exp\big{(}\frac{-T}{16f(T)}\big{)}-\exp\big{(}-\frac{f(T)-|Q_{T}|}{16}\big{)}$,
-$\prod_{t=1}^{T}\mu_{f,\mathbf{x}}(s_{t},a_{t})\leq\exp\left(-\frac{LT(f(T)-|Q_{T}|)}{2^{7}f(T)^{2}}\right)$
+Let $f,g:\mathbb{N}\to\mathbb{N}$.
+Suppose $|Q_{T}|\leq g(T)$ always and independently sample $\boldsymbol{a}\sim U(\{0,1\}^{f(T)})$ and $\boldsymbol{x}\sim U^{T}$.
+[^14] That is, the entire set $\{a_{1},\dots,a_{f(T)},x_{1},\dots,x_{T}\}$ is mutually independent.
+Then with probability at least $1-\exp\big{(}\frac{-T}{16f(T)}\big{)}-\exp\big{(}-\frac{f(T)-g(T)}{16}\big{)}$,
+$\prod_{t=1}^{T}\mu_{f,\boldsymbol{a}}(x_{t},y_{t})\leq\exp\left(-\frac{LT(f(T)-g(T))}{2^{7}f(T)^{2}}\right)$
 Proof.
 Part 1: setup.
-Let $J_{\neg Q}=\{j\in[f(T)]:s_{t}\not\in S_{j}\ \forall t\in Q_{T}\}$ be the set of sections that are never queried.
-Since each query appears in exactly one section (because each state appears in exactly one section), $|J_{\neg Q}|\geq f(T)-|Q_{T}|$.
-For each $j\in J_{\neg Q}$, let $y_{j}$ be the action taken most frequently among time steps in $V_{j}$:
-$y_{j}=\operatorname*{arg\,max}_{a\in\{0,1\}}\Big{|}\{t\in V_{j}:a=a_{t}\}\Big{|}$
-Let $\bar{J}=\{j\in J_{\neg Q}:x_{j}\neq y_{j}\}$.
-For each $j\in\bar{J}$, let $V_{j}^{\prime}=\{t\in V_{j}:a_{t}\neq x_{j}\}$ be the set of time steps where the agent chooses the wrong action (assuming payoff function $\mu_{f,\mathbf{x}}$).
+Let $J_{\neg Q}=\{j\in[f(T)]:x_{t}\not\in X_{j}\ \forall t\in Q_{T}\}$ be the set of sections that are never queried.
+Since each query appears in exactly one section (because each input appears in exactly one section), $|J_{\neg Q}|\geq f(T)-|Q_{T}|\geq f(T)-g(T)$.
+For each $j\in J_{\neg Q}$, let $y_{j}$ be the action taken most frequently among time steps in $M_{j}$:
+$y_{j}=\operatorname*{arg\,max}_{y\in\{0,1\}}\Big{|}\{t\in M_{j}:y=y_{t}\}\Big{|}$
+Let $\bar{J}=\{j\in J_{\neg Q}:a_{j}\neq y_{j}\}$.
+For each $j\in\bar{J}$, let $M_{j}^{\prime}=\{t\in M_{j}:y_{t}\neq a_{j}\}$ be the set of time steps where the agent chooses the wrong action (assuming payoff function $\mu_{f,\boldsymbol{a}}$).
 […]
-$\Pr\left[|\bar{J}|\leq\frac{f(T)-|Q_{T}|}{4}\right]\leq\exp\left(-\frac{f(T)-|Q_{T}|}{16}\right)$
-Part 3: $|V_{j}^{\prime}|\geq|V_{j}|/2$.
-Since $j\in J_{\neg Q}$, the mentor is not queried on any time step $t\in V_{j}$, so $a_{t}\in\{0,1\}$ for all $t\in V_{j}$.
-Since the agent chooses one of two actions for each $t\in V_{j}$, the more frequent action must be chosen chosen at least half of the time: $a_{t}=y_{j}$ for at least half of the time steps in $V_{j}$.
-Since $x_{j}\neq y_{j}$ for $j\in\bar{J}$, we have $a_{t}=y_{j}\neq x_{j}$ for those time steps, so $|V_{j}^{\prime}|\geq|V_{j}|/2$.
-Part 4: a bound in terms of $\bar{J}$ and $V_{j}$.
-Consider any $j\in\bar{J}$ and $t\in V_{j}^{\prime}\subseteq V_{j}$.
-By definition of $V_{j}$, we have $|m_{j}-s_{t}|\leq\frac{1}{4f(T)}$.
-Then by definition of $\mu_{f,\mathbf{x}}$,
-$\mu_{f,\mathbf{x}}(s_{t},a_{t})=$ $\ 1-L\left(\frac{1}{2f(T)}-|s_{t}-m_{j}|\right)$
+$\Pr\left[|\bar{J}|\leq\frac{f(T)-g(T)}{4}\right]\leq\exp\left(-\frac{f(T)-g(T)}{16}\right)$
+Part 3: $|M_{j}^{\prime}|\geq|M_{j}|/2$.
+Since $j\in J_{\neg Q}$, the mentor is not queried on any time step $t\in M_{j}$, so $y_{t}\in\{0,1\}$ for all $t\in M_{j}$.
+Since the agent chooses one of two actions for each $t\in M_{j}$, the more frequent action must be chosen chosen at least half of the time: $y_{t}=y_{j}$ for at least half of the time steps in $M_{j}$.
+Since $a_{j}\neq y_{j}$ for $j\in\bar{J}$, we have $y_{t}=y_{j}\neq a_{j}$ for those time steps, so $|M_{j}^{\prime}|\geq|M_{j}|/2$.
+Part 4: a bound in terms of $\bar{J}$ and $M_{j}$.
+Consider any $j\in\bar{J}$ and $t\in M_{j}^{\prime}\subseteq M_{j}$.
+By definition of $M_{j}$, we have $|m_{j}-x_{t}|\leq\frac{1}{4f(T)}$.
+Then by definition of $\mu_{f,\boldsymbol{a}}$,
+$\mu_{f,\boldsymbol{a}}(x_{t},y_{t})=$ $\ 1-L\left(\frac{1}{2f(T)}-|x_{t}-m_{j}|\right)$
 $\leq$ $\ 1-L\left(\frac{1}{2f(T)}-\frac{1}{4f(T)}\right)$
@@
 […]
-$\log\prod_{t=1}^{T}\mu_{f,\mathbf{x}}(s_{t},a_{t})\leq$ $\ \log\prod_{j\in\bar{J}}\left(1-\frac{L}{4f(T)}\right)^{|V_{j}|/2}$
-$=$ $\ \sum_{j\in\bar{J}}\frac{|V_{j}|}{2}\log\left(1-\frac{L}{4f(T)}\right)$
-$\leq$ $\ -\sum_{j\in\bar{J}}\frac{L|V_{j}|}{8f(T)}$
+$\log\prod_{t=1}^{T}\mu_{f,\boldsymbol{a}}(x_{t},y_{t})\leq$ $\ \log\prod_{j\in\bar{J}}\left(1-\frac{L}{4f(T)}\right)^{|M_{j}|/2}$
+$=$ $\ \sum_{j\in\bar{J}}\frac{|M_{j}|}{2}\log\left(1-\frac{L}{4f(T)}\right)$
+$\leq$ $\ -\sum_{j\in\bar{J}}\frac{L|M_{j}|}{8f(T)}$
 […]
-$\log\prod_{t=1}^{T}\mu_{f,\mathbf{x}}(s_{t},a_{t})\leq$ $\ -\sum_{j\in\bar{J}}\frac{L|V_{j}|}{8f(T)}$
+$\log\prod_{t=1}^{T}\mu_{f,\boldsymbol{a}}(x_{t},y_{t})\leq$ $\ -\sum_{j\in\bar{J}}\frac{L|M_{j}|}{8f(T)}$
 $\leq$ $\ -\frac{L}{8f(T)}\cdot\frac{T|\bar{J}|}{4f(T)}$
-$\leq$ $\ -\frac{L}{8f(T)}\cdot\frac{T}{4f(T)}\cdot\frac{f(T)-|Q_{T}|}{4}$
-$=$ $\ -\frac{LT(f(T)-|Q_{T}|)}{2^{7}f(T)^{2}}$
+$\leq$ $\ -\frac{L}{8f(T)}\cdot\frac{T}{4f(T)}\cdot\frac{f(T)-g(T)}{4}$
+$=$ $\ -\frac{LT(f(T)-g(T))}{2^{7}f(T)^{2}}$
 Exponentiating both sides proves the lemma. ∎
-Let $\alpha(T)=\exp\big{(}\frac{-T}{16f(T)}\big{)}+\exp\big{(}-\frac{f(T)-|Q_{T}|}{16}\big{)}$ for brevity. See 4.1
+Let $\alpha(T)=\exp\big{(}\frac{-T}{16f(T)}\big{)}+\exp\big{(}-\frac{f(T)-g(T)}{16}\big{)}$ for brevity (with $f$ and $g$ to be chosen later). See 4.1
 Proof.
-If the algorithm has sublinear queries, then there exists $g(T)\in o(T)$ such that $|Q_{T}|\leq g(T)$ for all $\mathbf{s},\mu,\pi^{m}$.
-Let $f(T)=\sqrt{(g(T)+1)T}$.
-Then by Lemma A.1, $g(T)\in o(f(T))$ and $f(T)\in o(T)$.
-Combining this with $|Q_{T}|\leq g(T)$, we get $\lim_{T\to\infty}\alpha(T)=0$.
-Also, since $|Q_{T}|\in o(f(T))$, there exists $T_{0}$ such that $|Q_{T}|\leq f(T)/2$ for all $T\geq T_{0}$.
-Combining this with Lemma A.4 and noting that $\prod_{t=1}^{T}\mu_{f,\mathbf{x}}(s_{t},a_{t})\leq 1$, we have
-$\ \operatorname*{\mathbb{E}}_{\mathbf{x}\sim U(\{0,1\}^{f(T)})}\ \,\operatorname*{\mathbb{E}}_{\mathbf{s}\sim U^{T},\,\mathbf{a}}\ \>\prod_{t=1}^{T}\mu_{f,\mathbf{x}}(s_{t},a_{t})$
-$\leq$ $\ \alpha(T)\cdot 1+\big{(}1-\alpha(T)\big{)}\exp\left(-\frac{LT(f(T)-|Q_{T}|)}{2^{7}f(T)^{2}}\right)$
+If the algorithm has sublinear queries, then there exists $g(T)\in o(T)$ such that $|Q_{T}|\leq g(T)$ always.
+Then by Lemma A.1, $g(T)\in o(\sqrt{g(T)T})$ and $\sqrt{g(T)T}\in o(T)$.
+Let $f(T)=\lceil\sqrt{g(T)T}\rceil$: then $f(T)\in\Theta(\sqrt{g(T)T})$, so $g(T)\in o(f(T))$ and $f(T)\in o(T)$.
+First, this implies that $\lim_{T\to\infty}\alpha(T)=0$.
+Next, $g(T)\in o(f(T))$ implies that exists $T_{0}$ such that $g(T)\leq f(T)/2$ for all $T\geq T_{0}$.
+Combining this with Lemma A.4 and noting that $\prod_{t=1}^{T}\mu_{f,\boldsymbol{a}}(x_{t},y_{t})\leq 1$, for all $T\geq T_{0}$ we have
+$\ \operatorname*{\mathbb{E}}_{\boldsymbol{a}\sim U(\{0,1\}^{f(T)})}\ \,\operatorname*{\mathbb{E}}_{\boldsymbol{x}\sim U^{T},\,\boldsymbol{y}}\ \>\prod_{t=1}^{T}\mu_{f,\boldsymbol{a}}(x_{t},y_{t})$
+$\leq$ $\ \alpha(T)\cdot 1+\big{(}1-\alpha(T)\big{)}\exp\left(-\frac{LT(f(T)-|g(T))}{2^{7}f(T)^{2}}\right)$
 $\leq$ $\ \alpha(T)+\big{(}1-\alpha(T)\big{)}\exp\left(-\frac{LTf(T)/2}{2^{7}f(T)^{2}}\right)$
 $=$ $\ \alpha(T)+\big{(}1-\alpha(T)\big{)}\exp\left(-\frac{LT}{2^{8}f(T)}\right)$
-whenever $T\geq T_{0}$. Since $\prod_{t=1}^{T}\mu_{f,\mathbf{x}}^{m}(s_{t})=1$ always, we have [^13] Fubini’s theorem means we need not worry about the order of the expectation operators.
-$\sup_{\mu}\ \operatorname*{\mathbb{E}}_{\mathbf{s}\sim U^{T},\,\mathbf{a}}\ R_{T}(\mathbf{s},\mathbf{a},\mu,\pi^{m})\geq$ $\ \operatorname*{\mathbb{E}}_{\mathbf{x}\sim U(\{0,1\}^{f(T)})}\ \,\operatorname*{\mathbb{E}}_{\mathbf{s}\sim U^{T},\,\mathbf{a}}\ R_{T}(\mathbf{s},\mathbf{a},\mu_{f,\mathbf{x}},\pi^{m})$
-$=\operatorname*{\mathbb{E}}_{\mathbf{x}\sim U(\{0,1\}^{f(T)})}\ \,\operatorname*{\mathbb{E}}_{\mathbf{s}\sim U^{T},\,\mathbf{a}}$ $\ \,\left[\prod_{t=1}^{T}\mu_{f,\mathbf{x}}^{m}(s_{t})-\prod_{t=1}^{T}\mu_{f,\mathbf{x}}(s_{t},a_{t})\right]$
+Since $\prod_{t=1}^{T}\mu_{f,\boldsymbol{a}}^{m}(x_{t})=1$ always, we have [^15] Fubini’s theorem means we need not worry about the order of the expectation operators.
+$\sup_{\boldsymbol{\mu}}\ \operatorname*{\mathbb{E}}_{\boldsymbol{x}\sim U^{T},\,\boldsymbol{y}}\ R_{T}(\boldsymbol{x},\boldsymbol{y},\boldsymbol{\mu},\pi^{m})\geq$ $\ \operatorname*{\mathbb{E}}_{\boldsymbol{a}\sim U(\{0,1\}^{f(T)})}\ \,\operatorname*{\mathbb{E}}_{\boldsymbol{x}\sim U^{T},\,\boldsymbol{y}}\ R_{T}(\boldsymbol{x},\boldsymbol{y},(\mu_{f,\boldsymbol{a}},\dots,\mu_{f,\boldsymbol{a}}),\pi^{m})$
+$=\operatorname*{\mathbb{E}}_{\boldsymbol{a}\sim U(\{0,1\}^{f(T)})}\ \,\operatorname*{\mathbb{E}}_{\boldsymbol{x}\sim U^{T},\,\boldsymbol{y}}$ $\ \,\left[\prod_{t=1}^{T}\mu_{f,\boldsymbol{a}}^{m}(x_{t})-\prod_{t=1}^{T}\mu_{f,\boldsymbol{a}}(x_{t},y_{t})\right]$
 $\geq$ $\ 1-\alpha(T)-\big{(}1-\alpha(T)\big{)}\exp\left(-\frac{LT}{2^{8}f(T)}\right)$
 Therefore
-$\lim_{T\to\infty}\sup_{\mu}\ \operatorname*{\mathbb{E}}_{\mathbf{s}\sim U^{T},\,\mathbf{a}}\ R_{T}(\mathbf{s},\mathbf{a},\mu,\pi^{m})\geq$ $\ 1-\lim_{T\to\infty}\alpha(T)-\big{(}1-\lim_{T\to\infty}\alpha(T)\big{)}\cdot\exp\left(\lim_{T\to\infty}-\frac{LT}{2^{8}f(T)}\right)$
+$\lim_{T\to\infty}\sup_{\boldsymbol{\mu}}\ \operatorname*{\mathbb{E}}_{\boldsymbol{x}\sim U^{T},\,\boldsymbol{y}}\ R_{T}(\boldsymbol{x},\boldsymbol{y},\boldsymbol{\mu},\pi^{m})\geq$ $\ 1-\lim_{T\to\infty}\alpha(T)-\big{(}1-\lim_{T\to\infty}\alpha(T)\big{)}\cdot\exp\left(\lim_{T\to\infty}-\frac{LT}{2^{8}f(T)}\right)$
 $=$ $\ 1-0-(1-0)\cdot\exp(-\infty)$
@@
 See 5.4
-Lemma 5.4 is a restatement and simplification of Lemma 3.5 in [33].
-First, [33] parametrize their algorithm by the expected number of queries $\hat{k}$ instead of the query probability $p=\hat{k}/T$.
-Second, [33] include a second parameter $k$, which is the eventual target number of queries for their unconditional query bound.
+Lemma 5.4 is a restatement and simplification of Lemma 3.5 in Russo et al.
+(2024).
+First, Russo et al.
+(2024) parametrize their algorithm by the expected number of queries $\hat{k}$ instead of the query probability $p=\hat{k}/T$.
+Second, Russo et al.
+(2024) include a second parameter $k$, which is the eventual target number of queries for their unconditional query bound.
 In our case, an expected query bound is sufficient, so we simply set $k=\hat{k}$.
-Third, [33] provide a second bound which is tighter for small $k$; that bound is less useful for us so we omit it.
+Third, Russo et al.
+(2024) provide a second bound which is tighter for small $k$; that bound is less useful for us so we omit it.
 Fourth, their number of actions $n$ is equal to $|\tilde{\Pi}|$ in our setting.
-(Their actions correspond to policies in $\tilde{\Pi}$, not our actions in $A$.) Since [33] set $\eta=\max\Big{(}\frac{1}{T}\sqrt{\frac{\hat{k}\log n}{2}},\frac{k\hat{k}}{\sqrt{2}T^{2}}\Big{)}$, we end up with $\eta=\max\Big{(}\sqrt{\frac{p\log|\tilde{\Pi}|}{2T}},\>\frac{p^{2}}{\sqrt{2}}\Big{)}$.
+(Their actions correspond to policies in $\tilde{\Pi}$, not our actions in $\mathcal{Y}$.) Since Russo et al.
+(2024) set $\eta=\max\Big{(}\frac{1}{T}\sqrt{\frac{\hat{k}\log n}{2}},\frac{k\hat{k}}{\sqrt{2}T^{2}}\Big{)}$, we end up with $\eta=\max\Big{(}\sqrt{\frac{p\log|\tilde{\Pi}|}{2T}},\>\frac{p^{2}}{\sqrt{2}}\Big{)}$.
 Algorithm 2 provides precise pseudocode for the HedgeWithQueries algorithm to which Lemma 5.4 refers.
@@
   2. 2.
-    Let $V_{T}=\{t\in[T]:\pi_{t}(s_{t})\neq\pi^{m}(s_{t})\}$ be the set of time steps where Hedge’s proposed action doesn’t match the mentor’s.
-Note that $|V_{T}|$ upper bounds the number of mistakes the algorithm makes (the number of mistakes could be smaller, since the algorithm sometimes queries instead of taking action $\pi_{t}(s_{t})$).
+    Let $M_{T}=\{t\in[T]:\pi_{t}(x_{t})\neq\pi^{m}(x_{t})\}$ be the set of time steps where Hedge’s proposed action doesn’t match the mentor’s.
+Note that $|M_{T}|$ upper bounds the number of mistakes the algorithm makes (the number of mistakes could be smaller, since the algorithm sometimes queries instead of taking action $\pi_{t}(x_{t})$).
   3. 3.
-    For $K\subseteq S$, let $\operatorname*{vol}(K)$ denote the $n$-dimensional Lebesgue measure of $K$.
+    For $X\subseteq\mathcal{X}$, let $\operatorname*{vol}(X)$ denote the $n$-dimensional Lebesgue measure of $X$.
   4. 4.
@@
   5. 5.
-    We will use “Case 1” to refer to finite VC dimension and $\sigma$-smooth $\mathbf{s}$ and “Case 2” to refer to finite Littlestone dimension.
-In Case 1, expectations are over the randomness of both $\mathbf{s}$ and the algorithm, while in Case 2, expectations are over just the randomness of the algorithm.
-When we need to distinction, we use $\operatorname*{\mathbb{E}}_{\mathbf{a}}$ to denote the expectation over randomness of the algorithm and $\operatorname*{\mathbb{E}}_{\mathbf{s}\sim\boldsymbol{\mathcal{D}}}$ to denote the expectation over $\mathbf{s}$.
+    We will use “Case 1” to refer to finite VC dimension and $\sigma$-smooth $\boldsymbol{x}$ and “Case 2” to refer to finite Littlestone dimension.
+In Case 1, expectations are over the randomness of both $\boldsymbol{x}$ and the algorithm, while in Case 2, expectations are over just the randomness of the algorithm.
+When we need to distinguish, we use $\operatorname*{\mathbb{E}}_{\boldsymbol{y}}$ to denote the expectation over randomness of the algorithm and $\operatorname*{\mathbb{E}}_{\boldsymbol{x}\sim\boldsymbol{\mathcal{D}}}$ to denote the expectation over $\boldsymbol{x}$.
 Lemma B.1.
 Under the conditions of Theorem 5.1, Algorithm 1 satisfies
-$\operatorname*{\mathbb{E}}[|V_{T}|]\in O\left(\frac{d}{\sigma}T\varepsilon\log(1/\varepsilon)\log T\right)$
+$\operatorname*{\mathbb{E}}[|M_{T}|]\in O\left(\frac{d}{\sigma}T\varepsilon\log(1/\varepsilon)\log T\right)$
 Proof.
-Define $\ell:[T]\times\tilde{\Pi}\to[0,1]$ by $\ell(t,\pi)=\mathbf{1}(\pi(s_{t})\neq\pi^{m}(s_{t}))$, and let $w^{h}$ and $\pi_{t}^{h}$ denote the values of $w$ and $\pi_{t}$ respectively in HedgeWithQueries, while $w$ and $\pi_{t}$ refer to the variables in Algorithm 1.
+Define $\ell:[T]\times\tilde{\Pi}\to[0,1]$ by $\ell(t,\pi)=\mathbf{1}(\pi(x_{t})\neq\pi^{m}(x_{t}))$, and let $w^{h}$ and $\pi_{t}^{h}$ denote the values of $w$ and $\pi_{t}$ respectively in HedgeWithQueries, while $w$ and $\pi_{t}$ refer to the variables in Algorithm 1.
 Then $w$ and $w^{h}$ evolve in the exact same way, so the distributions of $\pi_{t}$ and $\pi_{t}^{h}$ coincide.
 Thus by Lemma 5.4,
-$\operatorname*{\mathbb{E}}_{\mathbf{a}}\left[\sum_{t=1}^{T}\ell(t,\pi_{t})\right]-\min_{\tilde{\pi}\in\tilde{\Pi}}\sum_{t=1}^{T}\ell(t,\tilde{\pi})\leq\frac{2\log|\tilde{\Pi}|}{p^{2}}=2T\varepsilon\log|\tilde{\Pi}|$
-Since Lemma 5.4 holds for any loss function, the bound above holds for any $\mathbf{s}\in S^{T}$, so the bound also holds in expectation over $\mathbf{s}\sim\boldsymbol{\mathcal{D}}$ (which is needed for Case 1).
-Next, observe that $|V_{T}|=\sum_{t=1}^{T}\mathbf{1}(\pi_{t}(s_{t})\neq\pi^{m}(s_{t}))=\sum_{t=1}^{T}\ell(t,\pi_{t})$, so
-$\operatorname*{\mathbb{E}}_{\mathbf{a}}[|V_{T}|]\leq 2T\varepsilon\log|\tilde{\Pi}|+\min_{\tilde{\pi}\in\tilde{\Pi}}\sum_{t=1}^{T}\mathbf{1}(\tilde{\pi}(s_{t})\neq\pi^{m}(s_{t}))$
+$\operatorname*{\mathbb{E}}_{\boldsymbol{y}}\left[\sum_{t=1}^{T}\ell(t,\pi_{t})\right]-\min_{\tilde{\pi}\in\tilde{\Pi}}\sum_{t=1}^{T}\ell(t,\tilde{\pi})\leq\frac{2\log|\tilde{\Pi}|}{p^{2}}=2T\varepsilon\log|\tilde{\Pi}|$
+Since Lemma 5.4 holds for any loss function, the bound above holds for any $\boldsymbol{x}\in S^{T}$, so the bound also holds in expectation over $\boldsymbol{x}\sim\boldsymbol{\mathcal{D}}$ (which is needed for Case 1).
+Next, observe that $|M_{T}|=\sum_{t=1}^{T}\mathbf{1}(\pi_{t}(x_{t})\neq\pi^{m}(x_{t}))=\sum_{t=1}^{T}\ell(t,\pi_{t})$, so
+$\operatorname*{\mathbb{E}}_{\boldsymbol{y}}[|M_{T}|]\leq 2T\varepsilon\log|\tilde{\Pi}|+\min_{\tilde{\pi}\in\tilde{\Pi}}\sum_{t=1}^{T}\mathbf{1}(\tilde{\pi}(x_{t})\neq\pi^{m}(x_{t}))$
 Case 1: Since $\tilde{\Pi}$ is a smooth $\varepsilon$-cover of $\Pi$, we have
-$\operatorname*{\mathbb{E}}_{\mathbf{s}\sim\boldsymbol{\mathcal{D}}}\left[\min_{\tilde{\pi}\in\tilde{\Pi}}\sum_{t=1}^{T}\mathbf{1}(\tilde{\pi}(s_{t})\neq\pi^{m}(s_{t}))\right]\leq$ $\ \operatorname*{\mathbb{E}}_{\mathbf{s}\sim\boldsymbol{\mathcal{D}}}\left[\sup_{\pi\in\Pi}\min_{\tilde{\pi}\in\tilde{\Pi}}\sum_{t=1}^{T}\mathbf{1}(\tilde{\pi}(s_{t})\neq\pi(s_{t}))\right]$
+$\operatorname*{\mathbb{E}}_{\boldsymbol{x}\sim\boldsymbol{\mathcal{D}}}\left[\min_{\tilde{\pi}\in\tilde{\Pi}}\sum_{t=1}^{T}\mathbf{1}(\tilde{\pi}(x_{t})\neq\pi^{m}(x_{t}))\right]\leq$ $\ \operatorname*{\mathbb{E}}_{\boldsymbol{x}\sim\boldsymbol{\mathcal{D}}}\left[\sup_{\pi\in\Pi}\min_{\tilde{\pi}\in\tilde{\Pi}}\sum_{t=1}^{T}\mathbf{1}(\tilde{\pi}(x_{t})\neq\pi(x_{t}))\right]$
 $\in$ $\ O\left(\frac{1}{\sigma}T\varepsilon\log T\sqrt{d\log(1/\varepsilon)}\right)$
@@
 The last component we need is that $|\tilde{\Pi}|\leq(41/\varepsilon)^{d}$ by construction (and such a $\tilde{\Pi}$ is guaranteed to exist by Lemma 5.1).
-Combining the above inequalities and taking the expectation over $\mathbf{s}\sim\boldsymbol{\mathcal{D}}$ (in addition to the randomness of the algorithm), we get
-$\operatorname*{\mathbb{E}}_{\mathbf{s}\sim\boldsymbol{\mathcal{D}},\,\mathbf{a}}\,[|V_{T}|]\leq$ $\ 2T\varepsilon\log|\tilde{\Pi}|+\operatorname*{\mathbb{E}}_{\mathbf{s}\sim\boldsymbol{\mathcal{D}}}\left[\min_{\tilde{\pi}\in\tilde{\Pi}}\sum_{t=1}^{T}\mathbf{1}(\tilde{\pi}(s_{t})\neq\pi^{m}(s_{t}))\right]$
+Combining the above inequalities and taking the expectation over $\boldsymbol{x}\sim\boldsymbol{\mathcal{D}}$ (in addition to the randomness of the algorithm), we get
+$\operatorname*{\mathbb{E}}_{\boldsymbol{x}\sim\boldsymbol{\mathcal{D}},\,\boldsymbol{y}}\,[|M_{T}|]\leq$ $\ 2T\varepsilon\log|\tilde{\Pi}|+\operatorname*{\mathbb{E}}_{\boldsymbol{x}\sim\boldsymbol{\mathcal{D}}}\left[\min_{\tilde{\pi}\in\tilde{\Pi}}\sum_{t=1}^{T}\mathbf{1}(\tilde{\pi}(x_{t})\neq\pi^{m}(x_{t}))\right]$
 $\leq$ $\ 2dT\varepsilon\log(41/\varepsilon)+O\left(\frac{1}{\sigma}T\varepsilon\log T\sqrt{d\log(1/\varepsilon)}\right)$
 $\in$ $\ O\left(\frac{d}{\sigma}T\varepsilon\log(1/\varepsilon)\log T\right)$
-Case 2: Since $\tilde{\Pi}$ is an adversarial cover of $\Pi$ and $\pi^{m}\in\Pi$, there exists $\tilde{\pi}\in\tilde{\Pi}$ such that $\sum_{t=1}^{T}\mathbf{1}(\tilde{\pi}(s_{t})\neq\pi^{m}(s_{t}))=0$.
+Case 2: Since $\tilde{\Pi}$ is an adversarial cover of $\Pi$ and $\pi^{m}\in\Pi$, there exists $\tilde{\pi}\in\tilde{\Pi}$ such that $\sum_{t=1}^{T}\mathbf{1}(\tilde{\pi}(x_{t})\neq\pi^{m}(x_{t}))=0$.
 Since $|\tilde{\Pi}|\leq(eT/d)^{d}$ (with such a $\tilde{\Pi}$ guaranteed to exist by Lemma 5.2),
-$\operatorname*{\mathbb{E}}_{\mathbf{a}}[|V_{T}|]\leq$ $\ 2T\varepsilon\log|\tilde{\Pi}|+\min_{\tilde{\pi}\in\tilde{\Pi}}\sum_{t=1}^{T}\mathbf{1}(\tilde{\pi}(s_{t})\neq\pi^{m}(s_{t}))$
+$\operatorname*{\mathbb{E}}_{\boldsymbol{y}}[|M_{T}|]\leq$ $\ 2T\varepsilon\log|\tilde{\Pi}|+\min_{\tilde{\pi}\in\tilde{\Pi}}\sum_{t=1}^{T}\mathbf{1}(\tilde{\pi}(x_{t})\neq\pi^{m}(x_{t}))$
 $\leq$ $\ 2T\varepsilon d\ln(eT/d)$
@@
 Lemma B.2.
-For all $t\in[T]$, $\mu(s_{t},a_{t})\geq\mu^{m}(s_{t})-L\varepsilon^{1/n}$.
+For all $t\in[T]$, $\mu_{t}(x_{t},y_{t})\geq\mu_{t}^{m}(x_{t})-L\varepsilon^{1/n}$.
 Proof.
-Fix any $t\in[T]$.
-If $t\in Q_{T}$, then $\mu(s_{t},a_{t})=\mu^{m}(s_{t})$, so assume $t\not\in Q_{T}$.
-Let $(s^{\prime},a^{\prime})=\operatorname*{arg\,min}_{(s,a)\in X_{t}:\pi_{t}(s_{t})=a}||s_{t}-s||$.
-Since $t\not\in Q_{T}$, we must have $||s_{t}-s^{\prime}||\leq\varepsilon^{1/n}$.
-We have $a^{\prime}=\pi^{m}(s^{\prime})$ by construction of $X_{t}$ and $\pi_{t}(s_{t})=a^{\prime}$ by construction of $a^{\prime}$. Combining these with the local generalization assumption, we get
-$\mu(s_{t},a_{t})=\mu(s_{t},\pi_{t}(s_{t}))=\mu(s_{t},\pi^{m}(s^{\prime}))\geq\mu^{m}(s_{t})-L||s_{t}-s^{\prime}||\geq\mu^{m}(s_{t})-L\varepsilon^{1/n}$
+Consider an arbitrary $t\in[T]$.
+If $t\in Q_{T}$, then $\mu_{t}(x_{t},y_{t})=\mu_{t}^{m}(x_{t})$, so assume $t\not\in Q_{T}$.
+Let $(x^{\prime},y^{\prime})=\operatorname*{arg\,min}_{(x,y)\in X_{t}:\pi_{t}(x_{t})=y}||x_{t}-x||$.
+Since $t\not\in Q_{T}$, we must have $||x_{t}-x^{\prime}||\leq\varepsilon^{1/n}$.
+We have $y^{\prime}=\pi^{m}(x^{\prime})$ by construction of $X_{t}$ and $\pi_{t}(x_{t})=y^{\prime}$ by construction of $y^{\prime}$. Combining these with the local generalization assumption, we get
+$\mu_{t}(x_{t},y_{t})=\mu_{t}(x_{t},\pi_{t}(x_{t}))=\mu_{t}(x_{t},\pi^{m}(x^{\prime}))\geq\mu_{t}^{m}(x_{t})-L||x_{t}-x^{\prime}||\geq\mu_{t}^{m}(x_{t})-L\varepsilon^{1/n}$
 as required. ∎
 Lemma B.3.
-Assume $x_{1},\dots,x_{T},y_{1},\dots,y_{T}\in[0,1]$ and $x_{t}\geq y_{t}$ for all $t\in[T]$. Then
-$\prod_{t=1}^{T}x_{t}-\prod_{t=1}^{T}y_{t}\leq\sum_{t=1}^{T}x_{t}-\sum_{t=1}^{T}y_{t}$
+Assume $a_{1},\dots,a_{T},b_{1},\dots,b_{T}\in[0,1]$ and $a_{t}\geq b_{t}$ for all $t\in[T]$. Then
+$\prod_{t=1}^{T}a_{t}-\prod_{t=1}^{T}b_{t}\leq\sum_{t=1}^{T}a_{t}-\sum_{t=1}^{T}b_{t}$
 Proof.
 We proceed by induction on $T$.
-The claim is trivially satisfied for $T=1$, so suppose $T>1$ and assume that $\prod_{t=1}^{T-1}x_{t}-\prod_{t=1}^{T-1}y_{t}\leq\sum_{t=1}^{T-1}x_{t}-\sum_{t=1}^{T-1}y_{t}$.
+The claim is trivially satisfied for $T=1$, so suppose $T>1$ and assume that $\prod_{t=1}^{T-1}a_{t}-\prod_{t=1}^{T-1}b_{t}\leq\sum_{t=1}^{T-1}a_{t}-\sum_{t=1}^{T-1}b_{t}$.
 Then
-$\sum_{t=1}^{T}x_{t}-\sum_{t=1}^{T}y_{t}-\prod_{t=1}^{T}x_{t}+\prod_{t=1}^{T}y_{t}=$ $\ x_{T}\sum_{t=1}^{T-1}x_{t}-y_{T}\sum_{t=1}^{T-1}y_{t}-x_{T}\prod_{t=1}^{T-1}x_{t}+y_{T}\prod_{t=1}^{T-1}y_{t}$
-$=$ $\ x_{T}\left(\sum_{t=1}^{T-1}x_{t}-\prod_{t=1}^{T-1}x_{t}\right)-y_{T}\left(\sum_{t=1}^{T-1}y_{t}-\prod_{t=1}^{T-1}y_{t}\right)$
-Since $T>1$ and $x_{t}\in[0,1]$ for all $t\in[T]$, we have $\sum_{t=1}^{T-1}x_{t}\geq x_{1}\geq\sum_{t=1}^{T-1}x_{t}$.
-Thus $\sum_{t=1}^{T-1}x_{t}-\prod_{t=1}^{T-1}x_{t}\geq 0$.
-Combining this with $x_{T}\geq y_{T}$, we get
-$\sum_{t=1}^{T}x_{t}-\sum_{t=1}^{T}y_{t}-\prod_{t=1}^{T}x_{t}+\prod_{t=1}^{T}y_{t}=$ $\ x_{T}\left(\sum_{t=1}^{T-1}x_{t}-\prod_{t=1}^{T-1}x_{t}\right)-y_{T}\left(\sum_{t=1}^{T-1}y_{t}-\prod_{t=1}^{T-1}y_{t}\right)$
-$\geq$ $\ y_{T}\left(\sum_{t=1}^{T-1}x_{t}-\prod_{t=1}^{T-1}x_{t}\right)-y_{T}\left(\sum_{t=1}^{T-1}y_{t}-\prod_{t=1}^{T-1}y_{t}\right)$
-$=$ $\ y_{T}\left(\sum_{t=1}^{T-1}x_{t}-\prod_{t=1}^{T-1}x_{t}-\sum_{t=1}^{T-1}y_{t}+\prod_{t=1}^{T-1}y_{t}\right)$
+$\sum_{t=1}^{T}a_{t}-\sum_{t=1}^{T}b_{t}-\prod_{t=1}^{T}a_{t}+\prod_{t=1}^{T}b_{t}=$ $\ a_{T}\sum_{t=1}^{T-1}a_{t}-b_{T}\sum_{t=1}^{T-1}b_{t}-a_{T}\prod_{t=1}^{T-1}a_{t}+b_{T}\prod_{t=1}^{T-1}b_{t}$
+$=$ $\ a_{T}\left(\sum_{t=1}^{T-1}a_{t}-\prod_{t=1}^{T-1}a_{t}\right)-b_{T}\left(\sum_{t=1}^{T-1}b_{t}-\prod_{t=1}^{T-1}b_{t}\right)$
+Since $T>1$ and $a_{t}\in[0,1]$ for all $t\in[T]$, we have $\sum_{t=1}^{T-1}a_{t}\geq a_{1}\geq\sum_{t=1}^{T-1}a_{t}$.
+Thus $\sum_{t=1}^{T-1}a_{t}-\prod_{t=1}^{T-1}a_{t}\geq 0$.
+Combining this with $a_{T}\geq b_{T}$, we get
+$\sum_{t=1}^{T}a_{t}-\sum_{t=1}^{T}b_{t}-\prod_{t=1}^{T}a_{t}+\prod_{t=1}^{T}b_{t}=$ $\ a_{T}\left(\sum_{t=1}^{T-1}a_{t}-\prod_{t=1}^{T-1}a_{t}\right)-b_{T}\left(\sum_{t=1}^{T-1}b_{t}-\prod_{t=1}^{T-1}b_{t}\right)$
+$\geq$ $\ b_{T}\left(\sum_{t=1}^{T-1}a_{t}-\prod_{t=1}^{T-1}a_{t}\right)-b_{T}\left(\sum_{t=1}^{T-1}b_{t}-\prod_{t=1}^{T-1}b_{t}\right)$
+$=$ $\ b_{T}\left(\sum_{t=1}^{T-1}a_{t}-\prod_{t=1}^{T-1}a_{t}-\sum_{t=1}^{T-1}b_{t}+\prod_{t=1}^{T-1}b_{t}\right)$
 […]
-The last step is due to $y_{T}\geq 0$ and our assumption of $\prod_{t=1}^{T-1}x_{t}-\prod_{t=1}^{T-1}y_{t}\leq\sum_{t=1}^{T-1}x_{t}-\sum_{t=1}^{T-1}y_{t}$. ∎
+The last step is due to $b_{T}\geq 0$ and our assumption of $\prod_{t=1}^{T-1}a_{t}-\prod_{t=1}^{T-1}b_{t}\leq\sum_{t=1}^{T-1}a_{t}-\sum_{t=1}^{T-1}b_{t}$. ∎
 Lemma B.4.
@@
 $\operatorname*{\mathbb{E}}\left[R_{T}\right]\in$ $\ O\left(\frac{dL}{\sigma}T\varepsilon^{1+1/n}\log(1/\varepsilon)\log T\right)$
+$\operatorname*{\mathbb{E}}\left[\sum_{t=1}^{T}\mu_{t}^{m}(x_{t})-\sum_{t=1}^{T}\mu_{t}(x_{t},y_{t})\right]\in$ $\ O\left(\frac{dL}{\sigma}T\varepsilon^{1+1/n}\log(1/\varepsilon)\log T\right)$
 Proof.
-We first claim that $a_{t}=\pi^{m}(s_{t})$ for all $t\not\in V_{T}$.
-If $t\in Q_{T}$, the claim is immediate; if not, we have $a_{t}=\pi_{t}(s_{t})$, and $\pi_{t}(s_{t})=\pi^{m}(s_{t})$ due to $t\not\in V_{T}$.
-Thus $\min(\mu^{m}(s_{t}),\mu(s_{t},a_{t}))=\mu^{m}(s_{t})$ for $t\not\in V_{T}$.
-We next claim that $\mu^{m}(s_{t})-\min(\mu^{m}(s_{t}),\mu(s_{t},a_{t}))\leq L\varepsilon^{1/n}$ for all $t\in[T]$.
-If $\mu(s_{t},a_{t})\leq\mu^{m}(s_{t})$, this follows from Lemma B.2.
-If $\mu(s_{t},a_{t})>\mu^{m}(s_{t})$, then $\mu^{m}(s_{t})-\min(\mu^{m}(s_{t}),\mu(s_{t},a_{t}))=0\leq L\varepsilon^{1/n}$.
+We first claim that $y_{t}=\pi^{m}(x_{t})$ for all $t\not\in M_{T}$.
+If $t\in Q_{T}$, the claim is immediate; if not, we have $y_{t}=\pi_{t}(x_{t})$, and $\pi_{t}(x_{t})=\pi^{m}(x_{t})$ due to $t\not\in M_{T}$.
+Thus $\min(\mu_{t}^{m}(x_{t}),\mu_{t}(x_{t},y_{t}))=\mu_{t}^{m}(x_{t})$ for $t\not\in M_{T}$.
+We next claim that $\mu_{t}^{m}(x_{t})-\min(\mu_{t}^{m}(x_{t}),\mu_{t}(x_{t},y_{t}))\leq L\varepsilon^{1/n}$ for all $t\in[T]$.
+If $\mu_{t}(x_{t},y_{t})\leq\mu_{t}^{m}(x_{t})$, this follows from Lemma B.2.
+If $\mu_{t}(x_{t},y_{t})>\mu_{t}^{m}(x_{t})$, then $\mu_{t}^{m}(x_{t})-\min(\mu_{t}^{m}(x_{t}),\mu_{t}(x_{t},y_{t}))=0\leq L\varepsilon^{1/n}$.
 Therefore
-$\sum_{t=1}^{T}\Big{(}\mu^{m}(s_{t})-\min(\mu^{m}(s_{t}),\mu(s_{t},a_{t}))\Big{)}\leq$ $\ \sum_{t\in V_{T}}\Big{(}\mu^{m}(s_{t})-\min(\mu^{m}(s_{t}),\mu(s_{t},a_{t}))\Big{)}$
-$\leq$ $\ \sum_{t\in V_{T}}L\varepsilon^{1/n}$
-$=$ $\ |V_{T}|L\varepsilon^{1/n}$
-Now let $x_{t}=\mu^{m}(s_{t})$ and $y_{t}=\min(\mu^{m}(s_{t}),\mu(s_{t},a_{t}))$ for all $t\in[T]$. Then by Lemma B.3,
-$\prod_{t=1}^{T}\mu^{m}(s_{t})-\prod_{t=1}^{T}\min(\mu^{m}(s_{t}),\mu(s_{t},a_{t}))\leq$ $\ \sum_{t=1}^{T}\Big{(}\mu^{m}(s_{t})-\min(\mu^{m}(s_{t}),\mu(s_{t},a_{t}))\Big{)}$
-Since $\mu(s_{t},a_{t})\geq\min(\mu^{m}(s_{t}),\mu(s_{t},a_{t}))$ for all $t\in[T]$, we have
-$R_{T}=$ $\ \prod_{t=1}^{T}\mu^{m}(s_{t})-\prod_{t=1}^{T}\mu(s_{t},a_{t})$
-$\leq$ $\ \sum_{t=1}^{T}\Big{(}\mu^{m}(s_{t})-\min(\mu^{m}(s_{t}),\mu(s_{t},a_{t}))\Big{)}$
-$\leq$ $\ |V_{T}|L\varepsilon^{1/n}$
-Taking the expectation and applying Lemma B.1 completes the proof. ∎
+$\sum_{t=1}^{T}\Big{(}\mu_{t}^{m}(x_{t})-\min(\mu_{t}^{m}(x_{t}),\mu_{t}(x_{t},y_{t}))\Big{)}\leq$ $\ \sum_{t\in M_{T}}\Big{(}\mu_{t}^{m}(x_{t})-\min(\mu_{t}^{m}(x_{t}),\mu_{t}(x_{t},y_{t}))\Big{)}$
+$\leq$ $\ \sum_{t\in M_{T}}L\varepsilon^{1/n}$
+$=$ $\ |M_{T}|L\varepsilon^{1/n}$
+Now let $a_{t}=\mu_{t}^{m}(x_{t})$ and $b_{t}=\min(\mu_{t}^{m}(x_{t}),\mu_{t}(x_{t},y_{t}))$ for all $t\in[T]$. Then by Lemma B.3,
+$\prod_{t=1}^{T}\mu_{t}^{m}(x_{t})-\prod_{t=1}^{T}\min(\mu_{t}^{m}(x_{t}),\mu_{t}(x_{t},y_{t}))\leq$ $\ \sum_{t=1}^{T}\Big{(}\mu_{t}^{m}(x_{t})-\min(\mu_{t}^{m}(x_{t}),\mu_{t}(x_{t},y_{t}))\Big{)}$
+Since $\mu_{t}(x_{t},y_{t})\geq\min(\mu_{t}^{m}(x_{t}),\mu_{t}(x_{t},y_{t}))$ for all $t\in[T]$, we have
+$R_{T}=$ $\ \prod_{t=1}^{T}\mu_{t}^{m}(x_{t})-\prod_{t=1}^{T}\mu_{t}(x_{t},y_{t})$
+$\leq$ $\ \sum_{t=1}^{T}\Big{(}\mu_{t}^{m}(x_{t})-\min(\mu_{t}^{m}(x_{t}),\mu_{t}(x_{t},y_{t}))\Big{)}$
+$\leq$ $\ |M_{T}|L\varepsilon^{1/n}$
+Since we also have $\sum_{t=1}^{T}\mu_{t}^{m}(x_{t})-\sum_{t=1}^{T}\mu_{t}(x_{t},y_{t})\leq\sum_{t=1}^{T}(\mu_{t}^{m}(x_{t})-\min(\mu_{t}^{m}(x_{t}),\mu_{t}(x_{t},y_{t})))$, we get both of the following
+$R_{T}\leq$ $\ L\varepsilon^{1/n}|M_{T}|$
+$\sum_{t=1}^{T}\mu_{t}^{m}(x_{t})-\sum_{t=1}^{T}\mu_{t}(x_{t},y_{t})\leq$ $\ L\varepsilon^{1/n}|M_{T}|$
+Taking the expectation and applying Lemma B.1 to both of the above inequalities completes the proof. ∎
 Definition B.1.
 Let $(K,||\cdot||)$ be a normed vector space and let $\delta>0$.
-Then $W\subseteq K$ is a $\delta$-packing of $K$ if for all $x,y\in W$, $||x-y||>\delta$.
+Then $X\subseteq K$ is a $\delta$-packing of $K$ if for all $x,y\in X$, $||x-y||>\delta$.
 The $\delta$-packing number of $K$, denoted $\mathcal{M}(K,||\cdot||,\delta)$, is the maximum cardinality of any $\delta$-packing of $K$.
 In this paper, we only consider the Euclidean distance norm, so we just write $M(K,||\cdot||,\delta)=M(K,\delta)$.
-Lemma B.5 (Theorem 14.2 in [40]).
+Lemma B.5 (Theorem 14.2 in Wu (2020)).
 If $K\subset\mathbb{R}^{n}$ is convex, bounded, and contains a ball with radius $\delta>0$, then
@@
 where $B$ is a unit ball.
-Lemma B.6 (Jung’s Theorem [22]).
+Lemma B.6 (Jung’s Theorem (Jung, 1901)).
 […]
@@
 Under the conditions of Theorem 5.1, Algorithm 1 satisfies
-$\operatorname*{\mathbb{E}}[|Q_{T}|]\in O\left(\sqrt{\frac{T}{\varepsilon}}+\frac{d}{\sigma}T\varepsilon\log(1/\varepsilon)\log T+\frac{\operatorname*{diam}(\mathbf{s})^{n}}{\varepsilon}\right)$
+$\operatorname*{\mathbb{E}}[|Q_{T}|]\in O\left(\sqrt{\frac{T}{\varepsilon}}+\frac{d}{\sigma}T\varepsilon\log(1/\varepsilon)\log T+\frac{\operatorname*{diam}(\boldsymbol{x})^{n}}{\varepsilon}\right)$
 Proof.
 […]
-Next, fix an $a\in A$ and let $S_{a}=\{s\in\mathbf{s}:\pi^{m}(s)=a\}$ be the set of observed states which share a mentor action.
-We claim that $\hat{S}_{2}\cap S_{a}$ is a packing of $S_{a}$.
-Suppose instead that there exists $s,s^{\prime}\in\hat{S}_{2}\cap S_{a}$, with $||s-s^{\prime}||\leq\varepsilon^{1/n}$.
-WLOG assume $s$ was queried after $s^{\prime}$ and let $t$ be the time step on which $s$ was queried.
-Then $(s^{\prime},\pi^{m}(s^{\prime}))\in X_{t}$.
-Also, $s,s^{\prime}\in\hat{S}_{2}\cap S_{a}$ implies that and $\pi_{t}(s_{t})=\pi^{m}(s_{t})=a=\pi^{m}(s^{\prime})$.
+Next, fix an $y\in\mathcal{Y}$ and let $X_{y}=\{x\in\boldsymbol{x}:\pi^{m}(x)=y\}$ be the set of observed inputs which share a mentor action.
+We claim that $\hat{X}_{2}\cap X_{y}$ is a packing of $X_{y}$.
+Suppose instead that there exists $x,x^{\prime}\in\hat{X}_{2}\cap X_{y}$, with $||x-x^{\prime}||\leq\varepsilon^{1/n}$.
+WLOG assume $x$ was queried after $x^{\prime}$ and let $t$ be the time step on which $x$ was queried.
+Then $(x^{\prime},\pi^{m}(x^{\prime}))\in X_{t}$.
+Also, $x,x^{\prime}\in\hat{X}_{2}\cap X_{y}$ implies that and $\pi_{t}(x_{t})=\pi^{m}(x_{t})=y=\pi^{m}(x^{\prime})$.
 Therefore
-$\min_{(s^{\prime\prime},a^{\prime\prime})\in X_{t}:a^{\prime\prime}=\pi_{t}(s_{t})}||s_{t}-s^{\prime\prime}||\leq||s_{t}-s^{\prime}||\leq\varepsilon^{1/n}$
-which contradicts $s_{t}\in\hat{S}$. Thus $\hat{S}_{2}\cap S_{a}$ is a $\varepsilon^{1/n}$-packing of $S_{a}$.
-By Lemma B.6, there exists a ball $B_{1}$ of diameter $\operatorname*{diam}(\mathbf{s})\sqrt{\frac{n}{2(n+1)}}$ which contains $\mathbf{s}$.
-Let $R=\operatorname*{diam}(\mathbf{s})\sqrt{\frac{n}{8(n+1)}}$ denote the radius of $B_{1}$.
+$\min_{(x^{\prime\prime},y^{\prime\prime})\in X_{t}:y^{\prime\prime}=\pi_{t}(x_{t})}||x_{t}-x^{\prime\prime}||\leq||x_{t}-x^{\prime}||\leq\varepsilon^{1/n}$
+which contradicts $x_{t}\in\hat{X}$. Thus $\hat{X}_{2}\cap X_{y}$ is a $\varepsilon^{1/n}$-packing of $X_{y}$.
+By Lemma B.6, there exists a ball $B_{1}$ of diameter $\operatorname*{diam}(\boldsymbol{x})\sqrt{\frac{n}{2(n+1)}}$ which contains $\boldsymbol{x}$.
+Let $R=\operatorname*{diam}(\boldsymbol{x})\sqrt{\frac{n}{8(n+1)}}$ denote the radius of $B_{1}$.
 Let $B_{2}$ be the ball with the same center as $B_{1}$ but with radius $\max(R,\varepsilon^{1/n})$.
-Since $S_{a}\subset\mathbf{s}\subset B_{1}\subset B_{2}$, $\hat{S}_{2}\cap S_{a}$ is also a $\varepsilon^{1/n}$-packing of $B_{2}$.
+Since $X_{y}\subset\boldsymbol{x}\subset B_{1}\subset B_{2}$, $\hat{X}_{2}\cap X_{y}$ is also a $\varepsilon^{1/n}$-packing of $B_{2}$.
 Also, $B_{2}$ must contain a ball of radius $\varepsilon^{1/n}$, so Lemma B.5 implies that
-$|\hat{S}_{2}\cap S_{a}|\leq$ $\ \mathcal{M}(B_{2},\varepsilon^{1/n})$
+$|\hat{X}_{2}\cap X_{y}|\leq$ $\ \mathcal{M}(B_{2},\varepsilon^{1/n})$
 $\leq$ $\ \frac{3^{n}\operatorname*{vol}(B_{2})}{\varepsilon\operatorname*{vol}(B)}$
 $=$ $\ \big{(}\max(R,\varepsilon^{1/n})\big{)}^{n}\,\frac{3^{n}\operatorname*{vol}(B)}{\varepsilon\operatorname*{vol}(B)}$
-$=$ $\ \max\left(\operatorname*{diam}(\mathbf{s})^{n}\left(\frac{n}{8(n+1)}\right)^{n/2},\>\varepsilon\right)\frac{3^{n}}{\varepsilon}$
-$\leq$ $\ O\left(\frac{\operatorname*{diam}(\mathbf{s})^{n}}{\varepsilon}+1\right)$
-(The $+1$ is necessary for now since $\operatorname*{diam}(\mathbf{s})$ could theoretically be zero.) Therefore
-$\operatorname*{\mathbb{E}}[|Q_{T}|]=$ $\ \sqrt{\frac{T}{\varepsilon}}+\operatorname*{\mathbb{E}}[|\hat{S}|]$
-$=$ $\ \sqrt{\frac{T}{\varepsilon}}+\operatorname*{\mathbb{E}}[|\hat{S}_{1}|]+\operatorname*{\mathbb{E}}\left[\sum_{a\in A}|\hat{S}_{2}\cap S_{a}|\right]$
-$\leq$ $\ \sqrt{\frac{T}{\varepsilon}}+O\left(\frac{d}{\sigma}T\varepsilon\log(1/\varepsilon)\log T\right)+\sum_{a\in A}O\left(\frac{\operatorname*{diam}(\mathbf{s})^{n}}{\varepsilon}+1\right)$
-$\leq$ $\ \sqrt{\frac{T}{\varepsilon}}+O\left(\frac{d}{\sigma}T\varepsilon\log(1/\varepsilon)\log T\right)+|A|\cdot O\left(\frac{\operatorname*{diam}(\mathbf{s})^{n}}{\varepsilon}+1\right)$
-$\leq$ $\ O\left(\sqrt{\frac{T}{\varepsilon}}+\frac{d}{\sigma}T\varepsilon\log(1/\varepsilon)\log T+\frac{\operatorname*{diam}(\mathbf{s})^{n}}{\varepsilon}\right)$
+$=$ $\ \max\left(\operatorname*{diam}(\boldsymbol{x})^{n}\left(\frac{n}{8(n+1)}\right)^{n/2},\>\varepsilon\right)\frac{3^{n}}{\varepsilon}$
+$\leq$ $\ O\left(\frac{\operatorname*{diam}(\boldsymbol{x})^{n}}{\varepsilon}+1\right)$
+(The $+1$ is necessary for now since $\operatorname*{diam}(\boldsymbol{x})$ could theoretically be zero.) Therefore
+$\operatorname*{\mathbb{E}}[|Q_{T}|]=$ $\ \sqrt{\frac{T}{\varepsilon}}+\operatorname*{\mathbb{E}}[|\hat{X}|]$
+$=$ $\ \sqrt{\frac{T}{\varepsilon}}+\operatorname*{\mathbb{E}}[|\hat{X}_{1}|]+\operatorname*{\mathbb{E}}\left[\sum_{y\in\mathcal{Y}}|\hat{X}_{2}\cap X_{y}|\right]$
+$\leq$ $\ \sqrt{\frac{T}{\varepsilon}}+O\left(\frac{d}{\sigma}T\varepsilon\log(1/\varepsilon)\log T\right)+\sum_{y\in\mathcal{Y}}O\left(\frac{\operatorname*{diam}(\boldsymbol{x})^{n}}{\varepsilon}+1\right)$
+$\leq$ $\ \sqrt{\frac{T}{\varepsilon}}+O\left(\frac{d}{\sigma}T\varepsilon\log(1/\varepsilon)\log T\right)+|\mathcal{Y}|\cdot O\left(\frac{\operatorname*{diam}(\boldsymbol{x})^{n}}{\varepsilon}+1\right)$
+$\leq$ $\ O\left(\sqrt{\frac{T}{\varepsilon}}+\frac{d}{\sigma}T\varepsilon\log(1/\varepsilon)\log T+\frac{\operatorname*{diam}(\boldsymbol{x})^{n}}{\varepsilon}\right)$
 as required. ∎
@@
 and
-$\operatorname*{\mathbb{E}}[|Q_{T}|]\in$ $\ O\left(\sqrt{T^{1+\frac{2n}{2n+1}}}+\frac{d}{\sigma}T^{1-\frac{-2n}{2n+1}}\log(T^{\frac{2n}{2n+1}})\log T+T^{\frac{2n}{2n+1}}\operatorname*{diam}(\mathbf{s})^{n}\right)$
-$=$ $\ O\left(T^{\frac{2n+0.5}{2n+1}}+\frac{d}{\sigma}T^{\frac{1}{2n+1}}\log T+T^{\frac{2n}{2n+1}}\operatorname*{diam}(\mathbf{s})^{n}\right)$
-$\leq$ $\ O\left(T^{\frac{4n+1}{4n+2}}\Big{(}\frac{d}{\sigma}\log T+\operatorname*{diam}(\mathbf{s})^{n}\Big{)}\right)$
-See 5.2.1
-Proof.
-Let $\xi$ be the event that Algorithm 1 exceeds $T^{\frac{4n+1.5}{4n+2}}(\frac{d}{\sigma}\log T+\operatorname*{diam}(\mathbf{s})^{n})$ queries. Noting that $|Q_{T}|\leq T$ always, we have
-$\operatorname*{\mathbb{E}}\big{[}|Q_{T}|\big{]}=$ $\ \Pr[\xi]\,\operatorname*{\mathbb{E}}\big{[}|Q_{T}|\mid\xi\big{]}+\Pr[\neg\xi]\,\operatorname*{\mathbb{E}}\big{[}|Q_{T}|\mid\neg\xi\big{]}$
-$\geq$ $\ \Pr[\xi]\,T^{\frac{4n+1.5}{4n+2}}\left(\frac{d}{\sigma}\log T+\operatorname*{diam}(\mathbf{s})^{n}\right)$
-Therefore by Theorem 5.2,
-$\Pr[\xi]\leq\frac{O\left(T^{\frac{4n+1}{4n+2}}\left(\frac{d}{\sigma}\log T+\operatorname*{diam}(\mathbf{s})^{n}\right)\right)}{T^{\frac{4n+1.5}{4n+2}}\left(\frac{d}{\sigma}\log T+\operatorname*{diam}(\mathbf{s})^{n}\right)}\leq O\left(T^{\frac{-0.5}{4n+2}}\right)$
-If $\xi$ does not occur, then the regret bound from Theorem 5.2 holds. If not, we still have $R_{T}\leq 1$ unconditionally. Thus the regret of the modified version of Algorithm 1 satisfies
-$\operatorname*{\mathbb{E}}[R_{T}]\leq$ $\ \Pr[\xi]\,\operatorname*{\mathbb{E}}[R_{T}\mid\xi]+\Pr[\neg\xi]\,\operatorname*{\mathbb{E}}[R_{T}\mid\neg\xi]$
-$\leq$ $\ O\left(T^{\frac{-0.5}{4n+2}}\right)\cdot 1+1\cdot O\left(\frac{dL}{\sigma}T^{\frac{-1}{2n+1}}\log T\right)$
-$\leq$ $\ O\left(\frac{dL}{\sigma}T^{\frac{-1}{8n+4}}\log T\right)$
+$\operatorname*{\mathbb{E}}[|Q_{T}|]\in$ $\ O\left(\sqrt{T^{1+\frac{2n}{2n+1}}}+\frac{d}{\sigma}T^{1-\frac{-2n}{2n+1}}\log(T^{\frac{2n}{2n+1}})\log T+T^{\frac{2n}{2n+1}}\operatorname*{diam}(\boldsymbol{x})^{n}\right)$
+$=$ $\ O\left(T^{\frac{2n+0.5}{2n+1}}+\frac{d}{\sigma}T^{\frac{1}{2n+1}}\log T+T^{\frac{2n}{2n+1}}\operatorname*{diam}(\boldsymbol{x})^{n}\right)$
+$\leq$ $\ O\left(T^{\frac{4n+1}{4n+2}}\Big{(}\frac{d}{\sigma}\log T+\operatorname*{diam}(\boldsymbol{x})^{n}\Big{)}\right)$
+If we instead use the second bound from Lemma B.4, the same arithmetic gives us:
+See 5.3
 B.3 Adaptive adversaries
@@
 For randomized algorithms, an adaptive adversary can base the choice of $s_{t}$ on the results of randomization on previous time steps (but not on the current time step), while an oblivious adversary cannot.
-In the standard online learning model, Hedge guarantees sublinear regret against both oblivious and adaptive adversaries (Chapter 5 of [35] or Chapter 21 of [34]).
-However, [33] state their result only for oblivious adversaries.
-In order for our overall proof of Theorem 5.1 to hold for adaptive adversaries, Lemma 5.4 (Lemma 3.5 in [33]) must also hold for adaptive adversaries.
+In the standard online learning model, Hedge guarantees sublinear regret against both oblivious and adaptive adversaries (Chapter 5 of Slivkins et al.
+(2019) or Chapter 21 of Shalev-Shwartz & Ben-David (2014)).
+However, Russo et al.
+(2024) state their result only for oblivious adversaries.
+In order for our overall proof of Theorem 5.1 to hold for adaptive adversaries, Lemma 5.4 (Lemma 3.5 in Russo et al.
+(2024)) must also hold for adaptive adversaries.
 In this section, we argue why the proof of Lemma 5.4 (Lemma 3.5 in their paper) goes through for adaptive adversaries as well.
-For this rest of Section B.3, lemma numbers refer to the numbering in [33].
+For this rest of Section B.3, lemma numbers refer to the numbering in Russo et al.
+(2024).
 The importance of independent queries.
-Recall from Section B.1 that [33] allow two separate parameters $k$ and $\hat{k}$, which we unify for simplicity.
+Recall from Section B.1 that Russo et al.
+(2024) allow two separate parameters $k$ and $\hat{k}$, which we unify for simplicity.
 Recall also that Lemma 3.5 refers to the variant of Hedge which queries with probability $p=\hat{k}/T=k/T$ independently on each time step (Algorithm 2.
@@
 This insight immediately extends Observation 3.3 to adaptive adversaries (with the minor modification that queries are now issued independently with probability $p$ on each time step instead of issuing $k$ uniformly distributed queries).
-Specifically, using the notation from [33] where $i_{t}$ is the action chosen at time $t$, $i_{t}^{0}$ is the action chosen at time $t$ if a query is not issued, and $i_{t}^{*}$ is the optimal action at time $t$, we have
+Specifically, using the notation from Russo et al.
+(2024) where $i_{t}$ is the action chosen at time $t$, $i_{t}^{0}$ is the action chosen at time $t$ if a query is not issued, and $i_{t}^{*}$ is the optimal action at time $t$, we have
 $\operatorname*{\mathbb{E}}[\ell_{t}(i_{t})]=(1-p)\operatorname*{\mathbb{E}}[\ell_{t}(i_{t}^{0})]+p\operatorname*{\mathbb{E}}[\ell_{t}(i_{t}^{*})]=\left(1-\frac{k}{T}\right)\operatorname*{\mathbb{E}}[\ell_{t}(i_{t}^{0})]+\frac{k}{T}\operatorname*{\mathbb{E}}[\ell_{t}(i_{t}^{*})]$
@@
 Appendix C Generalizing Theorem 5.2 to many actions
-We use the standard “one versus rest” reduction (see, e.g., Chapter 29 of [34]).
-For each action $a$, we will learn a binary classifier which predicts whether action $a$ is the mentor’s action.
-Formally, for each $a\in A$, define the policy class $\Pi_{a}=\{\pi_{a}:\pi\in\Pi\text{ and }\pi_{a}(s)=\mathbf{1}(\pi(s)=a))\ \forall s\in S\}$.
-Informally, for each policy $\pi:S\to A$ in $\Pi$, there exists a policy $\pi_{a}:S\to\{0,1\}$ in $\Pi_{a}$ such that $\pi_{a}(s)=\mathbf{1}(\pi(s)=a)$ for all $s\in S$.
-Algorithm 3 runs one copy of our binary-action algorithm Algorithm 1 for each action $a\in A$.
-At each time step $t$, the copy for action $a$ returns an action $b_{t}^{a}$, with $b_{t}^{a}=1$ indicating a belief that $a=\pi^{m}(s_{t})$ and $b_{t}^{a}=0$ indicating a belief that $a\neq\pi^{m}(s_{t})$.
-(Note that $b_{t}^{a}=\hat{a}$ is also possible, indicating that the mentor was queried.)
-The key idea is that if $b_{t}^{a}$ is correct for each action $a$, there will be exactly one $a$ such that $b_{t}^{a}=1$, and specifically it will be $a=\pi^{m}(s_{t})$.
+We use the standard “one versus rest” reduction (see, e.g., Chapter 29 of Shalev-Shwartz & Ben-David (2014)).
+For each action $y$, we will learn a binary classifier which predicts whether action $y$ is the mentor’s action.
+Formally, for each $y\in\mathcal{Y}$, define the policy class $\Pi_{y}=\{\pi_{y}:\pi\in\Pi\text{ and }\pi_{y}(x)=\mathbf{1}(\pi(x)=y))\ \forall x\in\mathcal{X}\}$.
+In words, for each policy $\pi:\mathcal{X}\to\mathcal{Y}$ in $\Pi$, there exists a policy $\pi_{y}:\mathcal{X}\to\{0,1\}$ in $\Pi_{y}$ such that $\pi_{y}(x)=\mathbf{1}(\pi(x)=y)$ for all $x\in\mathcal{X}$.
+Algorithm 3 runs one copy of our binary-action algorithm Algorithm 1 for each action $y\in\mathcal{Y}$.
+At each time step $t$, the copy for action $y$ returns an action $b_{t}^{y}$, with $b_{t}^{y}=1$ indicating a belief that $y=\pi^{m}(x_{t})$ and $b_{t}^{y}=0$ indicating a belief that $y\neq\pi^{m}(x_{t})$.
+(Note that $b_{t}^{y}=\tilde{y}$ is also possible, indicating that the mentor was queried.)
+The key idea is that if $b_{t}^{y}$ is correct for each action $y$, there will be exactly one $y$ such that $b_{t}^{y}=1$, and specifically it will be $y=\pi^{m}(x_{t})$.
 Thus we are guaranteed to take the mentor’s action on such time steps.
-The analysis for Theorem 5.2 (specifically, Lemma B.1) bounds the number of time steps when a given copy of Algorithm 1 is incorrect, so by the union bound, the number of time steps where any copy is incorrect is $|A|$ times that bound.
+The analysis for Theorem 5.2 (specifically, Lemma B.1) bounds the number of time steps when a given copy of Algorithm 1 is incorrect, so by the union bound, the number of time steps where any copy is incorrect is $|\mathcal{Y}|$ times that bound.
 That in turn bounds the number of time steps where Algorithm 3 takes an action other than the mentor’s.
-Similarly, the number of queries made by Algorithm 3 is at most $|A|$ times the bound from Theorem 5.2.
+Similarly, the number of queries made by Algorithm 3 is at most $|\mathcal{Y}|$ times the bound from Theorem 5.2.
 The result is the following theorem:
 1:function AvoidCatastropheManyActions($T\in\mathbb{N},\>\varepsilon\in\mathbb{R}_{>0},\>d\in\mathbb{N},\,$ policy class $\Pi$)
-2:     for $a\in A$ do
+2:     for $y\in\mathcal{Y}$ do
 3:         if $\Pi$ has VC dimension $d$ then
-4:              $\tilde{\Pi}_{a}\leftarrow$ any smooth $\varepsilon$-cover of $\Pi$ of size at most $(41/\varepsilon)^{d}$
+4:              $\tilde{\Pi}_{y}\leftarrow$ any smooth $\varepsilon$-cover of $\Pi$ of size at most $(41/\varepsilon)^{d}$
 5:         else if $\Pi$ has Littlestone dimension $d$ then
-6:              $\tilde{\Pi}_{a}\leftarrow$ any adversarial $\varepsilon$-cover of size at most $(eT/d)^{d}$
+6:              $\tilde{\Pi}_{y}\leftarrow$ any adversarial $\varepsilon$-cover of size at most $(eT/d)^{d}$
 7:     for $t$ from $1$ to $T$ do
-8:         for $a\in A$ do
-9:              $b_{t}^{a}\leftarrow$ action from running one step of Algorithm 1 on $\Pi_{a}$ (with the same $T,\varepsilon,d$)
-10:         if $b_{t}^{a}\neq\hat{a}$ $\forall a\in A$ and $\exists a\in A:b_{t}^{a}=1$ then
-11:              Take any action $a$ with $b_{t}^{a}=1$
+8:         for $y\in\mathcal{Y}$ do
+9:              $b_{t}^{y}\leftarrow$ action from running one step of Algorithm 1 on $\Pi_{y}$ (with the same $T,\varepsilon,d$)
+10:         if $b_{t}^{y}\neq\tilde{y}$ $\forall y\in\mathcal{Y}$ and $\exists a\in\mathcal{Y}:b_{t}^{y}=1$ then
+11:              Take any action $y$ with $b_{t}^{y}=1$
 12:         else
@@
 Theorem C.1.
-Assume $\pi^{m}\in\Pi$ where either (1) $\Pi_{a}$ has finite VC dimension $d$ and $\mathbf{s}$ is $\sigma$-smooth or (2) $\Pi_{a}$ has finite Littlestone dimension $d$ for all $a\in A$.
+Assume $\pi^{m}\in\Pi$ where either (1) $\Pi_{y}$ has finite VC dimension $d$ and $\boldsymbol{x}$ is $\sigma$-smooth or (2) $\Pi_{y}$ has finite Littlestone dimension $d$ for all $y\in\mathcal{Y}$.
 Then for any $T\in\mathbb{N}$, Algorithm 3 with $T$ and $\varepsilon=T^{\frac{-2n}{2n+1}}$ satisfies
-$\operatorname*{\mathbb{E}}\left[R_{T}\right]\in$ $\ O\left(\frac{|A|dL}{\sigma}T^{\frac{-1}{2n+1}}\log T\right)$
-$\operatorname*{\mathbb{E}}[|Q_{T}|]\in$ $\ O\left(|A|T^{\frac{4n+1}{4n+2}}\left(\frac{d}{\sigma}\log T+\operatorname*{diam}(\mathbf{s})^{n}\right)\right)$
+$\operatorname*{\mathbb{E}}\left[R_{T}\right]\in$ $\ O\left(\frac{|\mathcal{Y}|dL}{\sigma}T^{\frac{-1}{2n+1}}\log T\right)$
+$\operatorname*{\mathbb{E}}[|Q_{T}|]\in$ $\ O\left(|\mathcal{Y}|T^{\frac{4n+1}{4n+2}}\left(\frac{d}{\sigma}\log T+\operatorname*{diam}(\boldsymbol{x})^{n}\right)\right)$
 We use the following terminology and notation in the proof of Theorem C.1:
   1. 1.
-    We refer to the copy of Algorithm 1 running on $\Pi_{a}$ as “copy $a$ of Algorithm 1”.
+    We refer to the copy of Algorithm 1 running on $\Pi_{y}$ as “copy $y$ of Algorithm 1”.
   2. 2.
-    Let $\pi_{t}^{a}$ and $X_{t}^{a}$ refer to the values of $\pi_{t}$ and $X_{t}$ for copy $a$ of Algorithm 1.
+    Let $\pi_{t}^{y}$ and $X_{t}^{y}$ refer to the values of $\pi_{t}$ and $X_{t}$ for copy $y$ of Algorithm 1.
   3. 3.
-    Let $\pi^{ma}:S\to\{0,1\}$ be the policy defined by $\pi^{ma}(s)=\mathbf{1}(\pi^{m}(s_{t})=a)$.
-Note that querying the mentor tells the agent $\pi^{m}(s_{t})$, which allows the agent to compute $\pi^{ma}(s_{t})$: this is necessary when Algorithm 1 queries while running on some $\Pi_{a}$.
+    Let $\pi^{my}:\mathcal{X}\to\{0,1\}$ be the policy defined by $\pi^{my}(x)=\mathbf{1}(\pi^{m}(x_{t})=y)$.
+Note that querying the mentor tells the agent $\pi^{m}(x_{t})$, which allows the agent to compute $\pi^{my}(x_{t})$: this is necessary when Algorithm 1 queries while running on some $\Pi_{y}$.
   4. 4.
-    Let $V_{T}^{a}=\{t\in[T]:b_{t}^{a}\neq\pi^{ma}(s_{t})\}$ be the set of time steps where $\pi_{t}^{a}$ does not correctly determine whether the mentor would take action $a$ and let $V_{T}=\{t\in[T]:a_{t}\neq\pi^{m}(s_{t})\}$ be the set of time steps where the agent’s action doesn’t match the mentor’s.
+    Let $M_{T}^{y}=\{t\in[T]:b_{t}^{y}\neq\pi^{my}(x_{t})\}$ be the set of time steps where $\pi_{t}^{y}$ does not correctly determine whether the mentor would take action $y$ and let $M_{T}=\{t\in[T]:y_{t}\neq\pi^{m}(x_{t})\}$ be the set of time steps where the agent’s action doesn’t match the mentor’s.
 Lemma C.1.
-We have $|V_{T}|\leq\sum_{a\in A}|V_{T}^{a}|$.
+We have $|M_{T}|\leq\sum_{y\in\mathcal{Y}}|M_{T}^{y}|$.
 Proof.
-We claim that $V_{T}\subseteq\cup_{a\in A}V_{T}^{a}$.
-Suppose the opposite: then there exists $t\in V_{T}$ such that $b_{t}^{a}=\pi^{ma}(s_{t})$ for all $a\in A$.
-Since $\pi^{m}(s_{t})\in A$, there is exactly one $a\in A$ such that $\mathbf{1}(\pi^{m}(s_{t})=a)=\pi^{ma}(s_{t})=b_{t}^{a}=1$.
-Specifically, this holds for $a=\pi^{m}(s_{t})$.
-But then Algorithm 3 takes action $\min\{a\in A:b_{t}^{a}=1\}=\pi^{m}(s_{t})$, which contradicts $t\in V_{T}$.
-Therefore $V_{T}\subseteq\cup_{a\in A}V_{T}^{a}$, and applying the union bound completes the proof.
+We claim that $M_{T}\subseteq\cup_{y\in\mathcal{Y}}M_{T}^{y}$.
+Suppose the opposite: then there exists $t\in M_{T}$ such that $b_{t}^{y}=\pi^{my}(x_{t})$ for all $y\in\mathcal{Y}$.
+Since $\pi^{m}(x_{t})\in\mathcal{Y}$, there is exactly one $y\in\mathcal{Y}$ such that $\mathbf{1}(\pi^{m}(x_{t})=y)=\pi^{my}(x_{t})=b_{t}^{y}=1$.
+Specifically, this holds for $y=\pi^{m}(x_{t})$.
+But then Algorithm 3 takes action $\min\{y\in\mathcal{Y}:b_{t}^{y}=1\}=\pi^{m}(x_{t})$, which contradicts $t\in M_{T}$.
+Therefore $M_{T}\subseteq\cup_{y\in\mathcal{Y}}M_{T}^{y}$, and applying the union bound completes the proof.
 ∎
 Lemma C.2.
-For all $t\in[T]$, $\mu^{m}(s_{t})-\mu(s_{t},a_{t})\leq L\varepsilon^{1/n}$.
+For all $t\in[T]$, $\mu_{t}^{m}(x_{t})-\mu_{t}(x_{t},y_{t})\leq L\varepsilon^{1/n}$.
 Proof.
 The argument is similar to the proof of Lemma B.2.
-If $\mu^{m}(s_{t})\neq\mu(s_{t},a_{t})$, then $a_{t}=a$ for some $a\in A$ where $b_{t}^{a}=1$.
-Therefore copy $a$ of Algorithm 1 did not query at time $t$ and $\pi_{t}^{a}(s_{t})=1$.
-Let $(s^{\prime},a^{\prime})=\operatorname*{arg\,min}_{(s,a)\in X_{t}^{a}:\pi_{t}^{a}(s_{t})=a}||s_{t}-s||$.
-Then $||s_{t}-s^{\prime}||\leq\varepsilon^{1/n}$ and $a^{\prime}=\pi_{t}^{a}(s_{t})=1$.
-By construction of $X_{t}^{a}$, $a^{\prime}=\pi^{ma}(s^{\prime})$ so $\pi^{ma}(s^{\prime})=1$ which implies $\pi^{m}(s^{\prime})=a$. Then by the local generalization assumption,
-$\mu(s_{t},a_{t})=\mu(s_{t},a)=\mu(s_{t},\pi^{m}(s^{\prime}))\geq\mu^{m}(s_{t})-L||s_{t}-s^{\prime}||\geq\mu^{m}(s_{t})-L\varepsilon^{1/n}$
+If $\mu_{t}^{m}(x_{t})\neq\mu_{t}(x_{t},y_{t})$, then $y_{t}=y$ for some $y\in\mathcal{Y}$ where $b_{t}^{y}=1$.
+Therefore copy $y$ of Algorithm 1 did not query at time $t$ and $\pi_{t}^{y}(x_{t})=1$.
+Let $(x^{\prime},y^{\prime})=\operatorname*{arg\,min}_{(x,y)\in X_{t}^{y}:\pi_{t}^{y}(x_{t})=y}||x_{t}-x||$.
+Then $||x_{t}-x^{\prime}||\leq\varepsilon^{1/n}$ and $y^{\prime}=\pi_{t}^{y}(x_{t})=1$.
+By construction of $X_{t}^{y}$, $y^{\prime}=\pi^{my}(x^{\prime})$ so $\pi^{my}(x^{\prime})=1$ which implies $\pi^{m}(x^{\prime})=y$. Then by the local generalization assumption,
+$\mu_{t}(x_{t},y_{t})=\mu_{t}(x_{t},y)=\mu_{t}(x_{t},\pi^{m}(x^{\prime}))\geq\mu_{t}^{m}(x_{t})-L||x_{t}-x^{\prime}||\geq\mu_{t}^{m}(x_{t})-L\varepsilon^{1/n}$
 as required. ∎
@@
 Proof.
-Theorem 5.2 implies that each copy of Algorithm 1 makes $O\big{(}T^{\frac{4n+1}{4n+2}}\big{(}\frac{d}{\sigma}\log T+\operatorname*{diam}(\mathbf{s})^{n}\big{)}\big{)}$ queries in expectation, so by linearity of expectation, the expected number of queries made by Algorithm 3 is $O\big{(}|A|T^{\frac{4n+1}{4n+2}}\big{(}\frac{d}{\sigma}\log T+\operatorname*{diam}(\mathbf{s})^{n}\big{)}\big{)}$.
+Theorem 5.2 implies that each copy of Algorithm 1 makes $O\big{(}T^{\frac{4n+1}{4n+2}}\big{(}\frac{d}{\sigma}\log T+\operatorname*{diam}(\boldsymbol{x})^{n}\big{)}\big{)}$ queries in expectation, so by linearity of expectation, the expected number of queries made by Algorithm 3 is $O\big{(}|\mathcal{Y}|T^{\frac{4n+1}{4n+2}}\big{(}\frac{d}{\sigma}\log T+\operatorname*{diam}(\boldsymbol{x})^{n}\big{)}\big{)}$.
 Using the same argument as in the proof of Lemma B.4 (with Lemma C.2 replacing Lemma B.2), we get
-$R_{T}=$ $\ \prod_{t=1}^{T}\mu^{m}(s_{t})-\prod_{t=1}^{T}\mu(s_{t},a_{t})$
-$\leq$ $\ \sum_{t=1}^{T}\Big{(}\mu^{m}(s_{t})-\min(\mu^{m}(s_{t}),\mu(s_{t},a_{t}))\Big{)}$
-$\leq$ $\ |V_{T}|L\varepsilon^{1/n}$
-Then by Lemma C.1, $R_{T}\leq L\varepsilon^{1/n}\sum_{a\in A}|V_{T}^{a}|$. Taking the expectation and applying Lemma B.1 to each $V_{T}^{a}$ gives us
-$\operatorname*{\mathbb{E}}[R_{T}]\leq L\varepsilon^{1/n}\sum_{a\in A}O\left(\frac{d}{\sigma}T\varepsilon\log(1/\varepsilon)\log T\right)=O\left(|A|L\varepsilon^{1/n}\frac{d}{\sigma}T\varepsilon\log(1/\varepsilon)\log T\right)$
-as required. ∎
+$R_{T}=$ $\ \prod_{t=1}^{T}\mu_{t}^{m}(x_{t})-\prod_{t=1}^{T}\mu_{t}(x_{t},y_{t})$
+$\leq$ $\ \sum_{t=1}^{T}\Big{(}\mu_{t}^{m}(x_{t})-\min(\mu_{t}^{m}(x_{t}),\mu_{t}(x_{t},y_{t}))\Big{)}$
+$\leq$ $\ |M_{T}|L\varepsilon^{1/n}$
+Then by Lemma C.1, $R_{T}\leq L\varepsilon^{1/n}\sum_{y\in\mathcal{Y}}|M_{T}^{y}|$. Taking the expectation and applying Lemma B.1 to each $M_{T}^{y}$ gives us
+$\operatorname*{\mathbb{E}}[R_{T}]\leq L\varepsilon^{1/n}\sum_{y\in\mathcal{Y}}O\left(\frac{d}{\sigma}T\varepsilon\log(1/\varepsilon)\log T\right)=O\left(|\mathcal{Y}|L\varepsilon^{1/n}\frac{d}{\sigma}T\varepsilon\log(1/\varepsilon)\log T\right)$
+Plugging in $\varepsilon=T^{\frac{-2n}{2n+1}}$ produces the desired bounds (see the arithmetic in the proof of Theorem 5.2 in Appendix B). ∎
+One can use the same argument to obtain a multi-action additive regret bound: simply modify the proof of Theorem C.1 to use $\sum_{t=1}^{T}\mu_{t}^{m}(x_{t})-\sum_{t=1}^{T}\mu_{t}(x_{t},y_{t})\leq|M_{T}|L\varepsilon^{1/n}$ instead of $R_{T}\leq|M_{T}|L\varepsilon^{1/n}$.
+Corollary C.1.1.
+Assume $\pi^{m}\in\Pi$ where either (1) $\Pi_{y}$ has finite VC dimension $d$ and $\boldsymbol{x}$ is $\sigma$-smooth or (2) $\Pi_{y}$ has finite Littlestone dimension $d$ for all $y\in\mathcal{Y}$.
+Then for any $T\in\mathbb{N}$, Algorithm 3 with $T$ and $\varepsilon=T^{\frac{-2n}{2n+1}}$ satisfies
+$\operatorname*{\mathbb{E}}\left[\sum_{t=1}^{T}\mu_{t}^{m}(x_{t})-\sum_{t=1}^{T}\mu_{t}(x_{t},y_{t})\right]\in$ $\ O\left(\frac{|\mathcal{Y}|dL}{\sigma}T^{\frac{-1}{2n+1}}\log T\right)$
 Appendix D There exist policy classes which are learnable in our setting but not in the standard online model
 This section presents another algorithm with subconstant regret and sublinear queries, but under different assumptions.
-The most important aspect of our algorithm is that it can handle the class of thresholds on $[0,1]$, which is known to have infinite Littlestone dimension and thus be hard in the standard online learning model.
-(Example 21.4 in [34]).
-Specifically, we assume a 1D state space and we allow the state sequence to be fully adversarial chosen. Instead of VC/Littlestone dimension, we consider the following notion of simplicity:
+The primary takeaway here is that our algorithm can handle the class of thresholds on $[0,1]$, which is known to have infinite Littlestone dimension and thus be hard in the standard online learning model.
+(Example 21.4 in Shalev-Shwartz & Ben-David (2014)).
+Specifically, we assume a 1D input space and we allow the input sequence to be fully adversarial chosen. Instead of VC/Littlestone dimension, we consider the following notion of simplicity:
 Definition D.1.
-Given a mentor policy $\pi^{m}$, partition the state space $S$ into intervals such that all states within each interval share the same mentor action.
-Let $\{S_{1},\dots,S_{k}\}$ be a partition that minimizes the number of intervals.
-We call each $S_{j}$ a segment.
-Let $f(\pi^{m})$ denote the number of segments in $\pi^{m}$.
-Bounding the number of segments is similar conceptually to VC dimension in that it limits the ability of the policy class to realize arbitrary combinations of labels (i.e., mentor actions) on $\mathbf{s}$.
+Given a mentor policy $\pi^{m}$, partition the input space $\mathcal{X}$ into intervals such that all inputs within each interval share the same mentor action.
+Let $\{X_{1},\dots,X_{k}\}$ be a partition that minimizes the number of intervals.
+We call each $X_{j}$ a segment.
+Let $S(\pi^{m})$ denote the number of segments in $\pi^{m}$.
+Bounding the number of segments is similar conceptually to VC dimension in that it limits the ability of the policy class to realize arbitrary combinations of labels (i.e., mentor actionx) on $\boldsymbol{x}$.
 For example, if $\Pi$ is the class of thresholds on $[0,1]$, every $\pi\in\Pi$ has at most two segments, and thus the positive result in this section will apply.
 This demonstrates the existence of policy classes which are learnable in our setting but not learnable in the standard online learning model, meaning that the two settings do not exactly coincide.
-We prove the following result:
+We prove the following regret bound. Like our other positive results, this bound applies to both multiplicative and additive regret.
 Theorem D.2.
-[…]
+For any $\boldsymbol{x}\in\mathcal{X}^{T}$, any $\pi^{m}$ with $S(\pi^{m})\leq K$, and any function $g:\mathbb{N}\to\mathbb{N}$, Algorithm 4 makes at most $(\operatorname*{diam}(\boldsymbol{x})+4)g(T)$ queries and satisfies
+$R_{T}\leq$ $\ \frac{2LKT}{g(T)^{2}}$
+$\sum_{t=1}^{T}\mu_{t}^{m}(x_{t})-\sum_{t=1}^{T}\mu_{t}(x_{t},y_{t})\leq$ $\ \frac{2LKT}{g(T)^{2}}$
 Choosing $g(T)=T^{c}$ for $c\in(1/2,1)$ is sufficient to subconstant regret and sublinear queries:
 Theorem D.3.
-For any $c\in(1/2,1)$, Algorithm 4 with $g(T)=T^{c}$ makes $O(T^{c}(\operatorname*{diam}(\mathbf{s})+1))$ queries and satisfies
-$\lim_{T\to\infty}\ \sup_{\mathbf{s}\in S^{T}}\ \sup_{\mu}\ \sup_{\pi^{m}:f(\pi^{m})\leq K}\ R_{T}=0$
+For any $c\in(1/2,1)$, Algorithm 4 with $g(T)=T^{c}$ makes $O(T^{c}(\operatorname*{diam}(\boldsymbol{x})+1))$ queries and satisfies
+$R_{T}\leq$ $\ 2LKT^{1-2c}$
+$\sum_{t=1}^{T}\mu_{t}^{m}(x_{t})-\sum_{t=1}^{T}\mu_{t}(x_{t},y_{t})\leq$ $\ 2LKT^{1-2c}$
 Our algorithm does not need to know $L$ or the number of segments; it only needs to know $T$.
 D.1 Intuition behind the algorithm
-The algorithm maintains a set of buckets which partition the observed portion of the state space.
-Each bucket’s length determines the maximum loss in payoff we will allow from that subset of the state space.
-As long as the bucket contains a query from a prior time step, local generalization allows us to bound $\mu^{m}(s_{t})-\mu(s_{t},a_{t})$ based on the length of the bucket containing $s_{t}$.
+The algorithm maintains a set of buckets which partition the observed portion of the input space.
+Each bucket’s length determines the maximum loss in payoff we will allow from that subset of the input space.
+As long as the bucket contains a query from a prior time step, local generalization allows us to bound $\mu_{t}^{m}(x_{t})-\mu_{t}(x_{t},y_{t})$ based on the length of the bucket containing $x_{t}$.
 We always query if the bucket does not contain a prior query
@@
 Regardless of the bucket length, the adversary can still place multiple segments in the same bucket $B$.
-A single query only tells us the optimal action for one of those segments, so we risk a payoff as bad as $\mu^{m}(s_{t})-O(\operatorname*{len}(B))$ whenever we choose not to query.
+A single query only tells us the optimal action for one of those segments, so we risk a payoff as bad as $\mu_{t}^{m}(x_{t})-O(\operatorname*{len}(B))$ whenever we choose not to query.
 We can endure a limited number of such payoffs, but if we never query again in that bucket, we may suffer $\Theta(T)$ such payoffs.
-Letting $\mu^{m}(s_{t})=1$ for simplicity, that would lead to $\prod_{t=1}^{T}\mu(s_{t},a_{t})\leq\big{(}1-\frac{1}{O(g(T))}\big{)}^{\Theta(T)}$, which converges to 0 (i.e., guaranteed catastrophe) when $g(T)\in o(T)$.
+Letting $\mu_{t}^{m}(x_{t})=1$ for simplicity, that would lead to $\prod_{t=1}^{T}\mu_{t}(x_{t},y_{t})\leq\big{(}1-\frac{1}{O(g(T))}\big{)}^{\Theta(T)}$, which converges to 0 (i.e., guaranteed catastrophe) when $g(T)\in o(T)$.
 This failure mode suggests a natural countermeasure: if we start to suffer significant (potential) losses in the same bucket, then we should probably query there again.
@@
 1:function AvoidCatastrophe($T\in\mathbb{N},\>g:\mathbb{N}\to\mathbb{N}$)
-2:     $S_{Q}\leftarrow\emptyset$triangleright Previously queried states
-3:     $\pi\leftarrow\emptyset$triangleright Records $\pi^{m}(s)$ for each $s\in S_{Q}$
+2:     $X_{Q}\leftarrow\emptyset$triangleright Previously queried inputs
+3:     $\pi\leftarrow\emptyset$triangleright Records $\pi^{m}(x)$ for each $x\in X_{Q}$
 4:     $\mathcal{B}\leftarrow\emptyset$triangleright The set of active buckets
 5:     for $t$ from $1$ to $T$ do
-6:         EvaluateState$(s_{t})$
-7:function EvaluateState($s\in S$)
-8:     if $s\not\in B$ for all $B\in\mathcal{B}$ then $\triangleright$ No bucket containing $s$: create a new bucket and try again
-9:         $B\leftarrow\left[\frac{j-1}{g(T)},\frac{j}{g(T)}\right]$ for $j\in\mathbb{Z}$ such that $s\in B$
+6:         EvaluateInput$(x_{t})$
+7:function EvaluateInput($x\in\mathcal{X}$)
+8:     if $s\not\in B$ for all $B\in\mathcal{B}$ then $\triangleright$ No bucket containing $x$: create a new bucket and try again
+9:         $B\leftarrow\left[\frac{j-1}{g(T)},\frac{j}{g(T)}\right]$ for $j\in\mathbb{Z}$ such that $x\in B$
 10:         $\mathcal{B}\leftarrow\mathcal{B}\cup\{B\}$
 11:         $n_{B}\leftarrow 0$triangleright Number of time steps that have used $B$
-12:         EvaluateState$(s)$
+12:         EvaluateInput$(x)$
 13:     else
-14:         $B\leftarrow$ any bucket containing $s$
-15:         if $S_{Q}\cap B=\emptyset$ then $\triangleright$ No queries in this bucket
-16:              Query mentor and observe $\pi^{m}(s)$
-17:              $\pi(s)\leftarrow\pi^{m}(s)$
-18:              $S_{Q}\leftarrow S_{Q}\cup\{s\}$
+14:         $B\leftarrow$ any bucket containing $x$
+15:         if $X_{Q}\cap B=\emptyset$ then $\triangleright$ No queries in this bucket
+16:              Query mentor and observe $\pi^{m}(x)$
+17:              $\pi(x)\leftarrow\pi^{m}(x)$
+18:              $X_{Q}\leftarrow X_{Q}\cup\{x\}$
 19:              $n_{B}\leftarrow n_{B}+1$
 20:         else if $n_{B}<T/g(T)$ then $\triangleright$ Bucket has a query and isn’t full: take that action
-21:              Let $s^{\prime}\in S_{Q}\cap B$
-22:              Take action $\pi(s^{\prime})$
+21:              Let $x^{\prime}\in X_{Q}\cap B$
+22:              Take action $\pi(x^{\prime})$
 23:              $n_{B}\leftarrow n_{B}+1$
 24:         $\triangleright$ Bucket is full: split bucket and try again
-25:              $B=[x,y]$
-26:              $(B_{1},B_{2})\leftarrow\Big{(}\left[x,\frac{x+y}{2}\right],\left[\frac{x+y}{2},y\right]\Big{)}$
-27:              $(s_{B_{1}},s_{B_{2}})\leftarrow(0,0)$
+25:              $B=[a,b]$
+26:              $(B_{1},B_{2})\leftarrow\Big{(}\left[a,\frac{a+b}{2}\right],\left[\frac{a+b}{2},b\right]\Big{)}$
+27:              $(x_{B_{1}},x_{B_{2}})\leftarrow(0,0)$
 28:              $\mathcal{B}\leftarrow\mathcal{B}\cup\{B_{1},B_{2}\}\setminus B$
-29:              EvaluateState$(s)$
+29:              EvaluateInput$(x)$
 Algorithm 4 achieves subconstant regret when the mentor’s policy has a bounded number of segments.
@@
 We will use the following notation throughout the proof of Theorem D.2:
-  • Let $V_{T}=\{t\in[T]:\mu(s_{t},a_{t})<\mu^{m}(s_{t})\}$ be the set of time steps with a suboptimal payoff.
+  1. 1.
+    Let $M_{T}=\{t\in[T]:\mu_{t}(x_{t},y_{t})<\mu_{t}^{m}(x_{t})\}$ be the set of time steps with a suboptimal payoff.
+  2. 2.
     Let $B_{t}$ be the bucket that is used on time step $t$ (as defined on line 14 of Algorithm 4).
+  3. 3.
     Let $d(B)$ be the depth of bucket $B$
+    1. (a)
       Buckets created on line 9 are depth 0.
+    2. (b)
       We refer to $B_{1},B_{2}$ created on line 26 as the children of the bucket $B$ defined on line 14.
+    3. (c)
       If $B^{\prime}$ is the child of $B$, $d(B^{\prime})=d(B)+1$.
+    4. (d)
 […]
+  4. 4.
     Viewing the set of buckets are a binary tree defined by the “child” relation, we use the terms “ancestor” and ”descendant” in accordance with their standard tree definitions.
-  • Let $\mathcal{B}_{V}=\{B:\exists t\in V_{T}\text{ s.t. }B_{t}=B\}$ be the set of buckets that ever produced a suboptimal payoff.
+  5. 5.
+    Let $\mathcal{B}_{V}=\{B:\exists t\in M_{T}\text{ s.t. }B_{t}=B\}$ be the set of buckets that ever produced a suboptimal payoff.
+  6. 6.
     Let $\mathcal{B}_{V}^{\prime}=\{B\in\mathcal{B}_{V}:\text{no descendant of $B$ is in $\mathcal{B}_{V}$}\}$.
@@
 Lemma D.1.
-Algorithm 4 performs at most $(\operatorname*{diam}(\mathbf{s})+4)g(T)$ queries.
+Algorithm 4 performs at most $(\operatorname*{diam}(\boldsymbol{x})+4)g(T)$ queries.
 Proof.
@@
 There are two ways to create a bucket: from scratch (line 9), or by splitting an existing bucket (line 26).
-Since depth 0 buckets overlap only at their boundaries, and each depth 0 bucket has length $1/g(T)$, at most $g(T)\max_{t,t^{\prime}\in[T]}|s_{t}-s_{t^{\prime}}|=g(T)\operatorname*{diam}(\mathbf{s})$ depth 0 buckets are subsets of the interval $[\min_{t\in[T]}s_{t},\max_{t\in[T]}s_{t}]$.
-At most two depth 0 buckets are not subsets of that interval (one at each end), so the total number of depth 0 buckets is at most $g(T)\operatorname*{diam}(\mathbf{s})+2$.
-We split a bucket $B$ when $n_{B}$ reaches $T/g(T)$, which creates two new buckets.
-Since each time step increments $n_{B}$ for a single bucket $B$, and there are a total of $T$ time steps, the total number of buckets created via splitting is at most
-$\frac{2T}{T/g(T)}=2g(T)$
-Therefore the total number of buckets ever in existence is $(\operatorname*{diam}(\mathbf{s})+2)g(T)+2\leq(\operatorname*{diam}(\mathbf{s})+4)g(T)$, so Algorithm 4 performs at most $(\operatorname*{diam}(\mathbf{s})+4)g(T)$ queries.
+Since depth 0 buckets overlap only at their boundaries, and each depth 0 bucket has length $1/g(T)$, at most $g(T)\max_{t,t^{\prime}\in[T]}|x_{t}-x_{t^{\prime}}|=g(T)\operatorname*{diam}(\boldsymbol{x})$ depth 0 buckets are subsets of the interval $[\min_{t\in[T]}x_{t},\max_{t\in[T]}x_{t}]$.
+At most two depth 0 buckets are not subsets of that interval (one at each end), so the total number of depth 0 buckets is at most $g(T)\operatorname*{diam}(\boldsymbol{x})+2$.
+[…]
+Lemma D.2.
+For each $t\in[T]$, $\mu_{t}(x_{t},y_{t})\geq\mu_{t}^{m}(x_{t})-L\operatorname*{len}(B_{t})$.
+Proof.
+If we query the mentor at time $t$, $\mu_{t}(x_{t},y_{t})=\mu_{t}^{m}(x_{t})$.
+Thus assume we do not query the mentor at time $t$: then there exists $x^{\prime}\in B_{t}$ (as defined on line 21 of Algorithm 4) such that $y_{t}=\pi(x^{\prime})=\pi^{m}(x^{\prime})$.
+Since $x_{t}$ and $x^{\prime}$ are both in $B_{t}$, $|x_{t}-x^{\prime}|\leq\operatorname*{len}(B_{t})$.
+Then by the local generalization assumption, $\mu_{t}(x_{t},y_{t})=\mu_{t}(x_{t},\pi^{m}(x^{\prime}))\geq\mu_{t}^{m}(x_{t})-L||x_{t}-x^{\prime}||\geq\mu_{t}^{m}(x_{t})-L\operatorname*{len}(B_{t})$.
 ∎
-Lemma D.2.
-For each $t\in[T]$, $\mu(s_{t},a_{t})\geq\mu^{m}(s_{t})-L\operatorname*{len}(B_{t})$.
-Proof.
-If we query the mentor at time $t$, $\mu(s_{t},a_{t})=\mu^{m}(s_{t})$.
-Thus assume we do not query the mentor at time $t$: then there exists $s^{\prime}\in B_{t}$ (as defined on line 21 of Algorithm 4) such that $a_{t}=\pi(s^{\prime})=\pi^{m}(s^{\prime})$.
-Since $s_{t}$ and $s^{\prime}$ are both in $B_{t}$, $|s_{t}-s^{\prime}|\leq\operatorname*{len}(B_{t})$.
-Then by the local generalization assumption,
-$\mu(s_{t},a_{t})=\mu(s_{t},\pi^{m}(s^{\prime}))\geq\mu^{m}(s_{t})-L||s_{t}-s^{\prime}||\geq\mu^{m}(s_{t})-L\operatorname*{len}(B_{t})$
-as required. ∎
 Lemma D.3.
@@
 Proof.
-If two buckets $B$ and $B^{\prime}$ overlap by more than at a single point, one must be the descendant of the other.
-Thus by definition of $\mathcal{B}_{V}^{\prime}$, any two buckets in $\mathcal{B}_{V}^{\prime}$ overlap by at most a single point.
 Now consider any $B\in\mathcal{B}_{V}^{\prime}$.
-By definition, there exists $t\in V_{T}$ such that $s_{t}\in B$.
-Then there exists $s^{\prime}\in B$ (as defined in Algorithm 4) such that $a_{t}=\pi(s^{\prime})=\pi^{m}(s^{\prime})$.
-Since $t\in V_{T}$, we have $\pi^{m}(s_{t})\neq a_{t}=\pi^{m}(s^{\prime})$.
-Thus $s_{t}$ and $s^{\prime}$ are in different segments; let $S_{B}$ be the segment containing $\min(s_{t},s^{\prime})$.
-Then $\min(s_{t},s^{\prime})\leq\max(S_{B})<\max(s_{t},s^{\prime})\leq\max(B)$.
-Consider any $B,B^{\prime}\in\mathcal{B}_{V}^{\prime}$ with $B\neq B^{\prime}$ and WLOG assume $\max(B)<\max(B^{\prime})$.
-Then $\max(S_{B})<\max(B)\leq\max(S_{B^{\prime}})$, so $S_{B}\neq S_{B^{\prime}}$.
-Thus the number of segments in $\pi^{m}$ is at least $\mathcal{B}_{V}^{\prime}$.
-Since the number of segments in $\pi^{m}$ is at most $K$, we must have $|\mathcal{B}_{V}^{\prime}|\leq K$.
+By definition of $\mathcal{B}_{V}^{\prime}$, there exists $t\in M_{T}$ such that $x_{t}\in B$.
+Then there exists $x^{\prime}\in B$ (as defined in Algorithm 4) such that $y_{t}=\pi(x^{\prime})=\pi^{m}(x^{\prime})$.
+Since $t\in M_{T}$, we have $\pi^{m}(x_{t})\neq y_{t}=\pi^{m}(x^{\prime})$.
+Thus $x_{t}$ and $x^{\prime}$ are in different segments, but are both in $B$.
+Therefore any $B\in\mathcal{B}_{V}^{\prime}$ must intersect at least two segments.
+Since $B$ is an interval, if it intersects two segments, it must intersect two adjacent segments $X_{j}$ and $X_{j+1}$.
+Furthermore, $B$ must contain an open neighborhood centered on the boundary between $X_{j}$ and $X_{j+1}$.
+Now consider some $B^{\prime}\in\mathcal{B}_{V}^{\prime}$ with $B\neq B^{\prime}$.
+We $|B\cap B^{\prime}|\leq 1$: otherwise one must be the descendant of the other, which contradicts the definition of $\mathcal{B}_{V}^{\prime}$.
+Suppose $B^{\prime}$ also intersects both $X_{j}$ and $X_{j+1}$: since $B^{\prime}$ is also an interval, $B^{\prime}$ must also contain an open neighborhood centered on the boundary between those two segments.
+But then $|B\cap B^{\prime}|>1$, which is a contradiction.
+Therefore any pair of adjacent segments $X_{j}$ and $X_{j+1}$, there is at most one bucket in $\mathcal{B}_{V}^{\prime}$ which contains an open neighborhood around their boundary.
+Since there are at most $K-1$ pairs of adjacent segments, we have $|\mathcal{B}_{V}^{\prime}|\leq K-1\leq K$.
 ∎
@@
 Proof.
-For every $t\in V_{T}$, we have $B_{t}=B$ for some $B\in\mathcal{B}_{V}$, so
-$\sum_{t\in V_{T}}\operatorname*{len}(B_{t})=\sum_{B\in\mathcal{B}_{V}}\ \sum_{t\in V_{T}:B=B_{t}}\operatorname*{len}(B_{t})$
+For every $t\in M_{T}$, we have $B_{t}=B$ for some $B\in\mathcal{B}_{V}$, so
+$\sum_{t\in M_{T}}\operatorname*{len}(B_{t})=\sum_{B\in\mathcal{B}_{V}}\ \sum_{t\in M_{T}:B=B_{t}}\operatorname*{len}(B_{t})$
 Next, observe that every $B\in\mathcal{B}_{V}\setminus\mathcal{B}_{V}^{\prime}$ must have a descendent in $\mathcal{B}_{V}^{\prime}$: otherwise we would have $B\in\mathcal{B}_{V}^{\prime}$.
@@
 Then we can write
-$\sum_{t\in V_{T}}$ $\ \operatorname*{len}(B_{t})\leq\sum_{B^{\prime}\in\mathcal{B}_{V}^{\prime}}\ \sum_{B\in\mathcal{A}(B^{\prime})}\ \sum_{t\in V_{T}:B=B_{t}}\operatorname*{len}(B_{t})$
-$=$ $\ \sum_{B^{\prime}\in\mathcal{B}_{V}^{\prime}}\ \sum_{B\in\mathcal{A}(B^{\prime})}\ |\{t\in V_{T}:B=B_{t}\}|\cdot\operatorname*{len}(B_{t})$
+$\sum_{t\in M_{T}}$ $\ \operatorname*{len}(B_{t})\leq\sum_{B^{\prime}\in\mathcal{B}_{V}^{\prime}}\ \sum_{B\in\mathcal{A}(B^{\prime})}\ \sum_{t\in M_{T}:B=B_{t}}\operatorname*{len}(B_{t})$
+$=$ $\ \sum_{B^{\prime}\in\mathcal{B}_{V}^{\prime}}\ \sum_{B\in\mathcal{A}(B^{\prime})}\ |\{t\in M_{T}:B=B_{t}\}|\cdot\operatorname*{len}(B_{t})$
 […]
-$\sum_{B\in\mathcal{A}(B^{\prime})}$ |$\ \frac{|\{t\in V_{T}:B=B_{t}\}|}{g(T)2^{d(B)}}\leq\frac{T}{g(T)^{2}}\sum_{B\in\mathcal{A}(B^{\prime})}\ \frac{1}{2^{d(B)}}$
+$\sum_{B\in\mathcal{A}(B^{\prime})}$ |$\ \frac{|\{t\in M_{T}:B=B_{t}\}|}{g(T)2^{d(B)}}\leq\frac{T}{g(T)^{2}}\sum_{B\in\mathcal{A}(B^{\prime})}\ \frac{1}{2^{d(B)}}$
 $=$ $\ \frac{T}{g(T)^{2}}\sum_{d=0}^{d(B^{\prime})}\frac{1}{2^{d}}\leq\frac{T}{g(T)^{2}}\sum_{d=0}^{\infty}\frac{1}{2^{d}}=\frac{2T}{g(T)^{2}}$
 Then by Lemma D.3,
-$\sum_{t\in V_{T}}\operatorname*{len}(B_{t})\leq\sum_{B^{\prime}\in\mathcal{B}_{V}^{\prime}}\frac{2T}{g(T)^{2}}=\frac{2T|\mathcal{B}_{V}^{\prime}|}{g(T)^{2}}\leq\frac{2KT}{g(T)^{2}}$
+$\sum_{t\in M_{T}}\operatorname*{len}(B_{t})\leq\sum_{B^{\prime}\in\mathcal{B}_{V}^{\prime}}\frac{2T}{g(T)^{2}}=\frac{2T|\mathcal{B}_{V}^{\prime}|}{g(T)^{2}}\leq\frac{2KT}{g(T)^{2}}$
 as claimed. ∎
@@
 Under the conditions of Theorem D.2, Algorithm 4 satisfies
-$\sum_{t=1}^{T}\Big{(}\mu^{m}(s_{t})-\min(\mu^{m}(s_{t}),\mu(s_{t},a_{t}))\Big{)}\leq\frac{2LKT}{g(T)^{2}}$
+$\sum_{t=1}^{T}\Big{(}\mu_{t}^{m}(x_{t})-\min(\mu_{t}^{m}(x_{t}),\mu_{t}(x_{t},y_{t}))\Big{)}\leq\frac{2LKT}{g(T)^{2}}$
 Proof.
-For $t\not\in V_{T}$ we have $\min(\mu^{m}(s_{t}),\mu(s_{t},a_{t}))=\mu^{m}(s_{t})$ by definition, and Lemma D.2 implies that $\min(\mu^{m}(s_{t}),\mu(s_{t},a_{t}))\geq L\operatorname*{len}(B_{t})$ for all $t\in[T]$.
+For $t\not\in M_{T}$ we have $\min(\mu_{t}^{m}(x_{t}),\mu_{t}(x_{t},y_{t}))=\mu_{t}^{m}(x_{t})$ by definition, and Lemma D.2 implies that $\min(\mu_{t}^{m}(x_{t}),\mu_{t}(x_{t},y_{t}))\geq L\operatorname*{len}(B_{t})$ for all $t\in[T]$.
 Thus
-$\sum_{t=1}^{T}\Big{(}\mu^{m}(s_{t})-\min(\mu^{m}(s_{t}),\mu(s_{t},a_{t}))\Big{)}\leq\sum_{t\in V_{T}}\Big{(}\mu^{m}(s_{t})-\min(\mu^{m}(s_{t}),\mu(s_{t},a_{t}))\Big{)}\leq L\sum_{t\in V_{T}}\operatorname*{len}(B_{t})$
+$\sum_{t=1}^{T}\Big{(}\mu_{t}^{m}(x_{t})-\min(\mu_{t}^{m}(x_{t}),\mu_{t}(x_{t},y_{t}))\Big{)}\leq$ $\ \sum_{t\in M_{T}}\Big{(}\mu_{t}^{m}(x_{t})-\min(\mu_{t}^{m}(x_{t}),\mu_{t}(x_{t},y_{t}))\Big{)}$
+$\leq$ $\ L\sum_{t\in M_{T}}\operatorname*{len}(B_{t})$
 Then by Lemma D.4,
-$\sum_{t=1}^{T}\Big{(}\mu^{m}(s_{t})-\min(\mu^{m}(s_{t}),\mu(s_{t},a_{t}))\Big{)}\leq\frac{2LKT}{g(T)^{2}}$
+$\sum_{t=1}^{T}\Big{(}\mu_{t}^{m}(x_{t})-\min(\mu_{t}^{m}(x_{t}),\mu_{t}(x_{t},y_{t}))\Big{)}\leq\frac{2LKT}{g(T)^{2}}$
 as required. ∎
 Lemma D.6.
-Under the conditions of Theorem D.2, Algorithm 4 satisfies $R_{T}\leq\frac{2LKT}{g(T)^{2}}$.
+Under the conditions of Theorem D.2, Algorithm 4 satisfies
+$R_{T}\leq$ $\ \frac{2LKT}{g(T)^{2}}$
+$\sum_{t=1}^{T}\mu_{t}^{m}(x_{t})-\sum_{t=1}^{T}\mu_{t}(x_{t},y_{t})\leq$ $\ \frac{2LKT}{g(T)^{2}}$
 Proof.
-Let $x_{t}=\mu^{m}(s_{t})$ and $y_{t}=\min(\mu^{m}(s_{t}),\mu(s_{t},a_{t}))$ for all $t\in[T]$. Then by Lemma B.3,
-$\prod_{t=1}^{T}\mu^{m}(s_{t})-\prod_{t=1}^{T}\min(\mu^{m}(s_{t}),\mu(s_{t},a_{t}))\leq$ $\ \sum_{t=1}^{T}\Big{(}\mu^{m}(s_{t})-\min(\mu^{m}(s_{t}),\mu(s_{t},a_{t}))\Big{)}$
-Since $\mu(s_{t},a_{t})\geq\min(\mu^{m}(s_{t}),\mu(s_{t},a_{t}))$ for all $t\in[T]$, we have
-$R_{T}=\prod_{t=1}^{T}\mu^{m}(s_{t})-\prod_{t=1}^{T}\mu(s_{t},a_{t})\leq\sum_{t=1}^{T}\Big{(}\mu^{m}(s_{t})-\min(\mu^{m}(s_{t}),\mu(s_{t},a_{t}))\Big{)}$
-Applying Lemma D.5 completes the proof. ∎
+Let $a_{t}=\mu_{t}^{m}(x_{t})$ and $b_{t}=\min(\mu_{t}^{m}(x_{t}),\mu_{t}(x_{t},y_{t}))$ for all $t\in[T]$. Then by Lemma B.3,
+$\prod_{t=1}^{T}\mu_{t}^{m}(x_{t})-\prod_{t=1}^{T}\min(\mu_{t}^{m}(x_{t}),\mu_{t}(x_{t},y_{t}))\leq$ $\ \sum_{t=1}^{T}\Big{(}\mu_{t}^{m}(x_{t})-\min(\mu_{t}^{m}(x_{t}),\mu_{t}(x_{t},y_{t}))\Big{)}$
+Since $\mu_{t}(x_{t},y_{t})\geq\min(\mu_{t}^{m}(x_{t}),\mu_{t}(x_{t},y_{t}))$ for all $t\in[T]$, we have
+$R_{T}=\prod_{t=1}^{T}\mu_{t}^{m}(x_{t})-\prod_{t=1}^{T}\mu_{t}(x_{t},y_{t})\leq\sum_{t=1}^{T}\Big{(}\mu_{t}^{m}(x_{t})-\min(\mu_{t}^{m}(x_{t}),\mu_{t}(x_{t},y_{t}))\Big{)}$
+Similarly,
+$\sum_{t=1}^{T}\mu_{t}^{m}(x_{t})-\sum_{t=1}^{T}\mu_{t}(x_{t},y_{t})\leq\sum_{t=1}^{T}\Big{(}\mu_{t}^{m}(x_{t})-\min(\mu_{t}^{m}(x_{t}),\mu_{t}(x_{t},y_{t}))$
+Applying Lemma D.5 to the above two inequalities completes the proof. ∎
 Theorem D.2 follows from Lemma D.1 and Lemma D.6.
 Appendix E Other proofs
+Proposition E.1 states that Lipschitz continuity implies local generalization when the mentor is optimal.
 Proposition E.1.
-Suppose $\mu$ is $L$-Lipschitz continuous and $\pi^{m}$ is optimal, i.e., $\mu(s,\pi^{m}(s))=\max_{a\in A}\mu(s,a)$ for all $s\in S$. Then $\pi^{m}$ satisfies local generalization with constant $2L$.
+Assume that $\mu$ satisfies Lipschitz continuity: for all $x,x^{\prime}\in\mathcal{X}$ and $y\in\mathcal{Y}$, $|\mu(x,y)-\mu(x^{\prime},y)|\leq L||x-x^{\prime}||$.
+Also assume that $\mu(x,\pi^{m}(x))=\max_{y\in\mathcal{Y}}\mu(x,y)$ for all $x\in\mathcal{X}$.
+Then $\mu$ satisfies local generalization with constant $2L$.
 Proof.
-For any $s,s^{\prime}\in S$, we have
-$\mu(s,\pi^{m}(s^{\prime}))\geq$ $\ \mu(s^{\prime},\pi^{m}(s^{\prime}))-L||s-s^{\prime}||$ (Lipschitz continuity of $\mu$)
-$\geq$ $\ \mu(s^{\prime},\pi^{m}(s))-L||s-s^{\prime}||$ ($\pi^{m}$ is optimal for $s^{\prime}$)
-$\geq$ $\ \mu(s,\pi^{m}(s))-2L||s-s^{\prime}||$ (Lipschitz continuity of $\mu$ again)
-$=$ $\ \mu^{m}(s)-2L||s-s^{\prime}||$ (Definition of $\mu^{m}(s)$)
-Since $\pi^{m}$ is optimal for $s$, we have
-$\mu^{m}(s)+2L||s-s^{\prime}||\geq\mu^{m}(s)\geq\mu(s,\pi^{m}(s^{\prime}))$
-Thus $-2L||s-s^{\prime}||\leq\mu(s,\pi^{m}(s^{\prime}))-\mu^{m}(s)\leq 2L||s-s^{\prime}||$.
-This is equivalent to $|\mu(s,\pi^{m}(s^{\prime}))-\mu^{m}(s)|\leq 2L||s-s^{\prime}||$, completing the proof.
+For any $x,x^{\prime}\in\mathcal{X}$, we have
+$\mu(x,\pi^{m}(x^{\prime}))\geq$ $\ \mu(x^{\prime},\pi^{m}(x^{\prime}))-L||x-x^{\prime}||$ (Lipschitz continuity of $\mu$)
+$\geq$ $\ \mu(x^{\prime},\pi^{m}(x))-L||x-x^{\prime}||$ ($\pi^{m}$ is optimal for $x^{\prime}$)
+$\geq$ $\ \mu(x,\pi^{m}(x))-2L||x-x^{\prime}||$ (Lipschitz continuity of $\mu$ again)
+$=$ $\ \mu^{m}(x)-2L||x-x^{\prime}||$ (Definition of $\mu^{m}(x)$)
+Since $\pi^{m}$ is optimal for $x$, we have
+$\mu^{m}(x)+2L||x-x^{\prime}||\geq\mu^{m}(x)\geq\mu(x,\pi^{m}(x^{\prime}))$
+Thus $-2L||x-x^{\prime}||\leq\mu(x,\pi^{m}(x^{\prime}))-\mu^{m}(x)\leq 2L||x-x^{\prime}||$.
+This is equivalent to $|\mu(x,\pi^{m}(x^{\prime}))-\mu^{m}(x)|\leq 2L||x-x^{\prime}||$, completing the proof.
+Proposition E.2 states that whether subconstant regret is achievable does not depend on whether we require expected sublinear queries or worst-case sublinear queries.
+Proposition E.2.
+Suppose an algorithm satisfies $\lim_{T\to\infty}\sup_{\boldsymbol{\mu},\pi^{m}}\operatorname*{\mathbb{E}}[R_{T}]=0$ and $\sup_{\boldsymbol{\mu},\pi^{m}}\operatorname*{\mathbb{E}}[|Q_{T}|]\in o(T)$.
+Then there exists $h:\mathbb{N}\to\mathbb{N}$ such that (1) $h(T)\in o(T)$ and (2) if the algorithm is modified to simply stop querying if the number of queries reaches $h(T)$, the algorithm still satisfies $\lim_{T\to\infty}\sup_{\boldsymbol{\mu},\pi^{m}}\operatorname*{\mathbb{E}}[R_{T}]=0$.
+Proof.
+We use $Q_{T},R_{T}$ to refer to the queries and regret of the original algorithm, and $Q_{T}^{\prime},R_{T}^{\prime}$ to refer to the queries and regret of the modified algorithm.
+Since $\sup_{\boldsymbol{\mu},\pi^{m}}\operatorname*{\mathbb{E}}[|Q_{T}|]\in o(T)$, there exists $g:\mathbb{N}\to\mathbb{N}$ such that $\sup_{\boldsymbol{\mu},\pi^{m}}\operatorname*{\mathbb{E}}[|Q_{T}|]\leq g(T)$ and $g(T)\in o(T)$.
+Let $h(T)=\sqrt{g(T)T}$; then $h(T)\in o(T)$ by Lemma A.1.
+Markov’s inequality implies that
+$\Pr\big{[}|Q_{T}|>h(T)\big{]}\leq\frac{\operatorname*{\mathbb{E}}[|Q_{T}|]}{h(T)}\leq\frac{g(T)}{\sqrt{g(T)T}}=\sqrt{\frac{g(T)}{T}}$
+Let $\xi$ denote the event that at some point, the original algorithm would query, but the modified algorithm cannot because $|Q_{T}^{\prime}|=h(T)$.
+Then $\Pr[\xi]\leq\Pr[|Q_{T}|>h(T)]$ (the inequality is because the modified algorithm might not want to query more anyway).
+Also note that the algorithms are equivalent if $\xi$ does not occur, so $\operatorname*{\mathbb{E}}[R_{T}^{\prime}\mid\neg\xi]=\operatorname*{\mathbb{E}}[R_{T}]$.
+Hence
+$\operatorname*{\mathbb{E}}[R_{T}^{\prime}]=$ $\ \operatorname*{\mathbb{E}}[R_{T}^{\prime}\mid\neg\xi]\Pr[\neg\xi]+\operatorname*{\mathbb{E}}[R_{T}^{\prime}\mid\xi]\Pr[\xi]$
+$\leq$ $\operatorname*{\mathbb{E}}[R_{T}]\cdot 1+1\cdot\Pr[\xi]$
+$\leq$ $\ \operatorname*{\mathbb{E}}[R_{T}]+\sqrt{\frac{g(T)}{T}}$
+Since $g(T)\in o(T)$, we get
+$\lim_{T\to\infty}\ \sup_{\boldsymbol{\mu},\pi^{m}}\ \operatorname*{\mathbb{E}}[R_{T}^{\prime}]\leq$ $\ \lim_{T\to\infty}\ \sup_{\boldsymbol{\mu},\pi^{m}}\ \left(\operatorname*{\mathbb{E}}[R_{T}]+\sqrt{\frac{g(T)}{T}}\right)$
+$=$ $\ \lim_{T\to\infty}\ \sup_{\boldsymbol{\mu},\pi^{m}}\ \operatorname*{\mathbb{E}}[R_{T}]+\lim_{T\to\infty}\ \sqrt{\frac{g(T)}{T}}$
+[…]
+as required. ∎
+Theorem E.3 shows that avoiding catastrophe is impossible without local generalization, even when $\boldsymbol{x}$ is $\sigma$-smooth and $\Pi$ has finite VC dimension.
+The first insight is that without local generalization, we can define $\mu(x,y)=\mathbf{1}(y=\pi^{m}(x))$ so that a single mistake causes $\prod_{t=1}^{T}\mu(x_{t},y_{t})=0$.
+To lower bound $\Pr\big{[}\prod_{t=1}^{T}\mu(x_{t},y_{t})=0\big{]}$, we use a similar approach to the proof of Theorem 4.1: divide $\mathcal{X}=[0,1]$ into $f(T)$ independent sections with $|Q_{T}|<<f(T)<<T$, so that the agent can only query a small fraction of these sections.
+However, the proof of Theorem E.3 is a bit easier, since we only need the agent to make a single mistake.
+Like the proof of Theorem 4.1, the proof of Theorem E.3 assumes worst-case sublinear queries.
+However, recall from Proposition E.2 that the distinction between worst-case sublinear queries and expected sublinear queries is not significant.
+Theorem E.3.
+Let $\mathcal{X}=[0,1]$ and $\mathcal{Y}=\{0,1\}$.
+Assume each input is sampled i.i.d.
+from the uniform distribution on $\mathcal{X}$ and define the mentor policy class by the set of intervals within $\mathcal{X}$, i.e., $\Pi=\{\pi:\exists a,b\in[0,1]\text{ s.t }\pi(x)=\mathbf{1}(x\in[a,b])\ \forall x\in\mathcal{X}\}$.
+Then without the local generalization assumption, any algorithm with sublinear queries satisfies $\lim_{T\to\infty}\sup_{\boldsymbol{\mu},\pi^{m}}\operatorname*{\mathbb{E}}[R_{T}]=1$.
+Proof.
+Part 1: Setup.
+Consider any algorithm which makes sublinear worst-case queries: then there exists $g:\mathbb{N}\to\mathbb{N}$ where $\sup_{\boldsymbol{\mu},\pi^{m}}|Q_{T}|\leq g(T)$ and $g(T)\in o(T)$.
+Define $f(T):=\sqrt{(g(T)+1)T}$; by Lemma A.1, $g(T)\in o(f(T))$ and $f(T)\in o(T)$.
+Divide $\mathcal{X}$ into $f(T)$ equally sized sections $X_{1},\dots,X_{f(T)}$ in the exactly the same way as in Section 4.2; see also Figure 2.
+Assume that each $x_{t}$ is in exactly one section: this assumption holds with probability 1, so it does not affect the regret.
+We use the probabilistic method: sample a segment $j^{m}\in[f(T)]$ uniformly at random, define $\pi^{m}$ by $\pi^{m}(x)=\mathbf{1}(x\in X_{j^{m}})$, and define $\mu$ by $\mu(x,y)=\mathbf{1}(y=\pi^{m}(x))$.
+In words, the mentor takes action 1 iff the input is in section $j^{m}$, and the agent receives payoff 1 if its action matches the mentor’s and zero otherwise.
+Since any choice of $j^{m}$ defines a valid $\mu$ and $\pi^{m}$, we have
+$\sup_{\boldsymbol{\mu},\pi^{m}}\ \operatorname*{\mathbb{E}}_{\boldsymbol{x},\boldsymbol{y}}\ [R_{T}(\boldsymbol{x},\boldsymbol{y},\boldsymbol{\mu},\pi^{m})]\geq\operatorname*{\mathbb{E}}_{j^{m}}\ \operatorname*{\mathbb{E}}_{\boldsymbol{x},\boldsymbol{y}}\ [R_{T}(\boldsymbol{x},\boldsymbol{y},(\mu,\dots,\mu),\pi^{m})]$
+Let $J_{\neg Q}=\{j\in[f(T)]:x_{t}\not\in X_{j}\ \forall t\in Q_{T}\}$ be the set of sections which are never queried.
+Let $j_{1},\dots,j_{k}$ be the sequence of sections queried by the agent: then $k\leq|Q_{T}|\leq g(T)$.
+Part 2: The agent is unlikely to determine $j^{m}$. By the chain rule of probability,
+$\Pr[j^{m}\in J_{\neg Q}]=\Pr\big{[}j_{i}\neq j^{m}\ \forall i\big{]}=\prod_{i=1}^{k}\Pr\big{[}j_{i}\neq j^{m}\mid j_{r}\neq j^{m}\ \forall r<i\big{]}$
+Now fix $i$ and assume $j_{r}\neq j^{m}\ \forall r<i$.
+Queries in sections other than $j^{m}$ provide no information about the value of $j^{m}$, so $j^{m}$ is uniformly distributed across the set of sections not yet queried, i.e., $\{j\in[f(T)]:j_{r}\neq j\ \forall r<i\}$.
+There are at least $f(T)-i+1$ such sections, since there are $i-1$ prior queries at this point.
+Thus $\Pr[j_{i}\neq j^{m}\mid j_{r}\neq j^{m}\ \forall r<i]\leq\frac{f(T)-i}{f(T)-i+1}$ (the inequality is because it could also be 0 if $j_{i}=j_{r}$ for some $i<r$).
+Therefore
+$\Pr\big{[}j^{m}\in J_{\neg Q}\big{]}\leq$ $\ \prod_{i=1}^{k}\frac{f(T)-i}{f(T)-i+1}$
+$=$ $\ \frac{f(T)-1}{f(T)}\cdot\frac{f(T)-2}{f(T)-1}\dots\frac{f(T)-k+1}{f(T)-k+2}\cdot\frac{f(T)-k}{f(T)-k+1}$
+$=$ $\ \frac{f(T)-k}{f(T)}$
+$\geq$ $\ 1-\frac{g(T)}{f(T)}$
+Part 3: If the agent fails to determine $j^{m}$, it is likely to make at least one mistake.
+For each $j\in J_{\neg Q}$, let $V_{j}=\{t\in[T]:x_{t}\in X_{j}\}$ be the set of time steps with inputs in section $j$.
+By Lemma A.3, $\Pr[|V_{j^{m}}|=0]\leq\exp\big{(}\frac{T}{16f(T)}\big{)}$.
+Then by the union bound, $\Pr[j^{m}\in J_{\neg Q}\text{ and }|V_{j^{m}}|>0]\geq 1-\frac{g(T)}{f(T)}-\exp\big{(}\frac{-T}{16f(T)}\big{)}$.
+For the rest of Part 3, assume $j^{m}\in J_{\neg Q}\text{ and }|V_{j^{m}}|>0$.
+Since $j^{m}\in J_{\neg Q}$, the agent has no information about $j^{m}$ other than that it is in $J_{\neg Q}$.
+This means that for all $j\in J_{\neg Q}$ and $t\in V_{j}$, $j^{m}$ is conditionally (under the condition of $j^{m}\in J_{\neg Q}$) independent of $y_{t}$.
+We proceed by case analysis.
+Case 1: For all $j\in J_{\neg Q},t\in V_{j}$, we have $y_{t}=0$.
+In particular, this holds for $j=j^{m}$, and we know there exists at least one $t\in V_{j^{m}}$ since $|V_{j^{m}}|>0$.
+Then $y_{t}\neq\pi^{m}(x_{t})$, so $\mu(x_{t},y_{t})=0$ and thus $\Pr\left[\prod_{r=1}^{T}\mu(x_{r},y_{r})=0\>\Big{|}\>j^{m}\in J_{\neg Q}\text{and }|V_{j^{m}}|>0\right]=1$.
+Case 2: There exists $j\in J_{\neg Q},t\in V_{j}$ with $y_{t}=1$. Then $\mu(x_{t},y_{t})=0$ unless $j=j^{m}$, so
+$\Pr\left[\prod_{r=1}^{T}\mu(x_{r},y_{r})=0\>\Big{|}\>j^{m}\in J_{\neg Q}\text{ and }|V_{j^{m}}|>0\right]\geq$ $\ \Pr\Big{[}\mu(x_{t},y_{t})=0\mid j^{m}\in J_{\neg Q}\text{ and }|V_{j^{m}}|>0\Big{]}$
+$=$ $\ \Pr\Big{[}j\neq j^{m}\mid j^{m}\in J_{\neg Q}\text{ and }|V_{j^{m}}|>0\Big{]}$
+Conditioned on $j^{m}\in J_{\neg Q}$, $j^{m}$ is uniformly distributed across $J_{\neg Q}$, so
+$\Pr\left[\prod_{r=1}^{T}\mu(x_{r},y_{r})=0\>\Big{|}\>j^{m}\in J_{\neg Q}\text{ and }|V_{j^{m}}|>0\right]\geq 1-\frac{1}{|J_{\neg Q}|}\geq 1-\frac{1}{f(T)-g(T)}$
+Combining Case 1 and Case 2, we get the overall bound of
+$\Pr\left[\prod_{t=1}^{T}\mu(x_{t},y_{t})=0\>\Big{|}\>j^{m}\in J_{\neg Q}\text{and }|V_{j^{m}}|>0\right]\geq 1-\frac{1}{f(T)-g(T)}$
+and thus
+$\Pr\left[\prod_{t=1}^{T}\mu(x_{t},y_{t})=0\right]\geq\Pr\left[\prod_{t=1}^{T}\mu(x_{t},y_{t})=0\text{ and }j^{m}\in J_{\neg Q}\text{ and }|V_{j^{m}}|>0\right]$
+$\quad\quad\quad=\Pr\left[\prod_{t=1}^{T}\mu(x_{t},y_{t})=0\>\Big{|}\>j^{m}\in J_{\neg Q}\text{ and }|V_{j^{m}}|>0\right]\cdot\Pr\Big{[}j^{m}\in J_{\neg Q}\text{ and }|V_{j^{m}}|>0\Big{]}$
+$\quad\quad\quad\geq\left(1-\frac{1}{f(T)-g(T)}\right)\left(1-\frac{g(T)}{f(T)}-\exp\left(\frac{-T}{16f(T)}\right)\right)$
+For brevity, let $\alpha(T)$ denote this final bound. Since $g(T)\in o(f(T))$ and $f(T)\in o(T)$, we have
+$\lim_{T\to\infty}\alpha(T)=$ $\ \lim_{T\to\infty}\left(1-\frac{1}{f(T)-g(T)}\right)\left(1-\frac{g(T)}{f(T)}-\exp\left(\frac{-T}{16f(T)}\right)\right)$
+$=$ $\ (1-0)(1-0-0)$
+[…]
+Part 4: Putting it all together. Since $\prod_{t=1}^{T}\mu(x_{t},y_{t})\leq 1$ always, we have
+$\operatorname*{\mathbb{E}}_{j^{m}}\ \operatorname*{\mathbb{E}}_{\boldsymbol{x},\boldsymbol{y}}\ \left[\prod_{t=1}^{T}\mu(x_{t},y_{t})\right]=$ $\ \operatorname*{\mathbb{E}}_{j^{m}}\ \operatorname*{\mathbb{E}}_{\boldsymbol{x},\boldsymbol{y}}\ \left[\prod_{t=1}^{T}\mu(x_{t},y_{t})\>\Big{|}\>\prod_{t=1}^{T}\mu(x_{t},y_{t})=0\right]\cdot\Pr\left[\prod_{t=1}^{T}\mu(x_{t},y_{t})=0\right]$
+$+$ $\ \operatorname*{\mathbb{E}}_{j^{m}}\ \operatorname*{\mathbb{E}}_{\boldsymbol{x},\boldsymbol{y}}\ \left[\prod_{t=1}^{T}\mu(x_{t},y_{t})\>\Big{|}\>\prod_{t=1}^{T}\mu(x_{t},y_{t})\neq 0\right]\cdot\Pr\left[\prod_{t=1}^{T}\mu(x_{t},y_{t})\neq 0\right]$
+$\leq 0$ $\cdot\Pr\left[\prod_{t=1}^{T}\mu(x_{t},y_{t})=0\right]+1\cdot\left(1-\Pr\left[\prod_{t=1}^{T}\mu(x_{t},y_{t})=0\right]\right)$
+$\leq 1$ $-\alpha(T)$
+Since $\prod_{t=1}^{T}\mu^{m}(x_{t})=1$ always, we have
+$\sup_{\boldsymbol{\mu},\pi^{m}}\ \operatorname*{\mathbb{E}}_{\boldsymbol{x},\boldsymbol{y}}\ [R_{T}(\boldsymbol{x},\boldsymbol{y},\boldsymbol{\mu},\pi^{m})]\geq$ $\ \operatorname*{\mathbb{E}}_{j^{m}}\ \operatorname*{\mathbb{E}}_{\boldsymbol{x},\boldsymbol{y}}\ [R_{T}(\boldsymbol{x},\boldsymbol{y},(\mu,\dots,\mu),\pi^{m})]$
+$=$ $\ 1-\operatorname*{\mathbb{E}}_{j^{m}}\operatorname*{\mathbb{E}}_{\boldsymbol{x},\boldsymbol{y}}\left[\prod_{t=1}^{T}\mu(x_{t},y_{t})\right]$
+$\geq$ $\ \alpha(T)$
+Therefore $\lim_{T\to\infty}\sup_{\boldsymbol{\mu},\pi^{m}}\operatorname*{\mathbb{E}}[R_{T}]\geq\lim_{T\to\infty}\alpha(T)=1$, as required. ∎

### v5

@@
 We propose an online learning problem where the goal is to minimize the chance of catastrophe.
-Specifically, we assume that the payoff in each round represents the chance of avoiding catastrophe that round and try to maximize the product of payoffs (the overall chance of avoiding catastrophe) while allowing a limited number of queries to a mentor.
-We first show that in general, any algorithm either constantly queries the mentor or is nearly guaranteed to cause catastrophe.
+Specifically, we assume that the payoff in each round represents the chance of avoiding catastrophe in that round and try to maximize the product of payoffs (the overall chance of avoiding catastrophe) while allowing a limited number of queries to a mentor.
+We also assume that the agent can transfer knowledge between similar inputs.
+We first show that in general, any algorithm either queries the mentor at a linear rate or is nearly guaranteed to cause catastrophe.
 However, in settings where the mentor policy class is learnable in the standard online model, we provide an algorithm whose regret and rate of querying the mentor both approach 0 as the time horizon grows.
+Although our focus is the product of payoffs, we provide matching bounds for the typical additive regret.
 Conceptually, if a policy class is learnable in the absence of catastrophic risk, it is learnable in the presence of catastrophic risk if the agent can ask for help.
+online learning, AI safety, asking for help, irreversibility
 1 Introduction
-There has been mounting concern over catastrophic risk from AI, including but not limited to autonomous weapon accidents (Abaimov & Martellini, 2020), bioterrorism (Mouton et al., 2024), and cyberattacks on critical infrastructure (Guembe et al., 2022).
+There has been mounting concern over catastrophic risk from AI, including but not limited to autonomous weapon accidents (Abaimov & Martellini, 2020), bioterrorism (Mouton et al., 2024), cyberattacks on critical infrastructure (Guembe et al., 2022), and loss of control (Bengio et al., 2024).
 See Critch & Russell (2023) and Hendrycks et al.
@@
 In this paper, we use “catastrophe” to refer to any kind of irreparable harm.
-This definition also covers smaller-scale (yet still unacceptable) incidents such as serious medical errors (Rajpurkar et al., 2022), crashing a robotic vehicle (Kohli & Chadha, 2020), and discriminatory sentencing (Villasenor & Foggo, 2020).
+In addition to the large-scale risks above, our definition also covers smaller-scale (yet still unacceptable) incidents such as serious medical errors (Rajpurkar et al., 2022), crashing a robotic vehicle (Kohli & Chadha, 2020), and discriminatory sentencing (Villasenor & Foggo, 2020).
 The gravity of these risks contrasts starkly with the dearth of theoretical understanding of how to avoid them.
@@
 More precisely, trial-and-error-style algorithms only work when catastrophe is assumed to be impossible.
-This assumption manifests differently in different subtypes of online learning.
-In the standard online learning model, the agent’s actions have no permanent effect on the environment.
-[^1] More precisely, the input can depend on the agent’s previous actions, but the agent’s performance is always evaluated relative to the optimal policy’s performance on the same sequence of inputs.
-Online reinforcement learning allows the agent’s actions to permanently affect the environment, but typically assumes that either no action has irreversible effects (e.g., Jaksch et al.
-(2010)) or that the agent is reset at the start of each “episode” (e.g., Azar et al.
-(2017)).
-One could train an agent entirely in a controlled lab setting where the above assumptions do hold, but we argue that sufficiently general agents will inevitably encounter novel scenarios when deployed in the real world.
-Machine learning models often behave unpredictably in unfamiliar environments (see, e.g., Quionero-Candela et al.
-(2009)), and we do not want AI biologists or robotic vehicles to behave unpredictably.
+This assumption can take multiple forms, such as that the agent’s actions do not affect future inputs (e.g., Slivkins, 2011), that no action has irreversible effects (e.g., Jaksch et al., 2010) or that the environment is reset at the start of each “episode” (e.g., Azar et al., 2017).
+One could train an agent entirely in a controlled lab setting where one of those assumptions does hold, but we argue that sufficiently general agents will inevitably encounter novel scenarios when deployed in the real world.
+Machine learning models often behave unpredictably in unfamiliar environments (see, e.g., Quiñonero-Candela et al., 2022), and we do not want AI biologists or robotic vehicles to behave unpredictably.
 The goal of this paper is to understand the conditions under which it is possible to formally guarantee avoidance of catastrophe in online learning.
-Certainly some conditions are necessary, because if the agent can only learn by trying actions directly, the problem is hopeless: any untried action could lead to paradise or disaster and the agent has no way to predict which.
-In the real world, however, one needn’t learn through pure trial-and-error: one can also ask for help.
+Certainly some conditions are necessary, because the problem is hopeless if the agent must rely purely on trial-and-error: any untried action could lead to paradise or disaster and the agent has no way to predict which.
+In the real world, however, one need not learn through pure trial-and-error: one can also ask for help.
 We think it is critical for high-stakes AI applications to employ a designated supervisor who can be asked for help.
 Examples include a human doctor supervising AI doctors, a robotic vehicle with a human driver who can take over in emergencies, autonomous weapons with a human operator, and many more.
-We hope that our work constitutes a small step in the direction of practical safety guarantees for such applications.
+We hope that our work constitutes a step in the direction of practical safety guarantees for such applications.
 1.1 Our model
 We propose an online learning model of avoiding catastrophe with mentor help.
-On each time step, the agent observes an input, selects an action (or queries the mentor), and obtains a payoff.
+On each time step, the agent observes an input, selects an action or queries the mentor, and obtains a payoff.
 Each payoff represents the probability of avoiding catastrophe on that time step (conditioned on no prior catastrophe).
@@
 As is standard in online learning, we consider the product of payoffs obtained while learning, not the product of payoffs of some final policy.
-The (possibly suboptimal) mentor has a fixed policy, and when queried, the mentor illustrates their policy’s action for the current input.
-We desire an agent whose regret – defined as the gap between the mentor’s performance and the agent’s performance – approaches zero as the time horizon $T$ grows.
+The (possibly suboptimal) mentor has a stationary policy, and when queried, the mentor illustrates their policy’s action for the current input.
+We want the agent’s regret – defined as the gap between the mentor’s performance and the agent’s performance – to go to zero as the time horizon $T$ grows.
 In other words, with enough time, the agent should avoid catastrophe nearly as well as the mentor.
-We also expect the agent to become self-sufficient over time: formally, the number of queries to the mentor should be sublinear in $T$, or equivalently, the rate of querying the mentor should go to zero.
+We also expect the agent to become self-sufficient over time: the number of queries to the mentor should be sublinear in $T$, or equivalently, the rate of querying the mentor should go to zero.
 1.2 Our assumptions
 The agent needs some way to make inferences about unqueried inputs in order to decide when to ask for help.
-Much past work has used Bayesian inference, which suffers tractability issues in complex environments.
+Much past work has used Bayesian inference, which suffers from tractability issues in complex environments.
 [^2] For the curious reader, Betancourt (2018) provides a thorough treatment.
 See also Section 2.
-We instead assume that the mentor policy satisfies what we call local generalization: informally, if the mentor told us that an action was safe for a similar input, then that action is probably also safe for the current input (see Section 3 for a formal definition and further discussion).
-This captures the intuition that one can transfer knowledge between similar situations.
-Unlike Bayesian inference, local generalization only requires computing distances and is compatible with any input space which admits a distance metric.
+We instead assume that the mentor policy satisfies a novel property that we call local generalization: informally, if the mentor told us that an action was safe for a similar input, then that action is probably also safe for the current input.
+For example, if it is safe to ignore a 3 mm spot on an X-ray, it is likely (but not certainly) also safe to ignore a 3.1 mm spot with the same density, location, etc.
+Unlike Bayesian inference, local generalization only requires computing distances and is compatible with any input space that admits a distance metric.
+See Section 5.2 for further discussion of local generalization.
 Unlike the standard online learning model, we assume that the agent does not observe payoffs.
@@
 [^3] One may be able to detect “close calls” in some cases, but observing the precise probability seems unrealistic.
-One consequence of this is that the payoff function does not affect the behavior of the algorithm, so our positive results hold simultaneously for all payoff functions which satisfy local generalization.
 1.3 Standard online learning
-An overview of standard online learning is in order before discussing our results.
+To properly understand our results, it is important to understand standard online learning.
 In the standard model, the agent observes an input on each time step and must choose an action.
@@
 The goal is sublinear regret with respect to the sum of payoffs, or equivalently, the average regret per time step should go to 0 as $T\to\infty$.
-Figure 1 delineates the precise differences between the standard model and our model.
+Table 1 delineates the precise differences between the standard model and our model.
 If the adversary’s choices are unconstrained, the problem is hopeless: if the adversary determines the correct action on each time step randomly and independently, the agent can do no better than random guessing.
 However, sublinear regret becomes possible if (1) the hypothesis class has finite Littlestone dimension (Littlestone, 1988), or (2) the hypothesis class has finite VC dimension (Vapnik & Chervonenkis, 1971) and the input is $\sigma$-smooth [^4] Informally, the adversary chooses a distribution over inputs instead of a precise input.
-See Section 3.
+See Section 3 for the formal definition.
 (Haghtalab et al., 2024).
@@
 Furthermore, standard online learning allows the agent to observe payoffs on every time step, while our agent only receives feedback on time steps with queries.
-However, access to a mentor (and local generalization) allows our agent to learn without trying actions directly, which is enough to offset all of the above disadvantages.
-Objective Regret goal Feedback Mentor Local generalization
-Standard model Sum of payoffs Sublinear Every time step No No
-Our model Product of payoffs Subconstant Only from queries Yes Yes
-Figure 1: Comparison between the standard online learning model and our model.
+However, the combination of a mentor and local generalization allows our agent to learn without trying actions directly, which is enough to offset all of the above disadvantages.
+Table 1: Comparison between the standard online learning model and our model.
+Standard model Our model
+Objective Sum of payoffs Product of payoffs
+Regret goal Sublinear Subconstant
+Feedback Every time step Only from queries
+Mentor No Yes
+Local generalization No Yes
 1.4 Our results
 At a high level, we show that avoiding catastrophe with the help of a mentor and local generalization is no harder than online learning without catastrophic risk.
-More precisely, we first show that in general, any algorithm with sublinear queries to the mentor has arbitrarily poor regret in the worst-case (Theorem 4.1).
-This means that even when the mentor can avoid catastrophe with certainty, any algorithm either needs excessive supervision or is nearly guaranteed to cause catastrophe.
-Unlike online learning where the general impossibility result is trivial (the agent might as well guess randomly given an unconstrained adversary), local generalization significantly limits the adversary’s power and necessitates a careful analysis.
-Next, we present a simple algorithm whose total regret and rate of querying the mentor both go to 0 as $T\to\infty$ when either (1) the mentor policy class has finite Littlestone dimension or (2) the mentor policy class has finite VC dimension and the input sequence is $\sigma$-smooth.
-Our algorithm can handle a unbounded input space and does not need detailed access to the feature embedding, instead relying on two simple operations.
-It does need to know the mentor policy class, as is standard in online learning.
-We initially prove the theorem for binary actions (Theorem 5.2) and then reduce learning with many actions to the binary action case (Theorem C.1).
-Along the way, we prove that the same subconstant bound holds for standard additive regret (Theorem 5.3).
-Essentially, our techniques are equally effective for maximizing the sum of payoffs and the product of payoffs.
-We emphasize the multiplicative objective due to our motivation of avoiding catastrophe, but our subconstant additive regret bound may also be of interest.
-In summary, the combination of a mentor and local generalization allows us to reduce the regret by an entire factor of $T$, resulting in subconstant regret (multiplicative or additive) instead of the typical sublinear regret.
+We first show that in general, any algorithm with sublinear queries to the mentor has unbounded regret in the worst-case (Theorem 4.1).
+As a corollary, even when the mentor can avoid catastrophe with certainty, any algorithm either needs extensive supervision or is nearly guaranteed to cause catastrophe (Corollary 4.1.1).
+Our primary result is a simple algorithm whose total regret and rate of querying the mentor both go to 0 as $T\to\infty$ when either (1) the mentor policy class has finite Littlestone dimension or (2) the mentor policy class has finite VC dimension and the input sequence is $\sigma$-smooth (Theorem 5.2).
+Conceptually, the algorithm has two components: (1) for “in-distribution” inputs, run a standard online learning algorithm (adjusted to account for only receiving feedback in response to queries), and (2) for “out-of-distribution” inputs, ask for help.
+Our algorithm can handle an unbounded input space and does not need to know the local generalization constant.
+Although we focus on the product of payoffs, we show that the results above (both positive and negative) hold for the typical additive regret as well.
+In fact, we show that multiplicative regret and additive regret are tightly related in our setting (Lemma A.1).
+In summary, the combination of local generalization and a mentor allows us to reduce the regret by an entire factor of $T$, resulting in subconstant regret (multiplicative or additive) instead of the typical sublinear regret.
 2 Related work
 Learning with irreversible costs.
-Despite the ubiquity of irreparable/irreversible costs in the real world, theoretical work on this topic remains limited.
-This may be due to the fundamental modeling question of how the agent should learn about novel inputs or actions without actually trying them.
+Despite the ubiquity of irreversible costs in the real world, theoretical work on this topic remains limited.
+This may be due to the fundamental modeling question of how the agent should learn about novel inputs or actions without trying them directly.
 The most common approach is to allow the agent to ask for help.
@@
 (2018).
-It is also worth mentioning Hadfield-Menell et al.
+We also mention Hadfield-Menell et al.
 (2017); Moldovan & Abbeel (2012); Turchetta et al.
 (2016), who utilize Bayesian inference in the context of safe (online) reinforcement learning without asking for help (and without regret bounds).
-We are only aware of two papers which theoretically address irreversibility without Bayesian inference: Grinsztajn et al.
+We are only aware of two papers that theoretically address irreversibility without Bayesian inference: Grinsztajn et al.
 (2021) and Maillard et al.
@@
 We are also not aware of prior results comparable to our negative result, including in the Bayesian regime.
-Constrained Markov Decision Processes (CMDPs).
-CMDPs (Altman, 2021) require the agent to maximize reward while also satisfying safety constraints.
+Safe reinforcement learning (RL).
+The safe RL problem is typically formulated as a constrained Markov Decision Process (CMDP) (Altman, 2021).
+In CMDPs, the agent must maximize reward while also satisfying safety constraints.
 See Gu et al.
@@
 (2024) for surveys.
-The two most relevant papers are Liu et al.
+The two most relevant safe RL papers are Liu et al.
 (2021) and Stradi et al.
@@
 See Cesa-Bianchi & Lugosi (2006) and Chapter 21 of Shalev-Shwartz & Ben-David (2014) for introductions to online learning.
-A classical result states that sublinear regret is possible iff the hypothesis class has finite Littlestone dimension (Littlestone, 1988).
-However, even some simple hypothesis classes have infinite Littlestone dimension, such as the class of thresholds on $[0,1]$ (Example 21.4 in Shalev-Shwartz & Ben-David (2014)).
+A classical result states that sublinear regret is possible if and only if the hypothesis class has finite Littlestone dimension (Littlestone, 1988).
+However, even some simple hypothesis classes have infinite Littlestone dimension, such as the class of thresholds on $[0,1]$ (Example 21.4 in Shalev-Shwartz & Ben-David, 2014).
 Recently, Haghtalab et al.
@@
 Multiplicative objectives.
-Although online learning traditionally studies the sum of payoffs, there is some work on maximizing the product of payoffs, or equivalently the sum of logarithms (Chapter 9 of Cesa-Bianchi et al.
-(2017)).
+Although online learning traditionally studies the sum of payoffs, there is some work which aims to maximize the product of payoffs (or equivalently, the sum of logarithms).
+See, e.g., Chapter 9 of Cesa-Bianchi & Lugosi (2006).
 However, these regret bounds are still sublinear in $T$, in comparison to our subconstant regret bounds.
-Also, that work still assumes that payoffs are observed on every time step, while our agent only receives feedback in response to queries (Figure 1).
+Also, like most online learning work, those results assume that payoffs are observed on every time step.
+In contrast, our agent only receives feedback in response to queries (Table 1) and never observes payoffs.
 Barman et al.
-(2023) recently provided regret bounds for a multiplicative objective in a multi-armed bandit problem, but their objective is the geometric mean of payoffs instead of the product.
+studied a multiplicative objective in a multi-armed bandit context, but their objective is the geometric mean of payoffs instead of the product.
 Interpreted in our context, their regret bounds imply that the average chance of catastrophe goes to zero, while we guarantee that the total chance of catastrophe goes to zero.
-This distinction is closely related to the difference between subconstant and sublinear regret discussed in Section 1.3.
+This distinction is closely related to the difference between subconstant and sublinear regret.
 Active learning and imitation learning.
@@
 This contrasts with passive learning, where the agent receives feedback automatically.
-Although ideas from active learning could be useful in our domain, we are not aware of any results from that literature which account for irreversible costs.
-The process of the agent learning from a mentor is also reminiscent of imitation learning (Osa et al., 2018), but we are not aware of any relevant technical implications.
+The way our agent learns from the mentor is also reminiscent of imitation learning (Osa et al., 2018).
+Although ideas from these areas could be useful in our setting, we are not aware of any results from that literature which account for irreversible costs.
 3 Model
 Inputs.
-Let $\mathbb{N}$ refer to the set of strictly positive integers and let $T\in\mathbb{N}$ be the time horizon.
-Let $\boldsymbol{x}=(x_{1},x_{2},\dots x_{T})\in\mathcal{X}^{T}$ be the sequence of inputs.
-In the adversarial setting, each $x_{t}$ can have arbitrary dependence on the events of prior time steps.
-In the smoothed setting, the adversary only chooses the distribution $x_{t}$ from which $x_{t}$ is sampled.
-Formally, a distribution $\mathcal{D}$ over $\mathcal{X}$ is $\sigma$-smooth if for any $X\subseteq\mathcal{X}$, $\mathcal{D}(X)\leq\frac{1}{\sigma}U(X)$.
-(In the smoothed setting, we assume that $\mathcal{X}$ supports a uniform distribution $U$ [^6] For example, $\mathcal{X}$ having finite Lebesgue measure is sufficient.
+Let $\mathbb{N}$ denote the strictly positive integers and let $T\in\mathbb{N}$ be the time horizon.
+Let $\boldsymbol{x}=(x_{1},x_{2},\dots,x_{T})\in\mathcal{X}^{T}$ be the sequence of inputs.
+In the fully adversarial setting, each $x_{t}$ can have arbitrary (possibly randomized) dependence on the events of prior time steps.
+In the smoothed setting, the adversary only chooses the distribution $\mathcal{D}_{t}$ from which $x_{t}$ is sampled.
+Formally, a distribution $\mathcal{D}$ over $\mathcal{X}$ is $\sigma$-smooth if for any $S\subseteq\mathcal{X}$, $\mathcal{D}(S)\leq\frac{1}{\sigma}U(S)$.
+(In the smoothed setting, we assume that $\mathcal{X}$ supports a uniform distribution $U$ [^5] For example, it suffices for $\mathcal{X}$ to have finite Lebesgue measure.
 Note that this does not imply boundedness.
-Alternatively, $\sigma$-smoothness can be defined with respect to a different distribution, as long as the Radon-Nikodym derivative is uniformly bounded; see Definition 1 of Block et al.
+Alternatively, $\sigma$-smoothness can be defined with respect to a different distribution; see Definition 1 of Block et al.
 (2022).) If each $x_{t}$ is sampled from a $\sigma$-smooth $\mathcal{D}_{t}$, we say that $\boldsymbol{x}$ is $\sigma$-smooth.
@@
 We allow the payoff function to vary between time steps: let $\boldsymbol{\mu}=(\mu_{1},\dots,\mu_{T})\in(\mathcal{X}\times\mathcal{Y}\to[0,1])^{T}$ be the sequence of payoff functions.
-[^7] Variation in $\mu_{t}$ is not a core part of our problem, but we allow it because the exact same proof holds.
-Since the mentor policy is fixed, local generalization limits how much $\mu_{t}$ can vary between time steps.
 Then $\mu_{t}(x_{t},y_{t})\in[0,1]$ is the agent’s payoff at time $t$.
 Like $\boldsymbol{\mathcal{D}}$, we allow $\boldsymbol{\mu}$ to be adaptive.
-Unless otherwise noted, all expectations are over the agent’s randomization (if any) and the randomization in $\boldsymbol{x}$ (if any).
+Unless otherwise noted, all expectations are over any randomization in the agent’s decisions, any randomization in $\boldsymbol{x}$, and any randomization in the adaptive choice of $\boldsymbol{\mu}$.
 Asking for help.
-The mentor is endowed with a (not necessarily optimal) policy $\pi^{m}:\mathcal{X}\to\mathcal{Y}$.
+The mentor is endowed with a (possibly suboptimal) policy $\pi^{m}:\mathcal{X}\to\mathcal{Y}$.
 When action $\tilde{y}$ is chosen, the mentor informs the agent of the action $\pi^{m}(x_{t})$ and the agent obtains payoff $\mu_{t}(x_{t},\pi^{m}(x_{t}))$.
@@
 Then we say that the (expected) number of queries is sublinear in $T$ if $\sup_{\boldsymbol{\mu},\pi^{m}}\operatorname*{\mathbb{E}}[|Q_{T}(\boldsymbol{\mu},\pi^{m})|]\in o(T)$.
-In other words, there must exist $g:\mathbb{N}\to\mathbb{N}$ such that $g(T)\in o(T)$ and $\sup_{\boldsymbol{\mu},\pi^{m}}\operatorname*{\mathbb{E}}[|Q_{T}(\boldsymbol{\mu},\pi^{m})|]\leq g(T)$.
-[^8] One could instead consider the worst-case number of queries, but this distinction does not affect whether subconstant regret is achievable (Proposition E.2).
+In other words, there must exist $g:\mathbb{N}\to\mathbb{N}$ which does not depend on $\boldsymbol{\mu}$ or $\pi^{m}$ such that $g(T)\in o(T)$ and $\sup_{\boldsymbol{\mu},\pi^{m}}\operatorname*{\mathbb{E}}[|Q_{T}(\boldsymbol{\mu},\pi^{m})|]\leq g(T)$.
 For brevity, we will usually write $Q_{T}=Q_{T}(\boldsymbol{\mu},\pi^{m})$.
 Local generalization.
-We assume that the mentor policy permits “local generalization”.
+We assume that $\boldsymbol{\mu}$ and $\pi^{m}$ satisfy local generalization.
 Informally, if the agent is given an input $x$, taking the mentor action for a similar input $x^{\prime}$ is almost as good as taking the mentor action for $x$.
 Formally, we assume $\mathcal{X}\subseteq\mathbb{R}^{n}$ and there exists $L>0$ such that for all $x,x^{\prime}\in\mathcal{X}$ and $t\in[T]$, $|\mu_{t}^{m}(x)-\mu_{t}(x,\pi^{m}(x^{\prime}))|\leq L||x-x^{\prime}||$, where |||||$\cdot$||| | ⋅ | | denotes Euclidean distance.
-[^9] One could also allow a generic metric space; our assumption of $\mathcal{X}\subseteq\mathbb{R}^{n}$ is only for convenience.
 This represents the ability to transfer knowledge between similar inputs:
-$|\underbrace{\mu_{t}(x,\pi^{m}(x))}_{\text{Taking the right action}}-\underbrace{\mu_{t}(x,\pi^{m}(x^{\prime}))}_{\text{Using what you learned in $x^{\prime}$}}|\ \ \ \leq\underbrace{L||x-x^{\prime}||}_{\text{Similarity between $x$ and $x^{\prime}$}}$
-Borrowing knowledge from similar experiences seems fundamental to learning and is well-understood in the psychology literature (Esser et al., 2023) and education literature (Hajian, 2019).
-Crucially, our input space can be any feature embedding of the agent’s situation, not just its physical positioning.
-Our algorithms will not require knowledge of the feature embedding and do not need to know $L$, so it suffices that there exists some feature embedding which satisfies local generalization.
-The agent does not even need to know which embedding it is.
-Finally, local generalization implies the more familiar Lipschitz continuity for an optimal mentor (Proposition E.1).
-Multiplicative objective and regret.
-If $\mu_{t}(x_{t},y_{t})\in[0,1]$ is the chance of avoiding catastrophe on time step $t$ (conditioned on no prior catastrophe), then $\prod_{t=1}^{T}\mu_{t}(x_{t},y_{t})$ is the agent’s overall chance of avoiding catastrophe by the chain rule of probability.
-For given $\boldsymbol{x},\boldsymbol{y},\boldsymbol{\mu},\pi^{m}$, the agent’s regret is
-$R_{T}(\boldsymbol{x},\boldsymbol{y},\boldsymbol{\mu},\pi^{m})=\prod_{t=1}^{T}\mu_{t}^{m}(x_{t})-\prod_{t=1}^{T}\mu_{t}(x_{t},y_{t})$
-We write $R_{T}=R_{T}(\boldsymbol{x},\boldsymbol{y},\boldsymbol{\mu},\pi^{m})$ for brevity.
-We desire subconstant worst-case regret: the total (not average) expected regret should go to 0 for any $\boldsymbol{\mu}$ and $\pi^{m}$; formally, $\lim_{T\to\infty}\sup_{\boldsymbol{\mu},\pi^{m}}\operatorname*{\mathbb{E}}[R_{T}]=0$.
-The interpretation of a bound on $\operatorname*{\mathbb{E}}[R_{T}]$ depends on the quality of the mentor.
-For example, subconstant regret becomes trivial if $\lim_{T\to\infty}\operatorname*{\mathbb{E}}\big{[}\prod_{t=1}^{T}\mu_{t}^{m}(x_{t})\big{]}=0$.
-However, we think that high-stakes AI applications should ensure the presence of a mentor who is almost always safe, i.e., $\operatorname*{\mathbb{E}}\big{[}\prod_{t=1}^{T}\mu_{t}^{m}(x_{t})\big{]}\approx 1$.
-If no such mentor exists for some application, perhaps it is better to avoid the application altogether.
-Also, our regret bounds include rates of convergence, so even if the mentor policy will eventually cause catastrophe, we can still bound how quickly the agent becomes unsafe.
+$\big{|}\!\!\underbrace{\mu_{t}(x,\pi^{m}(x))}_{\text{Taking the right action}}-\!\underbrace{\mu_{t}(x,\pi^{m}(x^{\prime}))}_{\text{Using what you learned in $x^{\prime}$}}\!\!\!\!\!\!\!\big{|}\,\leq\>\underbrace{L||x-x^{\prime}||}_{\text{Input similarity}}$
+This ability seems fundamental to intelligence and is well-understood in psychology (e.g., Esser et al., 2023) and education (e.g., Hajian, 2019).
+Note that the input space $\mathcal{X}\subseteq\mathbb{R}^{n}$ can be any encoding of the agent’s situation, not just its physical positioning.
+See Section 5.2 for further discussion.
+All suprema over $\boldsymbol{\mu},\pi^{m}$ pairs are assumed to be restricted to $\boldsymbol{\mu},\pi^{m}$ pairs which satisfy local generalization.
+Regret.
+If $\mu_{t}(x_{t},y_{t})\in[0,1]$ is the chance of avoiding catastrophe at time $t$ (conditioned on no prior catastrophe), then by the chain rule of probability, $\prod_{t=1}^{T}\mu_{t}(x_{t},y_{t})$ is the agent’s overall chance of avoiding catastrophe.
+For given $\boldsymbol{x},\boldsymbol{y},\boldsymbol{\mu},\pi^{m}$, the agent’s multiplicative regret [^6] One could also define the multiplicative regret as $R_{T}^{\prime}=\prod_{t=1}^{T}\mu_{t}^{m}(x_{t})-\prod_{t=1}^{T}\mu_{t}(x_{t},y_{t})$, but our definition is actually stricter: $\lim_{T\to\infty}R_{T}^{\times}\to 0$ implies $\lim_{T\to\infty}R_{T}^{\prime}\to 0$, while the reverse is not true.
+In particular, $\lim_{T\to\infty}R_{T}^{\prime}\to 0$ is trivial if $\lim_{T\to\infty}\prod_{t=1}^{T}\mu_{t}^{m}(x_{t})\to 0$.
+is
+$R_{T}^{\times}(\boldsymbol{x},\boldsymbol{y},\boldsymbol{\mu},\pi^{m})=\log\prod_{t=1}^{T}\mu_{t}^{m}(x_{t})-\log\prod_{t=1}^{T}\mu_{t}(x_{t},y_{t})$
+when all payoffs are strictly positive.
+To handle the case where some payoffs are zero, we assume the existence of $\mu_{0}^{m}>0$ such that $\mu^{m}(x_{t})\geq\mu_{0}^{m}$ always.
+Thus only the agent’s payoffs can be zero, so we can safely define $R_{T}^{\times}(\boldsymbol{x},\boldsymbol{y},\boldsymbol{\mu},\pi^{m})=\infty$ whenever $\mu_{t}(x_{t},y_{t})=0$ for some $t\in[T]$.
+We write $R_{T}^{\times}=R_{T}^{\times}(\boldsymbol{x},\boldsymbol{y},\boldsymbol{\mu},\pi^{m})$ for brevity.
+The assumption of $\mu^{m}(x_{t})\geq\mu_{0}^{m}>0$ means that the mentor cannot be abysmal.
+In fact, we argue that high-stakes AI applications should employ a mentor who is almost always safe, i.e., $\mu_{0}^{m}\approx 1$.
+If no such mentor exists for some application, perhaps that application should be avoided altogether.
+We also define the agent’s additive regret as
+$R_{T}^{+}(\boldsymbol{x},\boldsymbol{y},\boldsymbol{\mu},\pi^{m})=\sum_{t=1}^{T}\mu_{t}^{m}(x_{t})-\sum_{t=1}^{T}\mu_{t}(x_{t},y_{t})$
+and similarly write $R_{T}^{+}=R_{T}^{+}(\boldsymbol{x},\boldsymbol{y},\boldsymbol{\mu},\pi^{m})$ for brevity.
+For both objectives, we desire subconstant worst-case regret: the total (not average) expected regret should go to 0 for any $\boldsymbol{\mu}$ and $\pi^{m}$.
+Formally, we want $\lim_{T\to\infty}\sup_{\boldsymbol{\mu},\pi^{m}}\operatorname*{\mathbb{E}}[R_{T}^{\times}]=0$ and $\lim_{T\to\infty}\sup_{\boldsymbol{\mu},\pi^{m}}\operatorname*{\mathbb{E}}[R_{T}^{+}]=0$.
 VC and Littlestone dimensions.
 VC dimension (Vapnik & Chervonenkis, 1971) and Littlestone dimension (Littlestone, 1988) are standard measures of learning difficulty which capture the ability of a hypothesis class (in our case, a policy class) to realize arbitrary combinations of labels (in our case, actions).
-We omit the precise dimensions since we only utilize these concepts via existing results.
+We omit the precise definitions since we only utilize these concepts via existing results.
 See Shalev-Shwartz & Ben-David (2014) for a comprehensive overview.
-Misc.
-The diameter of a set $X\subseteq\mathcal{X}$ is defined by $\operatorname*{diam}(X)=\max_{x,x^{\prime}\in X}||x-x^{\prime}||$. All logarithms and exponents are base $e$ unless otherwise noted.
-4 Avoiding catastrophe is impossible in general
-We begin by showing that in general, any algorithm with sublinear mentor queries has arbitrary poor regret in the worst-case, even when inputs are i.i.d.
-on $[0,1]$.
-The result also holds even if the algorithm knows $L$ and $\boldsymbol{x}$ ahead of time.
-The proof can be found in Appendix A.
-Theorem 4.1.
-The worst-case expected regret of any algorithm with sublinear queries goes to 1 as $T$ goes to infinity.
-Formally, $\lim_{T\to\infty}\ \,\sup_{\boldsymbol{\mu},\pi^{m}}\,\operatorname*{\mathbb{E}}[R_{T}]=1$.
-4.1 Intuition
-We partition $\mathcal{X}$ into equally-sized sections that are “independent” in the sense that querying an input in section $i$ gives you no information about section $j$.
-The number of sections is determined by a function $f:\mathbb{N}\to\mathbb{N}$ that we will choose.
-If $|Q_{T}|\in o(f(T))$, most of these sections will never contain a query.
-When the agent sees an input in a section not containing a query, it essentially has to guess, meaning it will be wrong a constant fraction of the time.
-We then choose a payoff function (which is the same for all time steps) which makes the wrong guesses as costly as possible, subject to the local generalization constraint.
-Figure 2 fleshes out this idea.
-Picking $\boldsymbol{f(T)}$.
-A natural idea is to try $f(T)=T$, but this doesn’t quite work: even if the agent chooses wrong on every time step, the minimum payoff is still at least $1-\frac{L}{2T}$, and $\lim_{T\to\infty}\prod_{t=1}^{T}\left(1-\frac{L}{2T}\right)=\lim_{T\to\infty}\left(1-\frac{L}{2T}\right)^{T}=e^{-L/2}$.
-In order for the regret to approach 1, we need $f(T)$ to be asymptotically between $|Q_{T}|$ and $T$ (such $f$ must exist since $|Q_{T}|\leq g(T)\in o(T)$).
-This leads to the following bound: $\prod_{t=1}^{T}\mu(x_{t},y_{t})\leq\big{(}1-\frac{L}{\Theta(f(T))}\big{)}^{\Theta(T)}$.
-When $f(T)\in o(T)$, the right hand side converges to 0, while $\prod_{t=1}^{T}\mu^{m}(x_{t})=1$.
-In words, the agent is nearly guaranteed to cause catastrophe, despite the existence of a policy which is guaranteed to avoid catastrophe.
-$x$$\mu$  $\mu(x,0)$$\mu(x,1)$$\cfrac{L}{2f(T)}$$\frac{1}{f(T)}$$\frac{2}{f(T)}$$\frac{3}{f(T)}$$\frac{f(T)-1}{f(T)}$$\dots$1001111
+Misc. The diameter of a set $S\subseteq\mathcal{X}$ is defined by $\operatorname*{diam}(S)=\max_{x,x^{\prime}\in S}||x-x^{\prime}||$. All logarithms and exponents are base $e$ unless otherwise noted.
+input $x$payoff $\mu$   $\mu(x,0)$$\mu(x,1)$$\cfrac{L}{2f(T)}$$\frac{1}{f(T)}$$\frac{2}{f(T)}$$\frac{3}{f(T)}$$\frac{f(T)-1}{f(T)}$$\dots$1001111
 Figure 1: An illustration of the construction we use to prove Theorem 4.1 (not to scale).
-The horizontal axis indicates the input $x\in[0,1]$ and the vertical axis indicates the payoff $\mu(x,y)\in[0,1]$.
+The horizontal axis indicates the input $x\in[0,1]=\mathcal{X}$ and the vertical axis indicates the payoff $\mu(x,y)\in[0,1]$.
 The solid line represents $\mu(x,0)$ and the dotted line represents $\mu(x,1)$.
@@
 As a result, we can freely toggle the optimal action for each section independently.
+4 Avoiding catastrophe with sublinear queries is impossible in general
+We first show that in general, any algorithm with sublinear mentor queries has unbounded regret in the worst-case, even when inputs are i.i.d.
+on $[0,1]$ and $\boldsymbol{\mu}$ does not vary over time.
+The formal proofs are deferred to Appendix A, but we provide intuition and define the construction here.
+Theorem 4.1.
+Any algorithm with sublinear queries has unbounded worst-case regret (both multiplicative and additive) as $T\to\infty$. Specifically,
+$\sup_{\boldsymbol{\mu},\pi^{m}}\operatorname*{\mathbb{E}}[R_{T}^{\times}],\,\sup_{\boldsymbol{\mu},\pi^{m}}\operatorname*{\mathbb{E}}[R_{T}^{+}]\in\Omega\left(\!L\sqrt{\frac{T}{\sup_{\boldsymbol{\mu},\pi^{m}}\operatorname*{\mathbb{E}}[|Q_{T}|]\!+\!1}}\right)$
+Intuitively, the regret decreases as the number of queries increases. However, as long as the number of queries remains sublinear in $T$, the regret is unbounded as $T\to\infty$.
+We also have the following corollary of Theorem 4.1:
+Corollary 4.1.1.
+Even when $\mu_{t}^{m}(x)=1$ for all $t$ and $x$, any algorithm with sublinear queries satisfies
+$\lim_{T\to\infty}\,\sup_{\boldsymbol{\mu},\pi^{m}}\,\operatorname*{\mathbb{E}}\left[\prod_{t=1}^{T}\mu_{t}(x_{t},y_{t})\right]=0$
+In other words, even if the mentor never causes catastrophe, any algorithm with sublinear queries causes catastrophe with probability 1 as $T\to\infty$ in the worst case.
+4.1 Intuition
+We partition $\mathcal{X}$ into equally-sized sections that are “independent” in the sense that querying an input in section $i$ provides no information about section $j$.
+There will be $f(T)$ sections, where $f$ is a function that we will choose.
+If $|Q_{T}|\in o(f(T))$, most of these sections will never contain a query.
+When the agent sees an input in a section not containing a query, it essentially must guess, meaning it will be wrong about half the time.
+We then choose a payoff function (which is the same for all time steps) which makes the wrong guesses as costly as possible, subject to the local generalization constraint.
+Figure 1 fleshes out this idea.
+The choice of $f$ is crucial. One idea is $f(T)=T$. If the agent is wrong about half the time, and the average payoff for wrong actions is $1-\frac{L}{4T}$, we can estimate the regret as
+$R_{T}^{\times}\,=$ $\ \log\prod_{t=1}^{T}\mu_{t}^{m}(x_{t})-\log\prod_{t=1}^{T}\mu_{t}(x_{t},y_{t})$
+$\approx$ $\ \log 1-\log\left(1-\frac{L}{4T}\right)^{T/2}$
+$=$ $\ -\frac{T}{2}\log\left(1-\frac{L}{4T}\right)$
+$\approx$ $\ \frac{T}{2}\cdot\frac{L}{4T}$
+$=$ $\ \frac{L}{8}$
+Thus $f(T)=T$ can at best give us a constant lower bound on regret.
+Instead, we choose $f$ such that $|Q_{T}|\in o(f(T))$ and $f(T)\in o(T)$.
+Specifically, we choose $f(T)=\max(\sqrt{\sup_{\boldsymbol{\mu},\pi^{m}}\operatorname*{\mathbb{E}}[|Q_{T}|]T},1)$.
+Most sections still will not contain a query, so the agent is still wrong about half the time, but the payoff for wrong actions is worse.
+Then
+$R_{T}^{\times}\,\approx$ $\ \log 1-\log\left(1-\frac{L}{4f(T)}\right)^{T/2}$
+$\approx$ $\ \frac{LT}{8f(T)}$
+$\in$ $\ \Omega\left(L\sqrt{\frac{T}{\sup_{\boldsymbol{\mu},\pi^{m}}\operatorname*{\mathbb{E}}[|Q_{T}|]+1}}\right)$
+which produces the bound in Theorem 4.1.
 VC dimension.
-The class of mentor policies induced by our construction has VC dimension $f(T)$; considered over all possible values of $T$, this implies infinite VC (and Littlestone) dimension.
-This is necessary given our positive results in Section 5.
+The class of mentor policies in our construction has VC dimension $f(T)$; across all possible values of $T$, this implies infinite VC (and Littlestone) dimension.
+We know that this is necessary given our positive results.
 4.2 Formal definition of construction
-Let $\mathcal{X}=[0,1]$ and $\mathcal{D}_{t}=U$ for each $t\in[T]$.
+Let $\mathcal{X}=[0,1]$ and $\mathcal{D}_{t}=U(\mathcal{X})$ for each $t\in[T]$, where $U(\mathcal{X})$ is the uniform distribution on $\mathcal{X}$.
 Assume that $L\leq 1$; this will simplify the math and only makes the problem easier for the agent.
@@
 The bit $a_{j}$ will denote the optimal action in section $j$.
-Note that $f(T)\geq 1$ and since we defined $\mathbb{N}$ to exclude 0.
-[…]
+Note that $f(T)\geq 1$.
+For each $j\in[f(T)]$, we refer to $X_{j}=\left[\frac{j-1}{f(T)},\frac{j}{f(T)}\right]$ as the $j$th section.
+Let $m_{j}=\frac{j-0.5}{f(T)}$ be the midpoint of $X_{j}$.
+Assume that each $x_{t}$ belongs to exactly one $X_{j}$ (this happens with probability 1, so this assumption does not affect the expected regret).
+Let $j(x)$ denote the index of the section containing input $x$.
+Then $\mu_{f,\boldsymbol{a}}$ is defined by
 $\mu_{f,\boldsymbol{a}}(x,y)=\begin{cases}1&\textnormal{ if }y=a_{j(x)}\\ 1-L\left(\cfrac{1}{2f(T)}-|m_{j(x)}-x|\right)&\textnormal{ if }y\neq a_{j(x)}\\ \end{cases}$
-We will use this payoff function for all time steps: $\mu_{t}=\mu_{f,\boldsymbol{a}}$ for all $t\in[T]$.
-Let $\pi^{m}$ be any policy which is optimal for $\mu_{f,\boldsymbol{a}}$.
+We use this payoff function for all time steps: $\mu_{t}=\mu_{f,\boldsymbol{a}}$ for all $t\in[T]$.
+Let $\pi^{m}$ be any optimal policy for $\mu_{f,\boldsymbol{a}}$.
 Note that there is a unique optimal action for each $x_{t}$, since each $x_{t}$ belongs to exactly one $X_{j}$; formally, $\pi^{m}(x_{t})=a_{j(x_{t})}$.
@@
 5 Avoiding catastrophe given finite VC or Littlestone dimension
-Theorem 4.1 shows that avoiding catastrophe is impossible in general, which is also true in online learning.
+Theorem 4.1 shows that avoiding catastrophe is impossible in general.
 What if we restrict ourselves to settings where standard online learning is possible?
 Specifically, we assume that $\pi^{m}$ belongs to a policy class $\Pi$ where either (1) $\Pi$ has finite VC dimension $d$ and $\boldsymbol{x}$ is $\sigma$-smooth or (2) $\Pi$ has finite Littlestone dimension $d$ [^7] Recall from Section 1.3 that standard online learning becomes tractable under either of these assumptions.
-This section presents a simple algorithm which guarantees subconstant regret and sublinear queries under either of those assumptions.
-Our algorithm needs to know $\Pi$, as is standard in online learning.
-The algorithm does not need to know $\sigma$ (in the smooth case) or $L$, and can handle an unbounded input space (the number of queries simply scales with the maximum distance between observed inputs).
-For simplicity, we initially prove our result for $\mathcal{Y}=\{0,1\}$ (the proof can be found in Appendix B).
-Appendix C extends our result to many actions using the standard “one versus rest” reduction.
-[^11] For each action $y$, we learn a binary classifier which predicts whether $\pi^{m}(x)=y$.
-If every binary classifier is correct, we can correctly determine $\pi^{m}(x)$.
-See, e.g., Chapter 29 of Shalev-Shwartz & Ben-David (2014).
+This section presents a simple algorithm which guarantees subconstant regret (both multiplicative and additive) and sublinear queries under either of those assumptions.
+Formal proofs are deferred to Appendix B but we provide intuition and a proof sketch here.
 5.1 Intuition behind the algorithm
 Algorithm 1 has two simple components: (1) run a modified version of the Hedge algorithm for online learning, but (2) ask for help for unfamiliar inputs (specifically, when the input is very different from any queried input with the same action under the proposed policy).
-Hedge ensures that the number of time steps where the agent’s action doesn’t match the mentor’s is small, and asking for help for unfamiliar inputs ensures that when we do make a mistake, the cost isn’t too high.
+Hedge ensures that the number of mistakes (i.e., the number of time steps where the agent’s action doesn’t match the mentor’s) is small, and asking for help for unfamiliar inputs ensures that when we do make a mistake, the cost isn’t too high.
 This algorithmic structure seems quite natural: mostly follow a baseline strategy, but ask for help when out-of-distribution.
-Agnostic to feature embedding.
-The algorithm does not need to know how inputs are represented in $\mathbb{R}^{n}$, since the only nontrivial operation it performs on the input space is computing a nearest neighbor distance (line 21 of Algorithm 1).
-This could be modeled as an out-of-distribution detector score, for which many methods are available (see e.g., Yang et al.
-(2024)).
-This agnosticism reduces the burden on the local generalization assumption: it suffices that there exists some feature embedding satisfying this assumption.
 Hedge.
 Hedge (Freund & Schapire, 1997) is a standard online learning algorithm which ensures sublinear regret when the number of hypotheses (in our case, the number of policies in $\Pi$) is finite.
-[^12] See Chapter 5 of Slivkins et al.
-(2019) and Chapter 21 of Shalev-Shwartz & Ben-David (2014) for modern introductions to Hedge.
+[^8] Chapter 5 of Slivkins et al.
+(2019) and Chapter 21 of Shalev-Shwartz & Ben-David (2014) give modern introductions to Hedge.
 We would prefer not to assume that $\Pi$ is finite.
-Luckily, any policy $\Pi$ can be approximated within $\varepsilon$ when either (1) $\Pi$ has finite VC dimension and $\boldsymbol{x}$ is $\sigma$-smooth or (2) $\Pi$ has finite Littlestone dimension.
+Luckily, any policy in $\Pi$ can be approximated within $\varepsilon$ when either (1) $\Pi$ has finite VC dimension and $\boldsymbol{x}$ is $\sigma$-smooth or (2) $\Pi$ has finite Littlestone dimension.
 Thus we can run Hedge on this approximative policy class instead.
@@
 (2024) analyze exactly this modification of Hedge.
-1:function AvoidCatastrophe($T\in\mathbb{N},\>\varepsilon\in\mathbb{R}_{>0},\>d\in\mathbb{N},\,$ policy class $\Pi$)
-2:     if $\Pi$ has VC dimension $d$ then
-3:         $\tilde{\Pi}\leftarrow$ any smooth $\varepsilon$-cover of $\Pi$ of size at most $(41/\varepsilon)^{d}$triangleright See Definition 5.4
-4:     else
-5:         $\tilde{\Pi}\leftarrow$ any adversarial cover of size at most $(eT/d)^{d}$triangleright See Definition 5.5
-6:     $X\leftarrow\emptyset$
-7:     $w(\pi)\leftarrow 1$ for all $\pi\in\tilde{\Pi}$
-8:     $p\leftarrow 1/\sqrt{\varepsilon T}$
-9:     $\eta\leftarrow\max\big{(}\sqrt{\frac{p\log|\tilde{\Pi}|}{2T}},\>\frac{p^{2}}{\sqrt{2}}\big{)}$
-10:     for $t$
-from
-$1$
-to
-$T$ do $\triangleright$ Run one step of Hedge, which selects policy $\pi_{t}$
+5.2 Local generalization
+Local generalization is vital: this is what allows us to detect when an input is unfamiliar.
+Crucially, our algorithm does not need to know how inputs are encoded in $\mathbb{R}^{n}$ and does not need to know $L$: it only needs to be able to compute the nearest-neighbor distance $\min_{(x,y)\in S:y=\pi_{t}(x_{t})}||x_{t}-x||$.
+Thus we only need to assume that there exists some encoding satisfying local generalization.
+To elaborate, recall the example that a 3 mm spot and a 3.1 mm spot on X-rays likely have similar risk levels (assuming similar density, location, etc.).
+If the risk level abruptly increases for any spot over 3 mm, then local generalization may not hold for a naive encoding which treats size as a single dimension.
+However, a more nuanced encoding would recognize that these two situations – a 3 mm vs 3.1 mm spot – are in fact not similar.
+Constructing a suitable encoding may be challenging, but we do not require the agent to have explicit access to such an encoding: the agent only needs a nearest-neighbor distance oracle.
+  Inputs: $T\in\mathbb{N},\>\varepsilon\in\mathbb{R}_{>0},\>d\in\mathbb{N},\,$ policy class $\Pi$
+  if $\Pi$ has VC dimension $d$ then
+     $\tilde{\Pi}\leftarrow$ any smooth $\varepsilon$-cover of $\Pi$ of size at most $(41/\varepsilon)^{d}$  (see Definition 5.3)
+  else
+     $\tilde{\Pi}\leftarrow$ any adversarial cover of size at most $(eT/d)^{d}$  (see Definition 5.4)
+  $S\leftarrow\emptyset$
+  $w(\pi)\leftarrow 1$ for all $\pi\in\tilde{\Pi}$
+  $p\leftarrow 1/\sqrt{\varepsilon T}$
+  $\eta\leftarrow\max\big{(}\sqrt{\frac{p\log|\tilde{\Pi}|}{2T}},\>\frac{p^{2}}{\sqrt{2}}\big{)}$
+  for $t$ from $1$ to $T$ do
+     Run one step of Hedge, which selects policy $\pi_{t}$
 […]
-12:         if hedgeQuery then
-13:              Query mentor and observe $\pi^{m}(x_{t})$
-14:              $\ell(t,\pi)\leftarrow\mathbf{1}(\pi(x_{t})\neq\pi^{m}(x_{t}))$ for all $\pi\in\tilde{\Pi}$
-15:              $\ell^{*}\leftarrow\min_{\pi\in\tilde{\Pi}}\ell(t,\pi)$
-16:              $w(\pi)\leftarrow w(\pi)\cdot\exp(-\eta(\ell(t,\pi)-\ell^{*}))$ for all $\pi\in\tilde{\Pi}$
-17:              $\pi_{t}\leftarrow\operatorname*{arg\,min}_{\pi\in\tilde{\Pi}}\ell(t,\pi)$
-18:         else
-19:              $P(\pi)\leftarrow w(\pi)/\sum_{\pi^{\prime}\in\tilde{\Pi}}w(\pi^{\prime})$ for all $\pi\in\tilde{\Pi}$
-20:              Sample $\pi_{t}\sim P$
-21:         if $\min_{(x,y)\in X:y=\pi_{t}(x_{t})}||x_{t}-x||>\varepsilon^{1/n}$ then $\triangleright$ Ask for help if out-of-distribution
-22:              Query mentor and observe $\pi^{m}(x_{t})$
-23:              $X\leftarrow X\cup\{(x_{t},\pi^{m}(x_{t}))\}$
-24:         $\triangleright$ Otherwise, follow Hedge’s chosen policy
-25:              Take action $\pi_{t}(x_{t})$
+[…]
+     if hedgeQuery [^9] The reader may notice that we do not update $S$ in this case.
+This is simply because those updates are not necessary for the desired bounds and omitting these updates simplifies the analysis.
+then
+        Query mentor and observe $\pi^{m}(x_{t})$
+        $\ell(t,\pi)\leftarrow\mathbf{1}(\pi(x_{t})\neq\pi^{m}(x_{t}))$ for all $\pi\in\tilde{\Pi}$
+        $\ell^{*}\leftarrow\min_{\pi\in\tilde{\Pi}}\ell(t,\pi)$
+        $w(\pi)\leftarrow w(\pi)\cdot\exp(-\eta(\ell(t,\pi)-\ell^{*}))$ for all $\pi\in\tilde{\Pi}$
+        $\pi_{t}\leftarrow\operatorname*{arg\,min}_{\pi\in\tilde{\Pi}}\ell(t,\pi)$
+     else
+        $P(\pi)\leftarrow w(\pi)/\sum_{\pi^{\prime}\in\tilde{\Pi}}w(\pi^{\prime})$ for all $\pi\in\tilde{\Pi}$
+        Sample $\pi_{t}\sim P$
+     if $S=\emptyset$ or $\min_{(x,y)\in S:y=\pi_{t}(x_{t})}||x_{t}-x||>\varepsilon^{1/n}$ then
+        Ask for help if out-of-distribution
+        Query mentor (if not already queried this round) and observe $\pi^{m}(x_{t})$
+        $S\leftarrow S\cup\{(x_{t},\pi^{m}(x_{t}))\}$
+     else
+        Otherwise, follow Hedge’s chosen policy
+        Take action $\pi_{t}(x_{t})$
 Algorithm 1 successfully avoids catastrophe assuming finite VC or Littlestone dimension.
-We prove the following theorem parametrized by $\varepsilon$:
+Conceptually, the algorithm only needs to be able to detect when an input is unfamiliar.
+While this task remains far from trivial, we argue that it is more tractable than fully constructing a suitable encoding.
+See Section 6 for a discussion of potential future work on this topic.
+We note that these encoding-related questions apply similarly to the more standard assumption of Lipschitz continuity.
+In fact, Lipschitz continuity implies local generalization when the mentor is optimal (Proposition E.1).
+We also mention that without local generalization, avoiding catastrophe is impossible even when the mentor policy class has finite VC dimension and $\boldsymbol{x}$ is $\sigma$-smooth (Theorem E.2).
+5.3 Main result
+For simplicity, here we only state our results for $\mathcal{Y}=\{0,1\}$; Appendix C extends our result to many actions using the standard “one versus rest” reduction.
+We first prove regret and query bounds parametrized by $\varepsilon$:
 Theorem 5.1.
-Let $\mathcal{Y}=\{0,1\}$.
-Assume $\pi^{m}\in\Pi$ where either (1) $\Pi$ has finite VC dimension $d$, $\boldsymbol{x}$ is $\sigma$-smooth, and $\varepsilon T\log T>12\sigma d\log(4e^{2}/\varepsilon)$ or (2) $\Pi$ has finite Littlestone dimension $d$.
-Then for any $T\in\mathbb{N}$ and $\varepsilon>0$, Algorithm 1 satisfies
-$\operatorname*{\mathbb{E}}\left[R_{T}\right]\in$ $\ O\left(\frac{dL}{\sigma}T\varepsilon^{1+1/n}\log(1/\varepsilon)\log T\right)$
-$\operatorname*{\mathbb{E}}[|Q_{T}|]\in$ $\ O\left(\sqrt{\frac{T}{\varepsilon}}+\frac{d}{\sigma}T\varepsilon\log(1/\varepsilon)\log T+\frac{\operatorname*{diam}(\boldsymbol{x})^{n}}{\varepsilon}\right)$
+[…]
+$\operatorname*{\mathbb{E}}\left[R_{T}^{\times}\right]\in$ $\ O\left(\frac{dL}{\sigma\mu_{0}^{m}}T\varepsilon^{1+1/n}\log(T+1/\varepsilon)\right)$
+$\operatorname*{\mathbb{E}}\left[R_{T}^{+}\right]\in$ $\ O\left(\frac{dL}{\sigma}T\varepsilon^{1+1/n}\log(T+1/\varepsilon)\right)$
+$\operatorname*{\mathbb{E}}[|Q_{T}|]\in$ $\ O\left(\!\sqrt{\frac{T}{\varepsilon}}+\frac{d}{\sigma}T\varepsilon\log(T\!+\!1/\varepsilon)+\frac{\operatorname*{\mathbb{E}}[\operatorname*{diam}(\boldsymbol{x})^{n}]}{\varepsilon}\right)$
 In Case 1, the expectation is over the randomness of both $\boldsymbol{x}$ and the algorithm, while in Case 2, the expectation is over only the randomness of the algorithm.
-Also, $R_{T}$ and $Q_{T}$ clearly have no dependence on $\sigma$ in Case 2, but we include $\sigma$ anyway to avoid writing two separate bounds.
-To obtain subconstant regret and sublinear queries, we can choose $\varepsilon=T^{\frac{-2n}{2n+1}}$.
-This also satisfies the requirement of $\varepsilon T\log T>12\sigma d\log(4e^{2}/\varepsilon)$ for large enough $T$.
+The bounds clearly have no dependence on $\sigma$ in Case 2, but we include $\sigma$ anyway to avoid writing two separate sets of bounds.
+To obtain subconstant regret and sublinear queries, we can choose $\varepsilon=T^{\frac{-2n}{2n+1}}$. This also satisfies $2L\varepsilon^{1/n}\leq\mu_{0}^{m}$ for large enough $T$.
 Theorem 5.2.
@@
 Then for any $T\in\mathbb{N}$, Algorithm 1 with $\varepsilon=T^{\frac{-2n}{2n+1}}$ satisfies
-$\operatorname*{\mathbb{E}}\left[R_{T}\right]\in$ $\ O\left(\frac{dL}{\sigma}T^{\frac{-1}{2n+1}}\log T\right)$
-$\operatorname*{\mathbb{E}}[|Q_{T}|]\in$ $\ O\left(T^{\frac{4n+1}{4n+2}}\left(\frac{d}{\sigma}\log T+\operatorname*{diam}(\boldsymbol{x})^{n}\right)\right)$
-Although our focus is the product of payoffs, Algorithm 1 also guarantees subconstant additive regret:
-Theorem 5.3.
-Let $\mathcal{Y}=\{0,1\}$.
-Assume $\pi^{m}\in\Pi$ where either (1) $\Pi$ has finite VC dimension $d$ and $\boldsymbol{x}$ is $\sigma$-smooth or (2) $\Pi$ has finite Littlestone dimension $d$.
-Then for any $T\in\mathbb{N}$, Algorithm 1 with $\varepsilon=T^{\frac{-2n}{2n+1}}$ satisfies
-$\operatorname*{\mathbb{E}}\left[\sum_{t=1}^{T}\mu_{t}^{m}(x_{t})-\sum_{t=1}^{T}\mu_{t}(x_{t},y_{t})\right]\in O\left(\frac{dL}{\sigma}T^{\frac{-1}{2n+1}}\log T\right)$
+$\operatorname*{\mathbb{E}}\left[R_{T}^{\times}\right]\in$ $\ O\left(\frac{dL}{\sigma\mu_{0}^{m}}T^{\frac{-1}{2n+1}}\log T\right)$
+$\operatorname*{\mathbb{E}}\left[R_{T}^{+}\right]\in$ $\ O\left(\frac{dL}{\sigma}T^{\frac{-1}{2n+1}}\log T\right)$
+$\operatorname*{\mathbb{E}}[|Q_{T}|]\in$ $\ O\left(T^{\frac{4n+1}{4n+2}}\left(\frac{d}{\sigma}\log T+\operatorname*{\mathbb{E}}[\operatorname*{diam}(\boldsymbol{x})^{n}]\right)\right)$
+Before proceeding to the proof sketch, we highlight some advantages of our algorithm.
+Limited knowledge required.
+Our algorithm needs to know $\Pi$, as is standard.
+However, the algorithm does not need to know $\sigma$ (in the smooth case) or $L$.
+Although Algorithm 1 as written does require $T$ as an input, it can be converted into an infinite horizon/anytime algorithm via the standard “doubling trick” (see, e.g., Slivkins et al., 2019).
+Unbounded environment.
+Our algorithm can handle an unbounded input space: the number of queries simply scales with the maximum distance between observed inputs in the form of $\operatorname*{\mathbb{E}}[\operatorname*{diam}(\boldsymbol{x})^{n}]$.
+Simultaneous bounds for all $\boldsymbol{\mu}$.
+Recall that the agent never observes payoffs and only learns from mentor queries.
+This means that the agent’s behavior does not depend on $\boldsymbol{\mu}$ at all.
+In other words, the distribution of $(\boldsymbol{x},\boldsymbol{y})$ depends on $\pi^{m}$ but not $\boldsymbol{\mu}$.
+Consequently, a given $\pi^{m}$ induces a single distribution $(\boldsymbol{x},\boldsymbol{y})$ which satisfies the bounds in Theorem 5.2 simultaneously for all $\boldsymbol{\mu}$ satisfying local generalization.
 5.4 Proof sketch
 The formal proof of Theorem 5.1 can be found in Appendix B, but we outline the key elements here.
-The regret analysis consists of two ingredients: analyzing the Hedge component, and analyzing the “ask for help when out-of-distrubtion” component.
+The regret analysis consists of two ingredients: analyzing the Hedge component and analyzing the “ask for help when out-of-distribution” component.
 The former will bound the number of mistakes made by the algorithm (i.e., the number of time steps where the agent’s action doesn’t match the mentor’s), and the latter will bound the cost of any single mistake.
-We must also carefully show that the latter does not result in excessively many queries, which we do via a novel packing argument.
-We begin by formalizing two notion of approximating a policy class:
+We must also show that the latter does not result in excessively many queries, which we do via a novel packing argument.
+We begin by formalizing two notions of approximating a policy class:
 Definition 5.3.
 Let $U$ be the uniform distribution over $\mathcal{X}$.
-For $\varepsilon>0$, a policy class $\tilde{\Pi}$ is a smooth $\varepsilon$-cover of a policy class $\Pi$ is for every $\pi\in\Pi$, there exists $\tilde{\pi}\in\tilde{\Pi}$ such that $\Pr_{x\sim U}[\pi(x)\neq\tilde{\pi}(x)]\leq\varepsilon$.
+For $\varepsilon>0$, a policy class $\tilde{\Pi}$ is a smooth $\varepsilon$-cover of a policy class $\Pi$ if for every $\pi\in\Pi$, there exists $\tilde{\pi}\in\tilde{\Pi}$ such that $\Pr_{x\sim U}[\pi(x)\neq\tilde{\pi}(x)]\leq\varepsilon$.
 Definition 5.4.
-A policy class $\tilde{\Pi}$ is an adversarial cover of a policy class $\Pi$ is for every $\boldsymbol{x}\in\mathcal{X}^{T}$ and $\pi\in\Pi$, there exists $\tilde{\pi}\in\tilde{\Pi}$ such that $\pi(x_{t})=\tilde{\pi}(x_{t})$ for all $t\in[T]$.
+A policy class $\tilde{\Pi}$ is an adversarial cover of a policy class $\Pi$ if for every $\boldsymbol{x}\in\mathcal{X}^{T}$ and $\pi\in\Pi$, there exists $\tilde{\pi}\in\tilde{\Pi}$ such that $\pi(x_{t})=\tilde{\pi}(x_{t})$ for all $t\in[T]$.
+An adversarial cover is a perfect cover by definition.
+The idea of a smooth $\varepsilon$-cover is that if the probability of disagreement over the uniform distribution is small, then the probability of disagreement over a $\sigma$-smooth distribution cannot be too much larger.
+Lemma 5.1.
+Let $\tilde{\Pi}$ be a smooth $\varepsilon$-cover of $\Pi$ and let $\mathcal{D}$ be a $\sigma$-smooth distribution.
+Then for any $\pi\in\Pi$, there exists $\tilde{\pi}\in\tilde{\Pi}$ such that $\Pr_{x\sim\mathcal{D}}[\pi(x)\neq\tilde{\pi}(x)]\leq\varepsilon/\sigma$.
+Proof.
+Define $S(\tilde{\pi})=\{x\in\mathcal{X}:\pi(x)\neq\tilde{\pi}(x)\}$.
+By the definition of a smooth $\varepsilon$-cover, there exists $\tilde{\pi}\in\tilde{\Pi}$ such that $\Pr_{x\sim U}[x\in S(\tilde{\pi})]\leq\varepsilon$.
+Since $\mathcal{D}$ is $\sigma$-smooth, $\Pr_{x\sim\mathcal{D}}[\pi(x)\neq\tilde{\pi}(x)]=\Pr_{x\sim\mathcal{D}}[x\in S(\tilde{\pi})]\leq\\ \Pr_{x\sim U}[x\in S(\tilde{\pi})]/\sigma\leq\varepsilon/\sigma$, as claimed.
+∎
 The existence of small covers is crucial:
-Lemma 5.1 (Lemma 7.3.2 in Haghtalab (2018) [^13] See also Haussler & Long (1995) or Lemma 13.6 in Boucheron et al. (2013) for variants which are less convenient for our purposes.).
+Lemma 5.2 (Lemma 7.3.2 in Haghtalab (2018) [^10] See also Haussler & Long (1995) or Lemma 13.6 in Boucheron et al. (2013) for variants of this lemma.).
 For all $\varepsilon>0$, any policy class of VC dimension $d$ admits a smooth $\varepsilon$-cover of size at most $(41/\varepsilon)^{d}$.
@@
 Any policy class of Littlestone dimension $d$ admits an adversarial cover of size at most $(eT/d)^{d}$.
-An adversarial cover is a perfect cover by definition. The following lemma establishes that a smooth $\varepsilon$-cover is a good approximation for any sequence of $\sigma$-smooth distributions.
-Lemma 5.3 (Equation 2 and Lemma 3.3 in Haghtalab et al. (2024)).
-Let $\tilde{\Pi}$ be a finite smooth $\varepsilon$-cover of $\Pi$ and let $\boldsymbol{\mathcal{D}}=\mathcal{D}_{1},\dots,\mathcal{D}_{T}$ be a sequence of $\sigma$-smooth distributions.
-If $\varepsilon T\log T>12\sigma d\log(4e^{2}/\varepsilon)$, then $\operatorname*{\mathbb{E}}\limits_{\boldsymbol{x}\sim\boldsymbol{\mathcal{D}}}\left[\sup\limits_{\pi\in\Pi}\min\limits_{\tilde{\pi}\in\tilde{\Pi}}\sum_{t=1}^{T}\mathbf{1}(\pi(x_{t})\neq\tilde{\pi}(x_{t}))\right]\in O\left(\frac{1}{\sigma}T\varepsilon\log T\sqrt{d\log(1/\varepsilon)}\right)$.
 We will run a variant of Hedge on $\tilde{\Pi}$.
@@
 They proved a regret bound when a query is issued on each time step with fixed probability $p$.
-Lemma 5.4 restates their result in a form that is more convenient for us (see Appendix B for details).
-Although their result is stated for non-adaptive adversaries, we explain in Section B.3 how their argument easily generalizes to adaptive adversaries.
+Lemma 5.4 restates their result in a form that is more convenient for us.
+See Appendices B.1 and B.3 for details on our usage of results from Russo et al.
+(2024).
 Full pseudocode for HedgeWithQueries can also be found in the appendix (Algorithm 2).
-Lemma 5.4 (Lemma 3.5 in Russo et al. (2024)).
-Assume $\tilde{\Pi}$ is finite. Then for any loss function $\ell:[T]\times\tilde{\Pi}\to[0,1]$ and query probability $p$, HedgeWithQueries enjoys the regret bound
-$\sum_{t=1}^{T}\operatorname*{\mathbb{E}}[\ell(t,\pi_{t})]-\min_{\tilde{\pi}\in\tilde{\Pi}}\sum_{t=1}^{T}\ell(t,\pi)\leq\frac{2\log|\tilde{\Pi}|}{p^{2}}$
+Lemma 5.4 (Lemma 3.5 in Russo et al., 2024).
+Assume $\tilde{\Pi}$ is finite. Then for any loss function $\ell:[T]\times\tilde{\Pi}\to[0,1]$ and query probability $p>0$, HedgeWithQueries enjoys the regret bound
+$\sum_{t=1}^{T}\operatorname*{\mathbb{E}}[\ell(t,\pi_{t})]-\min_{\tilde{\pi}\in\tilde{\Pi}}\sum_{t=1}^{T}\operatorname*{\mathbb{E}}[\ell(t,\pi)]\leq\frac{\log|\tilde{\Pi}|}{p^{2}}$
 where $\pi_{t}$ is the policy chosen at time $t$ and the expectation is over the randomness of the algorithm.
-We apply Lemma 5.4 with $\ell(t,\pi)=\mathbf{1}(\pi(x_{t})\neq\pi^{m}(x_{t}))$ and combine this with Lemmas 5.1 and 5.3 (in the $\sigma$-smooth case) and with Lemma 5.2 (in the adversarial case).
+We apply Lemma 5.4 with $\ell(t,\pi)=\mathbf{1}(\pi(x_{t})\neq\pi^{m}(x_{t}))$ and combine this with Lemmas 5.2 and 5.1 (in the $\sigma$-smooth case) and with Lemma 5.3 (in the adversarial case).
 This yields a $O\left(\frac{d}{\sigma}T\varepsilon\log(1/\varepsilon)\log T\right)$ bound on the number of mistakes made by Algorithm 1 (Lemma B.1).
@@
 The trickier part is bounding the number of resulting queries.
-It is tempting to claim that the inputs queried in the out-of-distribution case must all be separated by at least $\varepsilon^{1/n}$ and thus form an $\varepsilon^{1/n}$-packing, but this is actually not true.
-Instead, we provide a novel method for bounding the number of data points (i.e., queries) needed to cover a set with respect to the realized actions of the algorithm (Lemma B.7).
-This is in contrast to vanilla packing arguments which consider all data points in aggregate.
-Our method may be useful in other contexts where a more refined packing argument is needed.
+It is tempting to claim that the inputs queried in the out-of-distribution case must all be separated by at least $\varepsilon^{1/n}$ and thus form an $\varepsilon^{1/n}$-packing, but this is actually false.
+Instead, we bound the number of data points (i.e., queries) needed to cover a set with respect to the realized actions of the algorithm (Lemma B.6).
+This contrasts with vanilla packing arguments which consider all data points in aggregate.
+The key to our analysis is that the number of mistakes made by the algorithm – which we already bounded in Lemma B.1 – gives us crucial information about how data points are distributed with respect to the actions of the algorithm.
+Our technique might be useful in other contexts where a more refined packing argument is needed and a bound on the number of mistakes already exists.
 6 Conclusion and future work
@@
 We showed that achieving subconstant regret in our problem (with the help of a mentor and local generalization) is no harder than achieving sublinear regret in standard online learning.
-There remain some technical questions within this paper’s model.
-One question is whether the time complexity of Algorithm 1 be improved, which currently stands at $\Omega(|\tilde{\Pi}|\cdot T)$ plus the time to compute the $\varepsilon$-cover.
-Also, we have not resolved whether our problem is tractable for finite VC dimension and fully adversarial inputs (although Appendix D shows that the problem is tractable for at least some classes with finite VC but infinite Littlestone dimension).
-We are also interested in alternatives to the local generalization assumption.
-We should expect some assumption to be necessary: if not, the payoff function $\mu(x,y)=\mathbf{1}(\pi^{m}(x)=y)$ means that the agent must make zero mistakes (with high probability), which turns out to be impossible even for $\sigma$-smooth $\boldsymbol{x}$ and finite VC dimension (Theorem E.3).
+Remaining technical questions.
+First, we have not resolved whether our problem is tractable for finite VC dimension and fully adversarial inputs (although Appendix D shows that the problem is tractable for at least some classes with finite VC but infinite Littlestone dimension).
+Second, the time complexity of Algorithm 1 currently stands at a hefty $\Omega(|\tilde{\Pi}|)$ per time step plus the time to compute $\tilde{\Pi}$.
+In the standard online learning setting, Block et al.
+(2022) and Haghtalab et al.
+(2022) show how to replace discretization approaches like ours with oracle-efficient approaches, where a small number of calls to an optimization oracle are made per round.
+We are optimistic about leveraging such techniques to obtain efficient algorithms in our setting.
+Local generalization.
+Our algorithm crucially relies on the ability to detect when an input is unfamiliar, i.e., differs significantly from prior observations in a metric space which satisfies local generalization.
+Without this ability, the practicality of our algorithm would be fundamentally limited.
+One option is to use out-of-distribution (OOD) detection, which is conceptually similar and well-studied (see Yang et al., 2024 for a survey).
+However, it is an open question whether standard OOD detection methods are measuring distance in a metric space which satisfies local generalization.
+We are also interested in alternatives to local generalization.
+Theorem E.2 shows that our positive result breaks down if local generalization is removed, so some sort of assumption is necessary.
 One possible alternative is Bayesian inference.
 We intentionally avoided Bayesian approaches in this paper due to tractability concerns, but it seems premature to abandon those ideas entirely.
+MDPs.
 Finally, we are excited to apply the ideas in this paper to Markov Decision Processes (MDPs): specifically, MDPs where some actions are irreversible (“non-communicating”) and the agent only gets one attempt (“single-episode”).
@@
 Is there an algorithm for non-communicating single-episode undiscounted MDPs which ensures that both the regret and the number of mentor queries are sublinear in $T$?
+Impact statement
+As AI systems become increasingly powerful, we believe that the safety guarantees of such systems should become commensurately robust.
+Irreversible costs are especially worrisome, and we hope that our work plays a small part in mitigating such risks.
+We do not believe that our work has any concrete potential risks that should be highlighted here.
+Author contributions
+B.
+Plaut conceived the project.
+B.
+Plaut designed the mathematical model with feedback from H.
+Zhu and S.
+Russell.
+B.
+Plaut proved all of the results.
+H.
+Zhu participated in proof brainstorming and designed counterexamples for several early conjectures.
+B.
+Plaut wrote the paper, with feedback from H.
+Zhu and S.
+Russell.
+S.
+Russell supervised the project and secured funding.
+Acknowledgements
+This work was supported in part by a gift from Open Philanthropy to the Center for Human-Compatible AI (CHAI) at UC Berkeley.
+This paper also benefited from discussions with many other researchers.
+We would like to especially thank (in alphabetical order) Aly Lidayan, Bhaskar Mishra, Cameron Allen, Juan Liévano-Karim, Matteo Russo, Michael Cohen, Nika Haghtalab, and Scott Emmons.
+We would also like to thank our anonymous reviewers for helpful feedback.
 Appendix A Proof of Theorem 4.1
-A.1 Proof roadmap
+A.1 Proof notation
+  1. 1.
 […]
+  2. 2.
+    Let $J_{\neg Q}=\{j\in[f(T)]:x_{t}\not\in X_{j}\ \forall t\in Q_{T}\}$ be the set of sections that are never queried.
+Since each query appears in exactly one section (because each input appears in exactly one section), $|J_{\neg Q}|\geq f(T)-|Q_{T}|$.
+  3. 3.
+    For each $j\in J_{\neg Q}$, let $y^{j}$ be the most frequent action among time steps in $M_{j}$: $y^{j}=\operatorname*{arg\,max}_{y\in\{0,1\}}|\{t\in M_{j}:y=y_{t}\}|$.
+  4. 4.
+    Let $J_{\neg Q}^{\prime}=\{j\in J_{\neg Q}:a_{j}\neq y^{j}\}$ be the set of sections where the more frequent action is wrong according to $\mu_{f,\boldsymbol{a}}$.
+  5. 5.
+    Let $M_{j}^{\prime}=\{t\in M_{j}:y_{t}\neq a_{j}\}$ be the set of time steps where the agent chooses the wrong action according to $\mu_{f,\boldsymbol{a}}$, and $x_{t}$ is close to the midpoint of section $j$.
+Since $\boldsymbol{x},\boldsymbol{y}$, and $\boldsymbol{a}$ are random variables, all variables defined on top of them (such as $M_{j}$) are also random variables.
+In contrast, the partition $\mathcal{X}=\{X_{1},\dots,X_{f(T)}\}$ and properties thereof (like the midpoints $m_{j}$) are not random variables.
+A.2 Proof roadmap
 The proof considers an arbitrary algorithm with sublinear queries, and proceeds via the following steps:
   1. 1.
-    Prove that $f(T)=\sqrt{(|Q_{T}|+1)T}$ is asymptotically between $|Q_{T}|$ and $T$ (Lemma A.1).
+    Show that multiplicative regret and additive regret are tightly related (Lemma A.1). We will also use this lemma for our positive results.
   2. 2.
+    Prove an asymptotic density lemma which we will use to show that $f(T)=\sqrt{|Q_{T}|T}$ is asymptotically between $|Q_{T}|$ and $T$ (Lemma A.2).
+  3. 3.
     Prove a simple variant of the Chernoff bound which we will apply multiple times (Lemma A.3).
-  3. 3.
-    Show that with high probability, $\sum_{j\in A}|M_{j}|$ is adequately large (Lemma A.3).
   4. 4.
+    Show that with high probability, $\sum_{j\in S}|M_{j}|$ is large for any subset of sections $S$ (Lemma A.4).
+  5. 5.
+    Prove that $|J_{\neg Q}^{\prime}|$ is large with high probability (Lemma A.5).
+  6. 6.
     The key lemma is Lemma A.6, which shows that a randomly sampled $\boldsymbol{a}$ produces poor agent performance with high probability.
@@
 As a result, the agent picks the wrong answer at least half the time on average (and at least a quarter of the time with high probability).
-Lemma A.3 implies that a constant fraction of those time steps will have quite suboptimal payoffs, again with high probability.
-  5. 5.
-    Finally, $\sup\limits_{\boldsymbol{\mu}}\ \operatorname*{\mathbb{E}}\limits_{\boldsymbol{x}\sim U^{T},\,\boldsymbol{y}}\ R_{T}(\boldsymbol{x},\boldsymbol{y},\boldsymbol{\mu},\pi^{m})\geq\operatorname*{\mathbb{E}}\limits_{\boldsymbol{a}\sim U(\{0,1\}^{f(T)})}\ \operatorname*{\mathbb{E}}\limits_{\boldsymbol{x}\sim U^{T},\,\boldsymbol{y}}\ R_{T}(\boldsymbol{x},\boldsymbol{y},\mu_{f,\boldsymbol{a}},\pi^{m})$, where $U(\{0,1\}^{f(T)})$ is the uniform distribution over bit strings of length $f(T)$.
-This is essentially an application of the probabilistic method: if a randomly chosen $\mu_{f,\boldsymbol{a}}$ has high expected regret, then the worst case $\boldsymbol{\mu}$ also has high expected regret.
-While reading the proof, one should keep in mind the following:
-  1. 1.
-    Since $\boldsymbol{x},\boldsymbol{y}$, and $\boldsymbol{a}$ are random variables, all variables defined on top of them (such as $M_{j}$) are also random variables.
-In contrast, the partition $\mathcal{X}=\{X_{1},\dots,X_{f(T)}\}$ and properties thereof (like the midpoints $m_{j}$) are not random variables.
-  2. 2.
-    We will assume that the worst-case (not expected) number of queries is sublinear in $T$. Proposition E.2 shows that relaxing this to expected sublinear queries would not affect our conclusion.
-  3. 3.
-    While the intuition provided in Section 4.1 is accurate, the analysis will mostly occur in log space, so the bounds will look different.
-However, bounds of the form discussed in Section 4.1 can still be found as an intermediate step in Part 4 of the proof of Lemma A.4.
-A.2 Proof
+Lemma A.4 implies that a constant fraction of those time steps will have significantly suboptimal payoffs, again with high probability.
+  7. 7.
+    Apply $\sup\limits_{\boldsymbol{\mu},\pi^{m}}\ \operatorname*{\mathbb{E}}\limits_{\boldsymbol{x},\boldsymbol{y}}\ R_{T}^{\times}(\boldsymbol{x},\boldsymbol{y},\boldsymbol{\mu},\pi^{m})\geq\operatorname*{\mathbb{E}}\limits_{\pi^{m},\boldsymbol{a}\sim U(\{0,1\}^{f(T)})}\ \operatorname*{\mathbb{E}}\limits_{\boldsymbol{x},\boldsymbol{y}}\ R_{T}^{\times}(\boldsymbol{x},\boldsymbol{y},\mu_{f,\boldsymbol{a}},\pi^{m})$.
+Here $U(\{0,1\}^{f(T)})$ is the uniform distribution over bit strings of length $f(T)$ and we write $\pi^{m},\boldsymbol{a}\sim U(\{0,1\}^{f(T)})$ with slight abuse of notation, since $\pi^{m}$ is not drawn from $U(\{0,1\}^{f(T)})$ but rather is determined by $\boldsymbol{a}$ which is drawn from $U(\{0,1\}^{f(T)})$.
+  8. 8.
+    The analysis above results in a lower bound on $R_{T}^{+}$. The last step is to use Lemma A.1 to obtain a lower bound on $R_{T}^{\times}$.
+Step 7 is essentially an application of the probabilistic method: if a randomly chosen $\mu_{f,\boldsymbol{a}}$ has high expected regret, then the worst-case $\boldsymbol{\mu}$ also has high expected regret.
+We have included subscripts in the expectations above to distinguish between the randomness over $\boldsymbol{a}$ and $\boldsymbol{x},\boldsymbol{y}$.
+When subscripts are omitted, the expected value is over all randomness, i.e., $\boldsymbol{a},\boldsymbol{x},$ and $\boldsymbol{y}$.
+A.3 Proof
 Lemma A.1.
-Let $a,b:\mathbb{N}\to\mathbb{N}$ be functions such that $a(x)\in o(b(x))$. Then $c(x)=\sqrt{a(x)b(x)}$ satisfies $a(x)\in o(c(x))$ and $c(x)\in o(b(x))$.
+[…]
+Proof.
+Recall the standard inequalities $11-\frac{1}{a}\leq\log a\leq a-$ 1 for any $a>0$.
+Part 1: $R_{T}^{+}\leq R_{T}^{\times}$.
+If $\mu_{t}(x_{t},y_{t})=0$ for any $t\in[T]$, then $R_{T}^{\times}=\infty$ and the claim is trivially satisfied.
+Thus assume $\mu_{t}(x_{t},y_{t})>0$ for all $t\in[T]$.
+$R_{T}^{+}=$ $\ \sum_{t=1}^{T}\mu_{t}^{m}(x_{t})-\sum_{t=1}^{T}\mu_{t}(x_{t},y_{t})$ (Definition of $(\text{Definition of $R_{T}^{+}$})$
+[…]
+[…]
+[…]
+$=$ $\ R_{T}^{\times}$ (Definition of $(\text{Definition of $R_{T}^{\times}$})$
+Part 2: $R_{T}^{\times}\leq R_{T}^{+}/\min_{t\in[T]}\mu_{t}(x_{t},y_{t})$. We have
+$R_{T}^{\times}=$ $\ \log\prod_{t=1}^{T}\mu_{t}^{m}(x_{t})-\log\prod_{t=1}^{T}\mu_{t}(x_{t},y_{t})$ (Definition of RT× given $(\text{Definition of $R_{T}^{\times}$ given $\mu_{t}(x_{t},y_{t})>0\ \forall t\in[T]$})$
+[…]
+[…]
+[…]
+[…]
+$=$ $\ \frac{R_{T}^{+}}{\min_{t\in[T]}\mu_{t}(x_{t},y_{t})}$ (Definition of $(\text{Definition of $R_{T}^{+}$})$
+as claimed. ∎
+Lemma A.2.
+Let $a,b:\mathbb{R}_{>0}\to\mathbb{R}_{>0}$ be functions such that $a(x)\in o(b(x))$. Then $c(x)=\sqrt{a(x)b(x)}$ satisfies $a(x)\in o(c(x))$ and $c(x)\in o(b(x))$.
 Proof.
@@
 Lemma A.4.
-let $A\subseteq[f(T)]$ be any nonempty subset of sections. Then
-$\Pr\left[\sum_{j\in A}|M_{j}|\leq\frac{T|A|}{4f(T)}\right]\leq\exp\left(\frac{-T}{16f(T)}\right)$
+Let $S\subseteq[f(T)]$ be any nonempty subset of sections. Then
+$\Pr\left[\sum_{j\in S}|M_{j}|\leq\frac{T|S|}{4f(T)}\right]\leq\exp\left(\frac{-T}{16f(T)}\right)$
 Proof.
-[…]
-$\Pr\left[\sum_{j\in A}|M_{j}|\leq\frac{T|A|}{4f(T)}\right]\leq\exp\left(\frac{-T|A|}{16f(T)}\right)\leq\exp\left(\frac{-T}{16f(T)}\right)$
-with the last step due to $|A|\geq 1$. ∎
+Fix any $j\in[f(T)]$.
+For each $t\in[T]$ , define the random variable $z_{t}$ by $z_{t}=1$ if $t\in M_{j}$ for some $j\in S$ and 0 otherwise.
+We have $t\in M_{j}$ iff $x_{t}$ falls within a particular interval of length $\frac{1}{2f(T)}$.
+Since these intervals are disjoint for different $j$’s, we have $z_{t}=1$ iff $x_{t}$ falls within a portion of the input space with total measure $\frac{|S|}{2f(T)}$.
+Since $x_{t}$ is uniformly random across $[0,1]$, we have $\operatorname*{\mathbb{E}}[z_{t}]=\frac{|S|}{2f(T)}$.
+Then $\operatorname*{\mathbb{E}}[\sum_{t=1}^{T}z_{t}]=\operatorname*{\mathbb{E}}[\sum_{j\in S}|M_{j}|]=\frac{T|S|}{2f(T)}$.
+Furthermore, since $x_{1},\dots,x_{T}$ are i.i.d., so are $z_{1},\dots,z_{T}$.
+Then by Lemma A.3,
+$\Pr\left[\sum_{j\in S}|M_{j}|\leq\frac{T|S|}{4f(T)}\right]\leq\exp\left(\frac{-T|S|}{16f(T)}\right)\leq\exp\left(\frac{-T}{16f(T)}\right)$
+with the last step due to $|S|\geq 1$. ∎
 Lemma A.5.
-Let $f,g:\mathbb{N}\to\mathbb{N}$.
-Suppose $|Q_{T}|\leq g(T)$ always and independently sample $\boldsymbol{a}\sim U(\{0,1\}^{f(T)})$ and $\boldsymbol{x}\sim U^{T}$.
+We have
+$\Pr\left[|J_{\neg Q}^{\prime}|\leq\frac{f(T)-\operatorname*{\mathbb{E}}[|Q_{T}|]}{4}\right]\leq\exp\left(-\frac{f(T)-\operatorname*{\mathbb{E}}[|Q_{T}|]}{16}\right)$
+Proof.
+Define a random variable $z_{j}=\mathbf{1}_{j\in J_{\neg Q}^{\prime}}$ for each $j\in J_{\neg Q}$.
+By definition, if $j\in J_{\neg Q}$, no input in $X_{j}$ is queried.
+Since queries outside of $X_{j}$ provide no information about $a_{j}$, the agent’s actions must be independent of $a_{j}$.
+In particular, the random variables $a_{j}$ and $y^{j}$ are independent.
+Combining that independence with $\Pr[a_{j}=0]=\Pr[a_{j}=1]=0.5$ yields $\Pr[z_{j}=1]=0.5$ for all $j\in J_{\neg Q}$.
+Then
+$\operatorname*{\mathbb{E}}\left[|J_{\neg Q}^{\prime}|\right]=$ $\ \operatorname*{\mathbb{E}}\left[\sum_{j\in J_{\neg Q}}z_{j}\right]$
+$=$ $\ |J_{\neg Q}|/2$
+$\geq$ $\ \frac{f(T)-\operatorname*{\mathbb{E}}[|Q_{T}|]}{2}$
+Furthermore, since $a_{1},\dots,a_{f(T)}$ are independent, the random variables $\{z_{j}:j\in J_{\neg Q}\}$ are also independent. Applying Lemma A.3 yields the desired bound. ∎
+Lemma A.6.
+Suppose $f:\mathbb{N}\to\mathbb{N}$ and independently sample $\boldsymbol{a}\sim U(\{0,1\}^{f(T)})$ and $\boldsymbol{x}\sim U(\mathcal{X})^{T}$.
 [^11] That is, the entire set $\{a_{1},\dots,a_{f(T)},x_{1},\dots,x_{T}\}$ is mutually independent.
-Then with probability at least $1-\exp\big{(}\frac{-T}{16f(T)}\big{)}-\exp\big{(}-\frac{f(T)-g(T)}{16}\big{)}$,
-$\prod_{t=1}^{T}\mu_{f,\boldsymbol{a}}(x_{t},y_{t})\leq\exp\left(-\frac{LT(f(T)-g(T))}{2^{7}f(T)^{2}}\right)$
+Then with probability at least $1-\exp\big{(}\frac{-T}{16f(T)}\big{)}-\exp\big{(}-\frac{f(T)-\operatorname*{\mathbb{E}}[|Q_{T}|]}{16}\big{)}$,
+$R_{T}^{+}\geq\frac{LT(f(T)-\operatorname*{\mathbb{E}}[|Q_{T}|])}{2^{7}f(T)^{2}}$
 Proof.
-Part 1: setup.
-Let $J_{\neg Q}=\{j\in[f(T)]:x_{t}\not\in X_{j}\ \forall t\in Q_{T}\}$ be the set of sections that are never queried.
-Since each query appears in exactly one section (because each input appears in exactly one section), $|J_{\neg Q}|\geq f(T)-|Q_{T}|\geq f(T)-g(T)$.
-For each $j\in J_{\neg Q}$, let $y_{j}$ be the action taken most frequently among time steps in $M_{j}$:
-$y_{j}=\operatorname*{arg\,max}_{y\in\{0,1\}}\Big{|}\{t\in M_{j}:y=y_{t}\}\Big{|}$
-Let $\bar{J}=\{j\in J_{\neg Q}:a_{j}\neq y_{j}\}$.
-For each $j\in\bar{J}$, let $M_{j}^{\prime}=\{t\in M_{j}:y_{t}\neq a_{j}\}$ be the set of time steps where the agent chooses the wrong action (assuming payoff function $\mu_{f,\boldsymbol{a}}$).
-[…]
-$\Pr\left[|\bar{J}|\leq\frac{f(T)-g(T)}{4}\right]\leq\exp\left(-\frac{f(T)-g(T)}{16}\right)$
-Part 3: $|M_{j}^{\prime}|\geq|M_{j}|/2$.
-Since $j\in J_{\neg Q}$, the mentor is not queried on any time step $t\in M_{j}$, so $y_{t}\in\{0,1\}$ for all $t\in M_{j}$.
-Since the agent chooses one of two actions for each $t\in M_{j}$, the more frequent action must be chosen chosen at least half of the time: $y_{t}=y_{j}$ for at least half of the time steps in $M_{j}$.
-Since $a_{j}\neq y_{j}$ for $j\in\bar{J}$, we have $y_{t}=y_{j}\neq a_{j}$ for those time steps, so $|M_{j}^{\prime}|\geq|M_{j}|/2$.
-Part 4: a bound in terms of $\bar{J}$ and $M_{j}$.
-Consider any $j\in\bar{J}$ and $t\in M_{j}^{\prime}\subseteq M_{j}$.
+Consider any $j\in J_{\neg Q}^{\prime}$ and $t\in M_{j}^{\prime}\subseteq M_{j}$.
 By definition of $M_{j}$, we have $|m_{j}-x_{t}|\leq\frac{1}{4f(T)}$.
-Then by definition of $\mu_{f,\boldsymbol{a}}$,
+Then by the definition of $\mu_{f,\boldsymbol{a}}$,
 $\mu_{f,\boldsymbol{a}}(x_{t},y_{t})=$ $\ 1-L\left(\frac{1}{2f(T)}-|x_{t}-m_{j}|\right)$
 $\leq$ $\ 1-L\left(\frac{1}{2f(T)}-\frac{1}{4f(T)}\right)$
-$=$ $\ =1-\frac{L}{4f(T)}$
-Now aggregating across time steps,
+$=$ $\ 1-\frac{L}{4f(T)}$
+Since $\mu_{f,\boldsymbol{a}}^{m}(x_{t})\geq\mu_{f,\boldsymbol{a}}(x_{t},y_{t})$ always, we can safely restrict ourselves to time steps $t\in M_{j}^{\prime}$ for some $j\in J_{\neg Q}^{\prime}$ and still obtain a lower bound:
+$R_{T}^{+}=$ $\ \sum_{t=1}^{T}(\mu_{f,\boldsymbol{a}}^{m}(x_{t})-\mu_{f,\boldsymbol{a}}(x_{t},y_{t}))$ (Definition of $(\text{Definition of $R_{T}^{+}$})$
+$\geq$ $\ \sum_{j\in J_{\neg Q}^{\prime}}\sum_{t\in M_{j}^{\prime}}\big{(}\mu_{f,\boldsymbol{a}}^{m}(x_{t})-\mu_{f,\boldsymbol{a}}(x_{t},y_{t})\big{)}$ $(\mu_{f,\boldsymbol{a}}^{m}(x_{t})\geq\mu_{f,\boldsymbol{a}}(x_{t},y_{t}))$
+$\geq$ $\ \sum_{j\in J_{\neg Q}^{\prime}}\sum_{t\in M_{j}^{\prime}}(1-\mu_{f,\boldsymbol{a}}(x_{t},y_{t}))$ $(\mu_{f,\boldsymbol{a}}^{m}(x_{t})=1\text{ always})$
+[…]
+[…]
+Since $j\in J_{\neg Q}$, the mentor is not queried on any time step $t\in M_{j}$, so $y_{t}\in\{0,1\}$ for all $t\in M_{j}$.
+Since the agent chooses one of two actions for each $t\in M_{j}$, the more frequent action must be chosen at least half of the time: $y_{t}=y^{j}$ for at least half of the time steps in $M_{j}$.
+Since $a_{j}\neq y^{j}$ for $j\in J_{\neg Q}^{\prime}$, we have $y_{t}=y^{j}\neq a_{j}$ for those time steps, so $|M_{j}^{\prime}|\geq|M_{j}|/2$.
+Thus
+$R_{T}^{+}\geq\sum_{j\in J_{\neg Q}^{\prime}}\frac{L|M_{j}|}{8f(T)}$
+[…]
+$R_{T}^{+}\geq$ $\ \sum_{j\in J_{\neg Q}^{\prime}}\frac{L|M_{j}|}{8f(T)}$
+$\geq$ $\ \frac{L}{8f(T)}\cdot\frac{T|J_{\neg Q}^{\prime}|}{4f(T)}$
+$\geq$ $\ \frac{L}{8f(T)}\cdot\frac{T}{4f(T)}\cdot\frac{f(T)-\operatorname*{\mathbb{E}}[|Q_{T}|]}{4}$
+$=$ $\ \frac{LT(f(T)-\operatorname*{\mathbb{E}}[|Q_{T}|])}{2^{7}f(T)^{2}}$
+as required. ∎
+For a given $f:\mathbb{N}\to\mathbb{N}$, define $\alpha_{f}(T)=\exp\big{(}\frac{-T}{16f(T)}\big{)}+\exp\big{(}-\frac{f(T)-\operatorname*{\mathbb{E}}[|Q_{T}|]}{16}\big{)}$ for brevity. See 4.1
+Proof.
+If the algorithm has sublinear queries, then there exists $g(T)\in o(T)$ such that $\sup_{\boldsymbol{\mu},\pi^{m}}\operatorname*{\mathbb{E}}_{\boldsymbol{x},\boldsymbol{y}}[|Q_{T}|]\leq g(T)$.
+Consider any such $g(T)$ satisfying $g(T)>0$.
+Since this holds for every $\boldsymbol{\mu}$, it also holds in expectation over $\boldsymbol{a}\sim U(\{0,1\})^{f(T)}$, so $\operatorname*{\mathbb{E}}_{\boldsymbol{a},\boldsymbol{x},\boldsymbol{y}}[|Q_{T}|]=\operatorname*{\mathbb{E}}[|Q_{T}|]\leq g(T)$.
+Next, Lemma A.2 gives us $g(T)\in o(\sqrt{g(T)T})$ and $\sqrt{g(T)T}\in o(T)$.
+Let $f(T)=\lceil\sqrt{g(T)T}\rceil$: then $f(T)\in\Theta(\sqrt{g(T)T})$, so $g(T)\in o(f(T))$ and $f(T)\in o(T)$.
+First, this implies that $\lim_{T\to\infty}\alpha_{f}(T)=0$.
+Second, $g(T)\in o(f(T))$ implies that exists $T_{0}$ such that $g(T)\leq f(T)/2$ for all $T\geq T_{0}$.
+We also have $R_{T}^{+}\geq 0$ always since $\mu_{f,\boldsymbol{a}}^{m}(x_{t})\geq\mu_{f,\boldsymbol{a}}(x_{t},y_{t})$ always.
+Then for all $T\geq T_{0}$ we have
+$\ \operatorname*{\mathbb{E}}_{\pi^{m},\boldsymbol{a}\sim U(\{0,1\}^{f(T)})}\ \,\operatorname*{\mathbb{E}}_{\boldsymbol{x},\boldsymbol{y}}\ \left[R_{T}^{+}(\boldsymbol{x},\boldsymbol{y},\boldsymbol{\mu},\pi^{m})\right]$
 […]
@@
 […]
+Since $\lim_{T\to\infty}\alpha_{f}(T)=0$,
+$\sup_{\boldsymbol{\mu},\pi^{m}}\ \operatorname*{\mathbb{E}}_{\boldsymbol{x},\boldsymbol{y}}\ \big{[}R_{T}^{+}(\boldsymbol{x},\boldsymbol{y},\boldsymbol{\mu},\pi^{m})\big{]}\geq$ $\ \operatorname*{\mathbb{E}}_{\pi^{m},\boldsymbol{a}\sim U(\{0,1\}^{f(T)})}\ \,\operatorname*{\mathbb{E}}_{\boldsymbol{x},\boldsymbol{y}}\ \big{[}R_{T}^{+}(\boldsymbol{x},\boldsymbol{y},(\mu_{f,\boldsymbol{a}},\dots,\mu_{f,\boldsymbol{a}}),\pi^{m})\big{]}$
+$\geq$ $\ \big{(}1-\alpha_{f}(T)\big{)}\left(\frac{LT}{2^{8}f(T)}\right)$
+$\in$ $\ \Omega\left(\frac{LT}{f(T)}\right)$
+$=$ $\ \Omega\left(L\sqrt{\frac{T}{g(T)}}\right)$
+This holds for any $g(T)\in o(T)$ such that $\sup_{\boldsymbol{\mu}}\operatorname*{\mathbb{E}}[|Q_{T}|]\leq g(T)$ and $g(T)>0$.
+Thus we can simply set $g(T)=\sup_{\boldsymbol{\mu},\pi^{m}}\operatorname*{\mathbb{E}}[|Q_{T}|]+1$, since $\sup_{\boldsymbol{\mu},\pi^{m}}\operatorname*{\mathbb{E}}[|Q_{T}|]$ is indeed a function of only $T$.
+Since $\mu_{f,\boldsymbol{a}}^{m}(x_{t})\geq\mu_{f,\boldsymbol{a}}(x_{t},y_{t})$ for all $t\in[T]$, Lemma A.1 implies that
+$\sup_{\boldsymbol{\mu},\pi^{m}}\ \operatorname*{\mathbb{E}}_{\boldsymbol{x},\boldsymbol{y}}\ \big{[}R_{T}^{\times}(\boldsymbol{x},\boldsymbol{y},\boldsymbol{\mu},\pi^{m})\big{]}\geq$ $\ \operatorname*{\mathbb{E}}_{\pi^{m},\boldsymbol{a}\sim U(\{0,1\}^{f(T)})}\ \,\operatorname*{\mathbb{E}}_{\boldsymbol{x},\boldsymbol{y}}\ \big{[}R_{T}^{\times}(\boldsymbol{x},\boldsymbol{y},(\mu_{f,\boldsymbol{a}},\dots,\mu_{f,\boldsymbol{a}}),\pi^{m})\big{]}$
+$\in$ $\ \Omega\left(L\sqrt{\frac{T}{\sup_{\boldsymbol{\mu},\pi^{m}}\operatorname*{\mathbb{E}}[|Q_{T}|]+1}}\right)$
+completing the proof.∎
+See 4.1.1
+Proof.
+We have $\prod_{t=1}^{T}\mu_{f,\boldsymbol{a}}^{m}(x_{t})=1$ from our construction. Then Lemma A.1 implies that $R_{T}^{\times}\geq R_{T}^{+}$, so
+$\exp(-R_{T}^{+})\geq$ $\ \exp(-R_{T}^{\times})$
+$=$ $\ \exp\left(\log\prod_{t=1}^{T}\mu_{f,\boldsymbol{a}}(x_{t},y_{t})-\log 1\right)$
+$=$ $\ \prod_{t=1}^{T}\mu_{f,\boldsymbol{a}}(x_{t},y_{t})$
+Then by Lemma A.6, with probability $1-\alpha_{f}(T)$,
+$\prod_{t=1}^{T}\mu_{f,\boldsymbol{a}}(x_{t},y_{t})\leq\exp\left(-\frac{LT(f(T)-\operatorname*{\mathbb{E}}[|Q_{T}|])}{2^{7}f(T)^{2}}\right)\leq\exp\left(-\frac{LT}{2^{8}f(T)}\right)$
+Since $\prod_{t=1}^{T}\mu_{f,\boldsymbol{a}}(x_{t},y_{t})\leq 1$ always,
+$\lim_{T\to\infty}\operatorname*{\mathbb{E}}_{\pi^{m},\boldsymbol{a}\sim U(\{0,1\}^{f(T)})}\ \operatorname*{\mathbb{E}}_{\boldsymbol{x},\boldsymbol{y}}\ \left[\prod_{t=1}^{T}\mu_{f,\boldsymbol{a}}(x_{t},y_{t})\right]\leq$ $\ \lim_{T\to\infty}\operatorname*{\mathbb{E}}_{\pi^{m},\boldsymbol{a}\sim U(\{0,1\}^{f(T)})}\ \operatorname*{\mathbb{E}}_{\boldsymbol{x},\boldsymbol{y}}\ \left[\prod_{t=1}^{T}\mu_{f,\boldsymbol{a}}(x_{t},y_{t})\right]$
+$\leq$ $\ \lim_{T\to\infty}\left(1\cdot\alpha_{f}(T)+(1-\alpha_{f}(T))\cdot\exp\left(-\frac{LT}{2^{8}f(T)}\right)\right)$
+$\leq$ $\ \lim_{T\to\infty}(1-\alpha_{f}(T))\cdot\lim_{T\to\infty}\exp\left(-\frac{LT}{2^{8}f(T)}\right)$
+$=$ $\ 1\cdot\exp(-\infty)$
 […]
-$\log\prod_{t=1}^{T}\mu_{f,\boldsymbol{a}}(x_{t},y_{t})\leq$ $\ \log\prod_{j\in\bar{J}}\left(1-\frac{L}{4f(T)}\right)^{|M_{j}|/2}$
-$=$ $\ \sum_{j\in\bar{J}}\frac{|M_{j}|}{2}\log\left(1-\frac{L}{4f(T)}\right)$
-$\leq$ $\ -\sum_{j\in\bar{J}}\frac{L|M_{j}|}{8f(T)}$
-[…]
-$\log\prod_{t=1}^{T}\mu_{f,\boldsymbol{a}}(x_{t},y_{t})\leq$ $\ -\sum_{j\in\bar{J}}\frac{L|M_{j}|}{8f(T)}$
-$\leq$ $\ -\frac{L}{8f(T)}\cdot\frac{T|\bar{J}|}{4f(T)}$
-$\leq$ $\ -\frac{L}{8f(T)}\cdot\frac{T}{4f(T)}\cdot\frac{f(T)-g(T)}{4}$
-$=$ $\ -\frac{LT(f(T)-g(T))}{2^{7}f(T)^{2}}$
-Exponentiating both sides proves the lemma. ∎
-Let $\alpha(T)=\exp\big{(}\frac{-T}{16f(T)}\big{)}+\exp\big{(}-\frac{f(T)-g(T)}{16}\big{)}$ for brevity (with $f$ and $g$ to be chosen later). See 4.1
-Proof.
-If the algorithm has sublinear queries, then there exists $g(T)\in o(T)$ such that $|Q_{T}|\leq g(T)$ always.
-Then by Lemma A.1, $g(T)\in o(\sqrt{g(T)T})$ and $\sqrt{g(T)T}\in o(T)$.
-Let $f(T)=\lceil\sqrt{g(T)T}\rceil$: then $f(T)\in\Theta(\sqrt{g(T)T})$, so $g(T)\in o(f(T))$ and $f(T)\in o(T)$.
-First, this implies that $\lim_{T\to\infty}\alpha(T)=0$.
-Next, $g(T)\in o(f(T))$ implies that exists $T_{0}$ such that $g(T)\leq f(T)/2$ for all $T\geq T_{0}$.
-Combining this with Lemma A.4 and noting that $\prod_{t=1}^{T}\mu_{f,\boldsymbol{a}}(x_{t},y_{t})\leq 1$, for all $T\geq T_{0}$ we have
-$\ \operatorname*{\mathbb{E}}_{\boldsymbol{a}\sim U(\{0,1\}^{f(T)})}\ \,\operatorname*{\mathbb{E}}_{\boldsymbol{x}\sim U^{T},\,\boldsymbol{y}}\ \>\prod_{t=1}^{T}\mu_{f,\boldsymbol{a}}(x_{t},y_{t})$
-$\leq$ $\ \alpha(T)\cdot 1+\big{(}1-\alpha(T)\big{)}\exp\left(-\frac{LT(f(T)-|g(T))}{2^{7}f(T)^{2}}\right)$
-$\leq$ $\ \alpha(T)+\big{(}1-\alpha(T)\big{)}\exp\left(-\frac{LTf(T)/2}{2^{7}f(T)^{2}}\right)$
-$=$ $\ \alpha(T)+\big{(}1-\alpha(T)\big{)}\exp\left(-\frac{LT}{2^{8}f(T)}\right)$
-Since $\prod_{t=1}^{T}\mu_{f,\boldsymbol{a}}^{m}(x_{t})=1$ always, we have [^15] Fubini’s theorem means we need not worry about the order of the expectation operators.
-$\sup_{\boldsymbol{\mu}}\ \operatorname*{\mathbb{E}}_{\boldsymbol{x}\sim U^{T},\,\boldsymbol{y}}\ R_{T}(\boldsymbol{x},\boldsymbol{y},\boldsymbol{\mu},\pi^{m})\geq$ $\ \operatorname*{\mathbb{E}}_{\boldsymbol{a}\sim U(\{0,1\}^{f(T)})}\ \,\operatorname*{\mathbb{E}}_{\boldsymbol{x}\sim U^{T},\,\boldsymbol{y}}\ R_{T}(\boldsymbol{x},\boldsymbol{y},(\mu_{f,\boldsymbol{a}},\dots,\mu_{f,\boldsymbol{a}}),\pi^{m})$
-$=\operatorname*{\mathbb{E}}_{\boldsymbol{a}\sim U(\{0,1\}^{f(T)})}\ \,\operatorname*{\mathbb{E}}_{\boldsymbol{x}\sim U^{T},\,\boldsymbol{y}}$ $\ \,\left[\prod_{t=1}^{T}\mu_{f,\boldsymbol{a}}^{m}(x_{t})-\prod_{t=1}^{T}\mu_{f,\boldsymbol{a}}(x_{t},y_{t})\right]$
-$\geq$ $\ 1-\alpha(T)-\big{(}1-\alpha(T)\big{)}\exp\left(-\frac{LT}{2^{8}f(T)}\right)$
-Therefore
-$\lim_{T\to\infty}\sup_{\boldsymbol{\mu}}\ \operatorname*{\mathbb{E}}_{\boldsymbol{x}\sim U^{T},\,\boldsymbol{y}}\ R_{T}(\boldsymbol{x},\boldsymbol{y},\boldsymbol{\mu},\pi^{m})\geq$ $\ 1-\lim_{T\to\infty}\alpha(T)-\big{(}1-\lim_{T\to\infty}\alpha(T)\big{)}\cdot\exp\left(\lim_{T\to\infty}-\frac{LT}{2^{8}f(T)}\right)$
-$=$ $\ 1-0-(1-0)\cdot\exp(-\infty)$
-[…]
-as required. ∎
+Since this upper bound holds for a randomly chosen $\mu_{f,\boldsymbol{a}},\pi^{m}$, the same upper bound holds for a worst-case choice of $\boldsymbol{\mu},\pi^{m}$ among $\boldsymbol{\mu},\pi^{m}$ which satisfy $\mu_{t}^{m}(x)=1$ for all $t\in[T],x\in\mathcal{X}$.
+Formally,
+$\lim_{T\to\infty}\ \sup_{\boldsymbol{\mu},\pi^{m}:\,\mu_{t}(x)=1\,\forall t,x}\operatorname*{\mathbb{E}}\left[\prod_{t=1}^{T}\mu_{t}(x_{t},y_{t})\right]\leq 0$
+Since $\prod_{t=1}^{T}\mu_{t}(x_{t},y_{t})\geq 0$ always, the inequality above holds with equality. ∎
 Appendix B Proof of Theorem 5.2
@@
 See 5.4
-Lemma 5.4 is a restatement and simplification of Lemma 3.5 in Russo et al.
-(2024).
-First, Russo et al.
-(2024) parametrize their algorithm by the expected number of queries $\hat{k}$ instead of the query probability $p=\hat{k}/T$.
-Second, Russo et al.
-(2024) include a second parameter $k$, which is the eventual target number of queries for their unconditional query bound.
+Lemma 5.4 is a restatement and simplification of Lemma 3.5 in Russo et al. (2024) with the following differences:
+  1. 1.
+    They parametrize their algorithm by the expected number of queries $\hat{k}$ instead of the query probability $p=\hat{k}/T$.
+  2. 2.
+    They include a second parameter $k$, which is the eventual target number of queries for their unconditional query bound.
 In our case, an expected query bound is sufficient, so we simply set $k=\hat{k}$.
-Third, Russo et al.
-(2024) provide a second bound which is tighter for small $k$; that bound is less useful for us so we omit it.
-Fourth, their number of actions $n$ is equal to $|\tilde{\Pi}|$ in our setting.
-(Their actions correspond to policies in $\tilde{\Pi}$, not our actions in $\mathcal{Y}$.) Since Russo et al.
+  3. 3.
+    They provide a second bound which is tighter for small $k$; that bound is less useful for us so we omit it.
+  4. 4.
+    Their “actions” correspond to policies in our setting, not actions in $\mathcal{Y}$. Their number of actions $n$ corresponds to $|\tilde{\Pi}|$.
+  5. 5.
+    We include an expectation over both loss terms, while they only include an expectation over the agent’s loss.
+This is because an adaptive adversary may choose the loss function for time $t$ in a randomized manner.
+Since we eventually set $\ell(t,\pi)=\mathbf{1}(\pi(x_{t})\neq\pi^{m}(x_{t}))$, the randomization in $\ell$ corresponds to the randomization in $x_{t}$.
+Altogether, since Russo et al.
 (2024) set $\eta=\max\Big{(}\frac{1}{T}\sqrt{\frac{\hat{k}\log n}{2}},\frac{k\hat{k}}{\sqrt{2}T^{2}}\Big{)}$, we end up with $\eta=\max\Big{(}\sqrt{\frac{p\log|\tilde{\Pi}|}{2T}},\>\frac{p^{2}}{\sqrt{2}}\Big{)}$.
 Algorithm 2 provides precise pseudocode for the HedgeWithQueries algorithm to which Lemma 5.4 refers.
-1:function HedgeWithQueries($p\in(0,1],\>\text{finite policy class }\tilde{\Pi},\>\text{unknown }\ell:[T]\times\tilde{\Pi}\to[0,1]$)
-2:     $w(\pi)\leftarrow 1$ for all $\pi\in\tilde{\Pi}$
-3:     $\eta\leftarrow\max\big{(}\sqrt{\frac{p\log|\tilde{\Pi}|}{2T}},\>\frac{p^{2}}{\sqrt{2}}\big{)}$
-4:     for $t$ from $1$ to $T$ do
+  function HedgeWithQueries$(p\in(0,1],\>\text{finite policy class }\tilde{\Pi},\>\text{unknown }\ell:[T]\times\tilde{\Pi}\to[0,1])$
+     $w(\pi)\leftarrow 1$ for all $\pi\in\tilde{\Pi}$
+     $\eta\leftarrow\max\big{(}\sqrt{\frac{p\log|\tilde{\Pi}|}{2T}},\>\frac{p^{2}}{\sqrt{2}}\big{)}$
+     for $t$ from $1$ to $T$ do
 […]
-6:         if hedgeQuery then
-7:              Query and observe $\ell(t,\pi)$ for all $\pi\in\tilde{\Pi}$
-8:              $\ell^{*}\leftarrow\min_{\pi\in\tilde{\Pi}}\ell(t,\pi)$
-9:              $w(\pi)\leftarrow w(\pi)\cdot\exp(-\eta(\ell(t,\pi)-\ell^{*}))$ for all $\pi\in\tilde{\Pi}$
-10:              Select policy $\arg\min_{\pi\in\tilde{\Pi}}\ell(t,\pi)$
-11:         else
-12:              $P(\pi)\leftarrow w(\pi)/\sum_{\pi^{\prime}\in\tilde{\Pi}}w(\pi^{\prime})$ for all $\pi\in\tilde{\Pi}$
-13:              Sample $\pi_{t}\sim P$
-14:              Select policy $\pi_{t}$
+[…]
+        if hedgeQuery then
+           Query and observe $\ell(t,\pi)$ for all $\pi\in\tilde{\Pi}$
+           $\ell^{*}\leftarrow\min_{\pi\in\tilde{\Pi}}\ell(t,\pi)$
+           $w(\pi)\leftarrow w(\pi)\cdot\exp(-\eta(\ell(t,\pi)-\ell^{*}))$ for all $\pi\in\tilde{\Pi}$
+           Select policy $\arg\min_{\pi\in\tilde{\Pi}}\ell(t,\pi)$
+        else
+           $P(\pi)\leftarrow w(\pi)/\sum_{\pi^{\prime}\in\tilde{\Pi}}w(\pi^{\prime})$ for all $\pi\in\tilde{\Pi}$
+           Sample $\pi_{t}\sim P$
+           Select policy $\pi_{t}$
 Algorithm 2 A variant of the Hedge algorithm which only observes losses in response to queries.
@@
   1. 1.
-    For each $t\in[T]$, let $X_{t}$ refer to the value of $X$ at the start of time step $t$.
+    For each $t\in[T]$, let $S_{t}$ refer to the value of $S$ at the start of time step $t$.
   2. 2.
@@
   3. 3.
-    For $X\subseteq\mathcal{X}$, let $\operatorname*{vol}(X)$ denote the $n$-dimensional Lebesgue measure of $X$.
+    For $S\subseteq\mathcal{X}$, let $\operatorname*{vol}(S)$ denote the $n$-dimensional Lebesgue measure of $S$.
   4. 4.
@@
     We will use “Case 1” to refer to finite VC dimension and $\sigma$-smooth $\boldsymbol{x}$ and “Case 2” to refer to finite Littlestone dimension.
-In Case 1, expectations are over the randomness of both $\boldsymbol{x}$ and the algorithm, while in Case 2, expectations are over just the randomness of the algorithm.
-When we need to distinguish, we use $\operatorname*{\mathbb{E}}_{\boldsymbol{y}}$ to denote the expectation over randomness of the algorithm and $\operatorname*{\mathbb{E}}_{\boldsymbol{x}\sim\boldsymbol{\mathcal{D}}}$ to denote the expectation over $\boldsymbol{x}$.
 Lemma B.1.
-Under the conditions of Theorem 5.1, Algorithm 1 satisfies
-$\operatorname*{\mathbb{E}}[|M_{T}|]\in O\left(\frac{d}{\sigma}T\varepsilon\log(1/\varepsilon)\log T\right)$
+Let $\mathcal{Y}=\{0,1\}$.
+Assume $\pi^{m}\in\Pi$ where either (1) $\Pi$ has finite VC dimension $d$ and $\boldsymbol{x}$ is $\sigma$-smooth, or (2) $\Pi$ has finite Littlestone dimension $d$.
+Then for any $T\in\mathbb{N}$ and $\varepsilon\geq 1/T$, [^12] Note that this lemma omits the assumption of $\varepsilon\leq(\frac{\mu_{0}^{m}}{2L})^{n}$, since we do not need it for this lemma, and we would like to apply this lemma in the multi-action case without that assumption.
+Algorithm 1 satisfies
+$\operatorname*{\mathbb{E}}[|M_{T}|]\in O\left(\frac{d}{\sigma}T\varepsilon\log(T+1/\varepsilon)\right)$
 Proof.
@@
 Then $w$ and $w^{h}$ evolve in the exact same way, so the distributions of $\pi_{t}$ and $\pi_{t}^{h}$ coincide.
+Also, $\varepsilon\geq 1/T$ implies that $p=1/\sqrt{\varepsilon T}\in(0,1]$.
 Thus by Lemma 5.4,
-$\operatorname*{\mathbb{E}}_{\boldsymbol{y}}\left[\sum_{t=1}^{T}\ell(t,\pi_{t})\right]-\min_{\tilde{\pi}\in\tilde{\Pi}}\sum_{t=1}^{T}\ell(t,\tilde{\pi})\leq\frac{2\log|\tilde{\Pi}|}{p^{2}}=2T\varepsilon\log|\tilde{\Pi}|$
-Since Lemma 5.4 holds for any loss function, the bound above holds for any $\boldsymbol{x}\in S^{T}$, so the bound also holds in expectation over $\boldsymbol{x}\sim\boldsymbol{\mathcal{D}}$ (which is needed for Case 1).
-Next, observe that $|M_{T}|=\sum_{t=1}^{T}\mathbf{1}(\pi_{t}(x_{t})\neq\pi^{m}(x_{t}))=\sum_{t=1}^{T}\ell(t,\pi_{t})$, so
-$\operatorname*{\mathbb{E}}_{\boldsymbol{y}}[|M_{T}|]\leq 2T\varepsilon\log|\tilde{\Pi}|+\min_{\tilde{\pi}\in\tilde{\Pi}}\sum_{t=1}^{T}\mathbf{1}(\tilde{\pi}(x_{t})\neq\pi^{m}(x_{t}))$
-Case 1: Since $\tilde{\Pi}$ is a smooth $\varepsilon$-cover of $\Pi$, we have
-$\operatorname*{\mathbb{E}}_{\boldsymbol{x}\sim\boldsymbol{\mathcal{D}}}\left[\min_{\tilde{\pi}\in\tilde{\Pi}}\sum_{t=1}^{T}\mathbf{1}(\tilde{\pi}(x_{t})\neq\pi^{m}(x_{t}))\right]\leq$ $\ \operatorname*{\mathbb{E}}_{\boldsymbol{x}\sim\boldsymbol{\mathcal{D}}}\left[\sup_{\pi\in\Pi}\min_{\tilde{\pi}\in\tilde{\Pi}}\sum_{t=1}^{T}\mathbf{1}(\tilde{\pi}(x_{t})\neq\pi(x_{t}))\right]$
-$\in$ $\ O\left(\frac{1}{\sigma}T\varepsilon\log T\sqrt{d\log(1/\varepsilon)}\right)$
-with the first step due to $\pi^{m}\in\Pi$ and the second step due to Lemma 5.3.
-The last component we need is that $|\tilde{\Pi}|\leq(41/\varepsilon)^{d}$ by construction (and such a $\tilde{\Pi}$ is guaranteed to exist by Lemma 5.1).
-Combining the above inequalities and taking the expectation over $\boldsymbol{x}\sim\boldsymbol{\mathcal{D}}$ (in addition to the randomness of the algorithm), we get
-$\operatorname*{\mathbb{E}}_{\boldsymbol{x}\sim\boldsymbol{\mathcal{D}},\,\boldsymbol{y}}\,[|M_{T}|]\leq$ $\ 2T\varepsilon\log|\tilde{\Pi}|+\operatorname*{\mathbb{E}}_{\boldsymbol{x}\sim\boldsymbol{\mathcal{D}}}\left[\min_{\tilde{\pi}\in\tilde{\Pi}}\sum_{t=1}^{T}\mathbf{1}(\tilde{\pi}(x_{t})\neq\pi^{m}(x_{t}))\right]$
-$\leq$ $\ 2dT\varepsilon\log(41/\varepsilon)+O\left(\frac{1}{\sigma}T\varepsilon\log T\sqrt{d\log(1/\varepsilon)}\right)$
-$\in$ $\ O\left(\frac{d}{\sigma}T\varepsilon\log(1/\varepsilon)\log T\right)$
+$\sum_{t=1}^{T}\operatorname*{\mathbb{E}}[\ell(t,\pi_{t})]-\min_{\tilde{\pi}\in\tilde{\Pi}}\sum_{t=1}^{T}\operatorname*{\mathbb{E}}[\ell(t,\tilde{\pi})]\leq$ $\ \frac{\log|\tilde{\Pi}|}{p^{2}}$
+$=$ $\ T\varepsilon\log|\tilde{\Pi}|$
+Since $|M_{T}|=\sum_{t=1}^{T}\mathbf{1}(\pi_{t}(x_{t})\neq\pi^{m}(x_{t}))=\sum_{t=1}^{T}\ell(t,\pi_{t})$, we have
+$\operatorname*{\mathbb{E}}[|M_{T}|]\leq T\varepsilon\log|\tilde{\Pi}|+\min_{\tilde{\pi}\in\tilde{\Pi}}\sum_{t=1}^{T}\operatorname*{\mathbb{E}}[\mathbf{1}(\tilde{\pi}(x_{t})\neq\pi^{m}(x_{t}))]$
+Case 1: Since $\tilde{\Pi}$ is a smooth $\varepsilon$-cover and $\pi^{m}\in\Pi$, Lemma 5.1 implies that $\operatorname*{\mathbb{E}}[\mathbf{1}(\tilde{\pi}(x_{t})\neq\pi^{m}(x_{t}))]\leq\varepsilon/\sigma$ for any $\tilde{\pi}\in\tilde{\Pi}$.
+Since $|\tilde{\Pi}|\leq(41/\varepsilon)^{d}$ by construction (and such a $\tilde{\Pi}$ is guaranteed to exist by Lemma 5.2), we get
+$\operatorname*{\mathbb{E}}[|M_{T}|]\leq$ $\ T\varepsilon\log((41/\varepsilon)^{d})+\min_{\tilde{\pi}\in\tilde{\Pi}}\sum_{t=1}^{T}\frac{\varepsilon}{\sigma}$
+$=$ $\ dT\varepsilon\log(41/\varepsilon)+\frac{T\varepsilon}{\sigma}$
+$\in$ $\ O\left(\frac{d}{\sigma}T\varepsilon\log(T+1/\varepsilon)\right)$
 Case 2: Since $\tilde{\Pi}$ is an adversarial cover of $\Pi$ and $\pi^{m}\in\Pi$, there exists $\tilde{\pi}\in\tilde{\Pi}$ such that $\sum_{t=1}^{T}\mathbf{1}(\tilde{\pi}(x_{t})\neq\pi^{m}(x_{t}))=0$.
 Since $|\tilde{\Pi}|\leq(eT/d)^{d}$ (with such a $\tilde{\Pi}$ guaranteed to exist by Lemma 5.3),
-$\operatorname*{\mathbb{E}}_{\boldsymbol{y}}[|M_{T}|]\leq$ $\ 2T\varepsilon\log|\tilde{\Pi}|+\min_{\tilde{\pi}\in\tilde{\Pi}}\sum_{t=1}^{T}\mathbf{1}(\tilde{\pi}(x_{t})\neq\pi^{m}(x_{t}))$
-$\leq$ $\ 2T\varepsilon d\ln(eT/d)$
-$\in$ $\ O\left(\frac{d}{\sigma}T\varepsilon\log(1/\varepsilon)\log T\right)$
+$\operatorname*{\mathbb{E}}[|M_{T}|]\leq$ $\ T\varepsilon\log|\tilde{\Pi}|+\min_{\tilde{\pi}\in\tilde{\Pi}}\sum_{t=1}^{T}\mathbf{1}(\tilde{\pi}(x_{t})\neq\pi^{m}(x_{t}))$
+$\leq$ $\ T\varepsilon d\ln(eT/d)$
+$\in$ $\ O\left(\frac{d}{\sigma}T\varepsilon\log(T+1/\varepsilon)\right)$
 as required. ∎
@@
 Consider an arbitrary $t\in[T]$.
-If $t\in Q_{T}$, then $\mu_{t}(x_{t},y_{t})=\mu_{t}^{m}(x_{t})$, so assume $t\not\in Q_{T}$.
-Let $(x^{\prime},y^{\prime})=\operatorname*{arg\,min}_{(x,y)\in X_{t}:\pi_{t}(x_{t})=y}||x_{t}-x||$.
+If $t\in Q_{T}$, then $\mu_{t}(x_{t},y_{t})=\mu_{t}^{m}(x_{t})$ trivially, so assume $t\not\in Q_{T}$.
+Let $(x^{\prime},y^{\prime})=\operatorname*{arg\,min}_{(x,y)\in S_{t}:\pi_{t}(x_{t})=y}||x_{t}-x||$.
 Since $t\not\in Q_{T}$, we must have $||x_{t}-x^{\prime}||\leq\varepsilon^{1/n}$.
-We have $y^{\prime}=\pi^{m}(x^{\prime})$ by construction of $X_{t}$ and $\pi_{t}(x_{t})=y^{\prime}$ by construction of $y^{\prime}$. Combining these with the local generalization assumption, we get
-$\mu_{t}(x_{t},y_{t})=\mu_{t}(x_{t},\pi_{t}(x_{t}))=\mu_{t}(x_{t},\pi^{m}(x^{\prime}))\geq\mu_{t}^{m}(x_{t})-L||x_{t}-x^{\prime}||\geq\mu_{t}^{m}(x_{t})-L\varepsilon^{1/n}$
+We have $y^{\prime}=\pi^{m}(x^{\prime})$ by construction of $S_{t}$ and $\pi_{t}(x_{t})=y^{\prime}$ by construction of $y^{\prime}$. Combining these with the local generalization assumption, we get
+$\mu_{t}(x_{t},y_{t})=$ $\ \mu_{t}(x_{t},\pi_{t}(x_{t}))$
+$=$ $\ \mu_{t}(x_{t},\pi^{m}(x^{\prime}))$
+$\geq$ $\ \mu_{t}^{m}(x_{t})-L||x_{t}-x^{\prime}||$
+$\geq$ $\ \mu_{t}^{m}(x_{t})-L\varepsilon^{1/n}$
 as required. ∎
 Lemma B.3.
-Assume $a_{1},\dots,a_{T},b_{1},\dots,b_{T}\in[0,1]$ and $a_{t}\geq b_{t}$ for all $t\in[T]$. Then
-$\prod_{t=1}^{T}a_{t}-\prod_{t=1}^{T}b_{t}\leq\sum_{t=1}^{T}a_{t}-\sum_{t=1}^{T}b_{t}$
-Proof.
-We proceed by induction on $T$.
-The claim is trivially satisfied for $T=1$, so suppose $T>1$ and assume that $\prod_{t=1}^{T-1}a_{t}-\prod_{t=1}^{T-1}b_{t}\leq\sum_{t=1}^{T-1}a_{t}-\sum_{t=1}^{T-1}b_{t}$.
-Then
-$\sum_{t=1}^{T}a_{t}-\sum_{t=1}^{T}b_{t}-\prod_{t=1}^{T}a_{t}+\prod_{t=1}^{T}b_{t}=$ $\ a_{T}\sum_{t=1}^{T-1}a_{t}-b_{T}\sum_{t=1}^{T-1}b_{t}-a_{T}\prod_{t=1}^{T-1}a_{t}+b_{T}\prod_{t=1}^{T-1}b_{t}$
-$=$ $\ a_{T}\left(\sum_{t=1}^{T-1}a_{t}-\prod_{t=1}^{T-1}a_{t}\right)-b_{T}\left(\sum_{t=1}^{T-1}b_{t}-\prod_{t=1}^{T-1}b_{t}\right)$
-Since $T>1$ and $a_{t}\in[0,1]$ for all $t\in[T]$, we have $\sum_{t=1}^{T-1}a_{t}\geq a_{1}\geq\sum_{t=1}^{T-1}a_{t}$.
-Thus $\sum_{t=1}^{T-1}a_{t}-\prod_{t=1}^{T-1}a_{t}\geq 0$.
-Combining this with $a_{T}\geq b_{T}$, we get
-$\sum_{t=1}^{T}a_{t}-\sum_{t=1}^{T}b_{t}-\prod_{t=1}^{T}a_{t}+\prod_{t=1}^{T}b_{t}=$ $\ a_{T}\left(\sum_{t=1}^{T-1}a_{t}-\prod_{t=1}^{T-1}a_{t}\right)-b_{T}\left(\sum_{t=1}^{T-1}b_{t}-\prod_{t=1}^{T-1}b_{t}\right)$
-$\geq$ $\ b_{T}\left(\sum_{t=1}^{T-1}a_{t}-\prod_{t=1}^{T-1}a_{t}\right)-b_{T}\left(\sum_{t=1}^{T-1}b_{t}-\prod_{t=1}^{T-1}b_{t}\right)$
-$=$ $\ b_{T}\left(\sum_{t=1}^{T-1}a_{t}-\prod_{t=1}^{T-1}a_{t}-\sum_{t=1}^{T-1}b_{t}+\prod_{t=1}^{T-1}b_{t}\right)$
-[…]
-The last step is due to $b_{T}\geq 0$ and our assumption of $\prod_{t=1}^{T-1}a_{t}-\prod_{t=1}^{T-1}b_{t}\leq\sum_{t=1}^{T-1}a_{t}-\sum_{t=1}^{T-1}b_{t}$. ∎
-Lemma B.4.
 Under the conditions of Theorem 5.1, Algorithm 1 satisfies
-$\operatorname*{\mathbb{E}}\left[R_{T}\right]\in$ $\ O\left(\frac{dL}{\sigma}T\varepsilon^{1+1/n}\log(1/\varepsilon)\log T\right)$
-$\operatorname*{\mathbb{E}}\left[\sum_{t=1}^{T}\mu_{t}^{m}(x_{t})-\sum_{t=1}^{T}\mu_{t}(x_{t},y_{t})\right]\in$ $\ O\left(\frac{dL}{\sigma}T\varepsilon^{1+1/n}\log(1/\varepsilon)\log T\right)$
+$\operatorname*{\mathbb{E}}\left[R_{T}^{\times}\right]\in$ $\ O\left(\frac{dL}{\sigma\mu_{0}^{m}}T\varepsilon^{1+1/n}\log(T+1/\varepsilon)\right)$
+$\operatorname*{\mathbb{E}}\left[R_{T}^{+}\right]\in$ $\ O\left(\frac{dL}{\sigma}T\varepsilon^{1+1/n}\log(T+1/\varepsilon)\right)$
 Proof.
 We first claim that $y_{t}=\pi^{m}(x_{t})$ for all $t\not\in M_{T}$.
-If $t\in Q_{T}$, the claim is immediate; if not, we have $y_{t}=\pi_{t}(x_{t})$, and $\pi_{t}(x_{t})=\pi^{m}(x_{t})$ due to $t\not\in M_{T}$.
-Thus $\min(\mu_{t}^{m}(x_{t}),\mu_{t}(x_{t},y_{t}))=\mu_{t}^{m}(x_{t})$ for $t\not\in M_{T}$.
-We next claim that $\mu_{t}^{m}(x_{t})-\min(\mu_{t}^{m}(x_{t}),\mu_{t}(x_{t},y_{t}))\leq L\varepsilon^{1/n}$ for all $t\in[T]$.
-If $\mu_{t}(x_{t},y_{t})\leq\mu_{t}^{m}(x_{t})$, this follows from Lemma B.2.
-If $\mu_{t}(x_{t},y_{t})>\mu_{t}^{m}(x_{t})$, then $\mu_{t}^{m}(x_{t})-\min(\mu_{t}^{m}(x_{t}),\mu_{t}(x_{t},y_{t}))=0\leq L\varepsilon^{1/n}$.
-Therefore
-$\sum_{t=1}^{T}\Big{(}\mu_{t}^{m}(x_{t})-\min(\mu_{t}^{m}(x_{t}),\mu_{t}(x_{t},y_{t}))\Big{)}\leq$ $\ \sum_{t\in M_{T}}\Big{(}\mu_{t}^{m}(x_{t})-\min(\mu_{t}^{m}(x_{t}),\mu_{t}(x_{t},y_{t}))\Big{)}$
+If $t\in Q_{T}$, the claim is immediate.
+If not, we have $y_{t}=\pi_{t}(x_{t})$ by the definition of the algorithm and $\pi_{t}(x_{t})=\pi^{m}(x_{t})$ by the definition of $t\not\in M_{T}$.
+Thus $\mu_{t}(x_{t},y_{t})=\mu_{t}^{m}(x_{t})$ for $t\not\in M_{T}$.
+For $t\in M_{T}$, Lemma B.2 implies that $\mu_{t}^{m}(x_{t})-\mu_{t}(x_{t},y_{t})\leq L\varepsilon^{1/n}$, so
+$R_{T}^{+}=$ $\ \sum_{t\in M_{T}}(\mu_{t}^{m}(x_{t})-\mu_{t}(x_{t},y_{t}))$
 $\leq$ $\ \sum_{t\in M_{T}}L\varepsilon^{1/n}$
-$=$ $\ |M_{T}|L\varepsilon^{1/n}$
-Now let $a_{t}=\mu_{t}^{m}(x_{t})$ and $b_{t}=\min(\mu_{t}^{m}(x_{t}),\mu_{t}(x_{t},y_{t}))$ for all $t\in[T]$. Then by Lemma B.3,
-$\prod_{t=1}^{T}\mu_{t}^{m}(x_{t})-\prod_{t=1}^{T}\min(\mu_{t}^{m}(x_{t}),\mu_{t}(x_{t},y_{t}))\leq$ $\ \sum_{t=1}^{T}\Big{(}\mu_{t}^{m}(x_{t})-\min(\mu_{t}^{m}(x_{t}),\mu_{t}(x_{t},y_{t}))\Big{)}$
-Since $\mu_{t}(x_{t},y_{t})\geq\min(\mu_{t}^{m}(x_{t}),\mu_{t}(x_{t},y_{t}))$ for all $t\in[T]$, we have
-$R_{T}=$ $\ \prod_{t=1}^{T}\mu_{t}^{m}(x_{t})-\prod_{t=1}^{T}\mu_{t}(x_{t},y_{t})$
-$\leq$ $\ \sum_{t=1}^{T}\Big{(}\mu_{t}^{m}(x_{t})-\min(\mu_{t}^{m}(x_{t}),\mu_{t}(x_{t},y_{t}))\Big{)}$
-$\leq$ $\ |M_{T}|L\varepsilon^{1/n}$
-Since we also have $\sum_{t=1}^{T}\mu_{t}^{m}(x_{t})-\sum_{t=1}^{T}\mu_{t}(x_{t},y_{t})\leq\sum_{t=1}^{T}(\mu_{t}^{m}(x_{t})-\min(\mu_{t}^{m}(x_{t}),\mu_{t}(x_{t},y_{t})))$, we get both of the following
-$R_{T}\leq$ $\ L\varepsilon^{1/n}|M_{T}|$
-$\sum_{t=1}^{T}\mu_{t}^{m}(x_{t})-\sum_{t=1}^{T}\mu_{t}(x_{t},y_{t})\leq$ $\ L\varepsilon^{1/n}|M_{T}|$
-Taking the expectation and applying Lemma B.1 to both of the above inequalities completes the proof. ∎
+$=$ $\ |M_{T}|L\varepsilon^{1/n}$ (1)
+[…]
+$R_{T}^{\times}\leq\frac{R_{T}^{+}}{\mu_{0}^{m}/2}\leq\frac{2|M_{T}|L\varepsilon^{1/n}}{\mu_{0}^{m}}$ (2)
+Taking the expectation and applying Lemma B.1 to Equations 1 and 2 completes the proof. ∎
 Definition B.1.
 Let $(K,||\cdot||)$ be a normed vector space and let $\delta>0$.
-Then $X\subseteq K$ is a $\delta$-packing of $K$ if for all $x,y\in X$, $||x-y||>\delta$.
+Then a multiset $S\subseteq K$ is a $\delta$-packing of $K$ if for all $a,b\in S$, $||a-b||>\delta$.
 The $\delta$-packing number of $K$, denoted $\mathcal{M}(K,||\cdot||,\delta)$, is the maximum cardinality of any $\delta$-packing of $K$.
-In this paper, we only consider the Euclidean distance norm, so we just write $M(K,||\cdot||,\delta)=M(K,\delta)$.
-Lemma B.5 (Theorem 14.2 in Wu (2020)).
+We only consider the Euclidean distance norm, so we just write $M(K,||\cdot||,\delta)=M(K,\delta)$.
+Lemma B.4 (Theorem 14.2 in (Wu, 2020)).
 If $K\subset\mathbb{R}^{n}$ is convex, bounded, and contains a ball with radius $\delta>0$, then
@@
 Lemma B.5 (Jung’s Theorem (Jung, 1901)).
-[…]
+If $S\subset\mathbb{R}^{n}$ is compact, then there exists a closed ball with radius at most $\operatorname*{diam}(S)\sqrt{\frac{n}{2(n+1)}}$ containing $S$.
 Lemma B.6.
 Under the conditions of Theorem 5.1, Algorithm 1 satisfies
-$\operatorname*{\mathbb{E}}[|Q_{T}|]\in O\left(\sqrt{\frac{T}{\varepsilon}}+\frac{d}{\sigma}T\varepsilon\log(1/\varepsilon)\log T+\frac{\operatorname*{diam}(\boldsymbol{x})^{n}}{\varepsilon}\right)$
+$\operatorname*{\mathbb{E}}[|Q_{T}|]\in O\left(\sqrt{\frac{T}{\varepsilon}}+\frac{d}{\sigma}T\varepsilon\log(T+1/\varepsilon)+\frac{\operatorname*{\mathbb{E}}[\operatorname*{diam}(\boldsymbol{x})^{n}]}{\varepsilon}\right)$
 Proof.
 […]
-Next, fix an $y\in\mathcal{Y}$ and let $X_{y}=\{x\in\boldsymbol{x}:\pi^{m}(x)=y\}$ be the set of observed inputs which share a mentor action.
-We claim that $\hat{X}_{2}\cap X_{y}$ is a packing of $X_{y}$.
+Next, fix a $y\in\mathcal{Y}$ and let $X_{y}=\{x\in\boldsymbol{x}:\pi^{m}(x)=y\}$ be the multiset of observed inputs whose mentor action is $y$.
+Also let $\hat{X}_{2}=\{x_{t}:t\in\hat{Q}_{2}\}$ be the multiset of inputs associated with time steps in $\hat{Q}_{2}$.
+Note that $|\hat{X}_{2}|=|\hat{Q}_{2}|$, since $\hat{X}_{2}$ is a multiset.
+We claim that $\hat{X}_{2}\cap X_{y}$ is a $\varepsilon^{1/n}$-packing of $X_{y}$.
 Suppose instead that there exists $x,x^{\prime}\in\hat{X}_{2}\cap X_{y}$, with $||x-x^{\prime}||\leq\varepsilon^{1/n}$.
 WLOG assume $x$ was queried after $x^{\prime}$ and let $t$ be the time step on which $x$ was queried.
-Then $(x^{\prime},\pi^{m}(x^{\prime}))\in X_{t}$.
-Also, $x,x^{\prime}\in\hat{X}_{2}\cap X_{y}$ implies that and $\pi_{t}(x_{t})=\pi^{m}(x_{t})=y=\pi^{m}(x^{\prime})$.
+Then $(x^{\prime},\pi^{m}(x^{\prime}))\in S_{t}$.
+Also, since $x,x^{\prime}\in\hat{X}_{2}$ we have $\pi_{t}(x_{t})=\pi^{m}(x_{t})=y=\pi^{m}(x^{\prime})$.
 Therefore
-$\min_{(x^{\prime\prime},y^{\prime\prime})\in X_{t}:y^{\prime\prime}=\pi_{t}(x_{t})}||x_{t}-x^{\prime\prime}||\leq||x_{t}-x^{\prime}||\leq\varepsilon^{1/n}$
-which contradicts $x_{t}\in\hat{X}$. Thus $\hat{X}_{2}\cap X_{y}$ is a $\varepsilon^{1/n}$-packing of $X_{y}$.
-By Lemma B.6, there exists a ball $B_{1}$ of diameter $\operatorname*{diam}(\boldsymbol{x})\sqrt{\frac{n}{2(n+1)}}$ which contains $\boldsymbol{x}$.
-Let $R=\operatorname*{diam}(\boldsymbol{x})\sqrt{\frac{n}{8(n+1)}}$ denote the radius of $B_{1}$.
+$\min_{(x^{\prime\prime},y^{\prime\prime})\in S_{t}:y^{\prime\prime}=\pi_{t}(x_{t})}||x_{t}-x^{\prime\prime}||\leq||x_{t}-x^{\prime}||\leq\varepsilon^{1/n}$
+which contradicts $t\in\hat{Q}$. Thus $\hat{X}_{2}\cap X_{y}$ is a $\varepsilon^{1/n}$-packing of $X_{y}$.
+By Lemma B.5, there exists a ball $B_{1}$ of radius $R:=\operatorname*{diam}(\boldsymbol{x})\sqrt{\frac{n}{2(n+1)}}$ which contains $\boldsymbol{x}$.
 Let $B_{2}$ be the ball with the same center as $B_{1}$ but with radius $\max(R,\varepsilon^{1/n})$.
-Since $X_{y}\subset\boldsymbol{x}\subset B_{1}\subset B_{2}$, $\hat{X}_{2}\cap X_{y}$ is also a $\varepsilon^{1/n}$-packing of $B_{2}$.
+Since $X_{y}\subset\boldsymbol{x}\subset B_{1}\subset B_{2}$ and $\hat{X}_{2}\cap X_{y}$ is a $\varepsilon^{1/n}$-packing of $X_{y}$, $\hat{X}_{2}\cap X_{y}$ is also a $\varepsilon^{1/n}$-packing of $B_{2}$.
 Also, $B_{2}$ must contain a ball of radius $\varepsilon^{1/n}$, so Lemma B.4 implies that
@@
 $=$ $\ \big{(}\max(R,\varepsilon^{1/n})\big{)}^{n}\,\frac{3^{n}\operatorname*{vol}(B)}{\varepsilon\operatorname*{vol}(B)}$
-$=$ $\ \max\left(\operatorname*{diam}(\boldsymbol{x})^{n}\left(\frac{n}{8(n+1)}\right)^{n/2},\>\varepsilon\right)\frac{3^{n}}{\varepsilon}$
+$=$ $\ \max\left(\operatorname*{diam}(\boldsymbol{x})^{n}\left(\frac{n}{2(n+1)}\right)^{n/2},\>\varepsilon\right)\frac{3^{n}}{\varepsilon}$
 $\leq$ $\ O\left(\frac{\operatorname*{diam}(\boldsymbol{x})^{n}}{\varepsilon}+1\right)$
-(The $+1$ is necessary for now since $\operatorname*{diam}(\boldsymbol{x})$ could theoretically be zero.) Therefore
-$\operatorname*{\mathbb{E}}[|Q_{T}|]=$ $\ \sqrt{\frac{T}{\varepsilon}}+\operatorname*{\mathbb{E}}[|\hat{X}|]$
-$=$ $\ \sqrt{\frac{T}{\varepsilon}}+\operatorname*{\mathbb{E}}[|\hat{X}_{1}|]+\operatorname*{\mathbb{E}}\left[\sum_{y\in\mathcal{Y}}|\hat{X}_{2}\cap X_{y}|\right]$
-$\leq$ $\ \sqrt{\frac{T}{\varepsilon}}+O\left(\frac{d}{\sigma}T\varepsilon\log(1/\varepsilon)\log T\right)+\sum_{y\in\mathcal{Y}}O\left(\frac{\operatorname*{diam}(\boldsymbol{x})^{n}}{\varepsilon}+1\right)$
-$\leq$ $\ \sqrt{\frac{T}{\varepsilon}}+O\left(\frac{d}{\sigma}T\varepsilon\log(1/\varepsilon)\log T\right)+|\mathcal{Y}|\cdot O\left(\frac{\operatorname*{diam}(\boldsymbol{x})^{n}}{\varepsilon}+1\right)$
-$\leq$ $\ O\left(\sqrt{\frac{T}{\varepsilon}}+\frac{d}{\sigma}T\varepsilon\log(1/\varepsilon)\log T+\frac{\operatorname*{diam}(\boldsymbol{x})^{n}}{\varepsilon}\right)$
+(The $+1$ is necessary for now since $\operatorname*{diam}(\boldsymbol{x})$ could theoretically be zero.) Since $\hat{X}_{2}\subseteq\{x_{1},\dots,x_{T}\}\subseteq\cup_{y\in\mathcal{Y}}X_{y}$, we have $|\hat{X}_{2}|\leq\sum_{y\in\mathcal{Y}}|\hat{X}_{2}\cap X_{y}|$ by the union bound.
+Therefore
+$\operatorname*{\mathbb{E}}[|Q_{T}|]\leq$ $\ \sqrt{\frac{T}{\varepsilon}}+1+\operatorname*{\mathbb{E}}[|\hat{Q}|]$
+$=$ $\ \sqrt{\frac{T}{\varepsilon}}+1+\operatorname*{\mathbb{E}}[|\hat{Q}_{1}|]+\operatorname*{\mathbb{E}}[|\hat{Q}_{2}|]$
+$=$ $\ \sqrt{\frac{T}{\varepsilon}}+1+\operatorname*{\mathbb{E}}[|\hat{Q}_{1}|]+\operatorname*{\mathbb{E}}[|\hat{X}_{2}|]$
+$\leq$ $\ \sqrt{\frac{T}{\varepsilon}}+1+\operatorname*{\mathbb{E}}[|\hat{Q}_{1}|]+\operatorname*{\mathbb{E}}\left[\sum_{y\in\mathcal{Y}}|\hat{X}_{2}\cap X_{y}|\right]$
+[…]
+$\leq$ $\ \sqrt{\frac{T}{\varepsilon}}+1+O\left(\frac{d}{\sigma}T\varepsilon\log(T+1/\varepsilon)\right)+|\mathcal{Y}|\cdot O\left(\frac{\operatorname*{\mathbb{E}}[\operatorname*{diam}(\boldsymbol{x})^{n}]}{\varepsilon}+1\right)$
+$\leq$ $\ O\left(\sqrt{\frac{T}{\varepsilon}}+\frac{d}{\sigma}T\varepsilon\log(T+1/\varepsilon)+\frac{\operatorname*{\mathbb{E}}[\operatorname*{diam}(\boldsymbol{x})^{n}]}{\varepsilon}\right)$
 as required. ∎
-Theorem 5.1 follows from Lemmas B.4 and B.7:
+Theorem 5.1 follows from Lemmas B.3 and B.6:
 See 5.1
-We then perform some arithmetic to get Theorem 5.2:
+We then perform some arithmetic to obtain Theorem 5.2:
 See 5.2
@@
 We have
-$\operatorname*{\mathbb{E}}\left[R_{T}\right]\in$ $\ O\left(\frac{dL}{\sigma}T^{1-\frac{2n}{2n+1}-\frac{2}{2n+1}}\log(1/\varepsilon)\log T\right)$
-$=$ $\ O\left(\frac{dL}{\sigma}T^{\frac{-1}{2n+1}}\log T\right)$
-and
-$\operatorname*{\mathbb{E}}[|Q_{T}|]\in$ $\ O\left(\sqrt{T^{1+\frac{2n}{2n+1}}}+\frac{d}{\sigma}T^{1-\frac{-2n}{2n+1}}\log(T^{\frac{2n}{2n+1}})\log T+T^{\frac{2n}{2n+1}}\operatorname*{diam}(\boldsymbol{x})^{n}\right)$
-$=$ $\ O\left(T^{\frac{2n+0.5}{2n+1}}+\frac{d}{\sigma}T^{\frac{1}{2n+1}}\log T+T^{\frac{2n}{2n+1}}\operatorname*{diam}(\boldsymbol{x})^{n}\right)$
-$\leq$ $\ O\left(T^{\frac{4n+1}{4n+2}}\Big{(}\frac{d}{\sigma}\log T+\operatorname*{diam}(\boldsymbol{x})^{n}\Big{)}\right)$
-If we instead use the second bound from Lemma B.4, the same arithmetic gives us:
-See 5.3
+$\operatorname*{\mathbb{E}}\left[R_{T}^{\times}\right]\in$ $\ O\left(\frac{dL}{\sigma\mu_{0}^{m}}T^{1-\frac{2n}{2n+1}-\frac{2}{2n+1}}\left(\log T+\log(T^{\frac{2n}{2n+1}})\right)\right)$
+$=$ $\ O\left(\frac{dL}{\sigma\mu_{0}^{m}}T^{\frac{-1}{2n+1}}\log T\right)$
+and similarly for $\operatorname*{\mathbb{E}}[R_{T}^{+}]$. For $\operatorname*{\mathbb{E}}[|Q_{T}|]$,
+$\operatorname*{\mathbb{E}}[|Q_{T}|]\in$ $\ O\left(\sqrt{T^{1+\frac{2n}{2n+1}}}+\frac{d}{\sigma}T^{1-\frac{-2n}{2n+1}}\left(\log T+\log(T^{\frac{2n}{2n+1}})\right)+T^{\frac{2n}{2n+1}}\operatorname*{\mathbb{E}}[\operatorname*{diam}(\boldsymbol{x})^{n}]\right)$
+$=$ $\ O\left(T^{\frac{2n+0.5}{2n+1}}+\frac{d}{\sigma}T^{\frac{1}{2n+1}}\log T+T^{\frac{2n}{2n+1}}\operatorname*{\mathbb{E}}[\operatorname*{diam}(\boldsymbol{x})^{n}]\right)$
+$\leq$ $\ O\left(T^{\frac{4n+1}{4n+2}}\Big{(}\frac{d}{\sigma}\log T+\operatorname*{\mathbb{E}}[\operatorname*{diam}(\boldsymbol{x})^{n}]\Big{)}\right)$
 B.3 Adaptive adversaries
@@
 (2024) state their result only for oblivious adversaries.
-In order for our overall proof of Theorem 5.1 to hold for adaptive adversaries, Lemma 5.4 (Lemma 3.5 in Russo et al.
-(2024)) must also hold for adaptive adversaries.
+In order for our overall proof of Theorem 5.1 to hold for adaptive adversaries, Lemma 5.4 (Lemma 3.5 in Russo et al., 2024) must also hold for adaptive adversaries.
 In this section, we argue why the proof of Lemma 5.4 (Lemma 3.5 in their paper) goes through for adaptive adversaries as well.
@@
 Recall also that Lemma 3.5 refers to the variant of Hedge which queries with probability $p=\hat{k}/T=k/T$ independently on each time step (Algorithm 2.
-More precisely, on each time step $t$, the algorithm samples a Bernoulli random variable $X_{t}\sim\text{Ber}(p)$ and queries if $X_{t}=1$.
+More precisely, on each time step $t$, the algorithm samples $X_{t}\sim\text{Bernoulli}(p)$ and queries if $X_{t}=1$.
 The key idea is that $X_{t}$ is independent of events on previous time steps.
@@
 (2024) where $i_{t}$ is the action chosen at time $t$, $i_{t}^{0}$ is the action chosen at time $t$ if a query is not issued, and $i_{t}^{*}$ is the optimal action at time $t$, we have
-$\operatorname*{\mathbb{E}}[\ell_{t}(i_{t})]=(1-p)\operatorname*{\mathbb{E}}[\ell_{t}(i_{t}^{0})]+p\operatorname*{\mathbb{E}}[\ell_{t}(i_{t}^{*})]=\left(1-\frac{k}{T}\right)\operatorname*{\mathbb{E}}[\ell_{t}(i_{t}^{0})]+\frac{k}{T}\operatorname*{\mathbb{E}}[\ell_{t}(i_{t}^{*})]$
+$\operatorname*{\mathbb{E}}[\ell_{t}(i_{t})]=$ $\ (1-p)\operatorname*{\mathbb{E}}[\ell_{t}(i_{t}^{0})]+p\operatorname*{\mathbb{E}}[\ell_{t}(i_{t}^{*})]$
+$=$ $\ \left(1-\frac{k}{T}\right)\operatorname*{\mathbb{E}}[\ell_{t}(i_{t}^{0})]+\frac{k}{T}\operatorname*{\mathbb{E}}[\ell_{t}(i_{t}^{*})]$
 The same logic applies to other statements like $\operatorname*{\mathbb{E}}[\hat{\ell}_{t}(i)\mid X_{\leq t-1},I_{\leq t-1}]=\ell_{t}(i)-\ell_{t}(i_{t}^{*})$ and immediately extends those statements to adaptive adversaries as well.
 Applying Observation 3.3.
-The other tricky part of the proof is applying Observation 3.3 using a new loss function $\hat{\ell}$ defined by ^$\hat{\ell}_{t}=\frac{T}{\hat{k}}(\ell_{t}(i)-\ell_{t}(i_{t}^{*}))\mathbf{1}(X_{t}=1)$.
+The other tricky part of the proof is applying Observation 3.3 using a new loss function $\hat{\ell}$ defined by ^$\hat{\ell}_{t}(i)=\frac{T}{\hat{k}}(\ell_{t}(i)-\ell_{t}(i_{t}^{*}))\mathbf{1}(X_{t}=1)$.
 To do so, we must argue that standard Hedge run on $\hat{\ell}$ is the “counterpart without queries” of HedgeWithQueries.
@@
 Thus standard Hedge run on $\hat{\ell}$ is the “counterpart without queries” of HedgeWithQueries.
+Note that since $\hat{\ell}$ is itself a random variable, the law of iterated expectation is necessary to formalize this.
 The rest of the proof.
@@
   2. 2.
-    Applying Lemma 3.1 to a $\hat{\ell}$.
+    Applying Lemma 3.1 to $\hat{\ell}$.
   3. 3.
@@
 Appendix C Generalizing Theorem 5.2 to many actions
-We use the standard “one versus rest” reduction (see, e.g., Chapter 29 of Shalev-Shwartz & Ben-David (2014)).
+We use the standard “one versus rest” reduction (see, e.g., Chapter 29 of Shalev-Shwartz & Ben-David, 2014).
 For each action $y$, we will learn a binary classifier which predicts whether action $y$ is the mentor’s action.
@@
 In words, for each policy $\pi:\mathcal{X}\to\mathcal{Y}$ in $\Pi$, there exists a policy $\pi_{y}:\mathcal{X}\to\{0,1\}$ in $\Pi_{y}$ such that $\pi_{y}(x)=\mathbf{1}(\pi(x)=y)$ for all $x\in\mathcal{X}$.
-Algorithm 3 runs one copy of our binary-action algorithm Algorithm 1 for each action $y\in\mathcal{Y}$.
+Algorithm 3 runs one copy of our binary-action algorithm (Algorithm 1) for each action $y\in\mathcal{Y}$.
 At each time step $t$, the copy for action $y$ returns an action $b_{t}^{y}$, with $b_{t}^{y}=1$ indicating a belief that $y=\pi^{m}(x_{t})$ and $b_{t}^{y}=0$ indicating a belief that $y\neq\pi^{m}(x_{t})$.
@@
 The result is the following theorem:
-1:function AvoidCatastropheManyActions($T\in\mathbb{N},\>\varepsilon\in\mathbb{R}_{>0},\>d\in\mathbb{N},\,$ policy class $\Pi$)
-2:     for $y\in\mathcal{Y}$ do
-3:         if $\Pi$ has VC dimension $d$ then
-4:              $\tilde{\Pi}_{y}\leftarrow$ any smooth $\varepsilon$-cover of $\Pi$ of size at most $(41/\varepsilon)^{d}$
-5:         else if $\Pi$ has Littlestone dimension $d$ then
-6:              $\tilde{\Pi}_{y}\leftarrow$ any adversarial $\varepsilon$-cover of size at most $(eT/d)^{d}$
-7:     for $t$ from $1$ to $T$ do
-8:         for $y\in\mathcal{Y}$ do
-9:              $b_{t}^{y}\leftarrow$ action from running one step of Algorithm 1 on $\Pi_{y}$ (with the same $T,\varepsilon,d$)
-10:         if $b_{t}^{y}\neq\tilde{y}$ $\forall y\in\mathcal{Y}$ and $\exists a\in\mathcal{Y}:b_{t}^{y}=1$ then
-11:              Take any action $y$ with $b_{t}^{y}=1$
-12:         else
-13:              Query the mentor
+  Inputs: $T\in\mathbb{N},\>\varepsilon\in\mathbb{R}_{>0},\>d\in\mathbb{N},\,$ policy class $\Pi$
+  for $y\in\mathcal{Y}$ do
+     if $\Pi_{y}$ has VC dimension $d$ then
+        $\tilde{\Pi}_{y}\leftarrow$ any smooth $\varepsilon$-cover of $\Pi$ of size at most $(41/\varepsilon)^{d}$
+     else if $\Pi$ has Littlestone dimension $d$ then
+        $\tilde{\Pi}_{y}\leftarrow$ any adversarial cover of size at most $(eT/d)^{d}$
+  for $t$ from $1$ to $T$ do
+     for $y\in\mathcal{Y}$ do
+        $b_{t}^{y}\leftarrow$ action at time $t$ from the copy of Algorithm 1 running on $\Pi_{y}$ (with the same $T,\varepsilon,d$)
+     if $b_{t}^{y}\neq\tilde{y}$ $\forall y\in\mathcal{Y}$ and $\exists a\in\mathcal{Y}:b_{t}^{y}=1$ then
+        Take any action $y$ with $b_{t}^{y}=1$
+     else
+        Take an arbitrary action in $\mathcal{Y}$
 Algorithm 3 extends Algorithm 1 to many actions.
@@
 Assume $\pi^{m}\in\Pi$ where either (1) $\Pi_{y}$ has finite VC dimension $d$ and $\boldsymbol{x}$ is $\sigma$-smooth or (2) $\Pi_{y}$ has finite Littlestone dimension $d$ for all $y\in\mathcal{Y}$.
-Then for any $T\in\mathbb{N}$, Algorithm 3 with $T$ and $\varepsilon=T^{\frac{-2n}{2n+1}}$ satisfies
-$\operatorname*{\mathbb{E}}\left[R_{T}\right]\in$ $\ O\left(\frac{|\mathcal{Y}|dL}{\sigma}T^{\frac{-1}{2n+1}}\log T\right)$
-$\operatorname*{\mathbb{E}}[|Q_{T}|]\in$ $\ O\left(|\mathcal{Y}|T^{\frac{4n+1}{4n+2}}\left(\frac{d}{\sigma}\log T+\operatorname*{diam}(\boldsymbol{x})^{n}\right)\right)$
+Then for any $T\in\mathbb{N}$, Algorithm 3 with $\varepsilon=T^{\frac{-2n}{2n+1}}$ satisfies
+$\operatorname*{\mathbb{E}}\left[R_{T}^{\times}\right]\in$ $\ O\left(\frac{|\mathcal{Y}|dL}{\sigma\mu_{0}^{m}}T^{\frac{-1}{2n+1}}\log T\right)$
+$\operatorname*{\mathbb{E}}\left[R_{T}^{+}\right]\in$ $\ O\left(\frac{|\mathcal{Y}|dL}{\sigma}T^{\frac{-1}{2n+1}}\log T\right)$
+$\operatorname*{\mathbb{E}}[|Q_{T}|]\in$ $\ O\left(|\mathcal{Y}|T^{\frac{4n+1}{4n+2}}\left(\frac{d}{\sigma}\log T+\operatorname*{\mathbb{E}}[\operatorname*{diam}(\boldsymbol{x})^{n}]\right)\right)$
 We use the following terminology and notation in the proof of Theorem C.1:
@@
   2. 2.
-    Let $\pi_{t}^{y}$ and $X_{t}^{y}$ refer to the values of $\pi_{t}$ and $X_{t}$ for copy $y$ of Algorithm 1.
+    Recall that $S_{t}$ refers to the value of $S$ at the start of time step $t$ in Algorithm 1.
+Let $\pi_{t}^{y}$ and $S_{t}^{y}$ refer to the values of $\pi_{t}$ and $S_{t}$ for copy $y$ of Algorithm 1.
   3. 3.
@@
   4. 4.
-    Let $M_{T}^{y}=\{t\in[T]:b_{t}^{y}\neq\pi^{my}(x_{t})\}$ be the set of time steps where $\pi_{t}^{y}$ does not correctly determine whether the mentor would take action $y$ and let $M_{T}=\{t\in[T]:y_{t}\neq\pi^{m}(x_{t})\}$ be the set of time steps where the agent’s action doesn’t match the mentor’s.
+    Let $M_{T}^{y}=\{t\in[T]:b_{t}^{y}\neq\pi^{my}(x_{t})\}$ be the set of time steps where $\pi_{t}^{y}$ does not correctly determine whether the mentor would take action $y$ and let $M_{T}=\{t\in[T]:y_{t}\neq\pi^{m}(x_{t})\}$ be the set of time steps where the agent’s action does not match the mentor’s.
 Lemma C.1.
@@
 Therefore copy $y$ of Algorithm 1 did not query at time $t$ and $\pi_{t}^{y}(x_{t})=1$.
-Let $(x^{\prime},y^{\prime})=\operatorname*{arg\,min}_{(x,y)\in X_{t}^{y}:\pi_{t}^{y}(x_{t})=y}||x_{t}-x||$.
+Let $(x^{\prime},y^{\prime})=\operatorname*{arg\,min}_{(x,y)\in S_{t}^{y}:\pi_{t}^{y}(x_{t})=y}||x_{t}-x||$.
 Then $||x_{t}-x^{\prime}||\leq\varepsilon^{1/n}$ and $y^{\prime}=\pi_{t}^{y}(x_{t})=1$.
-By construction of $X_{t}^{y}$, $y^{\prime}=\pi^{my}(x^{\prime})$ so $\pi^{my}(x^{\prime})=1$ which implies $\pi^{m}(x^{\prime})=y$. Then by the local generalization assumption,
+By construction of $S_{t}^{y}$, $y^{\prime}=\pi^{my}(x^{\prime})$ so $\pi^{my}(x^{\prime})=1$ which implies $\pi^{m}(x^{\prime})=y$. Then by the local generalization assumption,
 $\mu_{t}(x_{t},y_{t})=\mu_{t}(x_{t},y)=\mu_{t}(x_{t},\pi^{m}(x^{\prime}))\geq\mu_{t}^{m}(x_{t})-L||x_{t}-x^{\prime}||\geq\mu_{t}^{m}(x_{t})-L\varepsilon^{1/n}$
 as required. ∎
-See C.1
-Proof.
-Theorem 5.2 implies that each copy of Algorithm 1 makes $O\big{(}T^{\frac{4n+1}{4n+2}}\big{(}\frac{d}{\sigma}\log T+\operatorname*{diam}(\boldsymbol{x})^{n}\big{)}\big{)}$ queries in expectation, so by linearity of expectation, the expected number of queries made by Algorithm 3 is $O\big{(}|\mathcal{Y}|T^{\frac{4n+1}{4n+2}}\big{(}\frac{d}{\sigma}\log T+\operatorname*{diam}(\boldsymbol{x})^{n}\big{)}\big{)}$.
-Using the same argument as in the proof of Lemma B.4 (with Lemma C.2 replacing Lemma B.2), we get
-$R_{T}=$ $\ \prod_{t=1}^{T}\mu_{t}^{m}(x_{t})-\prod_{t=1}^{T}\mu_{t}(x_{t},y_{t})$
-$\leq$ $\ \sum_{t=1}^{T}\Big{(}\mu_{t}^{m}(x_{t})-\min(\mu_{t}^{m}(x_{t}),\mu_{t}(x_{t},y_{t}))\Big{)}$
-$\leq$ $\ |M_{T}|L\varepsilon^{1/n}$
-Then by Lemma C.1, $R_{T}\leq L\varepsilon^{1/n}\sum_{y\in\mathcal{Y}}|M_{T}^{y}|$. Taking the expectation and applying Lemma B.1 to each $M_{T}^{y}$ gives us
-$\operatorname*{\mathbb{E}}[R_{T}]\leq L\varepsilon^{1/n}\sum_{y\in\mathcal{Y}}O\left(\frac{d}{\sigma}T\varepsilon\log(1/\varepsilon)\log T\right)=O\left(|\mathcal{Y}|L\varepsilon^{1/n}\frac{d}{\sigma}T\varepsilon\log(1/\varepsilon)\log T\right)$
-Plugging in $\varepsilon=T^{\frac{-2n}{2n+1}}$ produces the desired bounds (see the arithmetic in the proof of Theorem 5.2 in Appendix B). ∎
-One can use the same argument to obtain a multi-action additive regret bound: simply modify the proof of Theorem C.1 to use $\sum_{t=1}^{T}\mu_{t}^{m}(x_{t})-\sum_{t=1}^{T}\mu_{t}(x_{t},y_{t})\leq|M_{T}|L\varepsilon^{1/n}$ instead of $R_{T}\leq|M_{T}|L\varepsilon^{1/n}$.
-Corollary C.1.1.
-Assume $\pi^{m}\in\Pi$ where either (1) $\Pi_{y}$ has finite VC dimension $d$ and $\boldsymbol{x}$ is $\sigma$-smooth or (2) $\Pi_{y}$ has finite Littlestone dimension $d$ for all $y\in\mathcal{Y}$.
-Then for any $T\in\mathbb{N}$, Algorithm 3 with $T$ and $\varepsilon=T^{\frac{-2n}{2n+1}}$ satisfies
-$\operatorname*{\mathbb{E}}\left[\sum_{t=1}^{T}\mu_{t}^{m}(x_{t})-\sum_{t=1}^{T}\mu_{t}(x_{t},y_{t})\right]\in$ $\ O\left(\frac{|\mathcal{Y}|dL}{\sigma}T^{\frac{-1}{2n+1}}\log T\right)$
+We now proceed to the proof of Theorem C.1.
+Proof of Theorem C.1.
+Theorem 5.2 implies that each copy of Algorithm 1 makes $O\big{(}T^{\frac{4n+1}{4n+2}}\big{(}\frac{d}{\sigma}\log T+\operatorname*{\mathbb{E}}[\operatorname*{diam}(\boldsymbol{x})^{n}]\big{)}\big{)}$ queries in expectation.
+Thus by linearity of expectation the expected number of queries made by Algorithm 3 is $O\big{(}|\mathcal{Y}|T^{\frac{4n+1}{4n+2}}\big{(}\frac{d}{\sigma}\log T+\operatorname*{\mathbb{E}}[\operatorname*{diam}(\boldsymbol{x})^{n}]\big{)}\big{)}$.
+[^13] This is an overestimate because the agent makes at most one query per time step, even if multiple copies request a query.
+Similar to the proof of Lemma B.3, we have
+[…]
+[…]
+[…]
+[…]
+Since each copy satisfies the conditions of Lemma B.1, we get
+$\operatorname*{\mathbb{E}}[R_{T}^{+}]\leq L\varepsilon^{1/n}\sum_{y\in\mathcal{Y}}O\left(\frac{d}{\sigma}T\varepsilon\log(1/\varepsilon)\log T\right)=O\left(|\mathcal{Y}|L\varepsilon^{1/n}\frac{d}{\sigma}T\varepsilon\log(1/\varepsilon)\log T\right)$
+Since $\lim_{T\to\infty}\varepsilon=0$, there exists $T_{0}$ such that $L\varepsilon^{1/n}\leq\mu_{0}^{m}/2$ for all $T\geq T_{0}$. Then by Lemma A.1,
+$\operatorname*{\mathbb{E}}[R_{T}^{\times}]\leq\frac{\operatorname*{\mathbb{E}}[R_{T}^{+}]}{\mu_{0}^{m}/2}\in O\left(|\mathcal{Y}|L\varepsilon^{1/n}\frac{d}{\sigma\mu_{0}^{m}}T\varepsilon\log(1/\varepsilon)\log T\right)$ (3)
+Plugging $\varepsilon=T^{\frac{-2n}{2n+1}}$ to the bounds above on $\operatorname*{\mathbb{E}}[R_{T}^{+}]$ and $\operatorname*{\mathbb{E}}[R_{T}^{\times}]$ yields the desired bounds (see the arithmetic in the proof of Theorem 5.2 in Appendix B).
+∎
 Appendix D There exist policy classes which are learnable in our setting but not in the standard online model
@@
 The primary takeaway here is that our algorithm can handle the class of thresholds on $[0,1]$, which is known to have infinite Littlestone dimension and thus be hard in the standard online learning model.
-(Example 21.4 in Shalev-Shwartz & Ben-David (2014)).
+(Example 21.4 in Shalev-Shwartz & Ben-David, 2014).
 Specifically, we assume a 1D input space and we allow the input sequence to be fully adversarial chosen. Instead of VC/Littlestone dimension, we consider the following notion of simplicity:
@@
 Let $S(\pi^{m})$ denote the number of segments in $\pi^{m}$.
-Bounding the number of segments is similar conceptually to VC dimension in that it limits the ability of the policy class to realize arbitrary combinations of labels (i.e., mentor actionx) on $\boldsymbol{x}$.
+Bounding the number of segments is similar conceptually to VC dimension in that it limits the ability of the policy class to realize arbitrary combinations of labels (i.e., mentor actions) on $\boldsymbol{x}$.
 For example, if $\Pi$ is the class of thresholds on $[0,1]$, every $\pi\in\Pi$ has at most two segments, and thus the positive result in this section will apply.
 This demonstrates the existence of policy classes which are learnable in our setting but not learnable in the standard online learning model, meaning that the two settings do not exactly coincide.
-We prove the following regret bound. Like our other positive results, this bound applies to both multiplicative and additive regret.
+Unlike our primary algorithm (Algorithm 1), this algorithm does require direct access to the input encoding.
+However, the point of this section is not to present a practical algorithm: it is simply to demonstrate that our setting and the standard online setting do not exactly coincide.
+We prove the following regret bound. Like our previous results, this bound applies to both multiplicative and additive regret.
 Theorem D.2.
-For any $\boldsymbol{x}\in\mathcal{X}^{T}$, any $\pi^{m}$ with $S(\pi^{m})\leq K$, and any function $g:\mathbb{N}\to\mathbb{N}$, Algorithm 4 makes at most $(\operatorname*{diam}(\boldsymbol{x})+4)g(T)$ queries and satisfies
-$R_{T}\leq$ $\ \frac{2LKT}{g(T)^{2}}$
-$\sum_{t=1}^{T}\mu_{t}^{m}(x_{t})-\sum_{t=1}^{T}\mu_{t}(x_{t},y_{t})\leq$ $\ \frac{2LKT}{g(T)^{2}}$
+For any $\boldsymbol{x}\in\mathcal{X}^{T}$, any $\pi^{m}$ with $S(\pi^{m})\leq K$, and any function $g:\mathbb{N}\to\mathbb{N}$ satisfying $g(T)\geq 2L/\mu_{0}^{m}$, Algorithm 4 satisfies
+$R_{T}^{\times}\leq$ $\ \frac{4LKT}{g(T)^{2}\mu_{0}^{m}}$
+$R_{T}^{+}\leq$ $\ \frac{2LKT}{g(T)^{2}}$
+$|Q_{T}|\leq$ $\ (\operatorname*{diam}(\boldsymbol{x})+4)g(T)$
 Choosing $g(T)=T^{c}$ for $c\in(1/2,1)$ is sufficient to subconstant regret and sublinear queries:
 Theorem D.3.
-For any $c\in(1/2,1)$, Algorithm 4 with $g(T)=T^{c}$ makes $O(T^{c}(\operatorname*{diam}(\boldsymbol{x})+1))$ queries and satisfies
-$R_{T}\leq$ $\ 2LKT^{1-2c}$
-$\sum_{t=1}^{T}\mu_{t}^{m}(x_{t})-\sum_{t=1}^{T}\mu_{t}(x_{t},y_{t})\leq$ $\ 2LKT^{1-2c}$
-Our algorithm does not need to know $L$ or the number of segments; it only needs to know $T$.
+For any $c\in(1/2,1)$, Algorithm 4 with $g(T)=T^{c}$ satisfies
+$R_{T}^{\times}\in$ $\ O\left(\frac{LKT^{1-2c}}{\mu_{0}^{m}}\right)$
+$R_{T}^{+}\in$ $\ O\left(LKT^{1-2c}\right)$
+$|Q_{T}|\in$ $\ O(T^{c}(\operatorname*{diam}(\boldsymbol{x})+1))$
+1:  function DBWRQ$(T\in\mathbb{N},\>g:\mathbb{N}\to\mathbb{N})$
+2:     $X_{Q}\leftarrow\emptyset$  (previously queried inputs)
+3:     $\pi\leftarrow\emptyset$  (records $\pi^{m}(x)$ for each $x\in X_{Q}$)
+4:     $\mathcal{B}\leftarrow\emptyset$  (The set of active buckets)
+5:     for $t$ from $1$ to $T$ do
+6:        EvaluateInput$(x_{t})$
+7:     end for
+8:  end function
+9:  function EvaluateInput$(x\in\mathcal{X})$
+10:     if $x\not\in B$ for all $B\in\mathcal{B}$ then
+11:        $B\leftarrow\left[\frac{j-1}{g(T)},\frac{j}{g(T)}\right]$ for $j\in\mathbb{Z}$ such that $x\in B$
+12:        $\mathcal{B}\leftarrow\mathcal{B}\cup\{B\}$
+13:        $n_{B}\leftarrow 0$
+14:        EvaluateInput$(x)$
+15:     else
+16:        $B\leftarrow$ any bucket containing $x$
+17:        if $X_{Q}\cap B=\emptyset$ then
+18:           Query mentor and observe $\pi^{m}(x)$
+19:           $\pi(x)\leftarrow\pi^{m}(x)$
+20:           $X_{Q}\leftarrow X_{Q}\cup\{x\}$
+21:           $n_{B}\leftarrow n_{B}+1$
+22:        else if $n_{B}<T/g(T)$ then
+23:           Let $x^{\prime}\in X_{Q}\cap B$
+24:           Take action $\pi(x^{\prime})$
+25:           $n_{B}\leftarrow n_{B}+1$
+26:        else
+27:           $B=[a,b]$
+28:           $(B_{1},B_{2})\leftarrow\Big{(}\left[a,\frac{a+b}{2}\right],\left[\frac{a+b}{2},b\right]\Big{)}$
+29:           $(x_{B_{1}},x_{B_{2}})\leftarrow(0,0)$
+30:           $\mathcal{B}\leftarrow\mathcal{B}\cup\{B_{1},B_{2}\}\setminus B$
+31:           EvaluateInput$(x)$
+32:        end if
+33:     end if
+34:  end function
+Algorithm 4 achieves subconstant regret when the mentor’s policy has a bounded number of segments.
 D.1 Intuition behind the algorithm
+We call our algorithm “Dynamic Bucketing With Routine Querying”, or DBWRQ (pronounced “DBWRQ”).
 The algorithm maintains a set of buckets which partition the observed portion of the input space.
@@
 As long as the bucket contains a query from a prior time step, local generalization allows us to bound $\mu_{t}^{m}(x_{t})-\mu_{t}(x_{t},y_{t})$ based on the length of the bucket containing $x_{t}$.
-We always query if the bucket does not contain a prior query
+We always query if the bucket does not contain a prior query; in this sense the querying is “routine”.
 The granularity of the buckets is controlled by a function $g$, with the initial buckets having length $1/g(T)$.
@@
 It turns out that splitting after $T/g(T)$ time steps is a sweet spot.
-1:function AvoidCatastrophe($T\in\mathbb{N},\>g:\mathbb{N}\to\mathbb{N}$)
-2:     $X_{Q}\leftarrow\emptyset$triangleright Previously queried inputs
-3:     $\pi\leftarrow\emptyset$triangleright Records $\pi^{m}(x)$ for each $x\in X_{Q}$
-4:     $\mathcal{B}\leftarrow\emptyset$triangleright The set of active buckets
-5:     for $t$ from $1$ to $T$ do
-6:         EvaluateInput$(x_{t})$
-7:function EvaluateInput($x\in\mathcal{X}$)
-8:     if $s\not\in B$ for all $B\in\mathcal{B}$ then $\triangleright$ No bucket containing $x$: create a new bucket and try again
-9:         $B\leftarrow\left[\frac{j-1}{g(T)},\frac{j}{g(T)}\right]$ for $j\in\mathbb{Z}$ such that $x\in B$
-10:         $\mathcal{B}\leftarrow\mathcal{B}\cup\{B\}$
-11:         $n_{B}\leftarrow 0$triangleright Number of time steps that have used $B$
-12:         EvaluateInput$(x)$
-13:     else
-14:         $B\leftarrow$ any bucket containing $x$
-15:         if $X_{Q}\cap B=\emptyset$ then $\triangleright$ No queries in this bucket
-16:              Query mentor and observe $\pi^{m}(x)$
-17:              $\pi(x)\leftarrow\pi^{m}(x)$
-18:              $X_{Q}\leftarrow X_{Q}\cup\{x\}$
-19:              $n_{B}\leftarrow n_{B}+1$
-20:         else if $n_{B}<T/g(T)$ then $\triangleright$ Bucket has a query and isn’t full: take that action
-21:              Let $x^{\prime}\in X_{Q}\cap B$
-22:              Take action $\pi(x^{\prime})$
-23:              $n_{B}\leftarrow n_{B}+1$
-24:         $\triangleright$ Bucket is full: split bucket and try again
-25:              $B=[a,b]$
-26:              $(B_{1},B_{2})\leftarrow\Big{(}\left[a,\frac{a+b}{2}\right],\left[\frac{a+b}{2},b\right]\Big{)}$
-27:              $(x_{B_{1}},x_{B_{2}})\leftarrow(0,0)$
-28:              $\mathcal{B}\leftarrow\mathcal{B}\cup\{B_{1},B_{2}\}\setminus B$
-29:              EvaluateInput$(x)$
-Algorithm 4 achieves subconstant regret when the mentor’s policy has a bounded number of segments.
-D.2 Notation for the proof
+D.2 Proof notation
 We will use the following notation throughout the proof of Theorem D.2:
@@
   2. 2.
-    Let $B_{t}$ be the bucket that is used on time step $t$ (as defined on line 14 of Algorithm 4).
+    Let $B_{t}$ be the bucket that is used on time step $t$ (as defined on line 16 of Algorithm 4).
   3. 3.
-    Let $d(B)$ be the depth of bucket $B$
+    Let $d(B)$ be the depth of bucket $B$.
     1. (a)
-      Buckets created on line 9 are depth 0.
+      Buckets created on line 11 are depth 0.
     2. (b)
-      We refer to $B_{1},B_{2}$ created on line 26 as the children of the bucket $B$ defined on line 14.
+      We refer to $B_{1},B_{2}$ created on line 28 as the children of the bucket $B$ defined on line 16.
     3. (c)
@@
   4. 4.
-    Viewing the set of buckets are a binary tree defined by the “child” relation, we use the terms “ancestor” and ”descendant” in accordance with their standard tree definitions.
+    Viewing the set of buckets as a binary tree defined by the “child” relation, we use the terms “ancestor” and ”descendant” in accordance with their standard tree definitions.
   5. 5.
@@
   3. 3.
-    Bound the sum of bucket lengths on time steps where we make a mistake (Lemma D.4), with Lemma D.3 as an intermediate step. This captures the total amount of suboptimality.
+    Bound the sum of bucket lengths on time steps where we make a mistake (Lemma D.4), with Lemma D.3 as an intermediate step. This captures the “total amount of suboptimality”.
   4. 4.
-    As in the proof of Theorem 5.2, Lemma B.3 transforms the multiplicative objective into an additive form. Lemma D.5 bounds the additive objective using Lemmas D.2 and D.4.
+    Lemma D.5 uses Lemma D.2 and Lemma D.4 to bound the regret.
   5. 5.
-    Combining Lemmas D.5 and B.3 bounds the regret (Lemma D.6).
-  6. 6.
-    Theorem D.2 directly follows from Lemmas D.1 and D.6.
+    Theorem D.2 directly follows from Lemmas D.1 and D.5.
 D.4 Proof
@@
 Algorithm 4 performs at most one query per bucket, so the total number of queries is bounded by the total number of buckets.
-There are two ways to create a bucket: from scratch (line 9), or by splitting an existing bucket (line 26).
+There are two ways to create a bucket: from scratch (line 11), or by splitting an existing bucket (line 28).
 Since depth 0 buckets overlap only at their boundaries, and each depth 0 bucket has length $1/g(T)$, at most $g(T)\max_{t,t^{\prime}\in[T]}|x_{t}-x_{t^{\prime}}|=g(T)\operatorname*{diam}(\boldsymbol{x})$ depth 0 buckets are subsets of the interval $[\min_{t\in[T]}x_{t},\max_{t\in[T]}x_{t}]$.
@@
 Proof.
-If we query the mentor at time $t$, $\mu_{t}(x_{t},y_{t})=\mu_{t}^{m}(x_{t})$.
-Thus assume we do not query the mentor at time $t$: then there exists $x^{\prime}\in B_{t}$ (as defined on line 21 of Algorithm 4) such that $y_{t}=\pi(x^{\prime})=\pi^{m}(x^{\prime})$.
+If we query at time $t$, then $\mu_{t}(x_{t},y_{t})=\mu_{t}^{m}(x_{t})$.
+Thus assume we do not query at time $t$: then there exists $x^{\prime}\in B_{t}$ (as defined on line 23 of Algorithm 4) such that $y_{t}=\pi(x^{\prime})=\pi^{m}(x^{\prime})$.
 Since $x_{t}$ and $x^{\prime}$ are both in $B_{t}$, $|x_{t}-x^{\prime}|\leq\operatorname*{len}(B_{t})$.
-Then by the local generalization assumption, $\mu_{t}(x_{t},y_{t})=\mu_{t}(x_{t},\pi^{m}(x^{\prime}))\geq\mu_{t}^{m}(x_{t})-L||x_{t}-x^{\prime}||\geq\mu_{t}^{m}(x_{t})-L\operatorname*{len}(B_{t})$.
+Then by local generalization, $\mu_{t}(x_{t},y_{t})=\mu_{t}(x_{t},\pi^{m}(x^{\prime}))\geq\mu_{t}^{m}(x_{t})-L||x_{t}-x^{\prime}||\geq\mu_{t}^{m}(x_{t})-L\operatorname*{len}(B_{t})$.
 ∎
@@
 Therefore any pair of adjacent segments $X_{j}$ and $X_{j+1}$, there is at most one bucket in $\mathcal{B}_{V}^{\prime}$ which contains an open neighborhood around their boundary.
-Since there are at most $K-1$ pairs of adjacent segments, we have $|\mathcal{B}_{V}^{\prime}|\leq K-1\leq K$.
-∎
+Since there are at most $K-1$ pairs of adjacent segments, we have $|\mathcal{B}_{V}^{\prime}|\leq K-1\leq K$.∎
 Lemma D.4.
@@
 $\sum_{t\in M_{T}}\operatorname*{len}(B_{t})=\sum_{B\in\mathcal{B}_{V}}\ \sum_{t\in M_{T}:B=B_{t}}\operatorname*{len}(B_{t})$
-Next, observe that every $B\in\mathcal{B}_{V}\setminus\mathcal{B}_{V}^{\prime}$ must have a descendent in $\mathcal{B}_{V}^{\prime}$: otherwise we would have $B\in\mathcal{B}_{V}^{\prime}$.
+Next, observe that every $B\in\mathcal{B}_{V}\setminus\mathcal{B}_{V}^{\prime}$ must have a descendant in $\mathcal{B}_{V}^{\prime}$: otherwise we would have $B\in\mathcal{B}_{V}^{\prime}$.
 Let $\mathcal{A}(B)$ denote the set of ancestors of $B$, plus $B$ itself.
@@
 […]
-$\sum_{B\in\mathcal{A}(B^{\prime})}$ |$\ \frac{|\{t\in M_{T}:B=B_{t}\}|}{g(T)2^{d(B)}}\leq\frac{T}{g(T)^{2}}\sum_{B\in\mathcal{A}(B^{\prime})}\ \frac{1}{2^{d(B)}}$
-$=$ $\ \frac{T}{g(T)^{2}}\sum_{d=0}^{d(B^{\prime})}\frac{1}{2^{d}}\leq\frac{T}{g(T)^{2}}\sum_{d=0}^{\infty}\frac{1}{2^{d}}=\frac{2T}{g(T)^{2}}$
+$\sum_{B\in\mathcal{A}(B^{\prime})}\frac{|\{t\in M_{T}:B=B_{t}\}|}{g(T)2^{d(B)}}\leq$ $\ \frac{T}{g(T)^{2}}\sum_{B\in\mathcal{A}(B^{\prime})}\ \frac{1}{2^{d(B)}}$
+$=$ $\ \frac{T}{g(T)^{2}}\sum_{d=0}^{d(B^{\prime})}\frac{1}{2^{d}}$
+$\leq$ $\ \frac{T}{g(T)^{2}}\sum_{d=0}^{\infty}\frac{1}{2^{d}}$
+$=$ $\ \frac{2T}{g(T)^{2}}$
 Then by Lemma D.3,
@@
 Under the conditions of Theorem D.2, Algorithm 4 satisfies
-$\sum_{t=1}^{T}\Big{(}\mu_{t}^{m}(x_{t})-\min(\mu_{t}^{m}(x_{t}),\mu_{t}(x_{t},y_{t}))\Big{)}\leq\frac{2LKT}{g(T)^{2}}$
+$R_{T}^{\times}\leq$ $\ \frac{2LKT}{\mu_{0}^{m}g(T)^{2}}$
+$R_{T}^{+}\leq$ $\ \frac{2LKT}{g(T)^{2}}$
 Proof.
-For $t\not\in M_{T}$ we have $\min(\mu_{t}^{m}(x_{t}),\mu_{t}(x_{t},y_{t}))=\mu_{t}^{m}(x_{t})$ by definition, and Lemma D.2 implies that $\min(\mu_{t}^{m}(x_{t}),\mu_{t}(x_{t},y_{t}))\geq L\operatorname*{len}(B_{t})$ for all $t\in[T]$.
-Thus
-$\sum_{t=1}^{T}\Big{(}\mu_{t}^{m}(x_{t})-\min(\mu_{t}^{m}(x_{t}),\mu_{t}(x_{t},y_{t}))\Big{)}\leq$ $\ \sum_{t\in M_{T}}\Big{(}\mu_{t}^{m}(x_{t})-\min(\mu_{t}^{m}(x_{t}),\mu_{t}(x_{t},y_{t}))\Big{)}$
-$\leq$ $\ L\sum_{t\in M_{T}}\operatorname*{len}(B_{t})$
-Then by Lemma D.4,
-$\sum_{t=1}^{T}\Big{(}\mu_{t}^{m}(x_{t})-\min(\mu_{t}^{m}(x_{t}),\mu_{t}(x_{t},y_{t}))\Big{)}\leq\frac{2LKT}{g(T)^{2}}$
-as required. ∎
-Lemma D.6.
-Under the conditions of Theorem D.2, Algorithm 4 satisfies
-$R_{T}\leq$ $\ \frac{2LKT}{g(T)^{2}}$
-$\sum_{t=1}^{T}\mu_{t}^{m}(x_{t})-\sum_{t=1}^{T}\mu_{t}(x_{t},y_{t})\leq$ $\ \frac{2LKT}{g(T)^{2}}$
-Proof.
-Let $a_{t}=\mu_{t}^{m}(x_{t})$ and $b_{t}=\min(\mu_{t}^{m}(x_{t}),\mu_{t}(x_{t},y_{t}))$ for all $t\in[T]$. Then by Lemma B.3,
-$\prod_{t=1}^{T}\mu_{t}^{m}(x_{t})-\prod_{t=1}^{T}\min(\mu_{t}^{m}(x_{t}),\mu_{t}(x_{t},y_{t}))\leq$ $\ \sum_{t=1}^{T}\Big{(}\mu_{t}^{m}(x_{t})-\min(\mu_{t}^{m}(x_{t}),\mu_{t}(x_{t},y_{t}))\Big{)}$
-Since $\mu_{t}(x_{t},y_{t})\geq\min(\mu_{t}^{m}(x_{t}),\mu_{t}(x_{t},y_{t}))$ for all $t\in[T]$, we have
-$R_{T}=\prod_{t=1}^{T}\mu_{t}^{m}(x_{t})-\prod_{t=1}^{T}\mu_{t}(x_{t},y_{t})\leq\sum_{t=1}^{T}\Big{(}\mu_{t}^{m}(x_{t})-\min(\mu_{t}^{m}(x_{t}),\mu_{t}(x_{t},y_{t}))\Big{)}$
-Similarly,
-$\sum_{t=1}^{T}\mu_{t}^{m}(x_{t})-\sum_{t=1}^{T}\mu_{t}(x_{t},y_{t})\leq\sum_{t=1}^{T}\Big{(}\mu_{t}^{m}(x_{t})-\min(\mu_{t}^{m}(x_{t}),\mu_{t}(x_{t},y_{t}))$
-Applying Lemma D.5 to the above two inequalities completes the proof. ∎
+We have
+[…]
+[…]
+[…]
+Since $g(T)\geq 2L/\mu_{0}^{m}$ and every bucket length is at most $\frac{1}{g(T)}$,
+$\mu_{t}(x_{t},y_{t})\geq$ $\ \mu_{t}^{m}(x_{t})-L\operatorname*{len}(B_{t})$
+$\geq$ $\ \mu_{0}^{m}-\frac{L}{g(T)}$
+$\geq$ $\ \mu_{0}^{m}-\mu_{0}^{m}/2$
+$\geq$ $\ \mu_{0}^{m}/2$
+[…]
+Invoking Lemma A.1 completes the proof:
+$R_{T}^{\times}\leq\frac{2R_{T}^{+}}{\mu_{0}^{m}}\leq\frac{4LKT}{g(T)^{2}\mu_{0}^{m}}$
 Theorem D.2 follows from Lemma D.1 and Lemma D.5.
-Appendix E Other proofs
+Appendix E Properties of local generalization
 Proposition E.1 states that Lipschitz continuity implies local generalization when the mentor is optimal.
@@
 $\mu^{m}(x)+2L||x-x^{\prime}||\geq\mu^{m}(x)\geq\mu(x,\pi^{m}(x^{\prime}))$
-Thus $-2L||x-x^{\prime}||\leq\mu(x,\pi^{m}(x^{\prime}))-\mu^{m}(x)\leq 2L||x-x^{\prime}||$.
-This is equivalent to $|\mu(x,\pi^{m}(x^{\prime}))-\mu^{m}(x)|\leq 2L||x-x^{\prime}||$, completing the proof.
-Proposition E.2 states that whether subconstant regret is achievable does not depend on whether we require expected sublinear queries or worst-case sublinear queries.
-Proposition E.2.
-Suppose an algorithm satisfies $\lim_{T\to\infty}\sup_{\boldsymbol{\mu},\pi^{m}}\operatorname*{\mathbb{E}}[R_{T}]=0$ and $\sup_{\boldsymbol{\mu},\pi^{m}}\operatorname*{\mathbb{E}}[|Q_{T}|]\in o(T)$.
-Then there exists $h:\mathbb{N}\to\mathbb{N}$ such that (1) $h(T)\in o(T)$ and (2) if the algorithm is modified to simply stop querying if the number of queries reaches $h(T)$, the algorithm still satisfies $\lim_{T\to\infty}\sup_{\boldsymbol{\mu},\pi^{m}}\operatorname*{\mathbb{E}}[R_{T}]=0$.
-Proof.
-We use $Q_{T},R_{T}$ to refer to the queries and regret of the original algorithm, and $Q_{T}^{\prime},R_{T}^{\prime}$ to refer to the queries and regret of the modified algorithm.
-Since $\sup_{\boldsymbol{\mu},\pi^{m}}\operatorname*{\mathbb{E}}[|Q_{T}|]\in o(T)$, there exists $g:\mathbb{N}\to\mathbb{N}$ such that $\sup_{\boldsymbol{\mu},\pi^{m}}\operatorname*{\mathbb{E}}[|Q_{T}|]\leq g(T)$ and $g(T)\in o(T)$.
-Let $h(T)=\sqrt{g(T)T}$; then $h(T)\in o(T)$ by Lemma A.1.
-Markov’s inequality implies that
-$\Pr\big{[}|Q_{T}|>h(T)\big{]}\leq\frac{\operatorname*{\mathbb{E}}[|Q_{T}|]}{h(T)}\leq\frac{g(T)}{\sqrt{g(T)T}}=\sqrt{\frac{g(T)}{T}}$
-Let $\xi$ denote the event that at some point, the original algorithm would query, but the modified algorithm cannot because $|Q_{T}^{\prime}|=h(T)$.
-Then $\Pr[\xi]\leq\Pr[|Q_{T}|>h(T)]$ (the inequality is because the modified algorithm might not want to query more anyway).
-Also note that the algorithms are equivalent if $\xi$ does not occur, so $\operatorname*{\mathbb{E}}[R_{T}^{\prime}\mid\neg\xi]=\operatorname*{\mathbb{E}}[R_{T}]$.
-Hence
-$\operatorname*{\mathbb{E}}[R_{T}^{\prime}]=$ $\ \operatorname*{\mathbb{E}}[R_{T}^{\prime}\mid\neg\xi]\Pr[\neg\xi]+\operatorname*{\mathbb{E}}[R_{T}^{\prime}\mid\xi]\Pr[\xi]$
-$\leq$ $\operatorname*{\mathbb{E}}[R_{T}]\cdot 1+1\cdot\Pr[\xi]$
-$\leq$ $\ \operatorname*{\mathbb{E}}[R_{T}]+\sqrt{\frac{g(T)}{T}}$
-Since $g(T)\in o(T)$, we get
-$\lim_{T\to\infty}\ \sup_{\boldsymbol{\mu},\pi^{m}}\ \operatorname*{\mathbb{E}}[R_{T}^{\prime}]\leq$ $\ \lim_{T\to\infty}\ \sup_{\boldsymbol{\mu},\pi^{m}}\ \left(\operatorname*{\mathbb{E}}[R_{T}]+\sqrt{\frac{g(T)}{T}}\right)$
-$=$ $\ \lim_{T\to\infty}\ \sup_{\boldsymbol{\mu},\pi^{m}}\ \operatorname*{\mathbb{E}}[R_{T}]+\lim_{T\to\infty}\ \sqrt{\frac{g(T)}{T}}$
-[…]
-as required. ∎
+So $-2L||x-x^{\prime}||\leq\mu(x,\pi^{m}(x^{\prime}))-\mu^{m}(x)\leq 2L||x-x^{\prime}||$ and therefore $|\mu(x,\pi^{m}(x^{\prime}))-\mu^{m}(x)|\leq 2L||x-x^{\prime}||$.∎
 Theorem E.2 shows that avoiding catastrophe is impossible without local generalization, even when $\boldsymbol{x}$ is $\sigma$-smooth and $\Pi$ has finite VC dimension.
@@
 However, the proof of Theorem E.2 is a bit easier, since we only need the agent to make a single mistake.
-Like the proof of Theorem 4.1, the proof of Theorem E.3 assumes worst-case sublinear queries.
-However, recall from Proposition E.2 that the distinction between worst-case sublinear queries and expected sublinear queries is not significant.
+Note that Theorem E.2 as stated only provides a bound on $R_{T}^{\times}$. A similar bound can be obtained for $R_{T}^{+}$, but it is more tedious and we do not believe it would add much to the paper.
 Theorem E.2.
 Let $\mathcal{X}=[0,1]$ and $\mathcal{Y}=\{0,1\}$.
-Assume each input is sampled i.i.d.
-from the uniform distribution on $\mathcal{X}$ and define the mentor policy class by the set of intervals within $\mathcal{X}$, i.e., $\Pi=\{\pi:\exists a,b\in[0,1]\text{ s.t }\pi(x)=\mathbf{1}(x\in[a,b])\ \forall x\in\mathcal{X}\}$.
-Then without the local generalization assumption, any algorithm with sublinear queries satisfies $\lim_{T\to\infty}\sup_{\boldsymbol{\mu},\pi^{m}}\operatorname*{\mathbb{E}}[R_{T}]=1$.
+Let each input be sampled i.i.d.
+from the uniform distribution on $\mathcal{X}$ and define the mentor policy class as the set of intervals within $\mathcal{X}$, i.e., $\Pi=\{\pi:\exists a,b\in[0,1]\text{ s.t }\pi(x)=\mathbf{1}(x\in[a,b])\ \forall x\in\mathcal{X}\}$.
+Then without the local generalization assumption, any algorithm with sublinear queries satisfies $\lim_{T\to\infty}\sup_{\boldsymbol{\mu},\pi^{m}}\operatorname*{\mathbb{E}}[R_{T}^{\times}]=\infty$.
 Proof.
 Part 1: Setup.
-Consider any algorithm which makes sublinear worst-case queries: then there exists $g:\mathbb{N}\to\mathbb{N}$ where $\sup_{\boldsymbol{\mu},\pi^{m}}|Q_{T}|\leq g(T)$ and $g(T)\in o(T)$.
-Define $f(T):=\sqrt{(g(T)+1)T}$; by Lemma A.1, $g(T)\in o(f(T))$ and $f(T)\in o(T)$.
+Consider any algorithm which makes sublinear worst-case queries: then there exists $g:\mathbb{N}\to\mathbb{N}$ where $\sup_{\boldsymbol{\mu},\pi^{m}}\operatorname*{\mathbb{E}}[|Q_{T}|]\leq g(T)$ and $g(T)\in o(T)$.
+WLOG assume $g(T)\geq 0$ for all $T$; if not, redefine $g(T)$ to be $\max(g(T),1)$.
+Define $f(T):=\lceil\sqrt{g(T)T}\rceil$; by Lemma A.2, $g(T)\in o(f(T))$ and $f(T)\in o(T)$.
 Divide $\mathcal{X}$ into $f(T)$ equally sized sections $X_{1},\dots,X_{f(T)}$ in the exactly the same way as in Section 4.2; see also Figure 1.
@@
 In words, the mentor takes action 1 iff the input is in section $j^{m}$, and the agent receives payoff 1 if its action matches the mentor’s and zero otherwise.
-Since any choice of $j^{m}$ defines a valid $\mu$ and $\pi^{m}$, we have
-$\sup_{\boldsymbol{\mu},\pi^{m}}\ \operatorname*{\mathbb{E}}_{\boldsymbol{x},\boldsymbol{y}}\ [R_{T}(\boldsymbol{x},\boldsymbol{y},\boldsymbol{\mu},\pi^{m})]\geq\operatorname*{\mathbb{E}}_{j^{m}}\ \operatorname*{\mathbb{E}}_{\boldsymbol{x},\boldsymbol{y}}\ [R_{T}(\boldsymbol{x},\boldsymbol{y},(\mu,\dots,\mu),\pi^{m})]$
+Since any choice of $j^{m}$ defines a valid $\mu$ and $\pi^{m}$,
+$\sup_{\boldsymbol{\mu},\pi^{m}}\ \operatorname*{\mathbb{E}}_{\boldsymbol{x},\boldsymbol{y}}\ [R_{T}^{+}(\boldsymbol{x},\boldsymbol{y},\boldsymbol{\mu},\pi^{m})]\geq\operatorname*{\mathbb{E}}_{j^{m}}\ \operatorname*{\mathbb{E}}_{\boldsymbol{x},\boldsymbol{y}}\ [R_{T}^{+}(\boldsymbol{x},\boldsymbol{y},(\mu,\dots,\mu),\pi^{m})]$
 Let $J_{\neg Q}=\{j\in[f(T)]:x_{t}\not\in X_{j}\ \forall t\in Q_{T}\}$ be the set of sections which are never queried.
-Let $j_{1},\dots,j_{k}$ be the sequence of sections queried by the agent: then $k\leq|Q_{T}|\leq g(T)$.
+Let $j_{1},\dots,j_{k}$ be the sequence of sections queried by the agent: then $k=|Q_{T}|\leq g(T)$.
 Part 2: The agent is unlikely to determine $j^{m}$. By the chain rule of probability,
@@
 There are at least $f(T)-i+1$ such sections, since there are $i-1$ prior queries at this point.
-Thus $\Pr[j_{i}\neq j^{m}\mid j_{r}\neq j^{m}\ \forall r<i]\leq\frac{f(T)-i}{f(T)-i+1}$ (the inequality is because it could also be 0 if $j_{i}=j_{r}$ for some $i<r$).
+Thus $\Pr[j_{i}\neq j^{m}\mid j_{r}\neq j^{m}\ \forall r<i]\geq\frac{f(T)-i}{f(T)-i+1}$ (the inequality is because this probability could also be 1 if $j_{i}=j_{r}$ for some $i<r$).
 Therefore
-$\Pr\big{[}j^{m}\in J_{\neg Q}\big{]}\leq$ $\ \prod_{i=1}^{k}\frac{f(T)-i}{f(T)-i+1}$
+$\Pr\big{[}j^{m}\in J_{\neg Q}\big{]}\geq$ $\ \prod_{i=1}^{k}\frac{f(T)-i}{f(T)-i+1}$
 $=$ $\ \frac{f(T)-1}{f(T)}\cdot\frac{f(T)-2}{f(T)-1}\dots\frac{f(T)-k+1}{f(T)-k+2}\cdot\frac{f(T)-k}{f(T)-k+1}$
@@
 For the rest of Part 3, assume $j^{m}\in J_{\neg Q}\text{ and }|V_{j^{m}}|>0$.
-Since $j^{m}\in J_{\neg Q}$, the agent has no information about $j^{m}$ other than that it is in $J_{\neg Q}$.
-This means that for all $j\in J_{\neg Q}$ and $t\in V_{j}$, $j^{m}$ is conditionally (under the condition of $j^{m}\in J_{\neg Q}$) independent of $y_{t}$.
-We proceed by case analysis.
-Case 1: For all $j\in J_{\neg Q},t\in V_{j}$, we have $y_{t}=0$.
+Case 1: For all $j\in J_{\neg Q}$ and $t\in V_{j}$, we have $y_{t}=0$.
 In particular, this holds for $j=j^{m}$, and we know there exists at least one $t\in V_{j^{m}}$ since $|V_{j^{m}}|>0$.
-Then $y_{t}\neq\pi^{m}(x_{t})$, so $\mu(x_{t},y_{t})=0$ and thus $\Pr\left[\prod_{r=1}^{T}\mu(x_{r},y_{r})=0\>\Big{|}\>j^{m}\in J_{\neg Q}\text{and }|V_{j^{m}}|>0\right]=1$.
-Case 2: There exists $j\in J_{\neg Q},t\in V_{j}$ with $y_{t}=1$. Then $\mu(x_{t},y_{t})=0$ unless $j=j^{m}$, so
-$\Pr\left[\prod_{r=1}^{T}\mu(x_{r},y_{r})=0\>\Big{|}\>j^{m}\in J_{\neg Q}\text{ and }|V_{j^{m}}|>0\right]\geq$ $\ \Pr\Big{[}\mu(x_{t},y_{t})=0\mid j^{m}\in J_{\neg Q}\text{ and }|V_{j^{m}}|>0\Big{]}$
-$=$ $\ \Pr\Big{[}j\neq j^{m}\mid j^{m}\in J_{\neg Q}\text{ and }|V_{j^{m}}|>0\Big{]}$
-Conditioned on $j^{m}\in J_{\neg Q}$, $j^{m}$ is uniformly distributed across $J_{\neg Q}$, so
-$\Pr\left[\prod_{r=1}^{T}\mu(x_{r},y_{r})=0\>\Big{|}\>j^{m}\in J_{\neg Q}\text{ and }|V_{j^{m}}|>0\right]\geq 1-\frac{1}{|J_{\neg Q}|}\geq 1-\frac{1}{f(T)-g(T)}$
+Then $y_{t}\neq\pi^{m}(x_{t})$, so $\mu(x_{t},y_{t})=0$ and thus $\Pr\left[\prod_{r=1}^{T}\mu(x_{r},y_{r})=0\>\Big{|}\>j^{m}\in J_{\neg Q}\text{and }|V_{j^{m}}|>0\text{ and }y_{t}=0\ \forall j\in J_{\neg Q},t\in V_{j}\right]=1$.
+Case 2: There exists $j\in J_{\neg Q}$ and $t\in V_{j}$ with $y_{t}=1$. Then $\mu(x_{t},y_{t})=0$ unless $j=j^{m}$, so
+$\ \Pr\left[\prod_{r=1}^{T}\mu(x_{r},y_{r})=0\>\Big{|}\>j^{m}\in J_{\neg Q}\text{ and }|V_{j^{m}}|>0\text{ and }\exists j\in J_{\neg Q},t\in V_{j}\text{ s.t. }y_{t}=1\right]$
+$\geq$ $\ \Pr\Big{[}\mu(x_{t},y_{t})=0\ \Big{|}\ j^{m}\in J_{\neg Q}\text{ and }|V_{j^{m}}|>0\text{ and }\exists j\in J_{\neg Q},t\in V_{j}\text{ s.t. }y_{t}=1\Big{]}$
+$=$ $\ \Pr\Big{[}j\neq j^{m}\ \Big{|}\ j^{m}\in J_{\neg Q}\text{ and }|V_{j^{m}}|>0\text{ and }\exists j\in J_{\neg Q},t\in V_{j}\text{ s.t. }y_{t}=1\Big{]}$
+Since $j^{m}\in J_{\neg Q}$, the agent has no information about $j^{m}$ other than that it is in $J_{\neg Q}$. This means that $j^{m}$ is uniformly distributed across $J_{\neg Q}$, so
+$\Pr\left[\prod_{r=1}^{T}\mu(x_{r},y_{r})=0\>\Big{|}\>j^{m}\in J_{\neg Q}\text{ and }|V_{j^{m}}|>0\text{ and }\exists j\in J_{\neg Q},t\in V_{j}\text{ s.t.
+}y_{t}=1\right]\geq 1-\frac{1}{|J_{\neg Q}|}\geq 1-\frac{1}{f(T)-g(T)}$
 Combining Case 1 and Case 2, we get the overall bound of
@@
 $\Pr\left[\prod_{t=1}^{T}\mu(x_{t},y_{t})=0\right]\geq$ $\ \Pr\left[\prod_{t=1}^{T}\mu(x_{t},y_{t})=0\text{ and }j^{m}\in J_{\neg Q}\text{ and }|V_{j^{m}}|>0\right]$
-$\quad\quad\quad=\Pr\left[\prod_{t=1}^{T}\mu(x_{t},y_{t})=0\>\Big{|}\>j^{m}\in J_{\neg Q}\text{ and }|V_{j^{m}}|>0\right]\cdot\Pr\Big{[}j^{m}\in J_{\neg Q}\text{ and }|V_{j^{m}}|>0\Big{]}$
-$\quad\quad\quad\geq\left(1-\frac{1}{f(T)-g(T)}\right)\left(1-\frac{g(T)}{f(T)}-\exp\left(\frac{-T}{16f(T)}\right)\right)$
+$=$ $\ \Pr\left[\prod_{t=1}^{T}\mu(x_{t},y_{t})=0\>\Big{|}\>j^{m}\in J_{\neg Q}\text{ and }|V_{j^{m}}|>0\right]\cdot\Pr\Big{[}j^{m}\in J_{\neg Q}\text{ and }|V_{j^{m}}|>0\Big{]}$
+$\geq$ $\ \left(1-\frac{1}{f(T)-g(T)}\right)\left(1-\frac{g(T)}{f(T)}-\exp\left(\frac{-T}{16f(T)}\right)\right)$
 For brevity, let $\alpha(T)$ denote this final bound. Since $g(T)\in o(f(T))$ and $f(T)\in o(T)$, we have
-$\lim_{T\to\infty}\alpha(T)=$ $\ \lim_{T\to\infty}\left(1-\frac{1}{f(T)-g(T)}\right)\left(1-\frac{g(T)}{f(T)}-\exp\left(\frac{-T}{16f(T)}\right)\right)$
-$=$ $\ (1-0)(1-0-0)$
-[…]
-Part 4: Putting it all together. Since $\prod_{t=1}^{T}\mu(x_{t},y_{t})\leq 1$ always, we have
-$\operatorname*{\mathbb{E}}_{j^{m}}\ \operatorname*{\mathbb{E}}_{\boldsymbol{x},\boldsymbol{y}}\ \left[\prod_{t=1}^{T}\mu(x_{t},y_{t})\right]=$ $\ \operatorname*{\mathbb{E}}_{j^{m}}\ \operatorname*{\mathbb{E}}_{\boldsymbol{x},\boldsymbol{y}}\ \left[\prod_{t=1}^{T}\mu(x_{t},y_{t})\>\Big{|}\>\prod_{t=1}^{T}\mu(x_{t},y_{t})=0\right]\cdot\Pr\left[\prod_{t=1}^{T}\mu(x_{t},y_{t})=0\right]$
-$+$ $\ \operatorname*{\mathbb{E}}_{j^{m}}\ \operatorname*{\mathbb{E}}_{\boldsymbol{x},\boldsymbol{y}}\ \left[\prod_{t=1}^{T}\mu(x_{t},y_{t})\>\Big{|}\>\prod_{t=1}^{T}\mu(x_{t},y_{t})\neq 0\right]\cdot\Pr\left[\prod_{t=1}^{T}\mu(x_{t},y_{t})\neq 0\right]$
-$\leq 0$ $\cdot\Pr\left[\prod_{t=1}^{T}\mu(x_{t},y_{t})=0\right]+1\cdot\left(1-\Pr\left[\prod_{t=1}^{T}\mu(x_{t},y_{t})=0\right]\right)$
-$\leq 1$ $-\alpha(T)$
+$\lim_{T\to\infty}\alpha(T)=\lim_{T\to\infty}\left(1-\frac{1}{f(T)-g(T)}\right)\left(1-\frac{g(T)}{f(T)}-\exp\left(\frac{-T}{16f(T)}\right)\right)=(1-0)(1-0-0)=1$
+Part 4: Putting it all together.
+Consider any $\varepsilon\in(0,1]$; to avoid dealing with infinite expectations, we will deal with $\Pr[\prod_{t=1}^{T}\mu(x_{t},y_{t})\leq\varepsilon]$ instead of $\Pr[\prod_{t=1}^{T}\mu(x_{t},y_{t})=0]$.
+Since $\prod_{t=1}^{T}\mu(x_{t},y_{t})\leq 1$ always, we have
+$\operatorname*{\mathbb{E}}_{j^{m}}\ \operatorname*{\mathbb{E}}_{\boldsymbol{x},\boldsymbol{y}}\ \left[\log\prod_{t=1}^{T}\mu(x_{t},y_{t})\right]=$ $\ \operatorname*{\mathbb{E}}_{j^{m}}\ \operatorname*{\mathbb{E}}_{\boldsymbol{x},\boldsymbol{y}}\ \left[\log\prod_{t=1}^{T}\mu(x_{t},y_{t})\>\Big{|}\>\prod_{t=1}^{T}\mu(x_{t},y_{t})\leq\varepsilon\right]\cdot\Pr\left[\prod_{t=1}^{T}\mu(x_{t},y_{t})\leq\varepsilon\right]$
+$+$ $\ \operatorname*{\mathbb{E}}_{j^{m}}\ \operatorname*{\mathbb{E}}_{\boldsymbol{x},\boldsymbol{y}}\ \left[\log\prod_{t=1}^{T}\mu(x_{t},y_{t})\>\Big{|}\>\prod_{t=1}^{T}\mu(x_{t},y_{t})>\varepsilon\right]\cdot\Pr\left[\prod_{t=1}^{T}\mu(x_{t},y_{t})>\varepsilon\right]$
+$\leq$ $\log\varepsilon\cdot\Pr\left[\prod_{t=1}^{T}\mu(x_{t},y_{t})\leq\varepsilon\right]+1\cdot\left(1-\Pr\left[\prod_{t=1}^{T}\mu(x_{t},y_{t})\leq\varepsilon\right]\right)$
+$\leq$ $\ 1-(1-\log\varepsilon)\Pr\left[\prod_{t=1}^{T}\mu(x_{t},y_{t})\leq\varepsilon\right]$
+Since $\varepsilon\in(0,1]$, we have $01-\log\varepsilon>$ 0. Also, $\Pr[\prod_{t=1}^{T}\mu(x_{t},y_{t})\leq\varepsilon]\geq\Pr[\prod_{t=1}^{T}\mu(x_{t},y_{t})=0]$, so
+$\operatorname*{\mathbb{E}}_{j^{m}}\ \operatorname*{\mathbb{E}}_{\boldsymbol{x},\boldsymbol{y}}\ \left[\log\prod_{t=1}^{T}\mu(x_{t},y_{t})\right]\leq$ $\ 1-(1-\log\varepsilon)\Pr\left[\prod_{t=1}^{T}\mu(x_{t},y_{t})\leq\varepsilon\right]$
+$\leq$ $\ 1-(1-\log\varepsilon)\Pr\left[\prod_{t=1}^{T}\mu(x_{t},y_{t})=0\right]$
+$\leq$ $\ 1-(1-\log\varepsilon)\alpha(T)$
 Since $\prod_{t=1}^{T}\mu^{m}(x_{t})=1$ always, we have
-$\sup_{\boldsymbol{\mu},\pi^{m}}\ \operatorname*{\mathbb{E}}_{\boldsymbol{x},\boldsymbol{y}}\ [R_{T}(\boldsymbol{x},\boldsymbol{y},\boldsymbol{\mu},\pi^{m})]\geq$ $\ \operatorname*{\mathbb{E}}_{j^{m}}\ \operatorname*{\mathbb{E}}_{\boldsymbol{x},\boldsymbol{y}}\ [R_{T}(\boldsymbol{x},\boldsymbol{y},(\mu,\dots,\mu),\pi^{m})]$
-$=$ $\ 1-\operatorname*{\mathbb{E}}_{j^{m}}\operatorname*{\mathbb{E}}_{\boldsymbol{x},\boldsymbol{y}}\left[\prod_{t=1}^{T}\mu(x_{t},y_{t})\right]$
-$\geq$ $\ \alpha(T)$
-Therefore $\lim_{T\to\infty}\sup_{\boldsymbol{\mu},\pi^{m}}\operatorname*{\mathbb{E}}[R_{T}]\geq\lim_{T\to\infty}\alpha(T)=1$, as required. ∎
+$\sup_{\boldsymbol{\mu},\pi^{m}}\ \operatorname*{\mathbb{E}}_{\boldsymbol{x},\boldsymbol{y}}\ [R_{T}^{+}(\boldsymbol{x},\boldsymbol{y},\boldsymbol{\mu},\pi^{m})]\geq$ $\ \operatorname*{\mathbb{E}}_{j^{m}}\ \operatorname*{\mathbb{E}}_{\boldsymbol{x},\boldsymbol{y}}\ [R_{T}^{+}(\boldsymbol{x},\boldsymbol{y},(\mu,\dots,\mu),\pi^{m})]$
+$=$ $\ \log 1-\operatorname*{\mathbb{E}}_{j^{m}}\operatorname*{\mathbb{E}}_{\boldsymbol{x},\boldsymbol{y}}\left[\log\prod_{t=1}^{T}\mu(x_{t},y_{t})\right]$
+$\geq$ $\ -1+(1-\log\varepsilon)\alpha(T)$
+Therefore
+$\lim_{T\to\infty}\sup_{\boldsymbol{\mu},\pi^{m}}\ \operatorname*{\mathbb{E}}_{\boldsymbol{x},\boldsymbol{y}}\ [R_{T}^{+}(\boldsymbol{x},\boldsymbol{y},\boldsymbol{\mu},\pi^{m})]\geq$ $\ -1+(1-\log\varepsilon)\lim_{T\to\infty}\alpha(T)$
+$\geq$ $\ -1+(1-\log\varepsilon)$
+$\geq$ $\ -\log\varepsilon$
+This holds for every $\varepsilon\in(0,1]$, which is only possible if $\lim_{T\to\infty}\sup_{\boldsymbol{\mu},\pi^{m}}\ \operatorname*{\mathbb{E}}_{\boldsymbol{x},\boldsymbol{y}}\ [R_{T}^{+}(\boldsymbol{x},\boldsymbol{y},\boldsymbol{\mu},\pi^{m})]=\infty$, as desired.
+∎

### v6

@@
 Barman et al.
-studied a multiplicative objective in a multi-armed bandit context, but their objective is the geometric mean of payoffs instead of the product.
+(2023) studied a multiplicative objective in a multi-armed bandit context, but their objective is the geometric mean of payoffs instead of the product.
 Interpreted in our context, their regret bounds imply that the average chance of catastrophe goes to zero, while we guarantee that the total chance of catastrophe goes to zero.
@@
 Formally, a distribution $\mathcal{D}$ over $\mathcal{X}$ is $\sigma$-smooth if for any $S\subseteq\mathcal{X}$, $\mathcal{D}(S)\leq\frac{1}{\sigma}U(S)$.
-(In the smoothed setting, we assume that $\mathcal{X}$ supports a uniform distribution $U$ [^5] For example, it suffices for $\mathcal{X}$ to have finite Lebesgue measure.
+(In the smoothed setting, we assume that $\mathcal{X}$ supports a uniform distribution UU.
+[^5] For example, it suffices for $\mathcal{X}$ to have finite Lebesgue measure.
 Note that this does not imply boundedness.
@@
 There also exists a special action $\tilde{y}$ which corresponds to querying the mentor.
-For $k\in\mathbb{N}$, let $[k]=\{1,2,\dots,k\}$.
+For $k\in\mathbb{N}$, let $k]=\{1,2,\dots,k\}$.
 On each time step $t\in[T]$, the agent must select action $y_{t}\in\ \mathcal{Y}\cup\{\tilde{y}\}$, which generates a payoff.
@@
 For brevity, we will usually write $Q_{T}=Q_{T}(\boldsymbol{\mu},\pi^{m})$.
-Local generalization.
-We assume that $\boldsymbol{\mu}$ and $\pi^{m}$ satisfy local generalization.
-Informally, if the agent is given an input $x$, taking the mentor action for a similar input $x^{\prime}$ is almost as good as taking the mentor action for $x$.
-Formally, we assume $\mathcal{X}\subseteq\mathbb{R}^{n}$ and there exists $L>0$ such that for all $x,x^{\prime}\in\mathcal{X}$ and $t\in[T]$, $|\mu_{t}^{m}(x)-\mu_{t}(x,\pi^{m}(x^{\prime}))|\leq L||x-x^{\prime}||$, where |||||$\cdot$||| | ⋅ | | denotes Euclidean distance.
-This represents the ability to transfer knowledge between similar inputs:
-$\big{|}\!\!\underbrace{\mu_{t}(x,\pi^{m}(x))}_{\text{Taking the right action}}-\!\underbrace{\mu_{t}(x,\pi^{m}(x^{\prime}))}_{\text{Using what you learned in $x^{\prime}$}}\!\!\!\!\!\!\!\big{|}\,\leq\>\underbrace{L||x-x^{\prime}||}_{\text{Input similarity}}$
+[…]
+[…]
 This ability seems fundamental to intelligence and is well-understood in psychology (e.g., Esser et al., 2023) and education (e.g., Hajian, 2019).
@@
 when all payoffs are strictly positive.
-To handle the case where some payoffs are zero, we assume the existence of $\mu_{0}^{m}>0$ such that $\mu^{m}(x_{t})\geq\mu_{0}^{m}$ always.
+To handle the case where some payoffs are zero, we assume the existence of $\mu_{0}^{m}>0$ such that $\mu_{t}^{m}(x_{t})\geq\mu_{0}^{m}$ always.
 Thus only the agent’s payoffs can be zero, so we can safely define $R_{T}^{\times}(\boldsymbol{x},\boldsymbol{y},\boldsymbol{\mu},\pi^{m})=\infty$ whenever $\mu_{t}(x_{t},y_{t})=0$ for some $t\in[T]$.
 We write $R_{T}^{\times}=R_{T}^{\times}(\boldsymbol{x},\boldsymbol{y},\boldsymbol{\mu},\pi^{m})$ for brevity.
-The assumption of $\mu^{m}(x_{t})\geq\mu_{0}^{m}>0$ means that the mentor cannot be abysmal.
+The assumption of $\mu_{t}^{m}(x_{t})\geq\mu_{0}^{m}>0$ means that the mentor cannot be abysmal.
 In fact, we argue that high-stakes AI applications should employ a mentor who is almost always safe, i.e., $\mu_{0}^{m}\approx 1$.
@@
 See Shalev-Shwartz & Ben-David (2014) for a comprehensive overview.
-Misc. The diameter of a set $S\subseteq\mathcal{X}$ is defined by $\operatorname*{diam}(S)=\max_{x,x^{\prime}\in S}||x-x^{\prime}||$. All logarithms and exponents are base $e$ unless otherwise noted.
-input $x$payoff $\mu$   $\mu(x,0)$$\mu(x,1)$$\cfrac{L}{2f(T)}$$\frac{1}{f(T)}$$\frac{2}{f(T)}$$\frac{3}{f(T)}$$\frac{f(T)-1}{f(T)}$$\dots$1001111
+[…]
+input xxpayoff $\mu$   $\mu(x,0)$$\mu(x,1)$$\cfrac{L}{2f(T)}$$\frac{1}{f(T)}$$\frac{2}{f(T)}$$\frac{3}{f(T)}$$\frac{f(T)-1}{f(T)}\dots$1011
 Figure 1: An illustration of the construction we use to prove Theorem 4.1 (not to scale).
@@
 Then $\mu_{f,\boldsymbol{a}}$ is defined by
-$\mu_{f,\boldsymbol{a}}(x,y)=\begin{cases}1&\textnormal{ if }y=a_{j(x)}\\ 1-L\left(\cfrac{1}{2f(T)}-|m_{j(x)}-x|\right)&\textnormal{ if }y\neq a_{j(x)}\\ \end{cases}$
+[…]
 We use this payoff function for all time steps: $\mu_{t}=\mu_{f,\boldsymbol{a}}$ for all $t\in[T]$.
@@
 What if we restrict ourselves to settings where standard online learning is possible?
-Specifically, we assume that $\pi^{m}$ belongs to a policy class $\Pi$ where either (1) $\Pi$ has finite VC dimension $d$ and $\boldsymbol{x}$ is $\sigma$-smooth or (2) $\Pi$ has finite Littlestone dimension $d$ [^7] Recall from Section 1.3 that standard online learning becomes tractable under either of these assumptions.
+Specifically, we assume that $\pi^{m}$ belongs to a policy class $\Pi$ where either (1) $\Pi$ has finite VC dimension $d$ and $\boldsymbol{x}$ is $\sigma$-smooth or (2) $\Pi$ has finite Littlestone dimension dd.
+[^7] Recall from Section 1.3 that standard online learning becomes tractable under either of these assumptions.
 This section presents a simple algorithm which guarantees subconstant regret (both multiplicative and additive) and sublinear queries under either of those assumptions.
@@
 5.2 Local generalization
-Local generalization is vital: this is what allows us to detect when an input is unfamiliar.
-Crucially, our algorithm does not need to know how inputs are encoded in $\mathbb{R}^{n}$ and does not need to know $L$: it only needs to be able to compute the nearest-neighbor distance $\min_{(x,y)\in S:y=\pi_{t}(x_{t})}||x_{t}-x||$.
-Thus we only need to assume that there exists some encoding satisfying local generalization.
+[…]
 To elaborate, recall the example that a 3 mm spot and a 3.1 mm spot on X-rays likely have similar risk levels (assuming similar density, location, etc.).
@@
 else
-     $\tilde{\Pi}\leftarrow$ any adversarial cover of size at most $(eT/d)^{d}$  (see Definition 5.4)
+  $\tilde{\Pi}\leftarrow$ any adversarial cover of $\Pi$ of size at most $(eT/d)^{d}$  (see Definition 5.4)
 $S\leftarrow\emptyset$
@@
 $\eta\leftarrow\max\big{(}\sqrt{\frac{p\log|\tilde{\Pi}|}{2T}},\>\frac{p^{2}}{\sqrt{2}}\big{)}$
-  for $t$ from $1$ to $T$ do
+for $t$ from 11 to $T$ do
   Run one step of Hedge, which selects policy $\pi_{t}$
-[…]
-[…]
+  with probability $p:$ $\texttt{hedgeQuery}\leftarrow\texttt{true}$
+  with probability 1p:1-p: $\texttt{hedgeQuery}\leftarrow\texttt{false}$
   if hedgeQuery [^9] The reader may notice that we do not update $S$ in this case.
@@
    Sample $\pi_{t}\sim P$
-     if $S=\emptyset$ or $\min_{(x,y)\in S:y=\pi_{t}(x_{t})}||x_{t}-x||>\varepsilon^{1/n}$ then
+[…]
    Ask for help if out-of-distribution
@@
 We showed that achieving subconstant regret in our problem (with the help of a mentor and local generalization) is no harder than achieving sublinear regret in standard online learning.
-Remaining technical questions.
-First, we have not resolved whether our problem is tractable for finite VC dimension and fully adversarial inputs (although Appendix D shows that the problem is tractable for at least some classes with finite VC but infinite Littlestone dimension).
-Second, the time complexity of Algorithm 1 currently stands at a hefty $\Omega(|\tilde{\Pi}|)$ per time step plus the time to compute $\tilde{\Pi}$.
-In the standard online learning setting, Block et al.
-(2022) and Haghtalab et al.
-(2022) show how to replace discretization approaches like ours with oracle-efficient approaches, where a small number of calls to an optimization oracle are made per round.
-We are optimistic about leveraging such techniques to obtain efficient algorithms in our setting.
+[…]
 Local generalization.
@@
 Proof.
-Recall the standard inequalities $11-\frac{1}{a}\leq\log a\leq a-$ 1 for any $a>0$.
+Recall the standard inequalities $1-\frac{1}{a}\leq\log a\leq a-1$ for any $a>0$.
 Part 1: $R_{T}^{+}\leq R_{T}^{\times}$.
@@
 Thus assume $\mu_{t}(x_{t},y_{t})>0$ for all $t\in[T]$.
-$R_{T}^{+}=$ $\ \sum_{t=1}^{T}\mu_{t}^{m}(x_{t})-\sum_{t=1}^{T}\mu_{t}(x_{t},y_{t})$ (Definition of $(\text{Definition of $R_{T}^{+}$})$
-[…]
-[…]
-[…]
-$=$ $\ R_{T}^{\times}$ (Definition of $(\text{Definition of $R_{T}^{\times}$})$
+$R_{T}^{+}=$ $\ \sum_{t=1}^{T}\mu_{t}^{m}(x_{t})-\sum_{t=1}^{T}\mu_{t}(x_{t},y_{t})$ $(\text{Definition of $R_{T}^{+}$})$
+$=$ $\ \sum_{t=1}^{T}\frac{\mu_{t}^{m}(x_{t})-\mu_{t}(x_{t},y_{t})}{\mu_{t}^{m}(x_{t})}$ $(\text{$\mu_{t}(x_{t},y_{t})\leq\mu_{t}^{m}(x_{t})$ and $0\leq\mu_{t}^{m}(x_{t})\leq 1$})$
+$\leq$ $\ \sum_{t=1}^{T}\log\left(\frac{\mu_{t}^{m}(x_{t})}{\mu_{t}(x_{t},y_{t})}\right)$ $\Big{(}\text{$1-\frac{1}{a}\leq\log a$ for any $a>0$}\Big{)}$
+$=$ $\ \log\prod_{t=1}^{T}\mu_{t}^{m}(x_{t})-\log\prod_{t=1}^{T}\mu_{t}(x_{t},y_{t})$ $(\text{Properties of logarithms})$
+$=$ $\ R_{T}^{\times}$ $(\text{Definition of $R_{T}^{\times}$})$
 Part 2: $R_{T}^{\times}\leq R_{T}^{+}/\min_{t\in[T]}\mu_{t}(x_{t},y_{t})$. We have
-$R_{T}^{\times}=$ $\ \log\prod_{t=1}^{T}\mu_{t}^{m}(x_{t})-\log\prod_{t=1}^{T}\mu_{t}(x_{t},y_{t})$ (Definition of RT× given $(\text{Definition of $R_{T}^{\times}$ given $\mu_{t}(x_{t},y_{t})>0\ \forall t\in[T]$})$
-[…]
-[…]
-[…]
-[…]
-$=$ $\ \frac{R_{T}^{+}}{\min_{t\in[T]}\mu_{t}(x_{t},y_{t})}$ (Definition of $(\text{Definition of $R_{T}^{+}$})$
+$R_{T}^{\times}=$ $\ \log\prod_{t=1}^{T}\mu_{t}^{m}(x_{t})-\log\prod_{t=1}^{T}\mu_{t}(x_{t},y_{t})$ $(\text{Definition of $R_{T}^{\times}$ given $\mu_{t}(x_{t},y_{t})>0\ \forall t\in[T]$})$
+$=$ $\ \sum_{t=1}^{T}\log\left(\frac{\mu_{t}^{m}(x_{t})}{\mu_{t}(x_{t},y_{t})}\right)$ $(\text{Properties of logarithms})$
+$\leq$ $\ \sum_{t=1}^{T}\left(\frac{\mu_{t}^{m}(x_{t})-\mu_{t}(x_{t},y_{t})}{\mu_{t}(x_{t},y_{t})}\right)$ $(\text{$\log a\leq a-1$ for any $a>0$})$
+$\leq$ $\ \sum_{t=1}^{T}\frac{\mu_{t}(x_{t},y_{t})}{\min_{i\in[T]}\mu_{i}(x_{i},y_{i})}\left(\frac{\mu_{t}^{m}(x_{t})-\mu_{t}(x_{t},y_{t})}{\mu_{t}(x_{t},y_{t})}\right)$ $(\text{$\mu_{t}(x_{t},y_{t})\geq\min_{i\in[T]}\mu_{i}(x_{i},y_{i})>0\ \forall t\in[T]$})$
+$=$ $\ \frac{1}{\min_{t\in[T]}\mu_{t}(x_{t},y_{t})}\sum_{t=1}^{T}(\mu_{t}^{m}(x_{t})-\mu_{t}(x_{t},y_{t}))$ $(\text{Arithmetic})$
+$=$ $\ \frac{R_{T}^{+}}{\min_{t\in[T]}\mu_{t}(x_{t},y_{t})}$ $(\text{Definition of $R_{T}^{+}$})$
 as claimed. ∎
@@
 binary variables, we have $\Pr[Z\leq\operatorname*{\mathbb{E}}[Z]/2]\leq\exp(-\operatorname*{\mathbb{E}}[Z]/8)$.
-Since $-\operatorname*{\mathbb{E}}[Z]\leq-W$ and $\exp$ is an increasing function, we have $\exp(-\operatorname*{\mathbb{E}}[Z]/8)\leq\exp(-W/8)$.
+Since $\operatorname*{\mathbb{E}}[Z]\leq-W$ and $\exp$ is an increasing function, we have $\exp(-\operatorname*{\mathbb{E}}[Z]/8)\leq\exp(-W/8)$.
 Also, $W/2\leq E[Z]/2$ implies $\Pr[Z\leq W/2]\leq\Pr[Z\leq\operatorname*{\mathbb{E}}[Z]/2]$.
@@
 Since $\mu_{f,\boldsymbol{a}}^{m}(x_{t})\geq\mu_{f,\boldsymbol{a}}(x_{t},y_{t})$ always, we can safely restrict ourselves to time steps $t\in M_{j}^{\prime}$ for some $j\in J_{\neg Q}^{\prime}$ and still obtain a lower bound:
-$R_{T}^{+}=$ $\ \sum_{t=1}^{T}(\mu_{f,\boldsymbol{a}}^{m}(x_{t})-\mu_{f,\boldsymbol{a}}(x_{t},y_{t}))$ (Definition of $(\text{Definition of $R_{T}^{+}$})$
+$R_{T}^{+}=$ $\ \sum_{t=1}^{T}(\mu_{f,\boldsymbol{a}}^{m}(x_{t})-\mu_{f,\boldsymbol{a}}(x_{t},y_{t}))$ $(\text{Definition of $R_{T}^{+}$})$
 $\geq$ $\ \sum_{j\in J_{\neg Q}^{\prime}}\sum_{t\in M_{j}^{\prime}}\big{(}\mu_{f,\boldsymbol{a}}^{m}(x_{t})-\mu_{f,\boldsymbol{a}}(x_{t},y_{t})\big{)}$ $(\mu_{f,\boldsymbol{a}}^{m}(x_{t})\geq\mu_{f,\boldsymbol{a}}(x_{t},y_{t}))$
 $\geq$ $\ \sum_{j\in J_{\neg Q}^{\prime}}\sum_{t\in M_{j}^{\prime}}(1-\mu_{f,\boldsymbol{a}}(x_{t},y_{t}))$ $(\mu_{f,\boldsymbol{a}}^{m}(x_{t})=1\text{ always})$
-[…]
-[…]
+$\geq$ $\ \sum_{j\in J_{\neg Q}^{\prime}}\sum_{t\in M_{j}^{\prime}}\left(1-1+\frac{L}{4f(T)}\right)$ $(\text{bound on $\mu_{f,\boldsymbol{a}}(x_{t},y_{t})$ for $t\in M_{j}^{\prime}$})$
+$=$ $\ \sum_{j\in J_{\neg Q}^{\prime}}\frac{L|M_{j}^{\prime}|}{4f(T)}$ $(\text{Simplifying inner sum})$
 Since $j\in J_{\neg Q}$, the mentor is not queried on any time step $t\in M_{j}$, so $y_{t}\in\{0,1\}$ for all $t\in M_{j}$.
@@
 […]
-[…]
-[…]
+$\geq$ $\ \big{(}1-\alpha_{f}(T)\big{)}\left(\frac{LT(f(T)-g(T))}{2^{7}f(T)^{2}}\right)$ $(\text{$\operatorname*{\mathbb{E}}[Q_{T}]\leq g(T)$})$
+$\geq$ $\ \big{(}1-\alpha_{f}(T)\big{)}\left(\frac{LT}{2^{8}f(T)}\right)$ $(\text{$g(T)\leq f(T)/2$})$
 Since $\lim_{T\to\infty}\alpha_{f}(T)=0$,
@@
 $=$ $\ 1\cdot\exp(-\infty)$
-[…]
+$=$ $\ 0$
 Since this upper bound holds for a randomly chosen $\mu_{f,\boldsymbol{a}},\pi^{m}$, the same upper bound holds for a worst-case choice of $\boldsymbol{\mu},\pi^{m}$ among $\boldsymbol{\mu},\pi^{m}$ which satisfy $\mu_{t}^{m}(x)=1$ for all $t\in[T],x\in\mathcal{X}$.
@@
   4. 4.
-    Their “actions” correspond to policies in our setting, not actions in $\mathcal{Y}$. Their number of actions $n$ corresponds to $|\tilde{\Pi}|$.
+[…]
   5. 5.
@@
 Algorithm 2 provides precise pseudocode for the HedgeWithQueries algorithm to which Lemma 5.4 refers.
-  function HedgeWithQueries$(p\in(0,1],\>\text{finite policy class }\tilde{\Pi},\>\text{unknown }\ell:[T]\times\tilde{\Pi}\to[0,1])$
+[…]
   $w(\pi)\leftarrow 1$ for all $\pi\in\tilde{\Pi}$
   $\eta\leftarrow\max\big{(}\sqrt{\frac{p\log|\tilde{\Pi}|}{2T}},\>\frac{p^{2}}{\sqrt{2}}\big{)}$
-     for $t$ from $1$ to $T$ do
-[…]
-[…]
+  for $t$ from 11 to $T$ do
+   with probability $p:$ $\texttt{hedgeQuery}\leftarrow\texttt{true}$
+   with probability 1p:1-p: $\texttt{hedgeQuery}\leftarrow\texttt{false}$
    if hedgeQuery then
@@
 Thus by Lemma 5.4,
-$\sum_{t=1}^{T}\operatorname*{\mathbb{E}}[\ell(t,\pi_{t})]-\min_{\tilde{\pi}\in\tilde{\Pi}}\sum_{t=1}^{T}\operatorname*{\mathbb{E}}[\ell(t,\tilde{\pi})]\leq$ $\ \frac{\log|\tilde{\Pi}|}{p^{2}}$
-$=$ $\ T\varepsilon\log|\tilde{\Pi}|$
+[…]
+[…]
 Since $|M_{T}|=\sum_{t=1}^{T}\mathbf{1}(\pi_{t}(x_{t})\neq\pi^{m}(x_{t}))=\sum_{t=1}^{T}\ell(t,\pi_{t})$, we have
@@
 Proof.
-Consider an arbitrary $t\in[T]$.
-If $t\in Q_{T}$, then $\mu_{t}(x_{t},y_{t})=\mu_{t}^{m}(x_{t})$ trivially, so assume $t\not\in Q_{T}$.
-Let $(x^{\prime},y^{\prime})=\operatorname*{arg\,min}_{(x,y)\in S_{t}:\pi_{t}(x_{t})=y}||x_{t}-x||$.
-Since $t\not\in Q_{T}$, we must have $||x_{t}-x^{\prime}||\leq\varepsilon^{1/n}$.
+[…]
 We have $y^{\prime}=\pi^{m}(x^{\prime})$ by construction of $S_{t}$ and $\pi_{t}(x_{t})=y^{\prime}$ by construction of $y^{\prime}$. Combining these with the local generalization assumption, we get
@@
 $=$ $\ \mu_{t}(x_{t},\pi^{m}(x^{\prime}))$
-$\geq$ $\ \mu_{t}^{m}(x_{t})-L||x_{t}-x^{\prime}||$
+[…]
 $\geq$ $\ \mu_{t}^{m}(x_{t})-L\varepsilon^{1/n}$
@@
 Let $(K,||\cdot||)$ be a normed vector space and let $\delta>0$.
-Then a multiset $S\subseteq K$ is a $\delta$-packing of $K$ if for all $a,b\in S$, $||a-b||>\delta$.
+Then a multiset $S\subseteq K$ is a $\delta$-packing of $K$ if for all $a,b\in S$, $a-b||>\delta$.
 The $\delta$-packing number of $K$, denoted $\mathcal{M}(K,||\cdot||,\delta)$, is the maximum cardinality of any $\delta$-packing of $K$.
@@
 Note that $|\hat{X}_{2}|=|\hat{Q}_{2}|$, since $\hat{X}_{2}$ is a multiset.
-We claim that $\hat{X}_{2}\cap X_{y}$ is a $\varepsilon^{1/n}$-packing of $X_{y}$.
-Suppose instead that there exists $x,x^{\prime}\in\hat{X}_{2}\cap X_{y}$, with $||x-x^{\prime}||\leq\varepsilon^{1/n}$.
+We claim that $\hat{X}_{2}\cap X_{y}$ is an $\varepsilon^{1/n}$-packing of $X_{y}$.
+Suppose instead that there exists $x,x^{\prime}\in\hat{X}_{2}\cap X_{y}$, with $x-x^{\prime}||\leq\varepsilon^{1/n}$.
 WLOG assume $x$ was queried after $x^{\prime}$ and let $t$ be the time step on which $x$ was queried.
-Then $(x^{\prime},\pi^{m}(x^{\prime}))\in S_{t}$.
+Since $x^{\prime}\in\hat{X}_{2}$, this implies $(x^{\prime},\pi^{m}(x^{\prime}))\in S_{t}$.
 Also, since $x,x^{\prime}\in\hat{X}_{2}$ we have $\pi_{t}(x_{t})=\pi^{m}(x_{t})=y=\pi^{m}(x^{\prime})$.
 Therefore
-$\min_{(x^{\prime\prime},y^{\prime\prime})\in S_{t}:y^{\prime\prime}=\pi_{t}(x_{t})}||x_{t}-x^{\prime\prime}||\leq||x_{t}-x^{\prime}||\leq\varepsilon^{1/n}$
-which contradicts $t\in\hat{Q}$. Thus $\hat{X}_{2}\cap X_{y}$ is a $\varepsilon^{1/n}$-packing of $X_{y}$.
+[…]
+which contradicts $t\in\hat{Q}$. Thus $\hat{X}_{2}\cap X_{y}$ is an $\varepsilon^{1/n}$-packing of $X_{y}$.
 By Lemma B.5, there exists a ball $B_{1}$ of radius $R:=\operatorname*{diam}(\boldsymbol{x})\sqrt{\frac{n}{2(n+1)}}$ which contains $\boldsymbol{x}$.
 Let $B_{2}$ be the ball with the same center as $B_{1}$ but with radius $\max(R,\varepsilon^{1/n})$.
-Since $X_{y}\subset\boldsymbol{x}\subset B_{1}\subset B_{2}$ and $\hat{X}_{2}\cap X_{y}$ is a $\varepsilon^{1/n}$-packing of $X_{y}$, $\hat{X}_{2}\cap X_{y}$ is also a $\varepsilon^{1/n}$-packing of $B_{2}$.
+Since $X_{y}\subset\boldsymbol{x}\subset B_{1}\subset B_{2}$ and $\hat{X}_{2}\cap X_{y}$ is an $\varepsilon^{1/n}$-packing of $X_{y}$, $\hat{X}_{2}\cap X_{y}$ is also an $\varepsilon^{1/n}$-packing of $B_{2}$.
 Also, $B_{2}$ must contain a ball of radius $\varepsilon^{1/n}$, so Lemma B.4 implies that
@@
 $\leq$ $\ O\left(\frac{\operatorname*{diam}(\boldsymbol{x})^{n}}{\varepsilon}+1\right)$
-(The $+1$ is necessary for now since $\operatorname*{diam}(\boldsymbol{x})$ could theoretically be zero.) Since $\hat{X}_{2}\subseteq\{x_{1},\dots,x_{T}\}\subseteq\cup_{y\in\mathcal{Y}}X_{y}$, we have $|\hat{X}_{2}|\leq\sum_{y\in\mathcal{Y}}|\hat{X}_{2}\cap X_{y}|$ by the union bound.
+(The +1+1 is necessary for now since $\operatorname*{diam}(\boldsymbol{x})$ could theoretically be zero.) Since $\hat{X}_{2}\subseteq\{x_{1},\dots,x_{T}\}\subseteq\cup_{y\in\mathcal{Y}}X_{y}$, we have $|\hat{X}_{2}|\leq\sum_{y\in\mathcal{Y}}|\hat{X}_{2}\cap X_{y}|$ by the union bound.
 Therefore
-$\operatorname*{\mathbb{E}}[|Q_{T}|]\leq$ $\ \sqrt{\frac{T}{\varepsilon}}+1+\operatorname*{\mathbb{E}}[|\hat{Q}|]$
-$=$ $\ \sqrt{\frac{T}{\varepsilon}}+1+\operatorname*{\mathbb{E}}[|\hat{Q}_{1}|]+\operatorname*{\mathbb{E}}[|\hat{Q}_{2}|]$
-$=$ $\ \sqrt{\frac{T}{\varepsilon}}+1+\operatorname*{\mathbb{E}}[|\hat{Q}_{1}|]+\operatorname*{\mathbb{E}}[|\hat{X}_{2}|]$
-$\leq$ $\ \sqrt{\frac{T}{\varepsilon}}+1+\operatorname*{\mathbb{E}}[|\hat{Q}_{1}|]+\operatorname*{\mathbb{E}}\left[\sum_{y\in\mathcal{Y}}|\hat{X}_{2}\cap X_{y}|\right]$
 […]
-$\leq$ $\ \sqrt{\frac{T}{\varepsilon}}+1+O\left(\frac{d}{\sigma}T\varepsilon\log(T+1/\varepsilon)\right)+|\mathcal{Y}|\cdot O\left(\frac{\operatorname*{\mathbb{E}}[\operatorname*{diam}(\boldsymbol{x})^{n}]}{\varepsilon}+1\right)$
+$=$ $\ \sqrt{\frac{T}{\varepsilon}}+\operatorname*{\mathbb{E}}[|\hat{Q}_{1}|]+\operatorname*{\mathbb{E}}[|\hat{Q}_{2}|]$
+$=$ $\ \sqrt{\frac{T}{\varepsilon}}+\operatorname*{\mathbb{E}}[|\hat{Q}_{1}|]+\operatorname*{\mathbb{E}}[|\hat{X}_{2}|]$
+$\leq$ $\ \sqrt{\frac{T}{\varepsilon}}+\operatorname*{\mathbb{E}}[|\hat{Q}_{1}|]+\operatorname*{\mathbb{E}}\left[\sum_{y\in\mathcal{Y}}|\hat{X}_{2}\cap X_{y}|\right]$
+$\ \leq$ $\ \sqrt{\frac{T}{\varepsilon}}+O\left(\frac{d}{\sigma}T\varepsilon\log(T+1/\varepsilon)\right)+\sum_{y\in\mathcal{Y}}O\left(\frac{\operatorname*{\mathbb{E}}[\operatorname*{diam}(\boldsymbol{x})^{n}]}{\varepsilon}+1\right)$
+$\leq$ $\ \sqrt{\frac{T}{\varepsilon}}+O\left(\frac{d}{\sigma}T\varepsilon\log(T+1/\varepsilon)\right)+|\mathcal{Y}|\cdot O\left(\frac{\operatorname*{\mathbb{E}}[\operatorname*{diam}(\boldsymbol{x})^{n}]}{\varepsilon}+1\right)$
 $\leq$ $\ O\left(\sqrt{\frac{T}{\varepsilon}}+\frac{d}{\sigma}T\varepsilon\log(T+1/\varepsilon)+\frac{\operatorname*{\mathbb{E}}[\operatorname*{diam}(\boldsymbol{x})^{n}]}{\varepsilon}\right)$
@@
 B.3 Adaptive adversaries
-If $s_{t}$ is allowed to depend on the events of prior time steps, we say that the adversary is adaptive.
+If $x_{t}$ is allowed to depend on the events of prior time steps, we say that the adversary is adaptive.
 In contrast, a non-adaptive or “oblivious” adversary must choose the entire input upfront.
@@
 In other words, the adversary gains no new information during the execution of the algorithm.
-For randomized algorithms, an adaptive adversary can base the choice of $s_{t}$ on the results of randomization on previous time steps (but not on the current time step), while an oblivious adversary cannot.
+For randomized algorithms, an adaptive adversary can base the choice of $x_{t}$ on the results of randomization on previous time steps (but not on the current time step), while an oblivious adversary cannot.
 In the standard online learning model, Hedge guarantees sublinear regret against both oblivious and adaptive adversaries (Chapter 5 of Slivkins et al.
@@
 In this section, we argue why the proof of Lemma 5.4 (Lemma 3.5 in their paper) goes through for adaptive adversaries as well.
-For this rest of Section B.3, lemma numbers refer to the numbering in Russo et al.
+For the rest of Section B.3, lemma numbers refer to the numbering in Russo et al.
 (2024).
@@
 (2024) allow two separate parameters $k$ and $\hat{k}$, which we unify for simplicity.
-Recall also that Lemma 3.5 refers to the variant of Hedge which queries with probability $p=\hat{k}/T=k/T$ independently on each time step (Algorithm 2.
+Recall also that Lemma 3.5 refers to the variant of Hedge which queries with probability $p=\hat{k}/T=k/T$ independently on each time step (Algorithm 2).
 More precisely, on each time step $t$, the algorithm samples $X_{t}\sim\text{Bernoulli}(p)$ and queries if $X_{t}=1$.
 The key idea is that $X_{t}$ is independent of events on previous time steps.
-Thus even conditioning on the history up to time $t$, for any for any random variable $Y_{t}$ we can write
+Thus even conditioning on the history up to time $t$, for any random variable $Y_{t}$ we can write
 $\operatorname*{\mathbb{E}}[Y_{t}]=(1-p)\operatorname*{\mathbb{E}}[Y_{t}\mid X_{t}=0]+p\operatorname*{\mathbb{E}}[Y_{t}\mid X_{t}=1]$
@@
 Applying Observation 3.3.
-The other tricky part of the proof is applying Observation 3.3 using a new loss function $\hat{\ell}$ defined by ^$\hat{\ell}_{t}(i)=\frac{T}{\hat{k}}(\ell_{t}(i)-\ell_{t}(i_{t}^{*}))\mathbf{1}(X_{t}=1)$.
+The other tricky part of the proof is applying Observation 3.3 using a new loss function $\hat{\ell}$ defined by $\hat{\ell}_{t}(i)=\frac{T}{\hat{k}}(\ell_{t}(i)-\ell_{t}(i_{t}^{*}))\mathbf{1}(X_{t}=1)$.
 To do so, we must argue that standard Hedge run on $\hat{\ell}$ is the “counterpart without queries” of HedgeWithQueries.
 Specifically, both algorithms must have the same weight vectors on every time step, and the only difference should be that HedgeWithQueries takes the optimal action on each time step independently with probability $p$ (and otherwise behaves the same as standard Hedge).
-On time steps with $X_{t}=0$, standard Hedge observes ^$\hat{\ell}_{t}(i)=0$ for all actions $i$ and thus makes no updates, and HedgeWithQueries makes no updates by definition.
+On time steps with $X_{t}=0$, standard Hedge observes $\hat{\ell}_{t}(i)=0$ for all actions $i$ and thus makes no updates, and HedgeWithQueries makes no updates by definition.
 On time steps with $X_{t}=1$, both algorithms perform the typical updates $w_{t+1}(i)=w_{t}(i)\cdot\exp(-\eta(\hat{\ell}_{t}(i)-\hat{\ell}_{t}(i_{t}^{*})))$.
@@
 For each action $y$, we will learn a binary classifier which predicts whether action $y$ is the mentor’s action.
-Formally, for each $y\in\mathcal{Y}$, define the policy class $\Pi_{y}=\{\pi_{y}:\pi\in\Pi\text{ and }\pi_{y}(x)=\mathbf{1}(\pi(x)=y))\ \forall x\in\mathcal{X}\}$.
+Formally, for each $y\in\mathcal{Y}$, define the policy class $\Pi_{y}=\{\pi_{y}:\pi\in\Pi\text{ and }\pi_{y}(x)=\mathbf{1}(\pi(x)=y)\ \forall x\in\mathcal{X}\}$.
 In words, for each policy $\pi:\mathcal{X}\to\mathcal{Y}$ in $\Pi$, there exists a policy $\pi_{y}:\mathcal{X}\to\{0,1\}$ in $\Pi_{y}$ such that $\pi_{y}(x)=\mathbf{1}(\pi(x)=y)$ for all $x\in\mathcal{X}$.
@@
   if $\Pi_{y}$ has VC dimension $d$ then
-        $\tilde{\Pi}_{y}\leftarrow$ any smooth $\varepsilon$-cover of $\Pi$ of size at most $(41/\varepsilon)^{d}$
-     else if $\Pi$ has Littlestone dimension $d$ then
-        $\tilde{\Pi}_{y}\leftarrow$ any adversarial cover of size at most $(eT/d)^{d}$
-  for $t$ from $1$ to $T$ do
+   $\tilde{\Pi}_{y}\leftarrow$ any smooth $\varepsilon$-cover of $\Pi_{y}$ of size at most $(41/\varepsilon)^{d}$
+  else if $\Pi_{y}$ has Littlestone dimension $d$ then
+   $\tilde{\Pi}_{y}\leftarrow$ any adversarial cover of $\Pi_{y}$ of size at most $(eT/d)^{d}$
+for $t$ from 11 to $T$ do
   for $y\in\mathcal{Y}$ do
@@
   3. 3.
-    Let $\pi^{my}:\mathcal{X}\to\{0,1\}$ be the policy defined by $\pi^{my}(x)=\mathbf{1}(\pi^{m}(x_{t})=y)$.
+    Let $\pi^{my}:\mathcal{X}\to\{0,1\}$ be the policy defined by $\pi^{my}(x)=\mathbf{1}(\pi^{m}(x)=y)$.
 Note that querying the mentor tells the agent $\pi^{m}(x_{t})$, which allows the agent to compute $\pi^{my}(x_{t})$: this is necessary when Algorithm 1 queries while running on some $\Pi_{y}$.
@@
 Proof.
-The argument is similar to the proof of Lemma B.2.
-If $\mu_{t}^{m}(x_{t})\neq\mu_{t}(x_{t},y_{t})$, then $y_{t}=y$ for some $y\in\mathcal{Y}$ where $b_{t}^{y}=1$.
-Therefore copy $y$ of Algorithm 1 did not query at time $t$ and $\pi_{t}^{y}(x_{t})=1$.
-Let $(x^{\prime},y^{\prime})=\operatorname*{arg\,min}_{(x,y)\in S_{t}^{y}:\pi_{t}^{y}(x_{t})=y}||x_{t}-x||$.
-Then $||x_{t}-x^{\prime}||\leq\varepsilon^{1/n}$ and $y^{\prime}=\pi_{t}^{y}(x_{t})=1$.
+[…]
 By construction of $S_{t}^{y}$, $y^{\prime}=\pi^{my}(x^{\prime})$ so $\pi^{my}(x^{\prime})=1$ which implies $\pi^{m}(x^{\prime})=y$. Then by the local generalization assumption,
-$\mu_{t}(x_{t},y_{t})=\mu_{t}(x_{t},y)=\mu_{t}(x_{t},\pi^{m}(x^{\prime}))\geq\mu_{t}^{m}(x_{t})-L||x_{t}-x^{\prime}||\geq\mu_{t}^{m}(x_{t})-L\varepsilon^{1/n}$
+[…]
 as required. ∎
@@
 Similar to the proof of Lemma B.3, we have
+$R_{T}^{+}=$ $\ \sum_{t\in M_{T}}(\mu_{t}^{m}(x_{t})-\mu_{t}(x_{t},y_{t}))$ $(\text{$\mu_{t}^{m}(x_{t})=\mu_{t}(x_{t},y_{t})$ for all $t\not\in M_{T}$})$
 […]
-[…]
-[…]
+$=$ $\ |M_{T}|L\varepsilon^{1/n}$ $(\text{Simplifying sum})$
 […]
@@
 $|Q_{T}|\leq$ $\ (\operatorname*{diam}(\boldsymbol{x})+4)g(T)$
-Choosing $g(T)=T^{c}$ for $c\in(1/2,1)$ is sufficient to subconstant regret and sublinear queries:
+Choosing $g(T)=T^{c}$ for $c\in(1/2,1)$ is sufficient for subconstant regret and sublinear queries:
 Theorem D.3.
@@
 $|Q_{T}|\in$ $\ O(T^{c}(\operatorname*{diam}(\boldsymbol{x})+1))$
-1:  function DBWRQ$(T\in\mathbb{N},\>g:\mathbb{N}\to\mathbb{N})$
+[…]
 2:  $X_{Q}\leftarrow\emptyset$  (previously queried inputs)
@@
 4:  $\mathcal{B}\leftarrow\emptyset$  (The set of active buckets)
-5:     for $t$ from $1$ to $T$ do
-6:        EvaluateInput$(x_{t})$
+5:  for $t$ from 11 to $T$ do
+[…]
 7:  end for
 8:end function
-9:  function EvaluateInput$(x\in\mathcal{X})$
+[…]
 10:  if $x\not\in B$ for all $B\in\mathcal{B}$ then
@@
 13:   $n_{B}\leftarrow 0$
-14:        EvaluateInput$(x)$
+14:   EvaluateInput(x)(x)
 15:  else
@@
 30:    $\mathcal{B}\leftarrow\mathcal{B}\cup\{B_{1},B_{2}\}\setminus B$
-31:           EvaluateInput$(x)$
+31:    EvaluateInput(x)(x)
 32:   end if
@@
   6. 6.
-    Let $\mathcal{B}_{V}^{\prime}=\{B\in\mathcal{B}_{V}:\text{no descendant of $B$ is in $\mathcal{B}_{V}$}\}$.
+[…]
 D.3 Proof roadmap
@@
 There are two ways to create a bucket: from scratch (line 11), or by splitting an existing bucket (line 28).
-Since depth 0 buckets overlap only at their boundaries, and each depth 0 bucket has length $1/g(T)$, at most $g(T)\max_{t,t^{\prime}\in[T]}|x_{t}-x_{t^{\prime}}|=g(T)\operatorname*{diam}(\boldsymbol{x})$ depth 0 buckets are subsets of the interval $[\min_{t\in[T]}x_{t},\max_{t\in[T]}x_{t}]$.
+Since depth 0 buckets overlap only at their boundaries, and each depth 0 bucket has length $1/g(T)$, at most $g(T)\max_{t,t^{\prime}\in[T]}|x_{t}-x_{t^{\prime}}|=g(T)\operatorname*{diam}(\boldsymbol{x})$ depth 0 buckets are subsets of the interval $\min_{t\in[T]}x_{t},\max_{t\in[T]}x_{t}]$.
 At most two depth 0 buckets are not subsets of that interval (one at each end), so the total number of depth 0 buckets is at most $g(T)\operatorname*{diam}(\boldsymbol{x})+2$.
@@
 Proof.
-If we query at time $t$, then $\mu_{t}(x_{t},y_{t})=\mu_{t}^{m}(x_{t})$.
-Thus assume we do not query at time $t$: then there exists $x^{\prime}\in B_{t}$ (as defined on line 23 of Algorithm 4) such that $y_{t}=\pi(x^{\prime})=\pi^{m}(x^{\prime})$.
-Since $x_{t}$ and $x^{\prime}$ are both in $B_{t}$, $|x_{t}-x^{\prime}|\leq\operatorname*{len}(B_{t})$.
-Then by local generalization, $\mu_{t}(x_{t},y_{t})=\mu_{t}(x_{t},\pi^{m}(x^{\prime}))\geq\mu_{t}^{m}(x_{t})-L||x_{t}-x^{\prime}||\geq\mu_{t}^{m}(x_{t})-L\operatorname*{len}(B_{t})$.
-∎
+[…]
 Lemma D.3.
@@
 Now consider some $B^{\prime}\in\mathcal{B}_{V}^{\prime}$ with $B\neq B^{\prime}$.
-We $|B\cap B^{\prime}|\leq 1$: otherwise one must be the descendant of the other, which contradicts the definition of $\mathcal{B}_{V}^{\prime}$.
+We have $|B\cap B^{\prime}|\leq 1$: otherwise one must be the descendant of the other, which contradicts the definition of $\mathcal{B}_{V}^{\prime}$.
 Suppose $B^{\prime}$ also intersects both $X_{j}$ and $X_{j+1}$: since $B^{\prime}$ is also an interval, $B^{\prime}$ must also contain an open neighborhood centered on the boundary between those two segments.
 But then $|B\cap B^{\prime}|>1$, which is a contradiction.
-Therefore any pair of adjacent segments $X_{j}$ and $X_{j+1}$, there is at most one bucket in $\mathcal{B}_{V}^{\prime}$ which contains an open neighborhood around their boundary.
-Since there are at most $K-1$ pairs of adjacent segments, we have $|\mathcal{B}_{V}^{\prime}|\leq K-1\leq K$.∎
+Therefore for any pair of adjacent segments $X_{j}$ and $X_{j+1}$, there is at most one bucket in $\mathcal{B}_{V}^{\prime}$ which contains an open neighborhood around their boundary.
+Since there are at most K1K-1 pairs of adjacent segments, we have $|\mathcal{B}_{V}^{\prime}|\leq K-1\leq K$.∎
 Lemma D.4.
@@
 We have
-[…]
+$R_{T}^{+}\leq$ $\ \sum_{t\in M_{T}}(\mu_{t}^{m}(x_{t})-\mu_{t}(x_{t},y_{t}))$ $(\text{$\mu_{t}(x_{t},y_{t})\geq\mu_{t}^{m}(x_{t})$ for $t\not\in M_{T}$})$
 […]
@@
 $\geq$ $\ \mu_{0}^{m}/2$
-[…]
+$>$ $\ 0$
 Invoking Lemma A.1 completes the proof:
@@
 Proposition E.1.
-Assume that $\mu$ satisfies Lipschitz continuity: for all $x,x^{\prime}\in\mathcal{X}$ and $y\in\mathcal{Y}$, $|\mu(x,y)-\mu(x^{\prime},y)|\leq L||x-x^{\prime}||$.
-Also assume that $\mu(x,\pi^{m}(x))=\max_{y\in\mathcal{Y}}\mu(x,y)$ for all $x\in\mathcal{X}$.
-Then $\mu$ satisfies local generalization with constant $2L$.
+[…]
 Proof.
 For any $x,x^{\prime}\in\mathcal{X}$, we have
-$\mu(x,\pi^{m}(x^{\prime}))\geq$ $\ \mu(x^{\prime},\pi^{m}(x^{\prime}))-L||x-x^{\prime}||$ (Lipschitz continuity of $\mu$)
-$\geq$ $\ \mu(x^{\prime},\pi^{m}(x))-L||x-x^{\prime}||$ ($\pi^{m}$ is optimal for $x^{\prime}$)
-$\geq$ $\ \mu(x,\pi^{m}(x))-2L||x-x^{\prime}||$ (Lipschitz continuity of $\mu$ again)
-$=$ $\ \mu^{m}(x)-2L||x-x^{\prime}||$ (Definition of $\mu^{m}(x)$)
+[…]
+[…]
+[…]
+[…]
 Since $\pi^{m}$ is optimal for $x$, we have
-$\mu^{m}(x)+2L||x-x^{\prime}||\geq\mu^{m}(x)\geq\mu(x,\pi^{m}(x^{\prime}))$
-So $-2L||x-x^{\prime}||\leq\mu(x,\pi^{m}(x^{\prime}))-\mu^{m}(x)\leq 2L||x-x^{\prime}||$ and therefore $|\mu(x,\pi^{m}(x^{\prime}))-\mu^{m}(x)|\leq 2L||x-x^{\prime}||$.∎
+[…]
+[…]
 Theorem E.2 shows that avoiding catastrophe is impossible without local generalization, even when $\boldsymbol{x}$ is $\sigma$-smooth and $\Pi$ has finite VC dimension.
@@
 Define $f(T):=\lceil\sqrt{g(T)T}\rceil$; by Lemma A.2, $g(T)\in o(f(T))$ and $f(T)\in o(T)$.
-Divide $\mathcal{X}$ into $f(T)$ equally sized sections $X_{1},\dots,X_{f(T)}$ in the exactly the same way as in Section 4.2; see also Figure 1.
+Divide $\mathcal{X}$ into $f(T)$ equally sized sections $X_{1},\dots,X_{f(T)}$ in exactly the same way as in Section 4.2; see also Figure 1.
 Assume that each $x_{t}$ is in exactly one section: this assumption holds with probability 1, so it does not affect the regret.
@@
 Queries in sections other than $j^{m}$ provide no information about the value of $j^{m}$, so $j^{m}$ is uniformly distributed across the set of sections not yet queried, i.e., $\{j\in[f(T)]:j_{r}\neq j\ \forall r<i\}$.
-There are at least $f(T)-i+1$ such sections, since there are $i-1$ prior queries at this point.
+There are at least f(T)i+1f(T)-i+1 such sections, since there are i1i-1 prior queries at this point.
 Thus $\Pr[j_{i}\neq j^{m}\mid j_{r}\neq j^{m}\ \forall r<i]\geq\frac{f(T)-i}{f(T)-i+1}$ (the inequality is because this probability could also be 1 if $j_{i}=j_{r}$ for some $i<r$).
@@
 $\leq$ $\ 1-(1-\log\varepsilon)\Pr\left[\prod_{t=1}^{T}\mu(x_{t},y_{t})\leq\varepsilon\right]$
-Since $\varepsilon\in(0,1]$, we have $01-\log\varepsilon>$ 0. Also, $\Pr[\prod_{t=1}^{T}\mu(x_{t},y_{t})\leq\varepsilon]\geq\Pr[\prod_{t=1}^{T}\mu(x_{t},y_{t})=0]$, so
+Since $\varepsilon\in(0,1]$, we have $1-\log\varepsilon>0$. Also, $\Pr[\prod_{t=1}^{T}\mu(x_{t},y_{t})\leq\varepsilon]\geq\Pr[\prod_{t=1}^{T}\mu(x_{t},y_{t})=0]$, so
 $\operatorname*{\mathbb{E}}_{j^{m}}\ \operatorname*{\mathbb{E}}_{\boldsymbol{x},\boldsymbol{y}}\ \left[\log\prod_{t=1}^{T}\mu(x_{t},y_{t})\right]\leq$ $\ 1-(1-\log\varepsilon)\Pr\left[\prod_{t=1}^{T}\mu(x_{t},y_{t})\leq\varepsilon\right]$